A Complete Bibliography of Publications in Limnology and Oceanography: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.05

Title word cross-reference

$102.65 [Edm82c]. $105.00 [Wan82]. $11.25 [Smi82a]. $110.00 [Lon85]. $120.00 [Mar83]. $137.00 [Hec86]. 15 [LK89]. $15.00 [Hal82a]. $17.95 [Thi83]. $18.00 [Fen84]. $189.50 [Pom89]. $19.95 [Wel88]. $20.00 [Pom80]. $21.55 [Thi83]. $21.95 [Mac80]. $22.95 [Smi82b]. $24.95 [Lon86]. $25.00 [Jah83]. $27.00 [Dug80]. $27.50 [Gar80, Gil80, Jum82, Pim83]. $28.50 [Pai84b]. $29.50 [Lis85, Smi82a]. $29.75 [Kru82]. $29.95 [Mor80b, Smi88]. $31.95 [Fre84, Mur83]. $32.50 [Lun84, Tut80, ZoB80]. $34.00 [Man83]. $35.00 [Den82a, EA84, Rzó81b]. $35.95 [Pai84a]. $36.00 [Fen84]. $36.95 [Mil84b]. $37.50 [Pic82, Pon82]. $39.80 [Ser80]. $39.95 [Ill80]. $42.50 [Now84]. $44.50 [Ril82]. $45.00 [Ano86j, Hrb82]. $46.00 [Bro82]. $46.95 [Cha85]. $47.50 [Koh82]. $48.75 [Kru80]. $49.00 [WK80]. $49.50 [Csa80, Hec84, Pra80, Smi86b]. $49.95 [Whi86, Bro80, Lec85, Whi86]. $55.00 [Bur86a, Mur84b, New86]. $59.50 [Edm82a]. $59.95 [Cam87, Pla86]. $64.00 [Bro83]. $65.00 [HG88]. $67.50 [Dav86]. $68.45 [Fry80]. $68.95 [Pic85]. $69.50 [Mac85, Wet82]. $69.95 [Lon84, Mil82, Mor88a]. $71.95 [Dav83]. $74.50 [Epp82]. $78.95 [Woo80]. $79.00 [Edm82d, Hyn81, Wet84]. $79.50 [MS84]. $85.00 [Ban84b]. $87.00 [Wet83]. $87.50 [Cat85]. $88.00 [Shi85]. $89.50 [Woo85]. $9.50 [Jah83]. $95.00 [Hor83]. $98.00 [McM86]. ^{0} [TAJ89]. ^{13} [HHH87, OSML87, POHMD84, ZFFC88]. ^{137} [MH86, SGMB87]. 14 [Col84a, DL83a, FH84, HD86, HS82, LDYY83, NO89, POHMD84, RJB81, RS81, SB80, TS85, TCN^{+}82]. ^{15} [DW86b, Gar84, GG84, GLMA85, KP82, Law84, OSML87]. ^{18} [POHMD84]. 2+ [KH80]. ^{210} [BH88a, BH88b, BB86, Mur85, SGMB87]. ^{222} [CA87, DQRM87]. ^{226} [BH88b]. ^{228} [TCN^{+}82]. ^{234} [CB85, CB87]. ^{238} [CB85, CB87]. ^{3} [RBNF87, RW89, THCP81]. ^{32} [AS87]. ^{35} [Kin83, TAJ89]. ^{55} [Kin83]. ^{7} [SGMB87]. _{18} [SH87a]. _{2} [DL83a, DL85b, Fra88, KM84, LDYY83, Ore83a, Pae85, PCP87, PM88, RJBC83, RAN88, RS88, SN85, TNEJ84, TAJ89, Unc81]. _{3} [HS82, LSK^{+}83, MD87, RJB81, WBEJ85]. _{4} [McK81]. _{4}^{+} [BM82, Law85]. β [Chr89, KGGW88]. : [AS83, CB85, CB87, Gro89a, PS89, SB89c]. G [WB84].

-carotene [KGGW88]. -Fe [Lyl83]. -glucosidase [Chr89].

/H [TNEJ84].

0 [SN85]. 0.1-1.0 [SM82a]. 0Mountain [Hyn81].

1 [Ano80a, Ano82a, Ano85a]. 1/2 [Mur84b]. 15^{o} [HF84, HFL84]. 1978 [Ano80a, Hrb82]. 1979 [Fen82]. 1980 [Ano82a, Pic82]. 1982 [Ano81e, Edm82e]. 1983 [Ano85a]. 1985 [Edm86b]. 1986 [BS89].

205 [Rzó81a]. 210 [Thi83]. 22.00 [Pai80]. 28 [WK80]. 286 [Man80]. 2nd [Bur86a, Edm81, Edm82d, Gil80].

36 [Fry80]. 3rd [Pai80, Pic82].

42nd [Ano80e]. 43rd [Ano81f]. 44th [Ano82h]. 45th [Ano83f]. 46th [Ano84h]. 47th [Ano85d].

5 [FW82]. 5’-monophosphate [FW82].

7 [Jum84, Ste84]. 7.5-m [LW81]. 79-E [MN86].

81-D [MN86]. ’89 [Ano88e, Ano88f, Ano89g].

93 [Mur84b].

Abarenicola [Tag88]. abilities [CK84a]. ability [SvdBR82]. Above [SH84, Sta80]. Above- [SH84]. Abrolhos [Smi81b]. absolute [BK82]. absorbance [SW80, SW81]. absorbed [KBO84]. Absorption [BMP81, DCV87, LWP85, WB86a, Ban88, HM87, IMK88, IS89, KS82a, KM83, LUP88, LE89, LD86, MH87, MCK89, PP89, PS81, RPC89, SAK89, SLP87, Til83, Vos89]. abundance [MP88, MH85, Pas80, SL89, Sul80, Two83, VL87]. abundant [SMR84]. Abyssal [SJC^{+}87]. abyssorum [GR84]. ac [CV82]. Academic [Bro82, Den82a, Fen82, Fen83, Ril82]. Academy [Lis85]. acantharian [Bra80]. Acanthaster [Ols87]. Acartia [DDSmV83, KM87, Rom84, Ros80]. Acceleration [ZKD87]. Access [BBH85]. acclimation [LHF^{+}84, SC81a]. according [KA89]. Accounting [Bay88]. accretion [HDP83]. accumulate [Whi81]. Accumulation [FTKB87, LL82, BZH^{+}85, Bru89b, ET80, Hal82b, HW85, HF84, Nür87, RKF86, WK86b]. Accuracy [MMR87, Ami88, Oud89]. accurate [WSU^{+}83]. acetate [MJSM89]. acetone [LD86]. Acid [Kil82, RWB87, BF84, DSH82, HHH87, IDB82, KH86a, RM81, WK86b]. acid-hydrolyzable [IDB82]. acidic [Col83]. Acidification [Mor84, DGN82, KRCS82, KRFS84]. acidified [CKST86, RKS^{+}86]. acids [CSHO89, Cof89, FS84, Fra87, FDK^{+}86, GCLC87, HFL84, Jør87, KW86, KB89b, ML82, RWB87, RKS^{+}86]. Acilius [AMP81]. Acoustic [Cha85]. acquire [CK84a]. across [CCW87, Pac86, RS86c, WAI89]. ACT [Koh84, Han84]. action [LWP85, LUP88]. active [CV82, Ger83a, Hun80, PB87]. activities [SF82, SRK86]. Activity [LB84, AFM86, BBG89, Chr89, Col84a, CS83, Cuk83, Duc87, FM82, GS86, HS80, HFFS85, KB89a, LMS87, Luc86, MB86, May89, PG83, SB83, SA85, VN87, War84, ZPVM83]. Acute [NDS86]. acutus [NO89]. adaptation [Dri81, GPHI83, SJR87]. adaptations [Nic89]. addition [WRS80]. additions [BKF^{+}87, FS84, HDLP85]. Adenine [BR89, FDKM86, FDK^{+}86, KW86]. adenosine [FW82]. adhesion [Jac89]. Adirondack [DGN82, VL89]. adjacent [Han81]. adjoining [WMSW89]. Adsorption [KB80a, BCY89, BM82, Ger83a, HASI80, MA84, ZoB80]. adult [DDSmV83, WBF85]. Adv [Thi83]. Advances [Ril82, Bro82]. advantage [TLH89]. Advantages [LH87]. advection [DG83]. Advective [GMK80, MBT80]. aerobic [HH80]. aerosols [Ger83b]. aeruginosa [FP87, HA81, NDS86, PTB83]. Affairs [Thi83]. affect [CSPV86, RJ84]. affected [WBV80]. affecting [Eck83, Leh88a, WBG^{+}87]. affects [RM81]. affinis [LE89]. Affinity [But86]. Africa [Rud80, Kli88, LHM84]. African [Hec84, KKH86, MM82a]. after [PA83]. agardhii [ZBM82]. ages [BT84, TH80, THCP81]. aggregate [Gar81a, JC80]. aggregates [ACC86, LH87]. Aggregation [PD80]. AGU [Ano81e]. AGU/ASLO [Ano81e]. aid [Ree81]. AIMS [Koh84]. Air [Csa80, DP81, DP86, Fra88, Mac85]. Air-sea [Csa80, Fra88, Mac85]. airborne [BSW81, Lew83a]. aircraft [HA85]. al [Boo80]. al. [Ami88, CP89, KW86]. Alan [Shi85]. Alaska [Ano88f, Ano89g, LBZC88, Mur84a]. albumin [ZST89]. alga [NO89, STC88, Str83, SH85]. algae [Ban82, CK79, COL84b, CK84a, DeM88, Dri81, FB80, GG81, Gro89b, Han88, IL85, Kan83, KB82, KC85, Leh80, LS85, MM80, MB89a, Pae89, SCH89, SMK81, Sta86, WMM83, Lun84]. Algal [ADK87, BK87, KDC86, AK89, Bel80, BS80b, Bel83, Bor87, CP86, CW88, Car85, Cuk83, EEMC88, EDKS86, FP87, Gro89a, HF81, KB89a, KH86b, KJP89, MB89b, PU82, PTB83, PM84a, Pri89, RS81, Smi82c, SF82, VP85, VG80]. algal-bacterial [Bel83]. Alkaline [Rue83, Har87]. Alkalinity [EH88, LSK^{+}83, Pse88, Sto87, UB89, CS88, ES83, GB80, Kel88, KM84, Sch88]. alkalization [Kil82]. Allanson [Fry80]. allocation [HH86, SCH89, THGD83]. allometric [MF89]. Alona [Fre88]. alone [Sch80]. along [PGB88, TML81]. aloricate [SSFN86]. alpine [GR84, Sto87]. alteration [Now84]. altered [Pet86c]. alternatives [Joh89]. alterniflora [LHH80]. altitude [Kir84, VWVR84]. aluminum [HDLP85]. Amazon [DQRM87, EHD^{+}86, FDR^{+}88, HCQ^{+}86, RDW^{+}88, SM84]. ambigua [HG85a]. America [Mac80, SIH85]. American [Ano80e, Ano81f, Ano82h, Ano83f, Ano84h, Ano85d, Ano87h, Hrb82, Ano80a, Ano81a, Ano81b, Ano81c, Ano82a, Ano82b, Ano83a, Ano83b, Ano84a, Ano84b, Ano85a, Ano86a, Ano86b, Ano86l, Ano87a, Ano88a, Ano89b, Ano89c, Ano89a, Ano89m, BR86, GC81, Koo85]. americium [FBF83]. amino [Cof89, FS84, GCLC87, HHH87, HFL84, Jør87, KH86a, ML82, RWB87]. Aminot [Oud89]. ammonia [BBP82, Gar81b]. Ammonium [KH88a, MA84, SH88, WGM82, BSS^{+}84, Bow84, Brz88, CMR89, DC88b, GLMA82, GG82, GCD87, KB80b, ML82, Ols80a, Smi82d, TLH89, War84, WM82, WKLK89, Wil84, ZFFC88]. ammonium-oxidizing [War84]. among [KKH86, LS85, SK83b, SK85a]. amperometric [ZPVM83]. amphipod [LR85, WH80]. amphipods [LE89]. amplitude [CM82]. Amsterdam [Kru80, Pom89, Wan82]. Anabaena [OW84, SS88b]. Anaerobic [IJ85, Ben89, Bez89, CK88, Dev83, IB82a, IOK87, SSS87, SK82]. anaerobiosis [ATA87]. analogue [WM82]. Analyses [BS85b, CS81, HT85a, POHMD84]. Analysis [Fro80, BK82, Bur80, CCF89, DPF87, FN85, GK83, Ham81, Iba81, IMK88, JP83, KSP^{+}81, KMSF81, LGV85, MMH85, ML82, ML89, Ols80b, Pac86, PD84, PSGD86, RS88, SAJB86, SL87a, SA81, SH89, Smi82c, SK82, YHM^{+}83, YJ84, ZKD87]. Analytic [Law85]. analytical [PP89]. analyzed [BP87, GSBP87]. analyzer [RS89]. Analyzing [GWB89, Tay88]. Ancellix [Fow80]. anchovy [SE82]. ancient [Tut80]. Anderson [Sto89]. Andes [HLM86]. anemones [ZST89]. angle [Gor89b]. angular [SP89a]. animal [Car89, Luc86, MMM81, Now84]. animal-particle [Car89]. Animal-sediment [Now84]. animals [Doy80, OSML87, Wot82]. Ann [Ban84a]. Announcement [Ano81b, Ano82e, Ano82g, Ano84e, Ano84c, Ano84d, Ano86h, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano87b, Ano87c, Ano87d, Ano88d, Ano88b, Ano88c, Ano89d, Ano89e, Hob85, Ano82i, Ano88l]. Announcements [Ano80b, Ano83c, Ano83d, Ano85b, Ano87e, Ano89f]. Annual [Ano80e, Ano81f, Ano82h, Ano83f, Ano84h, Ano85d, Ano88m, Ano89n, PAGG87, WH80, KTM83, LCK89, PO80, RD89]. annularis [DL83b]. Anomalous [SP89b]. anoxic [CM83, DH80, DWR82, HL85, KB80a, KB80b, Lid83, MJSM89, Nür84, Nür87, SMSM86, SNL84, TNEJ84]. Antarctic [Lon86, Top87, WM87, Ber89, HHOC89, Ols80a, OBP87, PPVHW87, PP85, SM80, Som89a, WMSP86]. Antarctica [Hor81, PS85, RV87]. antennae [Lan80]. antimony [Kan83]. Antonio [Ano81e]. apex [Gar81b]. Aphanizomenon [HS84b]. Aphanizomenonflosaquae [HGS83]. Apparent [Gar84, JKMN87, Kir84]. appendicularian [All81]. appendix [Ban84b]. applicability [KR88]. Application [JD86a, New86, VWVR84, BW89, BH88a, BB86, CV82, Den82b, HW86, RF88, Iba81]. Applications [Cha85, Fow80, MHB84, MB89b]. applied [Gor89a, GBF86]. approach [Bel80, Kel89, MFHM80, MTL84, Pet88, RR80, TML81, WSU^{+}83, vdLAH^{+}84, Fen84, Pai84b, Pra80]. Approaches [Joh89, PG81]. April [Woo84]. aquae [HS84b, OW84]. Aquat [Rzó81a, Thi83]. Aquatic [LT88, Ano88k, Bro82, Bro83, CK79, CK80, EA84, FH84, Fly88, GFCP85, GG81, HG88, HB84, HF88, Jac89, KB89b, Koh82, Lam87, Lan82, Lec85, Mac80, MS87, Pae85, PF80, RW89, SP89a, TS85, Wil80b, YHM^{+}83, Mor88a]. aqueous [AM82]. Arafura [VN87]. Archipelago [POHMD84]. Arco [LS83]. Arctic [POHMD84, But82, CM81, Cuk83, GPHI83, Har83b, LO83, MRLD86, SCH89, WRS80]. Arctica [TCN^{+}82]. area [And85, BHK^{+}82, RE81]. areal [CR80]. areas [MMP86]. arguments [MJN84]. arising [Mor88b]. Arkansas [JKMN87]. arsenic [HASI80]. art [Mol89]. artifact [HM80]. Artificial [BHK^{+}82, CK79, DeM88, GG81, San88]. Ascophyllum [Str83]. ashfall [Abe88]. ASI [Mac85]. Asiatic [CDPC84]. ASLO [Ano81e, Ano88e, Ano88f, Ano89g]. ASLO-’89 [Ano88e, Ano88f, Ano89g]. asphyxiation [vdLAH^{+}84]. Asplanchna [GS84]. Asplanchna-induced [GS84]. assay [HS82, Kri83]. assemblage [Gro89a]. assemblages [BCL85, DSH82, HFA80, Moo88, SGSS84, Whi89]. assess [JD86b]. assessing [MH86]. Assessment [ZBM82, Bur86a, Lis85, Man80, NG83]. assimilated [MMTL88]. Assimilation [TS85, AGG82, Col82, Col80b, HGS83, Jør87, LHF^{+}84, NO89, PS85, PGO82, SM80, UEB88]. Assoc [Smi82b, Tut80]. associated [KC85, KDR89, MS85b, SP85, UB89]. associates [Har84]. Association [FH82]. associations [SJR87, SCR84, TQS86]. Asterionella [HA81]. asteroid [OBP87]. Asymmetrical [SW81]. asymptotic [Zan89]. ation [SGSS84]. Atlantic [BW86, GA81, McM86, SSS87, SC81b, BS89, CW83, CJTL83, Fis80, GK83, GMFLJ85, Han87, Le 82, LUP88, MH85, SHRW84, Wro89]. atmosphere [BK80, BK84, KA89]. atmosphere-ocean [KA89]. atmospheric [BCM86, DL85b, GPS85, HACF85, Kor80]. atoll [Smi82b, ASS81, SMK81]. ATP [Col84a, HL81, PME87]. attached [Kir83, MK84, Sim87, TS89]. attenuation [Gor89a, Sch85, SYB^{+}89]. attraction [Jac89]. attributable [Pai86]. August [Hrb82]. Auke [LBZC88]. Australia [Koh84, HH81, KM87, SA83]. Australian [Koh84, BB81]. Authors [Ano80c, Ano81d, Ano82c, Ano83e, Ano84f, Ano85c, Ano86i, Ano87f, Ano88g, Ano89h]. autocorrelation [LT88, Mac84]. automated [FNL85, GCAN87]. autoradiography [DM84, SK83a, War84]. Autotrophic [CCO83, Ano89i, Gop89]. Availability [TH80, WST80, HF88, LS83, PCP87, Pet81, Run85, Som84]. available [BF87, CA89, RO82]. avirostris [TTF88]. avoidance [Dod88, Fie82]. Award [Ano88n]. Awards [Ano88e, Ano88m, Ano89n].

B [Ano89i, Ben89, Bez89, Den82a, Edm87, Fen83, Fry80, Gop89, Har89b, Jum82, Koh84, Man80, Pon82, Ril82, Wet82, Le 82, SB89c]. Bacillariophyceae [SP85]. backswimmer [SM83]. bacteria [ACC86, BSW81, CFPW89, CLS82, Col80a, CCF89, CK84a, DM83, DH85a, DH85b, Edw87, FS84, GMM88, GCL89, GCD87, GMPA^{+}85, Har87, Hob88, Kir83, KCN82, LB84, LT88, PB81, RS86b, RBNF87, SLF86, Sca88, SM85, SSS87, Sim87, VJOR88, VHJ^{+}89, War84, WBS83, WK86a, vGMMG85, Ano89i, Gop89]. Bacterial [Bel80, BS80b, Bel83, BBG89, Bur86b, Cof89, FCC^{+}86, BLKH88, BSW81, CP86, CDG88, CCS^{+}89, Jac89, JL80, Jør87, KB89a, KRCS82, KH86b, KJP89, MH89, PB80a, PC84, RW89, RMN82, TQS86, VO89, WB84, WWS84, ZHO^{+}87, Duc82]. Bacterioplankton [SL87b, BRH^{+}88, CK84b, HFA80, SF82, WCEP82]. Bacteriorhodopsin [Ore83a]. Bacteriorhodopsin-mediated [Ore83a]. Bacterivory [AF85, MF86, EDKS86, SPBD89]. Baffin [Her83]. Baja [KS82a]. Baker [KR89]. balance [KM84, LWL86, Smi81a, SA83]. balances [Cor81]. Baldeggersee [Lem87, IJ84]. Balkema [Ban84b]. Baltic [ML82, Neh87, Rah87]. Banda [VN87]. Bank [DL83b, Fry88]. Barataria [HDP83]. Barcelona [Edm84a]. Barica [Wet83]. Baringo [BSB^{+}87]. Barnacle [Far88]. Barnes [Koh84, New86, Smi82a]. Barnett [RF89]. barrels [WD87]. Barrier [HH81, SA89, AS87, Smi82b]. Baruch [Gar80, Jum82]. based [BF84, BLKH88, DL83a, KTM83, KM83, KH86c, Li83, PS81, RBNF87]. bases [Thi83]. Basin [HDP83, MTL84, Cor81, DH80, HTE84, HI88, PB80b, RMB82, WS81, LL81]. basins [Kel88]. Bassa [SG86]. bathyal [Smi82d]. Baxter [Rzó81a]. Bay [DDSmV83, Han81, KM87, LBZC88, MRLD86, Mur85, SA83, SHDV89, HCC88a, HCC88b, Iba81, Duc82, Her83, MPsCS80, MMD^{+}80, MKN84, RI88, WB86a, WGM82, WM82]. be [Gor89a, SGMB87]. beaches [McL89]. beam [SYB^{+}89]. beams [SVG89]. Beckel [Bur88]. Bedford [MTL84]. beds [CK80, KCN82, Pet86c, Ste88b]. been [GLMA85]. Beer [Gor89a]. behavior [AG88, BMS83, Dod88, HSMH87, MMH85, NMM83, Pat89, PGO82, Pri89, Run80, Sul80, Wil80c, Wil87, ZF89]. Behavioral [Ger82, HD84, AFM86]. Belize [Smi82b]. bell [LSdKvD85]. Belle [Gar80, Jum82]. belowground [SH84]. Belt [Woo85]. benefits [Ste88a, Tag88]. Benguela [OH87]. Benoit [BB88]. Benthic [AHI^{+}86, Kor80, Pai86, AS83, Ber89, Bri82, BWJ80, CGCG89, CA87, Dri81, DMM82, Eck83, FBG89, HAvdL^{+}89, Hun83b, JRC83, JD88, Jum82, LE89, MMD^{+}80, OSML87, RJ83, Sel81, Smi82d, SS85].

Benthopelagic [CGCG89]. benthos [Wil80a]. Bera [Wet84]. Bering [AN81, HS86, SA81, Woo84]. Berkeley [Ano86j, Epp82, Fen84]. Berlin [Dug80, Edm81, Ser80]. Berner [Jah83]. beryllina [WL85b]. between [BF87, BRH^{+}88, BMPD84, BP87, BG86a, BG86b, CSPV86, Car85, CK79, CKT89, CG80, CPT86, DM83, DeM88, DL85b, EEMC88, Fee79, Fen80c, FM82, GG81, HB84, HM89, IL85, JBvdB89, Ker88, Kir84, Leh80, LKS89, MH84, MMD^{+}80, Pat80, RM86b, RM81, SG81, Smi87, Som85, Sta85b, SW81, SH83, WS81, WMSW89, Wil80a, Wil87, WLR88, WHM^{+}85, YP89, ZFFC88, ZF89, vdLAH^{+}84]. beware [Par82]. Bibliography [Ano88h]. Biennial [Fen82]. Big [CCO83, CCW87, IOK87, SO87, ZHO^{+}87]. bight [LB84, Bal84, DG80, Gar81b, GA81, HN81, SE82, TCN^{+}82]. binding [Kep85, FO82]. Binford [BH88a]. Binnengewässer [Edm82b]. Bio [BS82, ML89, NCY89, SPBB89]. Bio-optical [BS82, ML89, NCY89, SPBB89]. bioassay [PHCK83]. bioassays [VWVR84]. bioavailable [BP87]. Biochemical [CW84, DP89a, SC81b, Mor80b, Rue83, ZST89]. biodegradability [Gel86]. biodegradation [BMH85]. bioenergetics [TC85]. Biogenic [RDW^{+}88, Bal84, Ham81, Now84, SEK84, TS85]. Biogeochemical [BMH86, San88, SKW86, HMC88]. biogeochemistry [BMS^{+}88, CSF88, CB84, HKT89, Dav86]. Biokinetics [FBF83]. Biol [Edm82c, Hec86, Hyn81, Pim83, Wet84]. Biological [BCM86, HS80, Sel81, TNJ84, WKRE89, HDLP85, HGH82, Jum84, LMS87, McK81, Ste84]. Biologically [SB80, vdLAH^{+}84]. Biologists [Par82]. Biology [Fry80, Shi85, CRGC87, Fen80b, Har89a, HLS87, Pai84b, Pai80, Lon84, Far88, Ben89, Bez89]. Bioluminescence [LLG86, BS89, BMS83]. Biomass [Sim87, ADK87, AK89, BZH^{+}85, BB89b, BFKN84, CGCG89, CRGC87, DTS^{+}89, DK86, Duc82, Han88, HK81a, HLS87, HL81, LKS89, MRLD86, MB89b, RMN82, RM86b, SNR86, Smi82c, Thr88, VN87, Wof83, ZHO^{+}87]. Biosynthesis [SCH89, VAI85]. biota [KE80]. biotic [Luc86]. bioturbation [MSG86, SGMB87]. birgei [Wal85]. birth [BB89a, KN81, TS81a]. Bitton [ZoB80]. bivalve [TQS86, WCEP82]. Biwa [Nag88]. black [DC88a, MP88, RR80]. blackwater [CM89, Edw87, FF85, SB89c]. Blackwell [Bur86a, Fen80a, Smi82b, Tut80]. bleaching [KDR89]. Blelham [Haw80]. bloom [IDB82, LB84, LBZC88, Ore85, Pae89, PAGG87, RI88, SP85, Wro89]. bloom-forming [Pae89]. blooms [FP87, HL88, KB89a, PU82, PTB83, Pae88]. Blue [PU82, FB80, FP87, MM80, PTB83, Pae89, PM84a]. Blue-green [PU82, FB80, FP87, MM80, PTB83, Pae89, PM84a]. Boca [Bro83]. body [DP80, Han85, KH86b, KH86c, Mal83, PL89, RM86b]. body-length-based [KH86c]. bog [FH82]. bogs [UEB88]. Boje [Dug80]. Bolivia [VWVR84]. bomb [TCN^{+}82]. bomb-produced [TCN^{+}82]. Book [Ano85d, Ano86j, Ano89i, Ban84a, Ban84b, Ben89, Bez89, Bir88, Boe89, Bro80, Bro82, Bro83, Bur88, Bur80, Bur86a, Cam87, Cat85, Cha85, Csa80, Dav83, Dav86, Den82a, Duc87, Dug80, Edm81, Edm82a, Edm82b, Edm82c, Edm82d, Edm84a, Edm84b, Edm87, EA84, Epp82, Far88, Fen80a, Fen80b, Fen82, Fen83, Fen84, Fow80, Fre84, Fry80, Gar80, Gil80, Gop89, Hal82a, Har89a, Har89b, HG88, Hay87, Hec84, Hec86, Hen87, Hor83, Hrb82, Hyn81, Ill80, Jah83, Joh89, Jum82, Jum84, Jum87a, Kir88, Koh82, Koh84, Kru80, Kru82, Lam87, Lec85, Leh89, Lis85, Lon84, Lon85, Lon86, Lun84, Mac85, Mac80, Man80, Man83, Mar83, McM86, Mil82, Mil87, Mil84b, Mol89, MS84, MS87, Mor80b, Mor88a, Mur83, Mur84b, New86, Now84, Pae89, Pai80, Pai84a, Pai84b]. Book [Pai89, Pic82, Pic85, Pim83, Pla86, Pom80, Pom89, Pon82, Pra80, Ril82, Rzó81a, Rzó81b, Ser80, Shi85, Smi82a, Smi82b, Smi86b, Smi88, Ste84, Sto89, Thi83, Tip87, Tut80, Wan82, Wel88, WK80, Wet82, Wet83, Wet84, Wet87a, Wet87b, Whi86, Woo80, Woo85, ZoB80]. Books [Ano80d, Ano82d, Ano84g, Ano87g]. Borntraeger [Edm81]. Bosmina [BS85a, BG82, DeM82, Thr81, ZK80]. Bosminidae [SCR84]. Boston [Mac85]. Boström [Lon85]. both [RS88]. bottle [BKF^{+}87, Gil89, LDYY83]. bottles [SB80, WBV80]. bottom [FP81, Han84, IB82b, KE80, LSI^{+}85, LI87, MRLD86]. bottom-ice [MRLD86]. Bougis [Ban84a]. boulders [TS83]. bound [KDC86, SM85]. Boundary [RW87, BWJ80, CGCG89, CA87, FBG89, HAvdL^{+}89, JR85, Lyl83, SBN^{+}83, Smi82d, SS85, Wol87]. Boundary-layer [RW87, FBG89]. bovine [ZST89]. Bowden [Pic85]. Bowien [Ano89i, Gop89]. box [CW85, RR80]. Boyles [Cha85]. brachyuran [BT84, VFB87]. brackish [HH85]. brackish-water [HH85]. Brady [JERC85]. breakdown [TNJ84]. breaking [Den82b, Bur88]. Brenner [BH88a]. bridge [YP89]. brine [GSB89]. brines [Kru80]. Bristol [Unc81]. British [Ano85d, Bot85, HW85, WS81, JT84, PGB88]. Brodsky [Hyn81]. Brook [Ano80e]. Brooks [Den82a]. Brown [Fen83, Lyl83]. brown-green [Lyl83]. Bubble [BSW81, Ger83a, Ger83b, WBS83]. bubble-generated [Ger83b]. bubbles [DMM82, Gar81a, JC80, JGCS82]. budget [CCS^{+}89, CL80, DWO86, JV82, Lue88, SMP^{+}89, Top87, WMSP86, WM87, YJ84]. budgets [BSB^{+}87, DQRM87, EH88, Kor80, LW84, LL89, WL85a]. buffer [Fro88]. Buffering [MG80, Lew82a]. Bull [Rzó81a, Thi83]. Bults [Mil82]. Buoyancy [Kon89, vRS83, HL88, OW84]. Burgis [Smi88]. Burlington [McK81]. business [Hyn80].

C [Cha85, Dav86, Den82a, Fen80b, Fen83, Fre84, Gol86, GCD87, Har89a, HG88, Hay87, Hec86, Hrb82, Jum82, Jum87a, Koh82, Mac85, Mil82, Mil87, Pai80, Pon82, Tip87, Wet87a, Wet87b, Woo80, Woo85, ZoB80, JW84, AS83, Col84a, FH84, Fry88, HHH87, HD86, LDYY83, NO89, OJR^{+}86, OSML87, POHMD84, PHCT89, PH88b, RS81, SB80, SH87a, TS85, TCN^{+}82, WLB84]. C$2.50 [Rzó81a]. C$3.00 [Rzó81a]. C-labeled [FH84, HD86]. C-labeling [NO89]. Ca [KH80, SSGB89]. CaCO [LSK^{+}83, WBEJ85]. cadmium [FM82, JM84, Ols80b]. Cahora [SG86]. calanid [Hun81]. calanoid [CG80, DKMvL86, GM88, Lan83, PPS83, RI88, WB86b]. Calanoida [GC81, Wal85]. Calanus [BF89, Håk87, Hak84, Har83a, Lan80, LHF^{+}84, MB86, McL86, Run80, Run85, SLL85]. calcareous [SMK81, Sta85b]. Calcite [Sta86, MHY83]. calcium [Avn83, CW83, Moo83, OC82]. calculate [PG81]. calculated [OJR^{+}86]. Calculation [Cul80, RBNF87, MC82]. calculations [KA89, Kim87]. caldera [Óla80]. calibration [CCS^{+}89, LSB81]. California [Ano86j, Epp82, Fen84, Gol88, SE82, Che80b, CB85, Håk87, Han81, KBB^{+}89, KS82a, LK89, PM86, SHDV89, Sto87, VG80, ZR85]. California-Nevada [VG80]. californium [FBF83]. Calkins [MS84]. Call [Ano82g, Ano85e, Ano85f, Ano88i]. calories [SM82a]. Cambridge [Lon86, Lun84, Mor80b, Pon82, Smi82a, Smi86b, Smi88, Woo85]. Cameroon [Kli88]. cAMP [FW82]. Can [BFKN84, Gor89a, Rzó81a, Thi83, BSN83, Whi81]. Canada [Rzó81a, Thi83, CW83]. canadensis [LL89]. Canadian [MRLD86, Har83b, POHMD84]. Capabilities [RJ83]. capacity [KSH85, Lew82a]. Capblancq [Edm84b]. Cape [Fis80]. capitata [TH80, TC85]. Capitella [TH80, TC85]. Capture [PP86, CH87, KS81, PPS83, SF81, VP85]. captured [KC85]. Caraco [KJP89]. carbcn [OAI^{+}85]. carbcn-13 [OAI^{+}85]. Carbohydrate [vRS85, Bur86b]. carbohydrates [IDB82]. Carbon [BLKH88, CFPW89, CL80, DL85a, HS84a, HDK84, Joh82, LMS87, Smi81b, WL84, WWS84, BMH86, Ber89, BR80, CK88, CSF88, CLS82, CCS^{+}89, CAS83, DSH82, DQRM87, FHHB80, Fro80, Fry86, HS82, HF88, HL81, JEE82, KLB80, KTM83, KM81, LS82a, Le 82, LHM84, LSLW86, LGM80, Lis85, LWL86, MFHM80, Mal82, MH89, Mul81, MMM81, PS85, PH87, PS89, QEQD86, RDW^{+}88, RV87, RJ84, SL87b, SGSS84, SRW80, SM80, SA83, Smi87, SHDV89, SW81, TS82, VHJ^{+}89, WSU^{+}83]. Carbon-14 [WL84, CL80]. carbon-based [KTM83]. carbon-controlled [SHDV89]. carbon-nitrogen [KM81]. carbon-specific [Mal82]. Carbonate [DWO86, AHI^{+}86, Avn83, Ham81, HS84a, LSK^{+}83, Moo83, SB89b, dJV89]. carbonate-containing [HS84a]. Carboxylating [PG83]. carboxypeptidase [HFFS85]. care [GBB81]. Cariaco [HE88]. Caribbean [CWV^{+}88, GCAN87, HF89, LGM80, MSG81, Sau80]. carinata [GB81]. Carmans [CD85]. Carmouze [Hec86]. carnivorous [AFM87, Yen85]. Carolina [Gar80, Jum82, Ano83f, Kin88, SH81, SH87b]. carotene [KGGW88]. Carotenoid [PTB83, Rin81, Zül81]. Carpenter [Leh89, Pai89]. carried [Ger83b]. carterae [VAI85]. cartridges [SH87a]. cascade [Thr88]. case [Den85, SHDV89]. Castenholz [Shi85]. Castle [MTW^{+}88, PG83]. catabolism [CK88, CDG88]. Catalina [ZR85]. catch [DC87]. catchment [ADO85, BF84]. cathode [Rev89]. Cattaneo [GG81]. caught [Hak84]. Caulerpa [Wil84]. Caura [LSLW86]. cause [Pai86, Whi81]. caused [FCO87]. Causes [HL87, HWL83, SM82b]. cd [Hen87, LL82]. Cell [Ban82, YMN82, ADK87, CW81, HL88, NG83, OKG81, PPS83, RBNF87, Som89a, War84, WMM83, YHM^{+}83]. cells [Bjø88, BB84a, CI88, ET80, MC82, PP89, PP86, RS81, SCO89]. cellular [KH88a, SH83, SH85]. Cellulose [SMH84]. center [BT84]. centers [Lon85]. Central [Edm82c, Bru89a, GDS83, Neh87, RK87, SLB83, Wei80, WL85a]. centrifugal [BQ86]. centrifugation [OKG81, TD83]. Centropages [CS83, DG80]. Ceratium [Wei80]. cf [DC87]. CH [MD87]. Chad [Hec86]. chaetognath [Fei82]. Chaetognatha [Sul80]. chain [SNR86]. Chalk [FP81]. Champ [Edm84b]. change [Ger82, Gri88, Hyn80, Woo80]. Changes [BB89b, IVH^{+}86, KDR89, Lew82a, AHRF89, CR83, Cor89, CRGC87, Dug81, EL81, GB80, HL85, HLS87, Iba81, LMS87, MTW^{+}88, Pai86, Pat89, Run85, SF82, WLB84, Wil80c, Col82, Col80b]. changing [KM89, LSdKvD85, MLG87, Lis85]. Channel [Unc81]. Chaoborus [SF81, Gig80, HD84, HG85a, Lue88, Moo88, Pas80, VKW89]. Chaoborus-induced [HG85a]. character [Ban87]. characteristics [ALRP85, BSS^{+}84, HGH82, HFFS85, LM86, MTL84, Yen83]. Characterization [Chr89, CS83, DSB86, Ger83b, TB80, IS89]. charge [HL82, LL85b]. charged [NL87]. Charlottesville [Pom80]. charts [McM86]. chelates [RNVB81]. Chemical [HF89, Kan83, McK81, Mor88a, Sta86, VO89, Bur84, Dod88, GDS83, GM84, HDLP85, HGH82, HS80, LCPL88, MMP86, MB80, RE82, SBN^{+}83, SC82, THCP81, WLG85, Cam87]. chemically [BP87, DeM89, HK81b, Ste88a]. Chemistry [FP81, LW81, Ser80, AM82, HL87, KB89a, Lan82, MMD^{+}80, Mil87, Mor88a, Óla80, Ril82, SCC82, WS81, Lec85, Wan82]. chemocline [CCW87]. chemoheterotrophy [ES82]. chemoreception [Ano88k, ZF89]. chemosensory [Jac87]. chemostat [Bor83]. chemostats [VO89]. Chesapeake [Duc82, MKN84, WGM82, WM82]. Chilwa [Woo80]. Chironomidae [HW86, Men81]. chironomids [SB89c]. Chironomus [But82, JBvdB89]. Chlamydomonas [SH85]. Chlorella [QHHLS89, Dou87, Dou88, McA86]. Chloride [BSB^{+}87]. chlorinity [Mil84a]. Chlorophyll [BSN83, CEE86, SCO89, WL85a, APR80, ADP^{+}84, ADF^{+}82, CP86, CSPV86, CSHO89, Car80, CDRW86, Edm80a, FVH^{+}84, GK83, GPS85, HM87, HK81a, Her83, HL81, KCB89, KJP89, Lan82, Lor80, LD86, MP89, NCY89, PME87, PGB88, PG83, Sta82, WL84, YMBO87]. chlorophyll-a [CP86, FVH^{+}84, GK83, HM87, HL81, KJP89, Lan82, MP89, NCY89, WL84]. chlorophyll-derived [CDRW86]. chlorophyll-Secchi [Car80]. chlorophyllide [LD86]. chlorophylls [BKB85]. Chromatic [Dri81]. chromatographic [Jef81]. chromatography [BKB85, HHH87, HFF87]. chromatography-mass [HHH87]. Chromium [CM80]. chronology [ADO85, HW85]. Chrysophyceae [BK87]. chrysophyte [PNL84]. chrysophytes [CS88, SH89]. Chydoridae [Fre88]. cichlid [DHV^{+}87]. ciliate [Gil89, SNR86]. ciliated [BC82]. ciliates [Ban82, Fen80c, Nic89, PS89, SPBD89, SSFN86, Sto84]. Circulation [ASS81, CWD^{+}86, CB86, HS86, HS88, MBT80, Rah85]. Circumpolar [Lon86]. Cisó [PAGG87, vGMMG85]. Cl [SSGB89]. Cladocera [SCR84, Fre88]. Cladoceran [KH86b, TG87, Han85, KH86c, MD85, TG82, TTF88]. cladocerans [PM84a, SPBD89, Zar81]. claim [BWS^{+}85]. clam [CDPC84]. Clap [RE81]. clarification [MC82]. clarity [DC88a]. Classification [PHI84, BS82, PS81]. clausii [Ros80]. clavipes [CG80]. clay [Cuk87]. clean [FS84, TD83]. clear [BF87, LFRT86]. clear-water [LFRT86]. Clearance [BG82, Fen80c]. Cliffs [Mil84b, Mur83, Pai84a]. climate [Lis85, Smi86b]. closed [HI88, PB80b]. closed-basin [HI88, PB80b]. closure [Zan89]. Clouston [Koh84]. clupeid [TD87]. clutch [Mal83]. CO [DL83a, RJB81, Fre84, Kru80, Mac85, Mac80, Wan82, SK83b, DL85b, Fra88, HS82, KM84, LDYY83, Ore83a, RS88, TNEJ84]. co-occurring [SK83b]. coast [Båm85, DP89a, ER88, JM84, PGB88]. Coastal [Wri89, AM82, CDG88, CL81, CP83, CM83, Den82a, EHE^{+}83, Fen80a, Fen84, GPS85, HCC88a, HT85b, Hun83b, Hun80, INA89, KS82a, KM89, KW87, LRH^{+}84, Man87, Man88, MV88, Mur84a, PCP87, Pae88, Pic85, PH88b, PS81, PS89, RBNF87, San88, SH84, SNP84, SN85, Sei88, SC81b, SPBB89, SH81, TAJ89, WD87, ZKD87, Mor88b, Smi82a]. coastlines [Wri89]. coatings [LL85b]. coccoid [MG81]. coccolith [SRW80, SW82]. coccolithophores [Bra82]. cochlearis [BG86b, BG86a]. cod [Ree81, Fis80]. cod-end [Ree81]. coeffcient [PP89]. coefficient [AB80, Gor89a, HM87, Vos89]. coefficients [Ban88, Ker88, KB80b, LD86, UA82, WB86a, Wof83]. coexistence [Eva88]. Cohen [Shi85]. cohesive [BWJ80, Pat89]. cohort [HT85a, SL87a]. coincidence [Pai86]. cold [KBB^{+}89, RJ84]. Cole [Man80]. collapse [OW84]. collecting [VW80]. Collection [BS85a, Edm84b, Lew83a, BW89, MHB84, Ree81, RF88]. collectors [Han81]. College [Hrb82]. Colloq [Mil84b]. colonial [FP87, Swa83]. colonization [SB89a]. Color [Mor88b, AB80, CSPV86, GPS85, LO81, Lyl83, MLG87]. Colorado [LW84, Lew82a]. Columbia [Gar80, Jum82, Bot85, CM80, HW85, HTE84, PGB88, WS81, Ano85d]. column [And85, BRBL83, HE88, HCC88b, Nag88, Ore83b, Pai86]. combination [BSN83]. Combined [War84, Cof89, RWB87]. Comment [BH88a, Boo80, CR83, Gar81a, GG81, Pat80, Sel81, Top87, Ami88, BW89, BB88, CP89, CK81, Edm80b, Edm84d, Edm84c, Edm86a, JC81, Jum87b, Jum88a, Jum88b, KR89, KJP89, Lan83, Oud89, Pec81, Phi80, Pra90, RF89, WM87]. comments [Fee80, Gie82, Jum88b, KW86, LD86, Sto88b]. commercial [Ami88, OGMG88]. committee [Lis85]. communities [Atk87, Car85, DK86, DMM82, Hrb82, JRC83, JD88, LO83, Pat88, VO89, Van86, Hal82a, HG88, Leh89, Pai89]. community [Abe88, BGJ^{+}87, BCL85, BG82, Bur80, Cuk83, DdNK82, DMdK84, DHV^{+}87, EJ86, Gil88, Gil89, Ive88, JBvdB89, KBB^{+}89, KH86c, LN86, Leh88a, Loe81, LS81, Mac84, PO81, Pac86, Pai86, Pim83, RV87, SPBD89, SLB83, Smi87, SK85b, TD87]. Comparative [ES82, GG82, KH88b, Kli88, Lop88, Ste88b, Nix88, Sto89]. compared [Kel89]. comparing [Bay88]. Comparison [CK88, Gel86, Haw80, JS83, KLJ86, PG81, SG81, SE82, Som85, BGJ^{+}87, Bot85, CK84a, EJ86, Har83b, HB85, Hob88, MH84, MH87, Nür87]. Comparisons [JEE82, ZST89, DL85b, Kel88, Unc81]. Competition [SK83b, TML81, Van86, GS85, HA81, Pim83, Pra90, SK85a, Som85, SK85b, Som89b, SH87a]. competitive [Han88, SH83]. completed [Ano88j]. completely [SMR84]. complex [FH82, GB81, Leh89, Pai89]. Complexation [Bru89a, CB88, MM80, SB89b, vdBD84, Dav86]. complexing [SvdBR82]. complications [Car80]. component [MCK89, SSFN86]. components [BH84]. composition [Ban84b, BCL85, BB89b, CD85, Csa80, CW84, Cuk83, Dav86, Dug80, Fen83, FVM82, GDS83, HK81a, HB84, Hor83, IDB82, Jum82, LMS87, Lis85, Lon85, Mac85, Mac84, Mar83, MS84, Rzó81b, Söd87, VO89, Wet82, WWS84]. Compositions [HCQ^{+}86, CFPW89, Fry88]. compounds [GZG80, Har83b, Neh87, NG87]. compression [Zar81]. Computation [KM80, SP89a, Mob89]. Computer [GG84, GCAN87, RMJ86]. Computer-automated [GCAN87]. Concentration [Syz82, BS80a, BK80, BK84, Ber89, BR86, DW86a, DP80, Gri88, GPS85, Han85, IVH^{+}86, KCB89, KH80, LS83, PGO82, SH87a, TS82]. concentration-dependent [BR86]. Concentrations [Jør87, All81, CMR89, Cor89, FDR^{+}88, FSW85, GS86, KLC81, Mar83, Smi86a, TW83, TML^{+}88, WK86b]. concept [Fly88, SS88a]. conceptual [Hil85, Pra80]. concerning [Jum88b, KMK88]. concerns [Jum88a]. Concurrent [HK88b]. condition [CR83, EL81, Håk87, Hak84, MS85b]. conditions [AK89, CW85, CS83, DP81, Fra87, KMK88, Smi81a]. conductivity [Mil84a]. conductivity-density-salinity-chlorinity [Mil84a]. Conf [Csa80, Lon85, MS84]. Conference [Fen82]. Confidence [KN81, KM80]. configuration [Win81]. confined [CSW^{+}80]. Conflicts [WK86b]. confoederata [HMG86]. Confusion [Sch88]. Congr [Pic82]. Congressional [Har89b]. Connecticut [Hal82a]. Consequences [CDG88, LG89, CDRW86, HWL83, HDLP85, SFDK84, TNJ84]. Consideration [Li83]. considerations [DW86b, JW85, SCC82, ZST89]. conspecifics [SF89]. Constance [Sim87, Sta86, Til83]. Constant [OP83]. constituents [CSPV86, DM85, Gel86, HDW85]. Construction [KH86c]. consumers [Fry86]. consumption [BBP82, Bow84, CGCG89, CAS83, Lid83, Rah87, RMJ86, RKS^{+}86, Smi82d, SLB83, Smi85, Smi87]. containing [HS84a]. Containment [RR80, JD86b]. contamination [Dou87, Dou88, FKM82]. contemporary [TD87]. content [CKT89, Fra87, Hak84, HK81a, Paa80, RGS89, Rin81, SGSS84, TB80]. continental [Fis80, Mac84, YMBO87, Pom89]. continentales [Edm84b]. continuous [Bot85, Haw80, Iba81, LB80, PHCK83, RE82, SG84]. continuous-flow [Bot85, PHCK83]. continuously [Kim87]. contrasting [BCL85]. Contrasts [CK88]. Contrib [Koh84]. Contribution [GMM88, CK80, LHH80, MMM81, VCG^{+}87]. contributions [BMH86, BG82, KC81, Kor80]. Control [FCS84, GS85, LL85b, Fro88, GMK80, LFRT86]. controlled [HA85, Hun83a, SHDV89, lWJ82]. controlling [Cuk83, FDR^{+}88, FVH^{+}84, FSW85, HA81, HF88, KC81, Klo85, McK81, San88, SH89]. controls [BM82, HMC88, SL87b]. Convective [HS88]. conventional [MH84]. Conversion [RMN82, BLKH88]. Cooke [Gar81a]. cooling [HS88]. Copepod [DFW89, KS81, BF89, BMS83, BWG83, Che80a, Che80b, CG80, CS83, DW87, GM88, Hai80, HWL83, HM80, Hay81, HGM80, HH85, Hun81, KM87, KMSF81, Ohm87, PP86, RR80, Rom84, Ros80, SNR86, VP85, WC86, WB86b, Yen85]. Copepoda [GC81, GR84, LL89, Wal85]. copepodites [Har83a]. copepods [BSC80, DW86a, DKMvL86, DeM88, DeM89, GBB81, Her83, Lan83, MHB84, O’D85, PHCT89, PPS83, Rin81, RI88, SL87a, SF89, Wil87]. Copper [CB88, MM80, JM84, McK81, RM81, Rue83, SH83, SH87a]. Coprecipitation [MHY83]. Coral [Vic84, Ano86j, Atk87, BC80, Car85, DL83b, EBSW83, Koh84, Koh87, Lew87, MS85a, MMM81, PS84, RM83, Smi81b, Wol87]. Corals [TS83, KDR89, MS85b, SA89, Ano86j]. Corbicula [CDPC84]. core [BW89, BMH^{+}87, CRGC87, DH85a, DH85b, GBF86, Han87, HLS87, LW81, MN86, RF88, SHRW84, ADO85]. corer [CW85]. cores [Chr84, MH86, SRW84]. Corp [Hor83, Koh82, MS84, Now84, Wet82]. correction [GPS85, LDYY83]. correlates [DL83b]. Correlation [FW82, ADO85, ER88]. Correspondence [Han81]. cosmogenic [LCPL88]. COSPAR [Hor83]. COSPAR/SCOR/IUCRM [Hor83]. cost [PM84a]. costatum [SAK89]. costs [Ste88a, Tag88]. Coull [Jum82]. Council [Lis85]. counted [BK82]. counting [CW81, Har87, PF80, RS81]. Coupled [OAI^{+}85]. Coupling [ZFFC88, JK84, LB84, MDR88, Pai86]. course [Har83b]. cove [Men81]. cover [LSdKvD85, PH88a, SM82b]. covered [CA83, CA87, PPVHW87, Rah85, Top87, WMSP86, WM87]. covering [CM86]. crabs [BT84, VFB87]. Craigie [Mor80b]. Craspedacusta [DC83]. crassa [MG83]. crayfish [Cap82]. CRC [Bro83]. Creek [WH80, WMSW89]. crests [GB81]. Cricotopus [Men81]. critical [CW81, KSP^{+}81]. critique [HT85a]. crop [Jav83, JRG81, Loe81]. cross [KM83, KSP^{+}81, PP89]. cross-section [PP89]. crown [Ols87]. crown-of-thorns [Ols87]. Crustacea [GR84]. crustacean [FCS84, GSBP87]. crustaceans [PO81, ZF89]. Cryptomonas [PAGG87]. Cryptophyceae [BS85b]. cryptophytes [EHE^{+}83]. Crystal [Mur85]. crystallina [DP80]. crystallinus [VKW89]. Cs [MH86, SGMB87]. CTD [Atk88]. ctenophore [HSMH87, SRW81]. Cu [LL82, SB89b]. cultural [Wri89]. culture [Bor83, FOB80, GHK80, HA81, LB80, Ros81, SB80, WBV80]. cultured [GB87, VO89]. cultures [Som84]. culturing [Ols87]. Cummins [Mac80]. cupressoides [Wil84]. cupric [SH87a]. current [CRFB80, HT85a, HK80, HA85, CB85, Håk87, Jum88a, OH87, PM86]. Currents [BC86, Gey89, IB82b, KS81, LG88, LSI^{+}85, LI87, MPsCS80]. curves [PS81]. CuSO [McK81]. cutthroat [Lue88]. cyanobacteria [EHE^{+}83, Kon89, MG81, OW84, Pae89, vRS85]. cyanobacterial [FB88, JRC83, JD88]. cyanobacterium [CW84, NDS86, PL89].

Cyanophyta [BB84a]. Cyanophyte [HL88]. cycle [GH84, Kin88, LR85, PO80, PB81, QEQD86, RDE82, vGMMG85]. cycles [MG83, San88, Str80]. cyclic [FW82]. Cycling [VHJ^{+}89, Bal84, BMH86, BH83, DM85, FHHB80, KMD^{+}88, Leh80, SNL84, SHDV89, Sta85a, WKRE89]. Cyclomorphosis [Hav85, Rie84]. Cyclops [GR84]. cyst [ATA87]. cysts [AATC82, ALRP85]. cytometric [SCO89]. cytometry [BCY89, GSBP87, NCY89, PP89, WHM^{+}85, YHM^{+}83]. CZCS [PGB88, YMBO87]. CZCS-chlorophyll [YMBO87].

D [Bir88, Dav83, Edm82a, EA84, Fen83, Ill80, Koh82, Koh84, Lam87, Mac85, MN86, Mur84b, Pic82, Pim83, WK80]. D.C [Lis85]. daily [MMM81]. Dalhousie [Ano82e]. damage [SRW81]. Dame [Gar80]. dams [Rzó81a]. Danish [BH83]. Daphnia [AMP81, BH84, BG86a, BG86b, CM82, DD89, DeM82, Dod88, DH88, DC87, EL82, EJ86, Ger82, GS85, Gil88, Gil89, GS86, GB81, Hal82b, HD84, Hav85, HG85a, HGS83, HS84b, IL85, IA85, JJ87, Kir85, KD81, LS83, LWL86, MMH85, Mur81a, NO89, NDS86, OP83, PL89, PGO82, RD83, Rie84, SIH85, THGD83, Van86, VKW89, WK89, Wei84, YTW84]. daphnid [BG86b]. DAPI [PF80]. Dark [PM88, LDYY83, SF81]. darkness [ATA87]. Dartmouth [Hrb82]. data [AH87, DPF87, GBF86, HT85a, HBT87, RMJ86, SL87a, Unc81, Jum87a]. date [Nür88]. dated [MH86]. dating [Mur85]. Davy [Fen80a]. Dawson [Wan82]. Day [CP83]. Day-to-day [CP83]. DCMU [LDYY83]. de-venir [Fow80]. Deacon [Lon86]. Dead [GK89, GSB89, KSTGB81, OS82, Ore83a, Ore83b, Ore85, Sir87, Ste83, Ste85, SC84]. death [KN81, TS81a]. December [Ano80a, Ano82a, Ano85a]. decline [Ore85]. d’Ecologie [Edm84b]. decomposing [LHH80]. Decomposition [HD86, RD89, SSB82, BZH^{+}85, BWG82, FB80, KRFS84]. decrease [D’A80]. deduced [MMP86]. Deep [Lon84, ADP^{+}84, CGCG89, GB82, HK80, Jah85, LG89, Mil84b, Neh87, NJF84, PG83, RS86c, SBN^{+}83, Smi85, Smi87, SHRW84, SRW84, SJC^{+}87]. Deep-sea [Lon84, CGCG89, Jah85, NJF84, SBN^{+}83, SHRW84, SRW84, SJC^{+}87]. Defant [LM86]. defecation [BFKN84]. defense [HF89]. defenses [HF89, Ste88a, Wil80c]. Defensive [Ker88, VKW89]. deficiency [RM81]. deficient [RLG86]. deficit [CR80]. deficits [Lin87]. definition [Par82, SC82]. degradation [Bal84, BKB85, CEE86, HA83, HASI80]. degree [GC81]. dehydration [Joh82]. Dekker [Kru82]. Delaware [CSF88, LWF86, SCC82]. delicate [Ree81]. delta [LG88, Moo83]. deltaic [MDR88]. demand [Smi87]. demissa [JV82]. demographic [SL87a]. Denitrification [SNP84, Sei88, SHRW84, GNM87, JEM82, JK84, RPL89, SN85, SJC^{+}87, UEB88]. Denmark [Høj89, IJ85]. dense [TM84]. densities [Pet86c]. Density [Moo88, ADK87, AK89, BG86b, CM82, DM84, FVM82, KSP^{+}81, Mil84a, OKG81, PHI84, PB87, TC85, TD83]. Density-dependent [Moo88]. Dependence [AB80, Gor89b, Kir84, OJR^{+}86, AK89, NO89, Smi82c, SB89b, Som89a]. dependent [BR86, Gro89b, Moo88]. depletion [PB87, SWSC85]. depolarization [CMB87]. deposit [Ham81, LE89, MJN84, NJF84, Phi80, Tag88]. deposit-feeding [Ham81]. deposition [Hab85, Hun83a, JKMN87, LE83, SRW80, TCN^{+}82, WED85]. Depositional [HTWR84]. Dept [Rzó81a]. Depth [SJC^{+}87, AB80, BWS^{+}85, Hay81, Kli88, KLC81, LS83, ML82]. depths [LE89]. derivatives [GK83]. derived [CDRW86, MH89]. Description [GK88, KM83, MCK89, SAK89]. desert [HB84, HG85b, PPVHW87]. desiccation [HM82]. Design [CRFB80, Den82b, DPF87, WLR88]. destratification [Duc82]. destructive [SG81]. detailed [CCS^{+}89]. detect [TQS86]. detecting [Zül81]. Detection [Lan80, Car89, Col80a, Iba81]. determinants [Sel81]. Determination [Chr84, DH80, HHH87, HFF87, JP83, KS82b, PP89, SAJB86, SS80a, SS80b, TCN^{+}82, WL84, Ami88, BKB85, CV82, FNL85, HY89, JD86b, KH80, LM86, LD86, NG87, OC82, OGMG88, WC86]. determinations [BSN83, HS84a, Ven87]. determined [CA87, Hak84, ML82, PM86, PS89, QBHS80, RMJ86]. determining [BGJ^{+}87, CSAO86, SB83]. Detrital [MH87, IS89, LHH80, MH89]. detritivore [BBA84]. detritus [BLKH88, BBA84, JR85, KBO84, Man88, RD89, RMN82, Rom84, SSB82, TH80]. detritus-based [BLKH88]. Develop [Cat85]. development [Han88, Har83a, Lan83, MP89, Ros81, Sch85, ZEY89]. Developments [Dav86, Edm82a, Kru80, Mur84b, Wet83, Zül81, Edm82d]. device [CH87, Den83, FP82, Ger83a, HB81, JGCS82, KE80, RG83]. Diagenesis [AR88, KM84, Jah83]. Diagenetic [GB82]. dialysis [Car84, Kep85]. diaptomid [Mal83, Wil87]. diaptomids [GC81]. Diaptomus [CG80, Elm83, Kir85, RD83, VP85, WBF85, WB86b]. diatom [Abe88, AM82, BK82, FM82, HM86, Haw80, HW85, JRG81, McB88, MLG87, PO80, Paa80, PH88b, SAK89, SH83, WHFT89, Whi89, YMN82, ZFFC88]. diatoms [BF84, CS88, Col82, Col80b, CKT89, GBB81, Hab85, KKH86, MB80, Pat89, SGSS84]. dictyopterenes [JW84]. Diel [AFM86, KB82, KB89a, LL84a, MG83, ML82, OP83, SF82, WKLK89, WK86b, WL85b, vGMMG85, BF89, DFW89, FCS84, Hai80, Han85, NL89, SLL85]. diet [BBA84, Dug81]. Differences [CKT89, IL85, BU86, EHE^{+}83, WLR88]. different [BH88a, BB86, BH83, JS83, Nic89, Sto88b, TH80, WST80]. Differential [SA89, Pet86c]. differentiation [Bra82, GC81]. differently [KGGW88]. differing [FW82, Söd87]. diffuse [Gor89a, Gor89b]. Diffusion [HY89, UA82, KB80b, QBHS80, RW87, vdLAH^{+}84]. Diffusive [JR85]. diffusivities [Lew82b]. diffusivity [CA83, CA87]. digestion [SMH84]. Dilution [BH88a, BB86, Bow84, GLMA82, GLMA85, Law84]. dimethyl [FH84, TML^{+}88]. dimethylpropiothetin [VAI85]. Dimethylsulfide [And85, INA89, VAI85]. dimethylsulfoniopropionate [TML^{+}88]. dimethylsulfonium [INA89]. Dimethylsulfoxide [And80]. dimorphism [GC81]. Dinobryon [BK87]. dinoflagellate [AATC82, ALRP85, ATA87, BMS83, CS85, KMK88, Whi81]. dinoflagellates [BQ86]. Dinophyceae [Fie82]. dioxide [Joh82, Lis85]. dipstick [RE82]. Diptera [HW86, Men81]. Direct [BG86a, HG88, OW84, PM84b, DWR82, WLR88]. direction [APR80, Ano88j]. directory [Har89b]. disappearance [Ste85]. discharge [WS81]. Discoveries [Mor88b, LG89]. discrepancy [Smi87]. Discrimination [DeM88, WHM^{+}85, WJ89]. discussing [Eva88]. Discussion [Man80]. disequilibria [CB85, CB87]. Disjunct [Wil80b]. disk [BWS^{+}85, Car80, DC88a, Edm80a, Lor80, Pre86]. disks [MSBC80]. Dismal [Pom80]. dispersal [SA89]. Displacement [Cap82]. displays [LLG86]. Dissertations [Ano88j]. dissipative [McL89]. dissociation [MPF88]. dissolution [Gar81a, JC80, dJV89]. Dissolved [BB81, BR86, CS81, Car83, EHD^{+}86, JT84, MM82b, OK87, SSS87, Ami88, BK80, BK84, BMP81, Bur86b, CL81, Cof89, CLS82, FH82, Fro88, GCLC87, HA83, KLB80, KB89b, KBO84, LH80, MWB82, MTW^{+}88, ML82, MH89, OGMG88, Oud89, PS81, RS88, RWB87, Smi86a, SS80a, SS80b, SW80, SW81, TS82, WLB84, HFL84]. distance [IH81]. Distinguishing [HM89, vdLAH^{+}84]. Distribution [HFFS85, RSBL80, SK82, BCY89, Ber89, Brz88, CWD^{+}86, CGCG89, DM84, Ebe81, FVH^{+}84, GHC89, GMPA^{+}85, Han87, Hay81, HK81a, HB85, HB84, Her83, HLA86, HF88, IBC^{+}83, KB89b, MH80, MMTL88, MJSM89, MH85, PGB88, PAGG87, PPVHW87, SP89a, SNL84, SH89, Sta82, Sul80, Vos89, Woo84, WL85b]. distributional [AH87]. distributions [AATC82, BD81, Elm83, GWB89, HTE84, KM89, McB88, Mob89, Mul89, NSM87, Sel81, VFB87, Wil80b]. Diurnal [BC80, Imb85, SIR86]. Div [Thi83]. diver [CH87]. diver-operated [CH87]. diversity [Mil82, Pat88, Som84]. division [FO82, Wei80, WMM83, WK86b, YMN82]. DM108 [Edm82b]. DM238 [Edm81]. DMSO [FH84, PME87]. DNA [BR89, Col80a, FO82, PC84, RW89, WK86b]. DNA-binding [FO82]. Do [Cur84, GBB81, Bjø88]. DOC [KB82, ZD89]. Does [FDK^{+}86, GM85, GC81, KW86, Paf84, Wot82]. Dokulil [Edm82a]. domains [BMP81]. Dominance [Pae89]. dominant [GMFLJ85, HLA86]. dominated [CGCG89, KTM83]. dominating [LaZ80]. domination [EHE^{+}83]. Dordrecht [Dav86, Mac85]. dormancy [JRG81]. dose [SB80]. Dowden [Man83]. Downing [Bur86a]. downward [Dee88]. Dr [Cat85, Dav86, Edm82a, Edm82c, Edm82d, Fry80, Woo80]. drainage [HTE84]. Drever [Mur83]. drift [AFM86, KC85, MH80, WH80]. drifter [HA85]. drifters [WD87]. Droop [Bro82]. drops [BSW81]. drying [AHRF89]. Dubuque [Mac80]. due [AHRF89, Den82b, EHE^{+}83, HK81b, HS88]. Duinker [Dav86]. Duluth [JSS81]. Dumont [Edm82d]. Dunaliella [OS82]. Duncan [Hay87]. Durand [Hec86]. duration [SC84]. during [BS89, BCL85, CCS^{+}89, Col80b, CDRW86, DP89a, Duc82, DL85b, FHHB80, FP87, Gol88, IDB82, JRG81, KB89a, LB84, LBZC88, LMS87, PU82, Pai86, Pet86c, PA83, RW89, SNM^{+}87, SLL85, SG84, WBEJ85, lWR89]. Duursma [Wan82]. dwelling [O’D85]. dye [FO82, HA85]. Dykyjova [WK80]. dynamic [GH84, PHI84]. Dynamics [Bor87, LI87, Lem89, SL87b, SMK81, WC86, ADP^{+}84, Alt89, AC82, Bor83, Bur84, But82, CK88, Car89, CPT86, Csa80, DWR82, Duc82, GK88, GMK80, HT85b, IP85, LBZC88, LS82b, Lem87, MG83, MKN84, OS82, Ore83b, Sto87, Tay88, VWVR84, WD87, dJV89, Jum82]. Dytiscidae [AMP81].

early [Gol88, Moo88, RK87, SRW81, Jah83]. East [Rud80, VN87, DL83b]. eastern [AN81, BWS^{+}85, BRBL83, DL85b, Le 82, LK89, SLB83, SHRW84, Har83b]. eaux [Edm84b]. Echinodermata [Ham81]. echinoids [Ham81]. Ecological [Leh88a, Pai80, Sei88, WK80, Dod88, Dri81, Fen87, GC81, GK88, Hai88, Jum87a, LS84, MG81, New86, O’D85, Pai84b, Wri89, Fen80a]. Écologie [Edm84b]. Ecology [Fen84, Hec86, EA84, Epp82, Fly88, Hay87, Hob88, Hrb82, Joh89, KH88b, Kon82, KD81, Li83, Lop88, Man87, Men81, New86, Nix88, PAGG87, Pet88, Pra80, Pur81, Rzó81b, Ste88b, Tip87, Wet84, YP89, Ban84b, Fen84, Mur84b, Ban84a, Pai84a, Thi83, Bir88]. ecosystem [BSON84, Boo80, CCS^{+}89, CSW^{+}80, Fen80b, FE81, FF85, Hun83a, Jav83, LWF86, Mac84, McK81, RM86b, SH84, SNP84, TM84, VCG^{+}87, Wet87a, Wet87b, ZKD87, Hec86, Woo80]. ecosystem-level [CCS^{+}89]. Ecosystems [Pom89, BMH85, BMH86, HMLC88, HMC88, Hun83b, KR88, Man88, Nix88, Sei88, Sto88a, Wet83, Wet83, Cat85, Dug80, MS84, WK80, Lam87, MS87]. ectocarpene [JW84]. Ed [Ban84a, Ban84b, Cat85, Edm82c, Epp82, Fen82, Fry80, Gar80, Hec84, Hor83, Hrb82, Koh84, Kru80, Lon84, Lun84, Man83, MS84, Mur84b, Pom80, Ril82, Ser80, Smi86b, Thi83, Ben89, Bez89, Bur86a, Cam87, Edm81, Edm87, Far88, Gil80, Joh89, Kir88, Leh89, Mil87, Mor88a, Pae89, Pai80, Pai89]. edax [Wil80c]. eddies [YAL^{+}81]. eddy [CA87, Lew82b]. edge [AN81]. Ediciones [Edm84a]. Edinburgh [Smi82b, Tut80]. Éditions [Fow80]. Editorial [Edm80b, Edm84d, Edm84c, Edm86a, Jum87b, Jum88a, Jum88b]. Edmondson [CR83, Ano88h, Ano88j, Hut88, Leh88b]. Eds [Bro82, Bur86a, Csa80, Dav83, Dav86, Dug80, Edm82a, Edm82d, EA84, Fen80a, Fen80b, Fen83, Hec86, Jum82, Jum84, Koh82, Lon85, Mac85, Mac80, Mil82, Mil84b, Mor80b, New86, Now84, Pon82, Pra80, Rzó81b, Shi85, Ste84, Tut80, Wan82, WK80, Wet82, Wet83, Wet84, Woo80, ZoB80, Duc87, Har89b, HG88, Lam87, Mol89, Pic82, Pom89, Wet87a, Wet87b]. EDTA [RNVB81]. edulis [FBG89]. eelgrass [Bor87]. Effect [CW88, DPF87, GB80, KA89, LHF^{+}84, MP88, RFS84, SH83, Abe88, Bal85, CM82, CR83, CDPC84, DW86a, DGN82, DP80, EL81, FKM82, Gil89, HL85, JJ87, LO83, MV88, Mon85, MB80, PL89, PGO82, RG81, uRG81, RD83, SP89a, STC88, Sch80, ZD89]. effective [SB80]. Effects [BZH^{+}85, BMH85, BG86b, FP87, Fur87, Gro89a, HH81, Han87, HJP87, Han88, HDW85, KB89a, KRFS84, Lan82, MMTL88, MB89a, MJN84, Run80, SF89, SG84, TS89, WK89, Win81, ZR85, APR80, ATA87, BMS83, CSHO89, CB86, DS87, DH88, DHV^{+}87, DTS^{+}89, Dug81, FB88, Gil88, HK88a, HK88b, HM82, Jah85, JD86b, Kel89, Leh88c, Man80, NDS86, OP83, PB80a, Pas80, RW87, RR80, Rzó81a, SB89a, Sta82, TC85, TD87, Thr88, WD87]. efficiency [BLKH88, BMP83, GCD87, KM83, KOTeI86, LHF^{+}84, Le 82, NO89, SF81, TS85]. effort [Bay88, HH86]. egestion [DW87]. egg [BB89a, Che80a, Che80b, CG80, KN81, Run85]. egg-ratio [BB89a, KN81]. EHP [CSAO86]. Ehrlich [Kru82]. Eisler [Mar83]. electrification [DP86]. electrochemistry [Dav83]. Electron [IB82a]. electronically [BK82]. Electrophoretic [GB87]. elegans [Fei82, Sul80]. element [Bal84]. elements [Bra80, FBF83, San88]. elevation [PB87]. Ellis [Pic85]. Elmsford [Mar83]. elongata [Yen85]. elongatus [PP86]. Elsevier [Kru80, Pom89, Wan82]. elsewhere [Rzó81a, Thi83, Man80]. elucidated [CB87]. elutriation [BQ86]. Embryological [KD81]. emergent [SH84]. Emery [McM86]. Emiliania [Bra82]. emission [KP82]. emphasis [Wet82]. Empirical [PD84, CR80, Kel89, LW84, Pac86, Smi82c]. enclosure [Bur84, MMTL88]. enclosures [BRH^{+}88, IVH^{+}86]. encounter [SFDK84]. encountered [WC81]. encyclopedia [Ill80]. end [PG83, Ree81, RR80]. endogenous [ADF^{+}82]. endosymbionts [CCF89]. endosymbiotic [CCF89, WMM83]. energetic [PM84a, ZST89]. Energy [LL89, Ohm87, Smi81a, THGD83, DP86, KBO84, MLG87, Nix88, Phi80, Smi87]. Enewetak [ASS81, SMK81]. England [HG88, Hrb82, HKT89, JV82]. Englewood [Mil84b, Mur83, Pai84a]. Enhanced [Col80a]. Enhancement [Pet86c, Ger82, PTB83, Pae85]. enhancing [Gel86]. enriched [Har87, IVH^{+}86, LK89]. enrichment [BSW81, HK88a, LP81, LS81, SM84]. entrained [CSW^{+}80]. Entrainment [IB82b]. environment [DC88b, DTBB87, DR88, Gey89, HW85, Høj89, Jac89, Mal82, McC86, Mil84b, PC84, RS86b, Smi84, TS81b, YMN82, Har89b]. Environmental [DL83b, KMK88, MB80, PS84, RG81, uRG81, Rzó81a]. environments [Col80a, HK88a, JEE82, KB89b, Leh88a, LBB88, SMSM86, SNL84, Ste88b, Swa83, Wri89, Fen80a, Kru80]. enzymatically [CSAO86]. enzyme [May89, PG83]. eopepod [Hab85]. epibenthic [AFM86, CH87, SG81]. epifauna [Mul89]. epifluorescence [WHM^{+}85]. epilimnetic [GCLC87]. epilimnion [Col83, GMK80, WED85]. epilithic [Cuk83, Loe81, RLG86]. epipelic [CW85, CW88, Pat89]. epiphyte [SJR87]. epiphyte-macrophyte [SJR87]. epiphytes [CK79, CK80, Cat83, GG81, MK84]. epiphyton [Bor87, CK82]. Epischura [SF89]. episodic [HDLP85]. equations [MF89]. equatorial [MM82a, SHRW84]. Equilibria [Whi86]. equilibrium [BK80, BK84, Sta86]. ergosterol [LHH80]. Erie [IB82b, IP84]. Ernst [Mil84b]. erodibility [Pat89]. erosion [ADO85]. erosional [DF82b]. Errata [Ano82f]. Error [HK81b, BSN83, BF84, CS81]. errors [KSP^{+}81, Li83]. Escape [GK88]. essay [Boe89]. ester [UB89]. esters [CVG89]. Estim [SGSS84]. estimate [Fra87, RO82]. Estimated [BS89, KN81]. Estimates [DL83a, SBN^{+}83, CDG88, MD87, MMR87, OH87, WK86b]. Estimating [Ebe81, MMM81, OØ85, Rod87, TS81a, TG82, FO82, HH86, KM80, LHH80, Li83, Men80, SW89]. Estimation [Ban88, Dee88, KSP^{+}81, Lyn83, SMP^{+}89, MG83, Vic84, War84]. estimators [HL81]. estuaries [BR86, Fro88, MDR88, NMM83, PH87, dAL87]. Estuarine [Fen82, ADF^{+}82, Bru89b, Cof89, CFPW89, DTBB87, DMM82, DP89b, EHE^{+}83, GM84, HM82, HMLC88, HMC88, HL82, INA89, JK84, Jon82, KSP^{+}81, LL82, LL85b, Man88, Men81, MJSM89, Mil84a, Mil86, Mil87, Pae88, PS89, Ste88b, SH83, TC84, dJV89, vdBD84, Ril82, Gar80, Fen82, Wet82]. estuary [BD81, CD85, CSF88, CM80, Duc82, FSW85, GL81, HS80, HP87, IBC^{+}83, NL87, SCC82, SH87b, Cor81]. Ethological [HSMH87]. etude [Fow80]. Eucalanus [PP86]. eucaryotes [BMH86]. Euchaeta [Yen85]. Eukrohnia [Sul80]. Euphausia [BW86, HHOC89, Ros81, RDE82]. euphausiid [BW86, KMSF81, McC86]. euphotic [CB87, HK81a, MB89b]. europas [Edm81]. Europe [Edm82c]. eutrophic [AR88, BU86, BFKN84, CAS83, DW86b, ES82, HLA86, JBvdB89, LWF86, MHY83, OAI^{+}85, Sta82, VHJ^{+}89, Wof83, WL85b]. Eutrophication [SN85, Fee79, Gol88, Hec84, Pat80, Sch80, SWSC85]. Evaluating [Thr81]. Evaluation [Col84a, WM82, CB84, HK80, Kin83, Mor84, Pae84, SC82, VW80, WLG85]. evaporated [LSK^{+}83]. evaporilic [Kru80]. Evelyn [Ano82g, Ano84e, Ano85e, Ano85f, Ano86h, Ano88d, Ano89j]. event [PA83]. Evidence [CCF89, HW88, JW84, KK80, RM83, VG80, GH84, Gil88, HK88a, KCN82, OW84, SH88, Wol87]. evolution [GK89, Pon82, Hrb82]. evolving [Jum87b]. examination [Dri81, GS81, HS84b]. examined [OBP87]. example [CDG88]. exchange [DQRM87, Fra88, Mac85, PB80b, RJB81, SLB83, SMK81, WMSW89]. exchangeable [BM82]. exchanges [Boo80, HS80, VTV^{+}80]. excitation [CI88, MH87]. excretion [Båm85, BBP82, Smi82d, Ver85, ZD89]. Executive [Ano89a]. exerted [Den83]. exoproteolytic [SB83]. expenditures [Gig80]. Experience [Rzó81a]. experiment [Blo81, Cuk87, Gro89a, LMS87, SA89, SS88b]. Experimental [BZH^{+}85, DHV^{+}87, WBG^{+}87, BB89a, DW86b, Elm83, OØ85, Pet88, Sch80, SM84, SNR86, TD87, TML81]. experimentally [CKST86, PS89]. Experiments [Som85, ZFFC88, Fee79, GG84, GLMA85, Gil89, HM80, Law85, LCK89, MMTL88, Pat80, QBHS80, RR80, SB80, SK85b, Som88]. explanation [PU82]. exponential [McL86]. export [KCN82, LG89]. exposed [DH88, IP85]. Extended [SP89a]. extension [LM86]. extinction [AB80, Wof83]. extra [HM89]. extra- [HM89]. Extracellular [MFHM80, May89, Bel80, BS80b, Bel83]. Extraction [FH84, PME87, SF84, Fra87]. exudates [SvdBR82]. exudation [Bjø88, CC84]. Eyrolles [Fow80]. eysenhardti [Pur81].

F [Bur86a, Edm82b, Edm87, Gar80, Hor83, Lec85, Mol89, Pae89, Pic85]. F. [Edm82e, Fow80]. Fabricius [Men81]. Fact [HM80]. Factors [Elm83, FDR^{+}88, FSW85, Klo85, San88, Wal85, BK87, BMP83, FVH^{+}84, GMPA^{+}85, HF88, LO81, PAGG87, RG81, uRG81, SH89, SW89, WBG^{+}87, YTW84, ZBM82]. facultatively [SSS87]. Failure [RG83]. failures [LG88]. Fairbanks [Ano88f, Ano89g]. fall [Atk88, Her82, Nür88]. fallen [RD89]. falling [Jac89]. Fallout [SMSM86]. Farley [Har89b]. Fast [Atk88, Ano88k]. Fast-response [Atk88]. Fate [HACF85, FCC^{+}86, For85, GCLC87, LBZC88]. fauna [KC85]. faunas [RJ84]. Favre [Csa80]. Fayetteville [THCP81]. Fe [Lyl83, Kin83, Lyl83]. features [MP88]. February [Ano81e]. fecal [DW86a, DL85a, Hab85, KMSF81, TNJ84]. fecundity [RDE82]. fed [HGS83, HS84b]. Fee [Pat80]. feed [Paf84]. feeder [FBG89, NJF84]. feeders [PB87, Phi80]. Feeding [DeM82, Fei82, HGS83, LE89, Pur81, BG82, CS83, CM81, DG80, DFW89, DeM89, DL85a, DdNK82, DMdK84, DHV^{+}87, Fen80c, GSBP87, Håk87, Hak84, Ham81, Han85, HM80, Hun81, KS81, MB86, MTL84, McC86, MMH85, MJN84, Moo88, Mul89, PD84, PGO82, RFS84, RD83, Ros80, Run80, SM85, SC88, Sul80, Tag88, TD87, WB86b, Yen85, YTW84]. females [DG80]. femorata [LE89]. fermenting [KCN82]. ferrous [DH80]. fertilization [MMTL88, Sch80, SS88b]. FeS [TAJ89]. fiats [PB87]. Fiber [JD88, JD86a, Yen83]. Fiber-optic [JD88, JD86a]. Field [Cuk87, Fra88, HHOC89, WCEP82, Hak84, HF82, IP85, KA89, McK81, WM82, Yen85, lWJ82]. field-caught [Hak84]. fields [MMP86]. Fifth [Fen82]. filamentous [MM80, PL89]. filaments [PM84a]. film [DP81, VW83, Wot82]. Filmed [Ros80]. films [BHK^{+}82, DP86, VW80]. filter [BSN83, DdNK82, DMdK84, DHV^{+}87, HHH83, Paa80, Ros80, TD87, Ven87]. filter-feeding [DdNK82, DMdK84, DHV^{+}87, TD87]. filter-retained [Ven87]. filter-transfer-freeze [HHH83]. Filtering [GS86, DP80, Edw87, Han85, HK81b, KH86b, KH86c, PL89, PD84]. filters [HHH83, TC84, Yen83]. filtration [JW85, McL89, WCEP82]. Final [Edm86a]. Financial [Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89b]. fine [MDM80]. fine-structure [MDM80]. finmarchicus [SLL85]. first [Lan80]. Fischer [Den82a]. Fish [HLM86, Rzó81a, Thi83, BF89, BFKN84, CPT86, DMdK84, DTS^{+}89, FCO87, GR84, HJP87, KR88, LO81, MMTL88, MS85b, Pur81, SP85, TD87, TM84, Vic84, Whi81, WL85b]. Fish-flamingo-plankton [HLM86]. fisheries [Bay88]. fishes [FF85, MS85a]. fishponds [vRS83, vRS85]. fitness [WK89]. five [Paa80]. fixation [BCM86, DSH82, GMM88, HG85b, HMLC88, HMC88, Jon82, LDYY83, LL84a, LL85a, LL84b, Pae84, Pae85, PCP87, RJB81, SNM^{+}87]. fixed [Pet86c]. fjord [Lid83, WL85a]. fjords [TNEJ84]. flagellate [SNR86]. flagellate-ciliate-copepod [SNR86]. flagellates [AF85, SPBD89]. flamingo [HLM86]. flat [Atk87, AS87, GCAN87, Pet88, Smi81a]. Flathead [Moo83]. Flatworm [DC87]. flavicans [Lue88]. flavor [RFS84]. Fleeger [BW89]. flexible [Bur84]. flocculated [TS89]. flocculent [Hun83a, SRK86]. flooded [HB81]. flooding [HM82]. floodplain [FDR^{+}88, HL87, RDW^{+}88, SM84]. floodplains [Bay88]. Florida [Bro83, CS85, Dee88, Elm83, FE81, VCG^{+}87, Whi89]. flos [HS84b, OW84]. flos-aquae [HS84b, OW84]. Flow [BCY89, YHM^{+}83, AC82, Bot85, GSBP87, GCAN87, HH81, HF82, HDK84, IB82a, JP83, JW85, JV82, Mac86, NCY89, PP89, PHCK83, PH87, SAJB86, Som84, SCO89, SW89, SG84, VTV^{+}80, WHM^{+}85]. flow-measuring [Mac86]. flow-through [Som84]. Flower [DL83b]. flowing [HT85b, MPsCS80]. flows [Csa87, FBG89, Han84, Kir85]. fluctuating [GHK80, OP83]. Fluctuations [Som84, CM83, CM81, GJP89, KB82, KB89a, Two83, WL83, vRS85]. fluid [AC82, Csa87, GK88]. fluminea [CDPC84]. Fluorescence [SW80, ARP82, CI88, CMB87, Cru87, DS87, EHE^{+}83, KCB89, MH87, NCY89, SCO89, SW81, SG84, Yen83]. fluorescent [FO82, MF86]. fluorochrome [Col80a, Har87]. fluorochrome-staining [Har87]. fluorometer [Ano88k, BSN83]. Fluorometric [Car89, CS81, LD86, Ven87]. fluviatilis [LB80]. Flux [DL85b, CW88, DFW89, Hun83b, LMS87, Lew87, MM82b, RMB82, RD89, San84, SL87b, SKW86, Smi87]. Fluxes [HCC88a, Boo80, CCW87, Csa80, DWR82, HTWR84, HCQ^{+}86, KSP^{+}81, KM81, LCPL88, LSdKvD85, MMD^{+}80, RS86c, RW87, SEK84, SS85]. fly [MP88]. focusing [DF82a, Hil85, HLA86]. following [Col82]. Food [Che80b, DDSmV83, Fry88, KS81, Smi87, WBF85, All81, BLKH88, BS85a, BH84, Bor83, Che80a, CS83, DW86a, DeM89, DP80, Edw87, Han85, JJ87, LHF^{+}84, Leh88c, McC86, Nag88, NO89, OJR^{+}86, Pas80, PGO82, RFS84, RD83, RO82, SNR86, SS88a, SRW81, TC85, TG82, TTF88, VP85, WKLK89, Wot82]. food-limited [Bor83]. food-size [VP85]. foraging [DeM89, HSMH87, NJF84, ZF89]. Foraminifera [Ber89]. Foraminiferan [JERC85]. force [Den83]. forced [Mon85, Rah85]. Forces [Den82b, Den85]. forcing [GJP89]. Forested [LBB88]. forests [RD89]. Fork [MTW^{+}88]. formalism [Li83, SP89a]. Formation [Jah85, MH89, Mul81, DP81, Gar81a, HM84, JC80, Kin83, Pae85, RNVB81, SRW80, SW82, TAJ89, UB89, VKW89]. forming [Moo83, Pae89, PAGG87, YAL^{+}81, vRS85]. formosa [HA81]. forms [WST80]. Forth [HP87]. fossil [LCK89]. four [BGJ^{+}87, BBS80, GG82, HI88, JRC83]. fraction [RMN82]. fractional [Til83]. fractionated [Fur87, PP85]. fractionation [BK80, BK84, LMS87, RO82]. fractions [BP87, MC82, WLG85]. fragile [JZF80]. framework [Eva88, Hil85]. Framvaren [Lid83]. Francisco [MPsCS80]. Fraser [WS81]. Free [Hal82b, Jør87, Atk88, Cof89, GCLC87, Ham82, HFL84, Her82, ML82, RS89, SM85, Sim87, SH87a, TS89]. free-fall [Atk88]. free-living [Sim87]. free-swimming [RS89]. Freeze [CBH83, HHH83, Pae84]. Freeze-out [CBH83]. freezing [MMW86]. frequency [AH87, ADF^{+}82, GWB89, HH86, KM80, MPsCS80, WL83]. fresh [Bur86a, Hob88, LL81, San88, Gil80]. Fresh-water [Gil80]. Freshwater [MDR88, BZH^{+}85, BC82, BMH86, BK80, BK84, BSW81, CK88, Cat85, CVG89, CMK89, CMN82, CKT89, CK84a, CK84b, DCV87, DeM88, DC83, Gro89b, HWL83, Hal82a, Hal82b, HK88a, HMLC88, HMC88, JT84, JW84, KR88, KH88b, Kir83, KM84, LS82a, LS85, Leh88a, Lop88, LBB88, Man88, McK81, Mur84a, OKG81, PO81, PU82, SPBD89, SMH84, SM85, Sei88, SK83b, Ste88b, Sto88a, Tip87, VP85, Wet84, WLG85, WC81, WWS84]. freshwaters [And80, BR89, GB87]. frontal [YAL^{+}81]. Fryer [Edm82b]. Fryxell [Lun84]. FTF [HHH83]. fucoid [CC84]. Fuhrman [KW86]. fulvic [CSHO89]. function [GSB89, LS83]. functioning [Pat88]. Functions [SW82, NG83]. fungi [LHH80]. furciliae [HHOC89]. Furtado [Wet84]. further [Mal83].

G [Ano86j, Ano89i, Edm82b, Gop89, Joh89, Lam87, Lon84, Lon86, Lun84, Mac85, Mil84b, New86, Wet87a, Wet87b, ZoB80]. G. [Ano82g, Ano84e, Ano85e, Ano85f, Ano86h, Ano88d, Ano89j]. Galapagos [Ano86j, BT84]. Gambia [LHM84]. Gammarus [LR85, WH80]. Gardens [DL83b]. Garland [Fen80b]. Garrett [JC81]. Gas [CMK89, PB80b, DQRM87, DMM82, HHH87, HFF87, OW84, RS88, SNL84, vRS85]. gaseous [RS88]. gases [BK80, Mac85, RDW^{+}88]. gastropod [Har89a, HP89]. Gebruder [Edm81]. Geer [VKW89]. gelatinous [CGCG89]. Gelidium [AC82]. Gen [Fen83]. General [MP89, MF89, New86, SCC82]. generated [Ger83b, KB82, MBT80]. generation [HFF87, Kel88, Pet86b, Pse88, UB89]. Genetic [Bra82, BW86, PC84]. Geneva [BC86, LG88, Lem89]. genotype [DD89]. genotypes [Wei84]. Gent [Edm82d]. Geobiology [Now84]. Geochemical [RK87, Ban87, FTKB87, GK89, Sei88]. geochemistry [BRBL83, Hen87, JM85, Mur83]. geographic [MRE84]. geologic [Mur85]. geology [McM86, Ser80, Mol89]. Geomicrobiology [Kru82]. Geophysics [Ril82]. George [Ill80, Ill80]. Georges [Fry88]. Georgia [Fen82, BMH85, HG83, Kin83, PH87, SH84]. Geothermal [RJ84]. geothermally [PSGD86]. Gephyrocapsa [Bra82]. Germain [Fow80]. Germanic [DSH82]. germination [ATA87]. Geukensia [JV82]. giant [MD87, ZR85]. gills [TQS86]. Gilmer [Har89a]. Ginsburg [Smi82b]. Girard [Cap82]. gizzard [DdNK82]. glacial [SM82b]. glass [Ban88, WBV80, Yen83]. glass-fiber [Yen83]. Glaude [Rzó81a]. global [MD87, Smi86b]. Globigerinoides [JERC85]. Glucose [HS80]. glucosidase [Chr89]. Glynn [Ano86j]. goal [Leh86]. Goldman [Fre84]. Good [Leh88b]. gorgonians [HF89]. Gough [CK81]. Gower [Hor83]. gradient [OKG81, TML81, TD83]. gradients [HP87, Pac86]. grain [DM84, DM83, Ham81]. grass [Mor80a, Ste88b]. Gravity [Gey89]. Grazing [BSC80, Cat83, Cuk83, Mor80a, RE81, Ver85, All81, BCL85, BBG89, BG82, CRGC87, DHV^{+}87, Fie82, Gro89a, HJP87, HK88b, JRG81, LFRT86, LS85, MD85, Mur84a, NL89, RR80, RI88, SLL85, Som88, TM84, WC86, WL85a]. Great [Pom80, HH81, HI88, LL81, SA89, WST80]. Green [Edm82d, THCP81, FB80, FP87, Lyl83, McA86, MM80, PU82, PTB83, Pae89, PM84a, STC88, Wil84, Sta85a]. Greifen [LMS87]. gross [GCD87]. Groundwater [Lee80, LCP80, Boo80, Lew87, LKS89, VTV^{+}80, WLR88]. Groundwater-solute [Lee80]. growing [CK79, GG81]. Growth [FDW85, HT85b, KM87, MLG87, SRW81, Str80, VJOR88, AK89, Ban82, BC80, BS80a, BLKH88, Bor87, Bot85, BS85b, CPT86, CJTL83, CM81, DL83b, DH85a, DH85b, Ebe81, EEMC88, FCC^{+}86, GB80, Gol86, GCD87, Gro89b, Har83a, HF81, HW86, JRG81, JY84, KM83, Koh82, KW87, KDC86, LB80, LRH^{+}84, LDR87, Le 82, LS85, LH87, Mal82, MC82, McL86, MS85b, OSML87, PO80, RG81, uRG81, RM86a, RW89, Rod87, She84, Som89a, SYB^{+}89, SB89c, Str83, SH83, TG87, TLH89, TM84, Ver85, WL84, WL85a, ZBM82, ZR85]. Growth-irradiance [FDW85]. Grupenhoff [Har89b]. guard [Rev89]. Guegueniat [Fow80]. Gulf [DL83b, GBF86, Han87, HLS87, Sau80, SSS87, SG84, YAL^{+}81]. gut [CDRW86, DW87, WC86]. Gymnodinium [Fie82]. gypsum [KSTGB81]. gyre [KCB89, LDR87, Wei80, SKW86]. Gyres [WL85a].

H [Ano89i, Bro82, Bro83, Bur86a, Den82a, Duc87, Edm82a, Edm82c, Edm82d, Edm82e, EA84, Fen84, Gop89, Jum84, Kru82, Man80, Pom89, RW89, Shi85, Ste84, TNEJ84, THCP81, Wel88, PM88, RJB81, RJBC83, RBNF87]. habitat [Han81, HB84, HH85, KB89a, Wei84]. habitats [SB89c]. haemulid [MS85a]. Hague [Cat85, Edm82a, Edm82c, Edm82d, Fry80, Hec84, Hec86, Hyn81, Mur84b, Wet83, Wet84, Woo80]. halide [MD87]. Halimeda [SMK81]. halite [GK89, GSB89, KSTGB81]. Hall [Mil84b, Mur83, Pai84a]. halobacteria [Ore83b, Ore85]. Halodule [MK84]. Halvorson [Shi85]. hamata [Sul80]. Hamilton [Wet82]. Hampshire [Hrb82, DF82a, RK87]. Handbook [Bur86a, Fen80b, Mor80b, Mil87]. Hanover [HG88, Hrb82]. Harbor [JSS81]. hard [CBH83, Kon89]. hard-water [CBH83, Kon89]. hardcover [Fen84, Jah83, Man80, Smi82a]. hardmannianus [SP85]. harmful [JD86b]. harpacticoid [HH85]. Harpacticoida [LL89]. Harper [Pai84b]. Harrison [Hay87]. Hasselmann [Csa80]. Hatteras [Fis80]. Haven [Hal82a]. Hawaiian [LRH^{+}84]. Head [Mil87, WLR88]. headlands [ER88]. heads [RM83]. headspace [SRW84]. headwater [Mor84]. headwaters [Cor81]. healthy [Bjø88]. heat [Lew83b]. heavy [SSK87, TB80]. Heidelberg [Dug80, Ser80, WK80]. held [Fen82, Hrb82]. Helix [SA89]. Hellebust [Mor80b]. helmets [Rie84]. Hemidiscus [SP85]. hemipelagic [JEE82]. hemoglobin [LS83]. Hemond [BB88]. herbivore [CL80]. herbivores [Leh80, RO82]. herbivorous [DL85a, OSML87, TS85]. herbivory [Bor87]. hermatypic [DL83b]. Heterocope [HGM80, LO81, LO83]. heterogeneities [Iba81]. heterogeneity [BW86, LL85a, Smi86a, Wei84, Wil80a]. heterotrophic [ACC86, BRH^{+}88, DH85a, DH85b, LB84, LT88, SSFN86, SYB^{+}89, VHJ^{+}89, Duc87]. heterotrophy [BMS^{+}88, VG80, ZHO^{+}87]. High [ADF^{+}82, CMR89, LRH^{+}84, LDR87, MPsCS80, BKB85, CA89, GWB89, Gri88, HMG86, HW85, HM82, JD86a, Pai86, RJ83, RMJ86, Smi81b, VWVR84, WL83]. high-altitude [VWVR84]. high-frequency [GWB89]. high-performance [BKB85]. high-precision [Gri88]. higher [LaZ80, OAI^{+}85]. highly [GSB89, Sir87]. Hill [Fre84]. Historical [ZEY89]. History [Pic82, DH88, Edm84d, Jum87b, Mur85, O’D85, OP83, Smi88, SRW81]. Hobbie [Duc87, Man83]. Hole [Pic82]. Holocene [DWO86, RK87]. holoplanktonic [Har89a]. holothurians [Ham81]. Horizontal [BU86, CA83, AATC82, IH81, Lem89, LL85a, Unc81]. Horne [Fre84]. Horwood [Pic85]. host [CCF89, DC88b, Dou87, Dou88]. Houghton [Smi86b]. Houtman [Smi81b]. Howard [Woo80]. Howard-Williams [Woo80]. HPLC [GK83, ML82]. Hudson [MRLD86, RI88]. humic [CSHO89, CS81, EHD^{+}86, FH82, Fra87, MTW^{+}88, RGS89, TW83, TS89]. hunger [DD89, Pas80, Run80]. Hunt [Mac80]. Hutchinson [Ano82g, Ano84e, Ano85e, Ano85f, Ano86h, Ano88d, Ano89j, Man83]. huxleyi [Bra82]. hydra [McA86, CM81, SC81a]. hydration [Joh82]. hydraulic [WLR88]. hydride [HFF87]. Hydrobiol [Cat85]. Hydrobiologia [Edm82d, Mur84b]. Hydrobiology [Edm82a, Mur84b, Wet83, Edm82d]. hydrocarbon [BMS^{+}88, BR86]. hydrocarbons [Sau80]. Hydrodynamic [Eck83, Luc86, Ste88a]. Hydrogen [NSM87, PZM87, SNM^{+}87, CKST86, MPF88, SAJB86, SSS87, SNL84, WWS84]. hydrogen-producing [SSS87]. Hydrographic [HN81, MMP86]. hydrography [ZR85]. hydrologic [HI88, LW84, PM84b]. hydrolysis [RWB87]. hydrolyzable [CSAO86, HFL84, IDB82]. hydrothermal [BT84, Smi85, VFB87, Lon85]. Hymenomonas [VAI85]. Hynes [KM80, Men80]. hypereutrophic [BA87]. hypersaline [FVM82, GK89, LSK^{+}83, Kru80]. Hypertrophic [Wet83, Pra90, Som89b, ZBM82, Wet83]. Hypolimnetic [Leh88c, CR80, Cor89, LN86, Lin87, Nür87, PSGD86]. hypolimnia [Nür84, Nür87]. hypolimnion [GM85, IJ84, IB82a, MDM80, QBHS80]. hypotheses [Mor88b]. hypothesis [OBP87, Ols87, SNR86]. Hypothesized [KKH86]. Hysteresis [WS81].

IBP [Bur86a, Man83]. Ice [Hor81, AN81, BBH85, CA83, CA87, HHOC89, KE80, MRLD86, PS85, PH88a, PPVHW87, Rah85, RI88, SA81, SVG89, SCH89, SM82b, Top87, WMSP86, WM87]. ice-covered [CA83, CA87, PPVHW87, Rah85, Top87, WMSP86, WM87]. ice-edge [AN81]. Iceland [ES83, Óla80]. Ideas [Lon84]. Identification [CI88, HHH83, SF84]. identifying [PF80]. II [Lyl83, RNVB81, SB89b, TNEJ84, Car85]. III [Lyl83]. illuminated [LSdKvD85]. illustrated [Ill80]. imagery [Mor88b, PGB88, YMBO87]. Imberger [Den82a]. Immediate [Iba81]. immunofluorescence [CI88, War84]. impact [All81, DdNK82, DMdK84, LW84, Wro89]. impaction [JW85]. impacts [HG88]. Implication [PC84, CB86]. Implications [BLKH88, CEE86, DL85b, Kir85, Lan80, LWP85, Mul89, Rue83, WKLK89, LB80, MG81, SB89c]. Importance [FS84, JY84, Pat88, BK87, CM89, CK84b, FBG89, Hai88, HMLC88, KRCS82, MSG86, PO81, PCP87, Thr81, Til83, UEB88]. important [CLS82, EDKS86]. impoundments [Rzó81a]. Improved [BH88b, Jef81]. improving [Kel88]. in-lake [Pse88]. Inc. [Ano80a, Ano81a, Ano81b, Ano81c, Ano82a, Ano82b, Ano83a, Ano83b, Ano84a, Ano84b, Ano85a, Ano86a, Ano86b, Ano86l, Ano87a, Ano87h, Ano88a, Ano89b, Koo85]. include [SP89a]. including [DS87]. income [CR83, EL81]. Incorporation [Hun83a, BA87, BR89, CW84, DL83a, DH85b, FDK^{+}86, KW86, RBNF87, RW89, WGM82, ZFFC88]. incubation [MH84, Nür87]. incubations [SNM^{+}87]. incubator [ZPVM83]. independent [BB89a]. index [BS86, BS85b, DL85a, KA89]. India [SP85]. indications [LG88, RK87]. indicator [PS84]. indicators [DP89a, FP81, Whi89, ZBM82]. indices [KR88]. Indirect [DH88, HG88]. Individual [GSBP87, CI88, Doy80, IMK88, PP89, WK89, YJ84]. individuals [MSG86]. Indonesia [VN87]. Induced [HGM80, DP89b, GS84, HG85a, HMS86, LLG86, Ste88a, WK89, WL83, Wil80c]. induction [GB81, KD81, VKW89]. inert [MF86]. inexpensive [Jah88, RS89]. infer [CS88]. inferences [NCY89]. Inferring [Mur81a]. infiltration [HB81]. inflow [FS83]. Influence [BSF81, PB87, AK89, AM82, AMP81, BFKN84, DTBB87, Dou87, Dou88, GM88, HASI80, IBC^{+}83, Jon87, MG83, SA81]. influenced [PSGD86]. influences [Cuk87]. influencing [CWD^{+}86, Elm83]. influx [Lee80, SP89b]. Infrared [RS88]. Ingestion [D’A80, DW87, DeM82, DMdK84, MF89, PGO82]. inhabiting [ACC86]. inherent [Kir84, Zan89]. Inhibition [IA85, DSH82, PL89, Rue83, lWR89]. injection [JP83, SAJB86]. injuries [Mur81a]. inland [Den82a, Pae88, RPC89]. Inlet [HW85, HA85, Dev83]. inner [PO80]. Inorganic [AGG82, HW88, RLG86, RJ84, SRW80, DQRM87, Har83b, HS82, HF88, HY89, LH80, MM82b, Neh87, OAI^{+}85, RKF86, TS82, WK86a]. input [KC81]. inputs [Boo80, Nix88]. insects [HK88b, Mac80, EA84]. inshore [EJ86]. Insights [DC88b]. insolation [BF87]. insoluble [LH80]. Inst [Ban84a, Koh84, Thi83]. instantaneous [BB89a]. instar [Moo88]. instrument [LSI^{+}85]. Instrumentation [TML83]. Int [Dav86, Edm82d, Pic82, Smi82b, Tut80]. intact [Chr84]. intensity [DKMvL86, NG83, PB80a]. Interaction [DD89, Bel83, BH88a, Csa80, FM82, RM81, TC85]. Interactions [FBF83, HF88, JBvdB89, BB86, BRH^{+}88, CK79, CPT86, Csa87, Cuk83, GG81, Han88, HHTV81, HLM86, Lam87, Leh89, NSL86, Pai89, Pet88, RS86b, RG81, uRG81, Sel81, SH83, TB80, WKLK89, Wil87]. Interannual [GJP89, Hai88]. Intercolony [HF89]. Interdependence [DTS^{+}89]. interface [DP81, DP86, KA89, RPL89, RS86c, vdLAH^{+}84, Mor88a]. interfaces [TNEJ84]. interfacial [BMH^{+}87]. Interference [OC82, BG86a, BG86b, GS85, Gil89, Ols80b]. interior [Høj89]. intermittently [HB81]. Internal [SSGB89, WC87, FHHB80, HMS86, LaZ80, LSM81, LM86, Lem87, Nür84, Nür87, RW87, Sir87, Sta85a]. International [Fen82]. Interpretation [SCO89, Jum87a]. interpretative [GL81]. interpreting [Fee79, Pat80]. interreplicate [DPF87]. Interscience [Whi86]. intersite [Bot85]. Interstitial [Car84, FP81, GB82, vdBD84]. intertidal [Cor81, DM83, Den82b, Den83, Den85, HP87, HM82, Jon82, May89, Mur84a, Pai80, Pat89, SB89a]. intervals [KN81, KM80]. intracellular [HM89, SGSS84]. intraspecific [DM84]. introduction [Mac80, Man87, Pla86, Wri89, Dav83]. Intrusions [Chu85]. Inverse [Ker88, PM84b]. inversion [SW89, WC81, Zan89]. invertebrate [DC88b, Ive88, MH80, SC88]. invertebrates [Edw87, GFCP85, Gil80, HB84, Ill80, Sel81, SP85]. investigating [Ano88k]. investigation [SNR86]. involving [CS81, SH87a]. inwash [BF84]. iodine [BSF81, CL81, JT84]. ion [CKST86, EH88, KH80, SH87a]. ionic [HB84]. ionie [Hal82b]. ions [KB80b]. Iowa [Mac80]. Ireland [DF82b]. iron [AM82, BFKN84, BSG83, Col83, DWR82, FM82, FH82, GH84, HM86, HK86, HM89, MMG86, NL87, PCP87, STC88]. iron-rich [NL87]. Irondequoit [WB86a]. irradiance [FDW85, KGGW88, PM84b, SW89, Zan89]. irradiances [MLG87]. irradiances12 [RV87]. Island [CD85, Chu85, Fen82, HG83, PH87, Ril82, ZR85, Bri82, HH81, DDSmV83, MMD^{+}80]. islandica [SWSC85, TCN^{+}82]. islands [Fre88, Ano86j, Smi81b]. isochronal [Lan83]. isolated [RS81, ZST89]. isolating [TD83]. Isolation [BQ86, McA86, FW82, Ger83a, SB89a]. Isotope [GLMA82, Law84, Bow84, CSF88, CCF89, GLMA85, LS82a, LMS87, QEQD86, RJ84]. isotopes [MRE84, OAI^{+}85, PH87]. Isotopic [POHMD84, AHRF89, BK80, BK84, CFPW89, Fry88, NL89, WWS84]. Israel [MWB82, DB81, GK89, HB84, SSGB89]. issued [Man80]. issuing [BB84a]. IUCRM [Hor83]. ix [Bur80, Pai80].

J [Ban84b, Ben89, Bez89, Bro80, Bur86a, Dav86, Den82a, Duc87, Edm82d, Edm84b, Fen80a, Fow80, Fre84, Har89b, Hay87, Hec84, Hec86, Hor83, Ill80, Koh84, Lam87, Man83, Mil84b, MS84, MS87, Mor80b, Mur83, New86, Now84, Pai84a, Pai84b, Pat80, Pom89, Pra80, Rzó81b, Smi86b, Smi88, WK80, Wet83, Wet84, Wet87a, Wet87b, Woo80]. Jagner [Dav83]. James [Smi82b]. Jannasch [Bro82]. January [Ano80a, Ano82a, Ano85a]. jar [Han81, LSdKvD85]. jar/microelectrode [LSdKvD85]. Jefferies [Fen80a]. Jekyll [Fen82]. jellyfish [DC83, Har84]. Jersey [MMG86, Mor84, Mor88a, Wel88]. jet [BSW81]. Jewson [Edm82a]. John [Ano84h, Cam87, Cha85, Dav83, Ill80, Lec85, Lon84, Mil82, Pic85, Pla86, Pra80, ZoB80]. Johnson [Gar81a]. Joint [Ano81e]. Jorg. [RE81]. journal [Edm84c]. Jr [Bur80, Lon85, Pom80]. July [BS89]. June [Ano80e, Ano81f, Ano82h, Ano83f, Ano84h, Ano85d]. Junk [Cat85, Dav86, Edm82a, Edm82c, Edm82d, Fry80, Hec84, Hec86, Hyn81, Mur84b, Wet83, Wet84, Woo80]. just [Jan81]. juvenile [TG87].

Kalff [GG81]. Kalk [Woo80]. Kattegat [IJ85]. Kelly [Sto88b]. kelp [Dug81, KCN82, MD87, ZR85]. Kendall [Mac80]. Kendall/Hunt [Mac80]. Kennedy [Fen82]. Kent [Pai80]. Keratella [BG82, BG86a, BG86b, GS84, GS85, GK88, MG83, Ste88a]. Kerfoot [Hrb82, HG88]. Kiefer [Edm82b]. Kiel [Bal84]. Kilham [Ano89k]. kills [Whi81]. kind [NCY89]. Kinetic [Fro88, GS81, KM89, Bel80, BS80b, Li83].

Kinetics [Kep85, BR86, Gar84, Gro89b, HAvdL^{+}89, Joh82, NSL86, RNVB81, TML81, VO89, WGM82]. King [GB81]. Kinneret [MWB82, SSGB89, Avn83, BSS^{+}84, DB81]. Kirk [MS87, Pom80]. Knowlton [CP89]. Knoxville [Ano81f]. Koh [Den82a]. Koste [Edm81]. Kramer [Dav86]. krill [Pri89]. krohnii [BW86]. Krøyer [CS83]. Kvet [WK80].

L [Bur88, Fen80a, Kru82, Lon85, Now84, Wet83]. L. [Bor87, FH85, FBG89, JBvdB89, KLJ86, Str83]. label [HK81b]. labeled [FH84, HD86, KP82, OSML87, TAJ89]. labeling [HFA80, NO89, RW89, WL84]. labile [EHE^{+}83]. Laboratory [BW89, Che80a, Nür87, Ros81, RF88, SVG89, Yen85, DW87, Gey89, VW80, WCEP82]. Labrador [Hun81]. Lachm [RE81]. Lack [Bal85]. lacustrine [AHRF89, DR88, DWO86]. lacustris [SF89]. lagoon [ASS81, FVM82, HH81]. lagoonal [AM83]. lagoons [Bay88, Smi82a]. Lagrangian [CRFB80]. Lake [CR83, CCO83, CCW87, DGN82, DF82a, DWO86, FP81, IOK87, JS83, LS83, LSM81, MTW^{+}88, Mur85, SO87, Sta85a, THCP81, ZHO^{+}87, AR88, ADO85, AGG82, BF84, BA87, BBG89, BU86, BKF^{+}87, CF81, CS81, CW88, CWD^{+}86, CMN82, Chr89, CCS^{+}89, CA83, CAS83, CKST86, CPT86, CM81, Cuk83, Cuk87, DM85, DWR82, Edm82c, ES82, Fee79, FE81, FH82, GMM88, Gel86, HL87, HGH82, HL85, HLA86, HG85b, IP85, IVH^{+}86, JBvdB89, JL80, JW84, KRCS82, KRFS84, Kel88, Kil82, Kor80, LaZ80, LP81, LCK89, LS82b, LN86, Leh88c, Leh89, LI87, Lew82b, Lin87, MM82a, MH86, MHY83, Nix88, OAI^{+}85, Pac86, Pai89, PB81, Pat80, PH88a, Pra90, PSGD86, Pse88, Rah85, RGS89, RLG86, RW87, RKF86, SM84, Sir87, Som89b, Sta80, SM82b, SS88b, Sto87, TW83]. lake [Top87, Tut80, VHJ^{+}89, VWVR84, WRS80, WMSP86, WM87, WC81, WC87, WWS84, Wot82, WL85b, Zül81, APR80, ADP^{+}84, Abe88, Avn83, BSS^{+}84, BC86, BWG83, DB81, EL81, EL82, ES83, EJ86, FB80, FHHB80, FS83, Fry80, GNM87, GCLC87, GCL89, Gol88, Hab85, Ham82, HH80, HK81a, HF81, Hec84, Hec86, HMS86, IL85, IA85, IB82a, IB82b, IP84, JEE82, KM84, KMD^{+}88, LG88, LMS87, LE83, Leh88c, Lem89, LL84a, LL85a, LW81, Lew83b, LL84b, LKS89, MDM80, MWB82, MTW^{+}88, Moo83, Mor88b, Mur81b, Nag88, Óla80, PAGG87, Pet81, PG83, PSGD86, QEQD86, RNVB81, Rud80, SLF86, SL87b, Sch85, SEK84, SSK87, Sim87, SSGB89, SG86, Sta86, SK82, SWSC85, Thr81, Til83, Two83, VG80, VWVR84, WBEJ85, WED85, Woo80, YJ84, ZBM82, vGMMG85]. lakebed [LCP80]. lakebeds [Win81]. Lakes [LL81, BSB^{+}87, Bay88, BC82, BH88a, BB86, BBS80, BFKN84, CBH83, CS88, Col83, CA87, DCV87, DGN82, Dee88, FDR^{+}88, FW82, Fry86, GM85, GK89, GR84, GJP89, GDS83, GMPA^{+}85, HK80, Han88, Har87, HI88, Hil85, Jør87, KK80, Kli88, KLC81, Kon89, LSI^{+}85, LM86, Lew82a, Mac86, MP89, Mon85, Nür84, Nür87, Nür88, POHMD84, PB80a, PHI84, PB80b, PPVHW87, QBHS80, RM86a, RKS^{+}86, RK87, Sch88, Sch80, Smi82c, Söd87, Sta82, Sta85b, Sto88b, SH88, VL89, WOS89, Wot82, Ser80, WST80, Woo85, Cam87, Edm82a, Smi88, Edm87]. Lalli [Har89a]. Lambert [Gor89a].

Lambert-Beer [Gor89a]. Lamp [BSN83]. Lamp/filter [BSN83]. Langmuir [ET80, HS86]. Lankester [DC83]. lapse [BC80]. Large [MMTL88, Ree81, BS85a, GMK80, KH86b, Lew82b, MH80, Sim87, WD87, WL85b]. larvae [But82, Gig80, Han84, MP88, Pas80, Pur81, Vic84]. larval [FCS84, Han81, HW86, OBP87, Ols87]. larvalfishes [GHC89]. Lateral [Zar81]. latitude [Smi81b]. latitudes [CA89, Pai86]. Latitudinal [ER88]. latitudinally [LL89]. Laubier [Lon85]. law [Gor89a]. layer [BWJ80, CGCG89, CA87, FBG89, HAvdL^{+}89, Imb85, IP85, Jef81, Mon85, MB89b, RW87, SBN^{+}83, SLL85, Smi82d, SIR86, SS85]. layers [JR85, PG83]. leachate [FCC^{+}86]. lead [DGN82]. leaf [CSW^{+}80]. leakage [Dee88]. leaky [Paf84]. leaves [Bor87]. LeB [Duc87]. Lectures [Shi85]. Lehman [CR83]. Lemna [FH85]. length [KH86b, KH86c, Str83]. lentic [Loe81]. Lerman [Ser80]. Leucothea [HSMH87]. level [CCS^{+}89]. levels [Ber89, HMG86, HE88, OAI^{+}85]. Lévèque [Hec86]. Levinton [Pai84a]. Lewis [Bur80]. Leydig [VKW89]. Library [Gar80, Jum82]. life [AH87, DH88, HT85a, HBT87, Ill80, LR85, O’D85, OP83, SRW81]. ligand [SH87a]. ligands [Bru89a]. Light [DB81, KTM83, KOTeI86, Mal82, MSBC80, uRG81, TLH89, ARP82, Bal85, Cru87, DCV87, GHK80, Gol86, HF88, JD86a, Jon87, KA89, Kel89, Kir89, KK80, LLG86, LB80, LL84b, MV88, NG83, OW84, PB80a, QHHLS89, SAK89, SJR87, SP89a, Sch85, SF81, Til83, WL83, WHFT89, WJ89, ZD89, MS87]. Light-limited [TLH89, LB80]. light-mediated [OW84]. Light-saturated [Mal82]. light-scattering [WJ89]. lignocellulose [BMH85, BLKH88, MH89]. lignocellulosic [BMH86]. like [Han84]. likely [GMFLJ85]. Likens [Joh89]. Limitation [BSG83, PCP87, Bot85, Che80b, DDSmV83, GCL89, HK88a, HA81, JRG81, JJ87, Klo85, LP81, MRLD86, PHCK83, RG81, uRG81, RM86a, Smi84, SKW86, Som88, SH88, TG82, VO89, VWVR84, WBF85]. limitations [RJ83]. limited [Bor83, LB80, Str80, TLH89, EEMC88]. limiting [GMPA^{+}85, ZBM82]. limits [Str88]. limnetic [VJOR88, VO89]. Limnología [Edm84a]. limnological [CM86]. Limnology [Ano80a, Ano80e, Ano81a, Ano81b, Ano81c, Ano81f, Ano82a, Ano82b, Ano82h, Ano83a, Ano83b, Ano83f, Ano84a, Ano84b, Ano84h, Ano85a, Ano85d, Ano86a, Ano86b, Ano86l, Ano87a, Ano87h, Ano88a, Ano89b, Ano89c, Ano89a, Ano89m, Hrb82, Koo85, Edm82c, Leh86, Pet86a, Ano80c, Ano84f, Fre84, Man83, Ano81d, Ano82c, Ano83e, Ano85c, Ano86i, Ano88g, Ano89h]. linear [SNR86]. lingua [DC87]. linkage [Nag88]. lipid [BS86, Hak84, HS84b, RW89]. lipids [GFCP85]. Liquid [RS81, BKB85, Sau80]. Liss [Shi85, Mac85]. List [Den82a]. litter [CSW^{+}80]. Littlefield [DWO86]. Littoral [DK86, HS88, Str83, SRK86, WK80]. live [Ree81]. lived [ZFFC88]. Livia [Edm86b]. Living [Koh87, Sim87, Wot82]. load [Leh88c, Nür84]. loading [Sta85b]. loads [Nür87]. Lobban [Hay87]. local [PB87, ZR85]. localization [Hal82b, HF89, SJC^{+}87]. localized [SA89]. locations [VFB87]. Löffler [Edm82c]. logic [Doy80]. London [Bro82, Fen80b, Fen82, Fen83, Lon85, Lun84, Mac85, Man80, MS84, Mor80b, New86, Rzó81b, Smi82b, Tut80, WK80]. Long [IH81, Ril82, SG86, CMR89, Hak84, RKF86, Ste85, CD85, Chu85, Joh89]. Long-distance [IH81]. long-term [CMR89, Joh89]. Longhurst [Bir88]. longicornis [WC86]. longirostris [Thr81, ZK80]. longispina [JJ87]. longitudinal [JKMN87]. look [Ven84]. loop [TTF88]. Loosdrecht [RM86a]. lose [Rie84]. Loss [AHRF89, JRG81, CDRW86, HK81b, LS85, PAGG87]. losses [WHFT89]. lotic [Bot85]. Lough [DF82b]. Louisiana [HDP83]. Lovén [RS86a]. low [CS88, Fei82, FS84, GS86, Gol86, HE88, KS81]. low-alkalinity [CS88]. lower [SL89]. Ltd [Bro82, Fen83, Pai80]. Lune [Jon82]. luteus [HT85b]. lutocline [WAI89]. Lutz [Koh82]. Lynch [Pra80]. lysing [BB84a].

M [Ano86j, Boe89, Bro82, Bur80, Dav83, Dav86, Dug80, Edm82a, EA84, Fen83, Hal82a, Har89a, Hay87, Hen87, Kir88, Lec85, Mol89, Mur84b, Now84, Pic82, Pra80, Rzó81a, Smi88, Tut80, Whi86, Woo80, KM81, LW81, Unc81]. M. [Mol89]. Macdonald [Wet82]. macroalga [Wil84]. macroalgae [CC84, NG87, SvdBR82]. macroconsumer [TH80]. macrocopa [D’A80]. Macrocystis [MD87, ZR85]. macrofauna [Lop88]. macromolecular [CJTL83, RW89]. macromolecules [KH88a]. macrophyte [BB84b, CK80, DK86, FCC^{+}86, HF88, KTM83, KLJ86, LKS89, SJR87, SH84]. macrophyte-dominated [KTM83]. macrophytes [CK82, CK80, LS82a, Lan82, TS82]. macroscopic [ACC86]. macrotidal [DTBB87]. macrozooplankton [Båm85, SM85]. made [Hec84]. Madeira [SJC^{+}87]. Madison [HH80]. Maerozooplankton [Smi85]. magna [BH84, CM82, DC87, Hal82b, NO89, NDS86]. Magnetic [ADO85, AR88, DF82b, HL85]. maintain [Som84]. maintaining [PTB83]. major [CS81, EH88, LHM84, SSFN86]. making [Cru87]. Malawi [Two83]. maltose [Dou87, Dou88]. man [Hec84]. man-made [Hec84]. Manganese [RAN88, SS85, BSG83, CMN82, DWR82, JM85, Kep85, MMD^{+}80, Mul89, STC88, SH83, SH85, SH87b]. mangrove [BB81, RD89, Wet87a, Wet87b]. manifestations [Sir87]. manipulated [BRH^{+}88, KK80]. manipulation [BZH^{+}85, SS88b]. Mann [Jum84, Ste84, Fen84]. Mannitol [KCN82]. Mannitol-fermenting [KCN82]. manual [Bur86a]. Manuka [Koh84]. manuscript [Edm84d, Jum87b]. manuscripts [Edm80b]. maps [Ban84b]. Mar [Hor83, Koh84, Lon85, MS84, Pic85]. Marcel [Kru82]. Margalef [Edm84a]. margin [Smi82b]. marin [Fow80]. marina [Bor87]. Marine [CSHO89, Fen87, Gar80, Jum82, Pai80, Wet82, Whi81, ACC86, AG88, And80, AS83, Ban82, BMS^{+}88, BMH86, BM82, BF89, Bra82, BSG83, BMS83, CK88, CDG88, Che80a, Che80b, Chr84, Col82, Col80b, CKT89, CL80, CP83, CM83, CW84, CB84, DW86a, DH80, DeM88, Dri81, FS84, FOB80, FBF83, Fis86, FTKB87, Fle89, Fow80, Fro80, GPHI83, GLMA82, GG82, HHH87, HM86, HW88, HK88a, HCC88a, HCC88b, HB85, HD86, Høj89, HMLC88, HMC88, HM89, Hun83a, Hun83b, IJ85, Jac87, JEM82, Jef81, JD86a, JEE82, Jon87, KMK88, Kan83, KR88, KS82a, KH88b, KH86a, Kri83, KB80a, KB80b, Lan83, Leh88a, Lid83, LH87, Lop88, LGV85, Lyl83, MA84, MDR88, MFHM80, Man88, Mar83, MHB84, MS84, MG81, MCK89, MB80, NL89, NG87, Nix88, NSM87, Paa80]. marine [PCP87, PZM87, PC84, PHCT89, PA83, PH88b, PS89, QHHLS89, RS86b, RSBL80, RBNF87, RS88, RPC89, Ros80, SAK89, San88, SNL84, SNP84, SN85, Sei88, SSFN86, Smi84, Ste88b, Sto88a, SvdBR82, TAJ89, TLH89, TTF88, UA82, VL87, WHFT89, WSU^{+}83, WK86a, Wil84, WHM^{+}85, WCEP82, Yen85, YMN82, ZFFC88, Ban84a, Dav83, Ill80, Pai84a, Pai84b, Wan82, Boe89, Jum82]. marker [Lyl83, Rue83]. markers [TS85]. marsh [BZH^{+}85, BMH85, Boo80, Bow84, CB86, HS80, HF82, HKT89, HM84, HHTV81, HDK84, HDW85, IBC^{+}83, JV82, Kin83, Kin88, LR85, LL82, PH87, SH84, SGMB87, WMSW89, Gar80]. Marsh-estuarine [Gar80]. marsh-estuary [HS80]. Marshall [ZoB80]. marshes [GH84, HDP83, HG83, VTV^{+}80]. Martinus [Dav86]. Maryland [CDPC84]. Mass [Csa86, Pic82, Pon82, SA83, SP85, BWG82, CWD^{+}86, Cor81, HHH87, HH81, THCP81, WHFT89]. Massachusetts [EH88, Fei82, LR85, McK81]. Masson [Ban84a, Edm84b]. Mastigias [IH81]. mat [RJBC83]. material [DF82b, FH82, HMG86, HCQ^{+}86, KSP^{+}81, PC84]. materials [Bro83, Car83, IBC^{+}83, Lew83a, PS81, VW80, WST80]. Math [Mac85]. Mathematical [Jum84, Ste84]. Mating [Hay81, DKMvL86, GM88]. matrix [FV85, KA89]. mats [JRC83, JD88, Shi85]. matter [Bal84, Bjø88, BMP81, CCW87, Cur84, DMM82, Ger83a, Ger83b, HTE84, HCC88a, HCC88b, HF84, Hun83a, Hun80, HL82, IDB82, JEM82, KC81, KRFS84, KBO84, KM84, LMS87, LG89, MS85a, PCP87, PH87, PD80, PS81, RE81, SW80, TS89, WLB84, WSU^{+}83, WB84, Tut80]. Maximal [Som89a, AK89, Ban82]. maximization [Phi80]. maximum [ADP^{+}84, ADK87, Den83, DK86, GA81, Wal80]. maximum-minimum [Wal80]. mayfly [AFM86]. Mazama [Abe88]. MBL [Shi85]. McCall [Now84]. mccradyi [SRW81]. McGraw [Fre84]. McGraw-Hill [Fre84]. McIlwaine [Hec84]. McLachlan [Woo80]. McMurdo [PS85, RV87]. McN [Bro80]. McRoy [Fen80b]. Mead [FS83]. mean [AB80]. measure [BR89, DW86b, FDK^{+}86, KW86, LSdKvD85, PME87, Vos89]. Measured [SEK84, AHI^{+}86, BC80, HTWR84, NCY89, Paa80, RS86c, RSBL80, RJB81, RJ83, RBNF87, SJR87]. Measurement [BS86, FV85, HS82, JL80, LWL86, MHB84, OSML87, SvdBR82, SH87a, BH88b, BKF^{+}87, BR80, CP86, DWR82, Fle89, GPS85, HB81, HM84, KB89b, KSP^{+}81, KJP89, LW84, TML83, VTV^{+}80, WLR88, ZST89]. Measurements [CL81, Lew87, OKG81, WHFT89, Ami88, FKM82, Fra88, GCAN87, HS80, HDW85, JD86a, KTM83, MPsCS80, MH84, Oud89, RW89, SVG89, SRW84, SG84, VW83, WOS89, WSU^{+}83, WM82, WCEP82, Yen85]. Measuring [DC88a, LS84, CMB87, Col84a, Gri88, HB85, KLJ86, LSI^{+}85, Loe81, Mac86, ZPVM83]. measurmg [NL89]. mechanism [Fro88, SFDK84, TS83, WBS83]. Mechanisms [CWD^{+}86, CKST86, BG86a, Gel86, Gil88, RM81]. mechanistic [Kel89, TML81]. mechanosensory [Kir85]. Medal [Ano82g, Ano84e, Ano85e, Ano85f, Ano86h, Ano88d, Ano89j]. mediated [CMN82, DeM89, EEMC88, GB80, OW84, Ore83a, vdLAH^{+}84]. mediation [TNEJ84]. Mediterranean [BWS^{+}85, BCM86, MB89a, RE81]. medium [CWD^{+}86]. medusae [HS86, MM82b]. Meeting [Ano80e, Ano81e, Ano81f, Ano82h, Ano83f, Ano84h, Ano85d, Ano88m, Ano89n, Man80]. megalops [BW86]. meiobenthos [BW89, RF88]. Melbourne [Smi82b, Tut80]. Melosira [SWSC85]. Member [Ano84b, Ano86b, Ano89c, Koo85]. Members [Ano81c, Ano82b, Ano83b]. memoriam [Ano89k, Edm82e, Edm86b, Hut86]. Memphremagog [Pet81]. Mendota [FB80, HH80, FHHB80, IB82a]. Menidia [WL85b]. Menten [Li83]. Mer [Iba81]. mercedis [Mur81b]. Mercenaria [Pet86c]. meromictic [CCO83, Fry86, POHMD84, PB81, Ste85]. Meromixis [MM82a, SC84]. Merriman [Pic82]. mesh [KC85]. mesocosm [Pri89, TD87]. Mesocyclops [Wil80c]. mesoplanktonic [BS89]. Mesostoma [DC87]. Metabolic [Gig80, KH86a, SRK86, FW82]. metabolically [DMM82]. Metabolism [FE81, GCAN87, BR86, CW85, Leh88c, LWF86, RV87, Smi81b, vGMMG85]. Metal [FKM82, Kep85, Mar83, San84, SvdBR82]. Metalimnetic [Kon89, BK89, Kon82, PAGG87, PLN84, PNL84]. metals [Dav86, Fis86, HTWR84, HACF85, Hun83a, SSK87, TB80]. Meteorological [GJP89]. meters [SP89b]. Methane [Dev83, KMD^{+}88, Lid83, Rud80, WRS80, dAL87, BRBL83, CMK89, CM83, FHHB80, HH80, IJ85, IOK87, KM84, WWS84]. methane-oxidizing [HH80]. methane-saturated [CMK89]. methanogenesis [IB82a, IOK87, KC81]. method [BB89a, CSAO86, CW81, Cru87, HH86, HHH83, HW86, KM80, LDYY83, LSdKvD85, Loe81, NL89, NG87, Paa80, Pae84, RJ83, SC88, SB83, TS85, TQS86, TB84, TD83, ZPVM83]. method1 [Men80]. Methodol [Jum84, Ste84]. Methodological [Car84, NG83]. methodology [MH86]. methods [BGJ^{+}87, BH88b, Bur86a, CS88, HT85a, HBT87, HY89, KLJ86, Mor80b, RJB81, WLG85, ZST89, Mor80b, Wet87a, Wet87b]. Methyl [MD87, RW89]. Methyl- [RW89]. Methylamine [DC88b, Bal85]. Methylammonium [WM82]. Metridia [MB86]. Metz [Edm82a]. Mexican [FVM82]. Mexico [DL83b, Sau80, SSS87, SG84]. Meyer [Edm84b]. Meyers [Lam87]. Michaelis [Li83]. Michigan [DWO86, EJ86, GNM87, GCLC87, GCL89, Kil82, LS82b, LN86, MDM80, Mor88b, SLF86, SL87b, SEK84]. micro [HB85]. micro- [HB85]. microaerophilic [SK82]. microalgae [CMR89, DMM82, PS85, RJB81, RI88]. microalgal [JY84, LSdKvD85, MRLD86]. microautoradiography [Pae84]. microbes [LH87, SS88a, Bro80, Kir88]. Microbial [CJTL83, HA83, RKS^{+}86, SH87b, TNEJ84, VN87, BA87, BR89, CK88, FDKM86, FDK^{+}86, HP89, HASI80, JBvdB89, JRC83, KW86, Kep85, Li83, MMG86, RJBC83, TTF88, WKLK89, YP89, Pra80, Shi85]. Microbially [CMN82]. Microbiol [Fen83]. microbiological [Bur84]. microbiology [Bro82, Fen83]. microcalorimeter [SM82a]. microcosm [GCAN87, Som88]. microcrustacea [SMH84]. Microcystis [FP87, HA81, NDS86, PTB83]. Microelectrode [JRC83, RJBC83, LSdKvD85, RMJ86]. microelectrodes [JERC85, RSBL80, SJR87]. Microelectrophoretic [Hun80]. microenvironments [Jah85, RAN88]. microflagellate [Nag88, SSB82]. microflagellate-picoplankton [Nag88]. microflora [PF80, RJ83]. microhabitat [MP88]. Microhabitats [Nic89]. microheterotrophic [AF85]. microlayer [CL81, HACF85, LL82, Syz82]. microlayers [Car83]. Micromethod [GFCP85]. Microorganisms [BBA84, CW85, Jac87, SK82, TS89, ZPVM83, ZoB80, Ben89, Bez89]. Microphotometric [IMK88, IS89]. microplankton [MWB82, PPVHW87]. microprobe [JD86a]. microprocessor [lWJ82]. microprocessor-controlled [lWJ82]. microprofile [RJB81, RJ83]. Microscale [Cur84, LS84]. microscope [HHH83, McB88]. microscopy [CW81, Cru87, WHM^{+}85]. microsensor [Rev89]. microspheres [BK82]. microstructure [FCO87, TB83]. Microzone [Pae85]. microzooplankton [Ban84a]. midcontinental [UEB88]. Middle [GA81, SC81b]. midst [HMG86]. midsummer [CCS^{+}89]. Mierobial [BMS^{+}88]. Migrating [MS85a]. migration [BF89, CM82, DFW89, FCS84, Ger82, JJ87, MG83, OP83, SG86, Wei84, vRS85]. migrations [IH81, SLL85, WL85b]. migratory [Pat89]. Miklavatn [ES83]. milieu [Fow80]. Mill [McK81]. Milwaukee [Ano82h]. mineralization [GNM87, GCL89, JEM82]. mineralogical [RK87]. minerals [AR88]. miniature [HA85]. minielectrodes [HB85]. minimize [BSN83]. minimum [Wal80]. Minnesota [LS83, LSM81, Mur85, WH80]. Minnetonka [Mur85]. minor [FH85]. Minshall [New86]. Mirror [DF82a]. Mississippi [FSW85, GHC89, MDR88]. mitosis [MC82]. mitotic [BS85b]. Mittel [Edm81]. Mittel-europas [Edm81]. Mixed [IP85, GHK80, Imb85, Mon85, SIR86, WLB84]. mixed-layer [Mon85]. Mixing [ADP^{+}84, Fis80, WAI89, DG83, HN81, IJ84, IP84, Kli88, Lem89, Lew83b, Ven84, Vic84, Wol87, Den82a]. mixtures [KSTGB81]. Mn [Hun83a, Hun83b, TNEJ84]. Mnemiopsis [SRW81]. mobile [CH87]. mobility [CF81, GB87]. mode [WC87]. Model [OØ85, BS82, CS85, CR80, Dri81, GL81, HF82, IP84, JEM82, Jah85, KMK88, KH86c, Mob89, Nür88, OJR^{+}86, WB84, Wof83, Wro89]. modeled [HMS86]. Modeling [BM82, Kel89, LT88, MJSM89, RPC89, SIR86, GG84, KM89, NSL86, SC81b, SPBB89, ZKD87, Wet82]. Models [WOS89, CSPV86, GLMA82, GB82, Jum84, LW84, Law84, Law85, PM84b, Rod87, SW89, Ste84]. Modern [Tut80]. Modes [PPS83, VP85, LaZ80, SB89a]. modification [Ham81]. Moina [D’A80]. molecular [SW80, TB80, vdLAH^{+}84]. molecular-weight [SW80]. mollusks [Har89a]. molting [MHB84]. Molybdenum [Col85, PCP87]. Monogr [Edm82c, Hec86, Hyn81, Jum84, Pim83, Ste84, Wet84]. Monographs [Fry80]. monomictic [HLA86]. monophosphate [FW82]. Monosaccharide [IDB82]. monoxide [CAS83]. monsoon [VN87]. Montana [Moo83]. Montastrea [DL83b]. Monterey [Han81]. Moore [Mol89]. Morel [Lec85]. Morgan [Mur84b]. Mori [Wet84]. Morin [Mil84b]. Morphological [SIH85, DH88]. morphology [FP87, HG85a, SWSC85, Wil87]. morphometric [SGSS84]. morphometry [Cor89, Fee79, Pat80, RGS89]. morphs [GB81, HD84, Hav85, LO81]. Morris [Smi88, Epp82]. Morritt [Mac80]. mortality [Ebe81, HWL83, KM87, Lyn83, PG81, SP85]. motion [IBC^{+}83, RS89]. motions [Mon85]. mountain [VL89]. mounts [CW81, Cru87]. mouth [Her82]. move [TS83]. Mozambique [SG86]. Mt. [Abe88]. mucronatus [LR85]. mucus [HP89]. muddy [HP87]. mudflat [May89]. muds [Lop88]. Mueller [FV85]. multichannel [Ano88k]. multidepression [Lin87]. multiple [BBH85]. multispecies [Moo88]. Multivariate [GBF86, SH89, Iba81, MTL84]. Mur [Wet83]. murres [Woo84]. mussel [JV82]. Myriophyllum [BB84b, SA86]. mysid [Mur81a]. Mysidacea [Mur84b]. Mysis [RS86a, Thr81]. mystery [GLMA85, Law84]. Mytilus [FBG89]. Myxococcal [FB88].

N [AS83, Den82a, EEMC88, Gol86, Pra80, RLG86, Smi82b, SH88, DL83a, DP89a, DW86b, Fry88, Gar84, GG84, GLMA85, GCD87, KP82, Law84, LK89, OSML87, Pae85, PCP87, PH88b, SN85, SSGB89, SH88, WLB84, ZFFC88]. N- [EEMC88]. N-deficient [RLG86]. N-enriched [LK89]. N-labeled [KP82]. N.J [Dav83, Jah83, Lec85, Mil84b, Mur83, Pai84a, Pic85, Pim83, Pla86]. N.Y [EA84, Kru82, Lon86, Pai84b]. NaH [HS82]. Naivasha [BSB^{+}87]. name [Hyn80]. nannoplankton [LS82b]. nanoflagellates [BRH^{+}88, EDKS86]. nanoplankton [HHH83, SSFN86]. nanozooplankton [RS86b]. nanus [HH85]. Narragansett [DDSmV83, MMD^{+}80]. narrow [SM82b]. National [Lis85, Lis85]. NATO [Csa80, Lon85, Mac85, MS84, Thi83]. Natural [DG80, KCB89, All81, BS82, BS80b, BG82, Bru89a, CP86, CK79, CM86, CSAO86, Col80a, DSH82, Dav86, Edm84d, FN85, FNL85, Fro88, GB87, Gil89, Gor89b, GG81, Gri88, Gro89a, Han81, HFF87, Hen87, HFA80, Hun80, IJ84, Jum87b, KSH85, KSH88, KN81, KS82b, Kir89, KJP89, KMSF81, LWP85, LUP88, Mob89, Mur83, NSL86, Pre86, RM86a, RS81, San84, SGSS84, Smi88, SB89b, SS80a, SS80b, Som85, SK85b, SB83, SYB^{+}89, TCN^{+}82, TB80, WL83, Whi86, WMM83, WCEP82, vRS83]. naturally [CSW^{+}80, Lan82, MHY83, RE81]. nature [Han85]. naupliar [CG80, WBF85]. Neagh [DF82b]. near [BS89, BHK^{+}82, Han84, HF84, HFL84, Wol87]. near-bottom [Han84]. near-surface [BS89]. nearshore [Bal84, Chu85, Lem89, TML^{+}88, UA82]. Nedwell [Fen83]. Needlelike [BB84a]. Nematoscelis [BW86]. Neocalanus [DW87, Ohm87]. Neomysis [Mur81b]. neotropical [BBA84]. neptunium [FBF83]. neritic [RE81]. Net [Le 82, LN86, Her82, HM82, KC85, OSML87, SRW80, SKW86]. Netherlands [Dav86, SK82, RM86a]. Neusiedlersee. [Edm82c]. neutralization [CKST86]. Nevada [CCW87, IOK87, SO87, Sto87, VG80, ZHO^{+}87, CCO83, Gol88, HG85b]. Newell [Pai80]. Newfoundland [Ano84h]. newly [OH87]. NH [BM82, Law85]. niche [TML81]. Nickel [JM84, Hal82b]. Night [COL84b].

Nijhoff [Dav86]. Nijhoff/Dr [Dav86]. Niño [Pai86, ZR85]. Niskin [lWR89]. Nissenbaum [Kru80]. nitracline [WKRE89]. Nitrate [CWV^{+}88, Gar81b, Ols80a, BSS^{+}84, BMPD84, Col82, Col80b, CMR89, GB82, JP83, Kri83, LK89, Ols80b, uRG81, TLH89, TB84, UEB88, YMN82]. nitric [RKS^{+}86]. Nitrification [RPL89, JK84]. Nitrification-denitrification [RPL89]. nitrite [Col82, JP83, MKN84]. Nitrogen [BH83, GNM87, HG85b, HMLC88, HMC88, Jon82, KSH85, KSH88, KW87, LWF86, PP85, SH81, WMSW89, Alt89, AGG82, BCM86, Bow84, CSF88, Col80b, DW86b, EBSW83, GS81, GCL89, GB80, Har83b, HW88, HS84a, IVH^{+}86, JV82, KM81, Le 82, LHH80, LHM84, LL84a, LL85a, LL84b, MN86, MRE84, MM82b, Neh87, OAI^{+}85, PH87, PSGD86, RLG86, Sch80, SNM^{+}87, Smi82c, Smi84, SHDV89, SS80a, SS88b, TB80, WKRE89, WK86a]. nitrogen-15 [Bow84, OAI^{+}85]. Nitrogenous [CD85, KH88a, MWB82]. Nitrous [KLC81, LL81, SNP84]. NJ [Mil82]. nm [HM87]. No [Bur86a, Gar80, Jum82, Wan82, Her82]. nodosum [Str83]. nodules [Moo83, Mul89]. Nominations [Ano82g, Ano85e, Ano85f]. non [Som85]. non-steady [Som85]. nonconcordance [AFM86]. noncontaminating [MH84]. nondestructive [SG81]. nonequilibria [Whi86]. nonsaturated [Hun81]. Nonselective [Hun81]. Nonsphericity [Jon87]. nonsymbiotic [MM82b]. nordenskioldi [SF89]. North [Ano83f, Mac80, ES83, GDS83, RKS^{+}86, BS89, Bru89a, BW86, BRBL83, BR86, GC81, GK83, GMFLJ85, LB84, LDR87, LUP88, MRE84, MH85, SSS87, SIH85, SLB83, SKW86, SH81, SH87b, Wei80, Wro89]. north-central [GDS83]. northeast [Ban87, JM84, JM85, KM81, CB88, Col85, ML89]. northeastern [Her83]. Northern [DF82b, BB81, Cap82, HKT89, LR85, KBB^{+}89]. northwest [DL83b, Han87, VN87]. northwestern [CJTL83, Moo83]. note [GHW84, Men80, RWB87, Yen83]. notes [Car84]. Notice [Ano86k]. Notonecta [DH88, SM83]. Nova [MTL84, FVH^{+}84]. NTA [RNVB81]. nucleation [Sta86]. nucleic [FDK^{+}86, KW86, KB89b, WK86b]. Nuclepore [TC84]. nudifrons [AC82]. Nuisance [Pae88]. number [KS81]. numbers [SF82]. numerical [Mob89, PSGD86, Unc81]. Nutrient [Boo80, Cor81, FB88, HK88a, LB80, MRLD86, PLN84, Pra90, Som89b, Sta85a, BZH^{+}85, CD85, CR83, CW84, Cuk83, Cur84, Duc82, EL81, FDR^{+}88, GMK80, HJP87, HA81, JRG81, KM89, Kor80, Lan82, LP81, LS84, LN86, Leh88c, Mal82, MKN84, MF89, NG83, PHCK83, RG81, RM83, SFDK84, SM84, SLB83, Smi86a, SKW86, Str80, TML81, VWVR84, ZKD87]. Nutrient- [LB80]. nutrient-limited [Str80]. nutrient-rich [Mal82]. nutrients [AM83, BZH^{+}85, CL81, Gol88, Jav83, Kel89, Leh80, MV88, MMD^{+}80, MWB82, MS85a]. nutrition [CCF89, DC88b, Fry86, KSH88, Rue83, SA85]. nutritional [HS84b]. Nybakken [Pai84b]. nymphs [AFM86].

O [McM86, MS87, PM88, Shi85, Sto89, POHMD84, RJBC83, RAN88, TNEJ84]. Observations [BWG83, FS83, KMK88, TB83, BG86a, HSMH87, HHOC89, HN81, HD84, HM82, Lon84, Mal83, Ros80, SNR86, Yen85]. observed [APR80, BB84a, HMS86, Unc81]. observing [BCY89]. obstructions [Her82]. obtain [SL87a, Zan89]. Obtaining [AH87, HT85a, HBT87, HW86]. obtusa [SIH85]. occurrence [Hil85, PU82]. occurring [RE81, SK83b]. Ocean [Ano81e, BF87, CSPV86, CSHO89, DG83, Den82b, Gor89a, KA89, LCPL88, LB80, LG89, LWP85, She84, Smi87, SMP^{+}89, Vos89, Wro89, YP89, Col85, CJTL83, GK83, Han87, JM85, KSH85, KSH88, Le 82, Lon86, McM86, ML89, SSS87, SM80, Sul80]. Ocean12 [DW87]. Oceanic [CB87, Cha85, EDKS86, Hay81, PS81, RM86b, SH83, Swa83, WKLK89]. oceanica [Bra82]. Oceanogr [Ban84a, Jum84, Ste84, Wan82]. oceanographers [Pla86]. oceanographic [Iba81]. Oceanography [AN81, Ano80a, Ano80e, Ano81a, Ano81b, Ano81c, Ano81f, Ano82a, Ano82b, Ano82h, Ano83a, Ano83b, Ano83f, Ano84a, Ano84b, Ano84h, Ano85a, Ano85d, Ano86a, Ano86b, Ano86l, Ano87a, Ano87h, Ano88a, Ano89b, Ano89c, Ano89a, Ano89m, Hrb82, Koo85, Pic82, Pic85, Pla86, Ano80c, Ano84f, Hor83, Jum84, Pon82, Ste84, Ano81d, Ano82c, Ano83e, Ano85c, Ano86i, Ano88g, Ano89h, Pic82]. Oceans [Rzó81a, DW86b, Bir88]. Odontaster [OBP87]. odor [Ano88k]. off [BSC80, CS85, Che80b, Chu85, DP89a, FVH^{+}84, JM84, JM85, LK89, MSG81]. Offer [Ano89l]. offshore [Bur86b, EJ86, HK86]. ofgrowth [SAK89]. ofreduced [HY89]. Okefenokee [BMH85, FF85]. old [ZEY89]. oligotrichous [PS89]. oligotrophic [Brz88, CCS^{+}89, DM85, ES82, Gol88, HG85b, JL80, LRH^{+}84, RS86b, SH88, Swa83, Wei80]. Olisthodiscus [HT85b]. Omega [Edm84a]. omnivorous [SF89, TD87, Wil87]. Only [Som89a, CM86]. Onondaga [WED85, YJ84]. onset [Gol88]. Ontario [FP81, Ham82, RNVB81, SWSC85]. onto [Lew87, SH87a]. ontogeny [HHOC89]. Onychodiaptomus [Wal85]. opal [Ban88]. open [HSMH87, LWP85]. operated [CH87]. opportunistic [TC85]. optic [JD86a, JD88]. Optical [BMP83, JD88, BS82, Kir84, KP82, ML89, NCY89, PS81, SVG89, SPBB89, SW89, Vos89, WBEJ85, Zan89]. optics [PM84b, Pre86, Wro89, YP89]. Optimal [DeM89, HH86]. optimum [Kri83]. Orconectes [Cap82]. order [LaZ80]. Oregon [HTE84, dAL87]. Organic [Bal84, GCL89, Gar81a, HCC88b, HL82, JC80, KM84, LSLW86, vdBD84, ACC86, BH88a, Ber89, BB86, Bjø88, BMP81, Bru89a, BR80, Car83, CLS82, CCS^{+}89, Fro80, Ger83a, Ger83b, GZG80, Har83b, HTE84, HCQ^{+}86, HCC88a, HF84, HD86, Hun83a, Hun80, JEM82, JEE82, KLB80, KC81, KRFS84, KBO84, KM89, LG89, LL85b, LE89, LWCV83, MS85a, MH89, Mul81, NG87, OK87, PCP87, PH87, PD80, PS81, RKF86, SSB82, SKW86, Smi87, SW80, SW81, TB80, VHJ^{+}89, WLB84, Wan82, WB84, WK86a, Bro83, Hen87]. organic-rich [KM89]. organically [Har87]. organism [TS81b]. organisms [But86, CH87, CGCG89, Den82b, Den83, Den85, GB87, HH80, MF89, SG81, TS85, Koh82, Mar83, Pai80]. Organizer [Man80]. Origin [Söd87, Hai88, Moo83, MB80, PNL84, Sch88]. Orinoco [HL87, MSG81, SL89]. orthophosphate [KDC86]. oscillation [JSS81]. oscillations [Ham82, JS83, Sir87]. Oscillatoria [vRS83, IA85, Kon82, LUP88, LGM80, SNM^{+}87, ZBM82]. Öskjuvatn [Óla80]. Oslofjord [PO80]. other [Håk87, HKT89, Hun83a, PS81, Rie84, SMR84]. Otisco [WBEJ85]. Ottawa [Rzó81a]. Oudot [Ami88]. outdoor [HT85b]. outer [Chu85, YAL^{+}81]. outlets [Wot82]. ovary [BS86]. overview [Sto88a]. overwintering [MP88]. Oxford [Bro80, Bur86a, Fen80a, Smi82b, Tut80, Wan82]. oxic [SMSM86]. oxic-anoxic [SMSM86]. Oxidation [HHTV81, BR86, CMN82, Dev83, IJ85, IOK87, Kep85, MMG86, RAN88, RDW^{+}88, Rud80, SH87b]. Oxidation-reduction [HHTV81]. Oxidative [LH80]. oxide [KLC81, LL81, SNP84]. oxidized [Chr84]. oxidizing [HH80, War84]. Oxygen [BBP82, HAvdL^{+}89, Rah87, RMJ86, Top87, WMSP86, WM87, Ami88, Atk88, BK84, Ber89, BB81, CW88, CGCG89, CR80, Cor89, DQRM87, Gri88, HE88, HB85, HDK84, JEM82, JR85, KTM83, KK80, LS83, Lin87, LSdKvD85, OGMG88, Oud89, RS86c, RSBL80, RJB81, RJ83, Rev89, SJR87, Smi82d, SLB83, Smi85, Smi87, SRW84, ZPVM83, BK80]. oxygen- [KTM83]. oyster [CMR89]. oyster-pond [CMR89].

P [Ano86j, Ban84a, Duc87, Edm84b, EEMC88, Fen80b, Fow80, Hay87, Hec86, Lon85, Mac85, Mil87, Now84, OJR^{+}86, Pom80, Rzó81a, Smi82b, Smi88, Sul80, Wet82, Ban84b, Bro80, Bro82, Bro83, Bur86a, Cam87, Cat85, Cha85, Csa80, Dav83, Dav86, Den82a, Dug80, Edm81, Edm82a, Edm82b, Edm82c, Edm82d, EA84, Epp82, Fen83, Fen84, Fre84, Fry80, Gar80, Gil80, Hal82a, HG88, Hec84, Hor83, Hrb82, Hyn81, Ill80, Jah83, Jum82, Koh82, Kru80, Kru82, Lec85, Lis85, Lon84, Lon86, Lun84, Mac80, Man80, Man83, Mar83, McM86, Mil82, Mil84b, MS84, Mor80b, Mor88a, Mur83, Mur84b, New86, Pai80, Pai84a, Pai84b, Pic82, Pic85, Pim83, Pla86, Pom89, Pon82, Pra80, Ril82, Rzó81b, Ser80, Shi85, Smi82a, Smi86b, Thi83, Tut80, Wan82, Wel88, WK80, Wet83, Wet84, Whi86]. p [Woo80, Woo85, ZoB80, AS83, AS87, Cuk87, Gol86, Gro89a, Le 82, RM86a, SSGB89, SB89c, SH88, WLB84]. P-limited [EEMC88]. p. [Ban84a, Bur80, Edm84a, Fen80a, Fen80b, Fen82, Jum84, Koh84, Ste84]. Pablo [MPsCS80]. Pacific [Col85, ML89, SKW86, WL85a, Bru89a, BRBL83, CB88, DW87, DL85b, JM84, JM85, KSH85, KSH88, KCB89, KM81, LDR87, LK89, MRE84, SLB83, Wei80, WKLK89]. pacifica [MB86, Ros81, RDE82, Tag88]. pacificus [BF89, Håk87, Hak84, Lan80, LHF^{+}84, MB86, Run80, Run85]. PAK [SH87a]. Palaemonetes [Mor80a]. Palau [HGH82]. paleolimnology [CEE86]. pallidus [WBF85, WB86b]. paper [Ami88]. papers [Ano88i]. Paracalanus [Paf84]. paradox [Som84]. parameters [Col80b, CP83, Ebe81, GPHI83, Li83, OP83, Rod87, SW89]. Parangipettai [SP85]. parasitism [JRG81]. Paris [Ban84a, Edm84b, Fow80, Jum84, Ste84]. Paronychocamptus [HH85]. Parsons [Gie82]. Particle [JW85, San84, Sto84, WOS89, Alt89, Car89, DMdK84, Fen80c, Gar81a, IMK88, JC80, KDC86, Mor88a, Pet86b, SM85, YJ84]. particle-bound [KDC86, SM85]. particles [BCY89, BS85a, BR80, Car89, Csa86, DeM88, GB87, Han84, HL82, Jac89, JZF80, Jon87, KS82a, Kir83, KH86b, LL85b, LWCV83, MF86, MCK89, NL87, SSK87, YHM^{+}83, Yen83, Mac85]. Particulate [DMM82, MV88, WBEJ85, Boo80, CCW87, Gol86, HMG86, HCQ^{+}86, IDB82, LMS87, MMTL88, Mul81, PS81, RE81, RMN82, Smi87, SS80b, WSU^{+}83, WK86b]. partitioning [HW88, Wei84]. parva [OS82]. passage [CDRW86, DW87]. passive [Han84]. past [CS88, Pic82, Zül81]. Patalas [Fee80]. Patch [Vic84, AM83, SB89a]. patches [Cur84, LS84, Pri89, Woo84, YAL^{+}81]. patchiness [DR88, DP89b, SFDK84]. patchy [McC86]. Pathways [PS85, SM80, HDK84]. pattern [HH81, Pet86c]. Patterns [KLB80, Kin88, LDYY83, APR80, BG82, BSC80, Bra82, Bru89b, CS83, CW84, FP81, Han85, Han81, Her83, HHTV81, MP89, MTL84, PM86, PGO82, SA89, SPBD89, SCH89, SCR84, WMM83, Yen85]. Pauly [Bir88]. Pb [BH88a, BH88b, BB86, FTKB87, LL82, Mur85, SGMB87]. peak [BK89, PNL84]. peaks [PLN84]. peat [HF82, WLG85]. Peckarsky [Sel81]. Pegea [HMG86]. pelagial [Ban82]. pelagic [Ban84a, JEE82, PD80, SS88a, TS89, TTF88, IOK87, SO87, ZHO^{+}87, Har89a]. pellet [DW86a]. pellets [DL85a, Hab85, KMSF81, TNJ84]. penetration [Fle89]. Penilia [TTF88]. Pennak [Gil80]. Pennsylvania [Man83]. perch [MMTL88, FP81]. perennially [Top87, WMSP86, WM87]. perfoliatus [KLJ86]. performance [BKB85]. performed [CI88]. Pergamon [Mar83]. period [NO89]. periodic [Ebe81, Lyn83]. periodicities [ADF^{+}82]. periodicity [WKLK89, YMN82]. periphytic [Han88]. periphyton [Bot85, HK88b, Loe81, MP88, OSML87, PME87, PHCK83, RLG86, RW87, Rod87, Cat85]. permanent [CW81, Cru87]. permanently [PPVHW87, WLB84]. permeable [LH87]. peroxide [PZM87]. Perspective [HBT87, SL87a, Fen80b]. perspectives [Fen82, Koh84]. perturbation [Col82]. Peru [VWVR84, And85, BMPD84, BSC80, GZG80, HF84, HFL84, HK86]. Peruvian [HLM86]. pesticide [GS86]. Peter [Ano89k]. pH [BF84, BMH85, BB81, CS88, CW83, Cul80, Lew82a, Mil86, RJBC83, RAN88, RK87, STC88, TS82, Whi89]. Phagotrophic [BK89]. phagotrophy [BK87]. phase [KA89, LFRT86]. phaseolus [PAGG87]. phenology [Wal85]. phenols [LH80]. Phenotypic [SCR84]. pheophorbide [LD86, VL87]. pheophorbide-a [VL87]. pheophytin [VL87]. pheophytin-a [VL87]. pheopigment [DL85a]. pheopigments [DFW89, LD86]. Phil [Man80]. Phillips [Fen80b]. phosphatase [Rue83]. phosphatases [BH84]. Phosphate [Ols80b, Atk87, BBP82, CSAO86, FNL85, Fro88, KB80a, KB80b, MG80, MMTL88, MHY83, Neh87, RM86a, SH88, TB84, VO89, dJV89]. Phosphorus [Avn83, Bot85, CK82, CW88, EBSW83, Gro89b, LS82b, Pet81, Smi84, SK85a, SA86, WED85, BP87, CF81, Cor89, CK84a, CK84b, DTS^{+}89, FSW85, FH82, GM85, GMM88, HFF87, IVH^{+}86, Klo85, LW84, LCPL88, LSM81, Le 82, LE83, LHM84, MMTL88, Nür84, Nür87, OØ85, OJR^{+}86, OK87, PCP87, PSGD86, RW87, Rue83, Sch80, Smi82c, SK83b, SA83, SS80b, Som84, Sta85b, SS88b, VJOR88, VO89, WST80]. Phosphorus-dependent [Gro89b]. photic [VHJ^{+}89]. Photoadaptations [RV87]. photoassimilation [LGM80, Ore83a]. photography [BC80]. photoheterotrophy [ES82]. photoinhibition [JY84]. photolithotrophy [ES82]. photometer [Col84a]. photon [SMP^{+}89]. Photoreduction [Col83, MMG86]. photorespiration [FH85]. Photoresponses [FCS84]. Photosynthate [PPVHW87, FH84, SCH89]. Photosynthesis [FH85, GHK80, JRC83, MH82, RJ83, SJR87, SRW80, WL83, BK87, Dri81, JERC85, KCB89, Mal82, MS87, MG81, MSG81, NG83, PME87, RJBC83, RV87, Str80]. Photosynthetic [GPHI83, LP81, LUP88, PB80a, TS82, CP83, FN85, HM82, JD88, KMK88, LWP85, PS85, PG83, SPBB89, Til83, ZPVM83]. Photosynthetically [CA89, CLS82, BF87]. Phototaxis [DD89]. Phototoxicity [LO81]. Phototrophic [GMPA^{+}85, Sto88a, MH85, PB81, SK82, SF84, ZPVM83, vGMMG85]. phycobiliprotein [SF84]. phycoerythrin [EHE^{+}83]. phycological [Mor80b]. Phys [Mac85]. Physical [FVH^{+}84, HDLP85, Nix88, HGH82, Par82, Pon82, SC82, TS81b, WBG^{+}87, Wri89, Pic85]. physics [Ril82, Ser80]. Physiological [Kon82, Mor80b, TS81b, Epp82, Hay87, KSH88, Kon89, MFHM80, VWVR84, ZBM82, Thi83]. Physiology [LGM80, MG81]. phytoplankter [PZM87]. phytoplankters [BMP83, OKG81]. Phytoplankton [Bjø88, ET80, HLS87, LFRT86, LBZC88, Mal82, PM86, RM86a, She84, SK85b, Som88, ARP82, ADK87, AB80, ADF^{+}82, Bal85, BCL85, BSS^{+}84, BSON84, BK89, BMPD84, BB89b, BFKN84, BSG83, BR80, BSF81, CK82, CD85, CDPC84, CL80, CP83, CW81, Cru87, Cur84, CK84b, DM84, DSH82, DG83, DL83a, DP89a, DMdK84, DHV^{+}87, DTS^{+}89, DB81, DP89b, Eva88, FO82, FDW85, FN85, FOB80, FBF83, Fis86, FTKB87, FBG89, FV85, Fur87, GHK80, GPHI83, GB80, GG82, Gol86, GBF86, Håk87, HHH87, HTWR84, Har83b, Haw80, HK81a, HK88a, HM89, HL81, IVH^{+}86, IS89, INA89, Jef81, Jør87, KSH85, KH88a, KSH88, KM83, KH88b, KBO84, KOTeI86, KW87, KP82, Kri83, LB84, LB80, LRH^{+}84, LDR87, LWP85, LS81, MFHM80, MH82, MH87, MC82, MSG81]. phytoplankton [NG83, OH87, Pae84, Pae88, PP89, PLN84, Pra90, PS81, PG83, PP85, RMB82, RG81, uRG81, RAN88, RV87, RPC89, RO82, Run85, SLP87, SA81, Sch85, SF89, SM84, SM80, SK83b, SK83a, SK85a, Som84, Som85, Som89a, Som89b, Str80, TML83, Thi83, TLH89, Thr88, Til83, TM84, VJOR88, Ven84, WL83, WL84, WL85a, WGM82, WM82, WD87, Wof83, WJ89, YAL^{+}81, ZKD87, ZD89, Zül81, Epp82, Tip87]. phytoplankton-herbivore [CL80]. phytoplankton-mediated [GB80]. picoplankton [HW88, Nag88, Sto88a]. Piedmont [KB82]. Pielou [Jum87a]. Pierrot [Mil82]. Pierrot-Bults [Mil82]. pigment [PM86, WC86, SK82]. pigmentation [Hai80, KDR89]. pigments [CDRW86, FN85, Håk87, Jef81, LCK89, MB89b, PS81, SF84, Til83, Ven87, WHM^{+}85]. Pinelands [MMG86, Mor84]. pipes [BBH85].

Pisces [BBA84]. pits [SC88]. pkss [Yen83]. plain [MDR88, SJC^{+}87]. planci [Ols87]. plancton [Edm84b]. planktivore [Thr88]. planktivorous [GR84, MMTL88]. Planktivory [Thr88]. plankton [Cuk87, DHV^{+}87, Duc82, FBF83, GHW84, GLMA82, Gro89a, HH81, Her82, HLM86, IDB82, KC85, LP81, Mac84, MSG86, MKN84, MF89, PO80, Paa80, RM86b, Som84, SYB^{+}89, SF84, TD87, Vic84, lWR89, Mil82]. plankton-mesh [KC85]. Planktonic [FB80, Han84, Jav83, Abe88, BGJ^{+}87, BF89, Che80a, Che80b, CLS82, CS83, DC87, FS84, For85, FW82, Han88, Hob88, JL80, JERC85, KKH86, Leh80, MV88, MHB84, Ros80, RI88, SAK89, SPBD89, SLF86, Sto84]. plant [BMH85, RPL89]. plants [AS83, CK79, FH84, GG81]. plastic [Bur84]. platensis [BBA84]. plates [Edm81]. Platt [Jum84, Ste84, Thi83]. plea [RR80]. Pleistocene [DL85b]. plenty [HMG86]. Plenum [Csa80, Hor83, Koh82, Lon85, MS84, New86, Now84, Rzó81b, Wet82]. plumchrus [DW87]. plume [GHC89]. plumosus [JBvdB89]. plunging [FS83]. plutei [RFS84]. Plutonium [FOB80, FBF83, SMSM86]. Poland [JS83]. Polar [Lon86]. polarization [KA89]. Polarized [QHHLS89]. Polarized-light [QHHLS89]. Polarographic [LGV85]. polarography [CV82]. pollen [BD81]. Pollingher [Woo85]. pollutants [Man80]. Polyarthra [BG82]. Polychaete [Han81, Tag88, TH80, TC85]. polymerization [LH80]. polymorphism [GS84]. Pond [McK81, AMP81, BBP82, CMR89, HFFS85, Kir83, WK80]. ponds [LO83, Man83]. Pontoporeia [LE89]. Pontoporera [LE89]. Poole [Pra80]. Pop [Pim83]. Population [Bor83, OS82, Ore83b, Wei80, Boe89, CM81, HHH87, Kon82, LBZC88, LUP88, Lue88, MG83, PAGG87, SA81, SE82, SYB^{+}89, Tay88, TC85, vRS83]. populations [BS80b, Boe89, BG82, BW86, D’A80, FN85, FB88, GS85, Gil88, Har83b, HA81, KSH85, KSH88, KN81, Kim87, LWP85, Lyn83, MC82, PP85, RM86a, TG82, Thr81, TM84, WMM83]. pore [BMH^{+}87, Jah88, KM89, MMD^{+}80]. pore-water [BMH^{+}87, Jah88, KM89]. Porewater [GH84, RM83, BB84b, HDW85, Jah85]. porewaters [And85, LGV85]. portion [WMSW89]. posed [SBN^{+}83]. positive [LKS89]. Positively [NL87]. Possible [Ger82, dJV89, SC84]. Post [MHB84]. Post-collection [MHB84]. Postdepositional [CF81]. Postma [Pom89]. Poststarvation [MB86]. Potamogeton [KLJ86]. Potential [HP89, BBG89, KRCS82, MB89b, SSS87, TS85, VW83, VCG^{+}87]. potentially [McL86, TM84]. potentials [HHTV81]. potentiomanometer [WLR88]. Potomac [BD81, CDPC84]. Pourriot [Edm84b]. powered [HHS80]. practical [Dav83, Gie82, Mil87]. Praeger [EA84]. precautionary [RWB87]. precipitation [Boo80, Kil82, Sta86, Ste83, TNEJ84, WBEJ85]. Precise [OGMG88, lWJ82, Ami88]. Precision [Ami88, Oud89, Gri88, MMR87]. Predation [Har84, Lue88, LS81, WB86b, AMP81, FB88, Hai88, HD84, KK80, KD81, LO81, LO83, MMTL88, Moo88, Mor80a, Mur81a, Mur81b, RS86a, SF89, SM83, Yen85, Hal82a, HG88].

Predator [Dod88, GB81, RS86b, Wil80c, Wil87, DC87, Pas80, Pur81, WK89]. Predator-avoidance [Dod88]. predator-induced [WK89]. Predator-prey [RS86b, Wil87]. predators [Kir85, Wil80b]. predatory [HGM80, Wil80c]. predict [Wof83]. predicted [RS86c, SEK84]. Predicting [Cor89, Hil85, Nür88]. Prediction [VFB87, KH86c, MD85, Nür84, Pet86a, ZEY89, Csa80]. predictor [DeM89, DK86]. preferences [Wil80c]. Preliminary [HM82, RS89]. Prentice [Mil84b, Mur83, Pai84a]. Prentice-Hall [Mil84b, Mur83, Pai84a]. Preparation [KP82, SK83a]. Preparing [McB88, CW81]. Prepublication [Ano89l]. prescreening [Fur87]. presence [CMR89, HE88, LT88, PL89, TM84]. preservation [Pae84]. preserved [HK81b]. Press [Ano86j, Bro80, Bro82, Bro83, Csa80, Den82a, Epp82, Fen80b, Fen82, Fen83, Fen84, Gar80, Hal82a, HG88, Hrb82, Jah83, Jum82, Jum84, Lis85, Lon85, Lon86, Lun84, Mar83, Mor80b, New86, Pim83, Pom80, Pon82, Ril82, Rzó81b, Smi82a, Smi86b, Smi88, Ste84, Wel88, Woo85]. pressure [Som88, VW83]. Prevention [TM84]. Prey [AFM87, CPT86, DH88, DC87, Kir85, Lan80, Moo88, Pas80, RS86b, Sul80, SF81, Wil80c, Wil87, Woo84, WL85b]. Primary [CK79, Gol88, GG81, HF81, Hor81, KM81, Mur84a, PA83, RJB81, SE82, BBS80, BKF^{+}87, CD85, Car85, DM84, Fee79, FKM82, Fly88, For85, GJP89, HJP87, HHTV81, Hun83b, JD86b, KSH85, Kel89, LB84, Loe81, MH84, MN86, Pat80, SP89a, SK82, VCG^{+}87, WBV80]. primer [Fro88]. Princeton [Jah83, Pim83, Wel88]. principles [Leh88a, Lec85]. printed [Jum88b]. Probable [CDRW86]. problem [SW89, lWR89]. Problems [WSU^{+}83, MHB84, NG83]. Proc [Dav86, Edm82d, Hor83, Pic82]. procaryotes [BMH86]. procaryotic [LS82b]. procedure [Ami88, LM86, OGMG88]. procedures [JS83, MH84]. Proceedings [Ano86l, Ano89m, Fen82, Hrb82, Ano80e, Ano81f, Ano82h, Ano83f, Ano84h, Ano87h, Ano85d]. Process [BM82]. processes [Cam87, CCO83, Eck83, Fen80a, Fen87, Fis80, HN81, HLA86, KRCS82, Lon85, Luc86, McK81, Mor88a, Pse88, Wet82]. Processing [CSW^{+}80, Edm80b]. Prochilodontidae [BBA84]. Prochilodus [BBA84]. produced [CLS82, DMM82, Kir85, RAN88, TCN^{+}82]. producers [HJP87]. produces [BSW81]. producing [SSS87]. Production [ACC86, But82, CAS83, FB80, FF85, INA89, Man88, Men81, SLF86, AHI^{+}86, BR89, BGJ^{+}87, Bow84, BKF^{+}87, BFKN84, Bri82, Bur86a, CW88, CDG88, CD85, Car85, CEE86, CK79, CK80, Che80a, Che80b, CM83, Duc82, DDSmV83, FCC^{+}86, FKM82, Fly88, For85, FDKM86, FDK^{+}86, GJP89, GG81, HHH87, HF81, HH86, Her83, HH85, Hor81, HHTV81, Hun83b, HW86, Hyn80, Ive88, JD86b, JGCS82, JL80, Jør87, KSH85, KW86, KLJ86, KM87, Kim87, Kir83, KM81, KM80, KM84, LB84, LRH^{+}84, LDR87, LBZC88, Le 82, LG89, LN86, Leh88a, LWP85, MD87, MH84, MV88, MN86, Men80, MB89b, MMR87, Mur84a, OH87, OK87, PO80, PZM87, PM88, PB80a, PA83, RJB81, RMJ86, RBNF87, Run85, SP89a, Sca88, SSS87, SH84, SNM^{+}87, SNP84, SN85, Sim87]. production [SE82, SKW86, SPBB89, SK82, Sto84, SF82, Str88, TML83, VCG^{+}87, WH80]. productive [RPC89]. Productivity [MMP86, MK84, Ami88, BBS80, DM84, Fee79, FH85, Fur87, Gol88, Hec86, HLS87, HF88, Kel89, KTM83, LR85, Leh88c, Loe81, OGMG88, Pat80, SB80, Til83, WBV80, lWR89]. Products [MSG81, Bel80, BS80b, Bel83, BKB85, BWG82, PG83]. Professor [Leh88b]. profiler [HK80]. profiles [GB82, Jah85, MB89b, RE82, RJBC83]. progress [Lon84]. projectiles [BB84a]. proper [Rah87]. Properties [DP81, DP86, CM86, Hun80, IS89, JD88, Kir84, Mur81a, WBEJ85, WJ89, Zan89]. propionate [INA89]. prospective [HBT87]. prospects [Rzó81a]. Protein [ZST89, COL84b, DL83a, TB80]. proteins [HA83]. proteolytic [May89]. protistan [Gro89a]. Protozoa [Ban84a, SSB82]. protozoan [BBG89, GSBP87]. Protozoans [CM89, BC82, PO81]. protozoology [Sto89]. protozooplankton [HK81a]. provide [Wot82]. provinces [ML89]. provisional [ML89]. pseudolimnaeus [WH80]. pseudonana [PH88b, RM81]. Pseudopterogorgia [HF89]. Publ [Bur86a, Cat85, Edm82a, Edm82c, Edm82d, Fen80a, Fen83, Fry80, Hec86, Hor83, Koh82, Kru80, Mac85, Mac80, MS84, Now84, Pom89, Smi82b, Tut80, Wan82, Wet82, Wet83, Woo80]. Publication [Ano81b, Ano82e, Ano82i, Ano88l, Edm80b]. Publishers [Dav86]. Puccoon [KJP89]. Puget [RDE82, Run85]. pugio [Mor80a]. pulex [DH88, HD84, Hav85, HGS83, HS84b, KD81, MMH85, SIH85, Van86, VKW89, WK89]. pulsed [LLG86]. pure [KBO84]. purple [ER88, vGMMG85]. purse [KC85]. pycnocline [GA81]. Pyramid [HG85b]. pyrifera [ZR85]. Pyrite [HM84, Kin83]. Pyrolysis [BWG82]. Pytkowicz [Whi86].

quadricauda [WST80]. quality [Høj89, MH86, MP88, RD83, VL89]. quantifying [WLG85]. quantitation [SF84]. Quantitative [DM84, MH87, NG87, SK83a]. quantum [KM83, KOTeI86, LUP88, MLG87]. Québec [MP88]. quickly [BSW81].

R [Bir88, Bro82, Cat85, Den82a, Dug80, Edm82e, Edm84a, Edm84b, Fen80a, Fen80b, Fre84, Fry80, Gar80, Gie82, Gil80, Har89a, Hec86, Jah83, Jum82, Jum84, Koh82, Lam87, Leh89, Mac80, Mar83, New86, Pai80, Pai89, Rzó81a, Shi85, Smi82a, Smi82b, Ste84, Sto89, Wan82, Wet83, Whi86]. Ra [BH88b, FTKB87, TCN^{+}82]. radar [MPsCS80]. Radertiere [Edm81]. radiance [Car85, KA89, Mob89, Vos89]. radiant [KBO84]. radiation [BF87, CA89, Fle89, JY84, MS84]. radiative [DS87]. radio [HA85]. radio-controlled [HA85]. Radioactively [HFA80]. radiocarbon [BKF^{+}87]. Radioecologie [Fow80]. radioisotope [ZFFC88]. radioisotopes [San88]. radioisotopic [HK81b]. Radiolaria [Swa83]. radiolarians [SA85]. Radiometric [BT84]. radionucleides [Fow80]. radionuclide [SMK81]. radioprotection [Fow80]. radiotracer [LCPL88]. Radium [SC84]. Radon [GM84, IJ84]. Radon-222 [GM84, IJ84]. Radunia [JS83]. rainfall [CB86].

Raleigh [Ano83f]. random [McB88]. range [CM86, SM82a, SG86]. Rapid [BKB85, MTW^{+}88, SB89c, ARP82, FNL85, GG82, Paa80, RE82]. raschii [McC86]. Rate [Edm80b, BB89a, D’A80, DP80, Ger82, GB80, HHH87, HW85, KRCS82, KH86b, MD85, MS85b, MLG87, OJR^{+}86, PL89, RE81, RW89, SN85, SG84, TC85, WK86b, ZBM82]. Rates [Atk87, Bru89b, Båm85, Ban82, BMH85, BG82, Bot85, Bra82, BSG83, BS85b, DG80, DeM82, Dev83, FO82, Gol86, Gri88, Han85, HF81, HP89, HASI80, HGS83, HM82, HK81b, HW86, IJ85, JKMN87, JRG81, Jør87, KN81, KC81, KRFS84, KH86c, KMSF81, LRH^{+}84, LDR87, LS85, LWCV83, Lyn83, Mal82, McC86, MC82, MHB84, MF89, NL89, OØ85, OSML87, PB80b, PD84, QBHS80, SM85, SSS87, SF89, She84, Smi82d, Smi85, SB89a, Som89a, SB89c, SH85, SH88, TS81a, TCN^{+}82, Ver85, WC86, Wei80, WL84, Yen85, DP81, HMLC88]. ratio [BB89a, DKMvL86, GCD87, Gro89a, KN81, LS82a, OJR^{+}86, PS89, TS85]. ratios [AHRF89, AS83, DL85a, Gol86, GM88, SW80, SB89c, SS88b, SH88]. Raton [Bro83]. Rayleigh [KA89]. re [CB84, Mor84]. re-evaluation [CB84, Mor84]. reaches [Mur84a]. reaction [SRW80]. reactive [GM85, MMW86]. reactivities [HCC88a]. reactivity [Fis86]. readers [Jum88a]. real [LM86]. Reassessment [CC84]. received [Ano80d, Ano82d, Ano84g, Ano87g]. Reciprocal [CPT86]. recolonization [Thr81]. Reconciling [RS86c]. reconstruction [BF84]. record [BWS^{+}85, LCK89]. recording [Den83]. recordings [Iba81]. Records [LG88, Haw80]. recovering [HHH83]. recovery [Hec84]. Recruitment [MSG86, CPT86, Ebe81, Eck83, Ohm87, SA89]. rectangula [Fre88]. recycling [LCPL88, OJR^{+}86, SH81]. red [CS85, TS81b, VCG^{+}87]. red-tide [CS85]. redistribution [Hil85]. Redox [HK86, HL85, Lyl83]. reduced [Fie82, GHW84, KS82b, RV87, Thr81]. reductase [BMPD84, Kri83]. reduction [Chr84, HE88, HKT89, HG83, HM84, HHTV81, HDK84, IB82a, JEM82, Kin83, Kin88, Ols80b, RDW^{+}88, RKF86, SO87, TAJ89, WB84]. reef [AM83, Atk87, AS87, Car85, CWV^{+}88, EBSW83, GCAN87, HH81, Lew87, PS84, Vic84, SA89]. reefs [Ano86j, Koh87, MS85a, Smi81b, Wol87, Koh84, Smi82b]. reference [VHJ^{+}89]. reflectance [CS85, CSPV86, Gor89b, VL89]. reflected [BB89b]. reflection [TB83]. refractive [KA89]. refractory [Gel86]. regeneration [GCD87, KM89, Law85, LN86, RM83, WKLK89]. region [BMPD84, DWO86, GZG80, HF84, HFL84, VL89]. Regional [KSH85]. regions [EJ86, HB84, KW87]. Regularly [HS86]. regulate [SKW86]. regulated [Rzó81b]. Regulating [BK87, KRCS82, Wal85]. Regulation [GCD87, Han85, SH85, Boe89, KH86a, Kon89, OW84, Sta85a, vRS83]. Reidel [Mac85]. reinterpretation [SK85b]. reinvestigated [MB89b]. rejetes [Fow80]. related [Pat89, SWSC85]. Relation [BF87, Fen80c, RM86b, Car83, Che80a, Fee79, Håk87, Jør87, MMD^{+}80, Pat80, RI88, SF82, Sul80, VAI85, VP85, WL85b]. relations [Now84]. Relationship [BS80a, BMPD84, KCB89, Run85, BBS80, BP87, Car80, CG80, HHOC89, Ker88, Kir84, LKS89, RGS89, ZF89]. Relationships [CSPV86, Car85, DM83, HB84, MB89b, Sta85b, BB81, DKMvL86, DC83, FDW85, GHK80, Ive88, KKH86, KH86b, Lor80, Mal83, Mil84a, SW81, WS81, Wil80a]. Relative [BMH86, Bor87, Leh88c, PCP87, SM85, BK87, CK80, CK84b, DeM82, MTL84, PO81, UEB88, VL87]. Release [Leh80, CK82, CWV^{+}88, Dou87, Dou88, GM85, GMM88, GS81, HJP87, MFHM80, OØ85, OJR^{+}86]. released [BH84]. releases [DC87]. relevance [NSL86]. relevant [SNR86]. reliable [Cru87, Jah88]. relicta [RS86a, Thr81]. remineralization [Gar84, GG84, GLMA82, HF84]. remote [CS85, CSPV86, CSHO89, KE80, MB89b, Pla86]. remote-sensing [CS85, CSPV86, MB89b, Pla86]. removal [TC84]. replace [LDYY83]. Reply [BB88, CK81, Fee80, JC81, KR89, Pec81, RF89, WM87, CP89, FDKM86, Gie82, Oud89, SK85a, Sto88b]. Report [Lis85, Edm86a]. Reprint [Man80]. Reprinted [Edm82d, Mur84b]. reproducing [Kim87]. reproduction [MG83]. Reproductive [RDE82, Ste88a, Bra82, BSG83, THGD83, Wal85]. Res [Lon86]. Research [Fen82, Lis85, Wet87a, Wet87b]. reserves [HS84b, THGD83]. Reservoir [McK81, ES82, EH88, GMK80, JKMN87, SF82]. reservoirs [LW84, Ree81, Edm87]. Resh [EA84]. residence [CWD^{+}86, HACF85, San84, WC81]. residence-time [CWD^{+}86]. residual [MBT80, TD87]. resistance [SBN^{+}83]. resolution [JD86a, MH86, RJ83, RMJ86]. resolved [GLMA85]. Resonant [JSS81]. Resource [PB87, KKH86, Som88, Pim83]. resources [FP87]. respect [Ber89]. respiration [AHI^{+}86, BA87, CCS^{+}89, MF89, MMM81, PGO82, RD83, Ver85]. Respiratory [WHFT89, VN87]. respirometer [Gri88]. respirometry [GCAN87]. respond [KGGW88]. Response [HM86, SS88b, ARP82, Atk88, BC82, Ger82, GK88, HMG86, KW86, KDC86, LP81, LL84b, MH82, McK81, MLG87, MB80, PM84a, Pri89, SYB^{+}89, TS82]. Responses [SM84, DH88, Fie82, Jac87, RV87, TS81b]. resulting [Gar81a, Gar84, Hun83b, JC80, LMS87]. Results [Kep85, WBV80, RS89]. Resuspension [DTBB87, DMM82]. retain [CK84a, Whi81]. retained [Ven87]. retention [Cor89]. revealed [GK83, Lue88, San88, SGMB87]. Reversal [FM82]. Review [Ano85d, Ano86j, Ano89i, Ban84a, Ban84b, Ben89, Bez89, Bir88, Boe89, Bro80, Bro82, Bro83, Bur88, Bur80, Bur86a, Cam87, Cat85, Cha85, Csa80, Dav83, Dav86, Den82a, Duc87, Dug80, Edm81, Edm82a, Edm82b, Edm82c, Edm82d, Edm84a, Edm84b, Edm87, EA84, Epp82, Far88, Fen80a, Fen80b, Fen82, Fen83, Fen84, Fow80, Fre84, Fry80, Gar80, Gil80, Gop89, Hal82a, Har89a, Har89b, HG88, Hay87, Hec84, Hec86, Hen87, Hor83, Hrb82, Hyn81, Ill80, Jah83, Joh89, Jum82, Jum84, Jum87a, Kir88, Koh82, Koh84, Kru80, Kru82, Lam87, Lec85, Leh89, Lis85, Lon84, Lon85, Lon86, Lun84, Mac85, Mac80, Man80, Man83, Mar83, McM86, Mil82, Mil87, Mil84b, Mol89, MS84, MS87, Mor80b, Mor88a, Mur83, Mur84b, New86, Now84, Pae89, Pai80, Pai84a, Pai84b].

Review [Pai89, Pic82, Pic85, Pim83, Pla86, Pom80, Pom89, Pon82, Pra80, Ril82, Rzó81a, Rzó81b, Ser80, Shi85, Smi82a, Smi82b, Smi86b, Smi88, Ste84, Sto89, Thi83, Tip87, Tut80, Wan82, Wel88, WK80, Wet82, Wet83, Wet84, Wet87a, Wet87b, Whi86, Woo80, Woo85, ZoB80, Gil88, HK88a]. revised [SS88a]. Reynolds [Tip87, KS81]. Rhizophysa [Pur81]. Rhoads [Koh82]. Rhode [DDSmV83, MMD^{+}80, Cor81]. Rhone [LG88]. rhythm [SLL85]. ribbed [JV82]. rich [KM89, Mal82, NL87]. Rigler [Bur86a, Edm82e]. ring [CRGC87, Han87, HLS87]. rings [DH85a, DH85b, GBF86, MN86]. rise [Ore85]. rising [WBS83]. River [Bot85, CDPC84, Cor81, Duc82, FP81, LHM84, LSLW86, SL89, WS81, Bay88, HFFS85, KTM83, MH80, MG80, NL87, SH81, SB89c, CD85, CM80, DQRM87, EHD^{+}86, GHC89, HL87, HTE84, HCQ^{+}86, Jon82, LWF86, MDR88, MSG81, RDW^{+}88]. river-floodplains [Bay88]. riverborne [LG88]. riverine [BBA84, Pat88]. rivers [CM89, CW83, Edw87, Fro88, dAL87]. Rn [CA87, DQRM87]. RNA [BS80a]. roach [CPT86]. Robinson [Pla86]. Rockies [Lew82a]. Role [HS86, HA81, MBT80, SS88a, YAL^{+}81, Ban82, CCF89, DQRM87, Dod88, FHHB80, Hab85, HL88, Kil82, MS84, OW84, PTB83, Pae85, PB81, Pet86a, RKF86, Sca88, SSB82, SK82, Swa83, WB84, Wil87, dJV89]. roles [Bor87, Luc86]. Rona [Lon85]. root [BB84b, RPL89]. Rosenberg [EA84]. Ross [Man83, Hor81]. Rotatoria. [Edm81, Edm82d]. Rotifer [Edm82d, GS84, GK88, Gil88, Gil89, MG83]. rotifers [Bor83, PO81, SPBD89, WB86b, Wil87]. Rotoiti [PSGD86]. Rotterdam [Ban84b]. roughened [Mob89]. Row [Pai84b]. Rowe [Lon84]. rows [HS86]. Roy [Man80]. Royal [Man80]. rubescens [Kon82]. Rubey [Mil84b]. rule [Lan83]. Rupert [HW85]. rusticus [Cap82]. Rutilus [CPT86]. Rutledge [BW89].

S [Fen82, Hay87, Leh89, Mac85, Mil82, Mor80b, Now84, Pai84a, Pai89, Pla86, Smi82a, Tip87, Wet84, Fry88, HF84, HFL84, Kin83, RJBC83, TNEJ84, TAJ89]. S-labeled [TAJ89]. S.A [Edm84a]. Saanich [Dev83]. sacculifer [JERC85]. Sagitta [Fei82, Sul80]. saline [GSB89, HGH82, Har87, HG85b, Sir87, TB84]. Salinity [GA81, Gie82, Mil84a, Par82, SC82, VAI85, Wal80]. salp [HMG86]. salt [BMH85, Boo80, CB86, GH84, HS80, HF82, HKT89, Hob88, HG83, HM84, HHTV81, HDK84, HDW85, IBC^{+}83, JV82, Kin83, Kin88, LR85, LCP80, LL82, LBB88, PH87, SGMB87, Ste83, VTV^{+}80, WMSW89]. salt-marsh [PH87]. salt-water [LBB88]. saltern [Jav83]. salts [CBH83]. Saltzman [Ril82]. Sample [SK83a, GL81]. sampler [BB84b, BWJ80, HHS80, Jah88, RE82, Wal80, lWR89]. samples [AF85, BS86, HK81b, KP82, MMW86, RS81, RS88, vdBD84]. Sampling [VW80, BBH85, BW89, BMH^{+}87, Car84, DPF87, FP82, HDW85, JD86b, KE80, Lyn83, RF88, SG81]. San [Ano81e, MPsCS80, MPsCS80]. sand [SC88]. sandy [LCP80, SRK86]. Santa [ZR85]. Sapelo [HG83, PH87]. Sargasso [Alt89, CI88, IS89, PM88, SMP^{+}89]. Sars [Fre88]. satellite [PGB88, PM86, Pla86]. saturated [CMK89, Mal82]. scale [Gie82, SFDK84, Smi86a]. scaled [SH89]. scales [DG83, GPHI83, YMBO87]. Scaling [MJN84]. scanner [GPS85, Mor88b]. scanning [WHM^{+}85]. scattering [Ban88, Jon87, KA89, QHHLS89, SLL85, WB86a, WJ89]. scavenging [BSW81, NSL86, WBS83]. Scenedesmus [NO89, WST80]. Schacht [CG80]. Schindler [Fee80, Sto88b]. Schlegel [Ano89i, Gop89]. schooling [HHOC89, MS85b]. Schweizerbartsche [Edm82b]. Sci [Bur86a, Fen80a, Hor83, Koh84, Kru80, Lon85, Mac85, MS84, Pic85, Pom89, Rzó81a, Smi82b, Thi83, Tut80, Wan82]. Science [Gar80, Jum82, Wet82, Pre86, Cha85]. Sciences [Ano81e]. scientists [Pla86]. scintillation [RS81]. Scirpus [BWG82]. Scope [Edm84c]. SCOR [Hor83]. Scotia [FVH^{+}84, MTL84]. Scotland [HP87]. Scottish [ADO85]. Scottolana [LL89]. scum [vRS85]. scum-forming [vRS85]. scums [PU82]. Scyphomedusae [IH81]. Sea [Dug81, GK89, BMP81, BHK^{+}82, Car83, CL81, Che80b, CGCG89, Csa80, Csa86, DS87, ER88, Fra88, GB82, HSMH87, HACF85, Jah85, KOTeI86, Lon84, Man80, NJF84, PS85, RS86c, SBN^{+}83, SVG89, Smi85, SHRW84, SRW84, SJC^{+}87, VW80, WL83, ZST89, Zan89, AN81, Alt89, Ban84b, BCM86, Bro80, CI88, GSB89, GK83, HS86, Hun81, IS89, KSTGB81, LB84, LGM80, MB89a, ML82, MSG81, OS82, Ore83a, Ore83b, Ore85, PM88, Sau80, SA81, Sir87, SMP^{+}89, Ste83, Ste85, SC84, VN87, Woo84, Ill80, Mac85, Mil84b, Wel88, Duc87, Kir88]. sea-ice [PS85]. sea-surface [HACF85]. seabird [Han87]. seabirds [Har84]. seafloor [Lon85]. seagrass [Fen80b, MK84, Pet86c, Smi81a]. Searching [Jan81]. Sears [Pic82]. seas [Lon84]. season [DKMvL86, Pet86c, Run80]. Seasonal [Båm85, BG82, JKMN87, MTL84, PGB88, SPBD89, TW83, Two83, VWVR84, WLB84, BC80, Eva88, Hai88, MP89, Run85, TML^{+}88]. Seasonality [CCW87, HL87, KM89, LE83]. seasonally [DH80, DWR82, HL85]. seaward [Smi82b]. Seawater [JW85, AF85, AC82, BK84, BRH^{+}88, BSF81, CMB87, CV82, Cul80, DSB86, DP81, DP86, Gar81a, Ger83a, Ger83b, GZG80, HA83, HS82, Hun80, JC80, Joh82, JGCS82, JP83, KSTGB81, KH80, LSK^{+}83, LK89, MF86, MPF88, OC82, RWB87, SAJB86, SB89b, SH87a, VW83, WSU^{+}83]. seawaters [HFFS85, OK87]. seaweed [RMN82]. seaweeds [Hay87]. Secchi [BWS^{+}85, Car80, Edm80a, MB89a, Pre86, Lor80, MSBC80]. second [WC87]. Secondary [Bri82, HH85, Ive88, Bur86a, HH86, HW86, Hyn80, KM87, Kim87, KM80, Leh88a, Men80, MMR87, Sca88, Str88, THCP81]. Secretary [Ano89a]. section [KM83, KSP^{+}81, PP89]. Sediment [Blo81, DF82a, Luc86, SLB83, Tut80, And85, Bal84, BF84, BA87, BH88a, BB86, BB84b, Bow84, Bru89b, BWJ80, CW85, CDG88, DGN82, DWR82, DL85b, GM84, HTWR84, Hil85, HLA86, JZF80, JD86a, KC81, KM89, LMS87, LW81, LSB81, MJN84, MH86, NSL86, RPL89, RS86c, RG83, Smi87, SB89a, SHRW84, SRW84, SJC^{+}87, TS83, Wil84, vdLAH^{+}84, Fen83, Now84]. sediment-based [BF84]. sediment-trapping [RG83]. sediment-water [BH88a, BB86, GM84, NSL86]. sedimentary [DR88, Fra87, GH84, HTE84, HKT89, RMB82, WB84, WST80].

Sedimentation [HDP83, SGMB87, BH88a, BB86, BBS80, BU86, BB89b, CEE86, DWO86, FB80, JRG81, Sta86]. sedimenting [DF82b]. Sedimentol [Smi82b]. sedimentological [TNJ84]. Sedimentology [Kru80]. sediments [AHRF89, AATC82, AR88, BMS^{+}88, Ber89, BH83, BM82, BD81, CK88, CF81, CW88, CB86, CMK89, Chr84, CM83, DM85, DM83, Dev83, DL85b, EBSW83, FP81, Fro80, GMM88, GNM87, GB82, Ham81, HP87, Haw80, HW85, HCC88b, HB85, HF84, HFL84, HD86, HL85, HASI80, HM82, HM84, HDK84, HDW85, HY89, IJ85, JEM82, Jah85, JK84, JEE82, JBvdB89, Jon82, JR85, KRFS84, Kep85, KE80, KK80, Klo85, KB80a, KB80b, KM84, KMD^{+}88, LSdKvD85, Lyl83, MA84, May89, MJSM89, NSM87, Pat89, RE82, RSBL80, RJB81, RMJ86, RKF86, SN85, SA86, SWSC85, SRK86, TAJ89, UA82, WSU^{+}83, YJ84, Zül81, Now84, Tut80]. seep [BMS^{+}88, KBB^{+}89]. seepage [FP81, Lew87, SP89b, Win81]. seiche [LaZ80, LM86]. seiches [HMS86, LM86]. seine [KC85]. Seki [Bro83]. Selection [TS81b, AFM87, DeM89, Fen80c, Kir85, Pas80, RFS84, VP85]. Selective [DdNK82, DMdK84, LO81, Mur81b, SM83, Yen85, DHV^{+}87, Gro89a, KK80, SSB82]. selectivities [BG82, DeM82]. selenium [CB84]. Sell [Pec81]. SEM [JRC83]. semicontinuous [HA81]. senescing [Lan82]. sensing [CS85, CSPV86, CSHO89, MB89b, Pla86]. sensitive [FNL85, FH82]. sensor [Atk88]. SEP [SH87a]. SEP-PAK [SH87a]. separate [McM86]. separated [LL89]. separation [SIH85]. September [Pic82]. septentrionalis [HGM80]. sequencing [JZF80]. Ser [Csa80, Lon85, Mac85, Man80, Man83, MS84, Pic85, Wan82]. series [Alt89, Iba81, JZF80, McC86, SE82, Sta80]. serious [lWR89]. Serratia [Syz82]. Serruya [Woo85]. serum [ZST89]. sessile [SG81]. seston [DM85, GM85, MTL84, MP88, TC84, WBG^{+}87]. Sestonic [Edw87]. settlement [Han81, Mul89, Pet86c]. settling [AG88, BR80, GM85, SSK87]. Sex [GM88, DKMvL86, HWL83]. sex-specific [HWL83]. sexual [GC81]. sexuality [D’A80]. shad [DdNK82]. Shagawa [LSM81]. shallow [Ber89, DTBB87, Edm82c, FP82, Hec86, JBvdB89, Men81, OAI^{+}85, Pra90, PG83, Smi81a, Som89b, TM84, WWS84, Edm82a]. shallow-water [Ber89]. Shan [Hyn81]. Shape [SW89, ZK80]. Shark [SA83]. shear [Mon85]. Shelf [Hor81, Chu85, Fis80, Mac84, VCG^{+}87, YAL^{+}81, YMBO87]. shell [TCN^{+}82]. shelves [Pom89]. Shifts [BCL85, EJ86, HG85a]. Shipboard [HB85, Ami88, ML82, OGMG88, WD87, lWJ82]. Short [CRGC87, Gil89, HLS87, Pat89, WGM82, Bur84, CW84, Dug81, Hak84, SP89b, Two83, WHFT89, WC81, ZFFC88]. short-lived [ZFFC88]. Short-term [CRGC87, Gil89, HLS87, Pat89, SP89b, WHFT89]. shown [PGB88]. shrimp [Mor80a]. Si [Gro89a]. Sibaya [Fry80]. sicilis [VP85]. Sida [DP80]. side [Rie84]. siderophores [MM80]. Sieburth [Bro80]. Sierra [Sto87]. Sierran [Sto88b]. sieve [Paf84]. signals [SCO89]. Significance [CK82, Chr89, DF82b, Dri81, FTKB87, GC81, GK88, JV82, LS84, O’D85, RW89, SL87b, Sei88]. Sih [HG88]. SIL [Wet83]. silica [CKT89, MG80, PO80, SEK84, SGSS84, SWSC85, TS85]. silicate [MMW86]. silicic [RM81]. Silicon [NMM83, Paa80]. Silicon-32 [NMM83]. simple [Cru87, Den83, FP82, GHW84, Jah88, JD86a, KM83, Nür88, Rod87, TQS86, Wal80, Wof83]. Simulating [Jac87, WL83]. Simulation [Jac89, JS83, PHI84, PH88a, RMJ86, Unc81, Gar80]. Simultaneous [BKF^{+}87, MMG86, NG83, Som88]. simultaneously [OSML87, PME87]. Sinclair [Boe89]. single [RS81, SCO89, War84]. Sinking [ALRP85, LWCV83, KMSF81]. sinks [MKN84, SSGB89]. siphonophore [Pur81]. site [DTS^{+}89]. sites [BCY89]. situ [ARP82, All81, AG88, Blo81, BS85b, Bur84, Bur86b, BR80, CW85, Cul80, DP80, Fra87, HW86, Kep85, Loe81, LSB81, MC82, NL89, Nür87, Ols87, Pet86b, RPC89, SNM^{+}87, SLL85, Smi82d, Smi85, Ste83, Sul80, WC86, Wei80]. situations [Col82, Col80b]. Size [KK80, RO82, ADK87, AK89, AFM87, BS80a, BCL85, BG86b, CG80, DW86a, DKMvL86, DP80, DHV^{+}87, Ebe81, ER88, Fen80c, Fur87, GC81, GM88, GWB89, Ham81, Han85, HW88, HH86, Ker88, KM80, Lyn83, Mal83, MMTL88, NG83, Pac86, PL89, PP85, RFS84, SB89a, Som89a, TB80, VP85, Vic84]. size-fractionated [Fur87, PP85]. size-frequency [HH86, KM80]. Size-selective [KK80, DHV^{+}87]. size-specific [Lyn83]. Skagerrak [IJ85]. Skeletal [Koh82]. Skeletonema [SAK89]. skeletons [Bra80]. skies [BF87]. slacki [GS84]. Sleigh [Kir88]. slicks [Söd87]. slides [McB88]. Slinn [Mac85]. Slope [BW86, Chu85, DK86, KBB^{+}89, LG88]. Slow [AS87]. Small [SFDK84, SSFN86, Smi86a, AMP81, BS85a, CA83, CPT86, Csa86, Cuk87, Gri88, Hil85, HLA86, HT85b, JGCS82, LaZ80, LSI^{+}85, LI87, Mac86, PP86, Rah85, RS88, Sim87, Van86]. Small-scale [SFDK84, Smi86a]. small-volume [RS88]. Smith [Lon85]. snails [Har89a]. Snedaker [Wet87a, Wet87b]. Snell [Jan81]. snow [ACC86, AG88, PA83]. Soc [Fen83, Man80]. Society [Ano80a, Ano80e, Ano81a, Ano81b, Ano81c, Ano81f, Ano82a, Ano82b, Ano82h, Ano83a, Ano83b, Ano83f, Ano84a, Ano84b, Ano84h, Ano85a, Ano85d, Ano86a, Ano86b, Ano86l, Ano87a, Ano87h, Ano88a, Ano89b, Ano89c, Ano89a, Ano89m, Hrb82, Koo85, Man80]. Soda [CCO83, CCW87, IOK87, SO87, ZHO^{+}87, MM82a]. soft [Kor80, Pse88, TW83]. soft-water [Kor80, Pse88, TW83]. Softcover [Edm84b, Jah83, Lis85, Fen84, Smi82a, Smi82b]. soils [Kin83, RK87]. solar [Kir84, MS84]. Solemya [CCF89]. solids [HS84a]. solitary [SA85]. Solubilities [KSTGB81, Avn83]. solubility [DH80, GSB89, LSK^{+}83]. soluble [FSW85, GM85, LH80]. solute [Lee80]. solutes [LHM84]. solutions [LSK^{+}83]. Some [DH88, Mal83, MG81, MRE84, SFDK84, WC81, Wil80a, Wol87, BWG82, BMP83, But82, BR86, CDRW86, DKMvL86, LG89, SNR86, SH88, SA85]. Somerset [Dav83, Lec85, Mil82, Mor88a, Pic85, Pla86]. Sommer [Pra90]. Sons [Cam87, Cha85, Dav83, Gil80, Ill80, Lec85, Lon84, Mil82, Mor88a, Pic85, Pla86, Pra80, ZoB80]. sorption [SH87a]. sorting [SMR84, YHM^{+}83]. Sound [Fei82, PS85, RV87, RDE82, Run85, GWB89, TB83, Ril82]. source [BF84, BMH85, CS81, CLS82, Edw87, GB80, LK89, MS85a, VAI85, Wot82].

Sources [Fry86, GCLC87, HTE84, MN86, ES83, HCC88b, Ohm87, RJ84, SRW80, SSGB89, MKN84]. South [Ban87, Gar80, Jum82, RMB82, KCB89, Kin88, MTW^{+}88]. Southeast [Mur84a, SA81, Sta85b, VN87]. southeastern [CM89, Edw87, EJ86, MRLD86, MB89a, Mul81, RI88, Woo84, YAL^{+}81, YMBO87]. southern [Che80b, LB84, LK89, MP88, Mor88b, OH87, Pet86b, SP85, SE82]. Southward [Far88]. southwestern [FVH^{+}84]. sowerbyi [DC83]. sp [HSMH87]. Space [AM83, DG83, Hor83]. Space-time [AM83]. spaced [HS86]. Spain [vGMMG85, PAGG87]. Spanish [GMPA^{+}85]. Sparkling [LKS89]. Spartina [LHH80]. Spatial [BD81, CM83, DR88, HHTV81, Lew87, Lin87, Mac84, Wei84, YMBO87, APR80, BSON84, Bra82, Car83, GHC89, KSP^{+}81, LT88, RJ83, RMJ86, Smi86a, Yen85]. spatially [IP85]. spatio [HK81a]. spatio-temporal [HK81a]. Spec [Fen83, Smi82b, Tut80]. Special [Ano89l, Hrb82, VHJ^{+}89]. speciation [BSF81, Kan83, Boe89]. Species [HK81a, LS85, AB80, BWG82, BR80, CD85, CVG89, CL81, CM80, Eva88, GG82, Gro89b, HASI80, HY89, Hun81, IL85, JT84, LH80, LGV85, Paa80, Pet88, SC81a, SH83, Van86, RV87]. Species-specific [LS85, RV87]. Specific [HM87, Nic89, HWL83, LS85, LD86, Lyn83, Mal82, PS81, RV87, SLP87, WL84]. specificity [Bel83]. spectra [IMK88, LWP85, LUP88, MH87, MD85, RPC89]. Spectral [EHE^{+}83, KS82a, VL89, AB80, MCK89, PS81, SLP87, SP89a]. spectrometry [BWG82, HHH87, KP82, WHFT89]. spectrophotometric [FN85]. spectroscopy [CI88, WHM^{+}85]. spectrum [RE81]. spicatum [BB84b, SA86]. spine [VKW89]. spined [HD84, Hav85]. Spines [GBB81, Ker88, WK89]. Spirit [MTW^{+}88]. splendens [Fie82]. splitter [GHW84]. Spoel [Mil82]. sponges [CWV^{+}88]. spp [CI88, GMFLJ85, HF89, WHM^{+}85, vRS83]. spreading [BT84, Lon85]. spreading-center [BT84]. spring [AN81, CP83, Duc82, Ive88, LFRT86, LB84, LBZC88, SA81, Wro89]. spring-summer [CP83]. Springer [Bur80, Dug80, McM86, Pic82, Ser80, WK80]. Springer-Verlag [Bur80, Dug80, McM86, Pic82, Ser80, WK80]. springs [RJ84]. Spumellaria [SA85]. squeezer [BMH^{+}87]. St [Ano84h]. stabile [EHE^{+}83]. stability [EHE^{+}83, HB84, Kli88, Luc86, RNVB81]. Stable [CSF88, LS82a, MRE84, CCF89, Fry88, OAI^{+}85, QEQD86]. stage [AH87, Bor83]. stage-frequency [AH87]. stages [SRW81]. staghorn [BC80]. staining [Col80a, Har87]. standing [Jav83, Loe81]. Stanford [Rzó81b]. Stanier [Edm82e]. starfish [Ols87]. Starvation [HMG86, THGD83, Col80b, OBP87, Ols87, Wil80c]. starvation-induced [Wil80c]. state [Car83, DP81, HK86, KSH88, KM83, Kon89, RM86b, SAK89, Som85, Sta85b, Whi89, Fry86]. Statement [Ano80a, Ano81a, Ano82a, Ano83a, Ano84a, Ano85a, Ano86a, Ano87a, Ano88a, Ano89b]. states [BH88a, BB86, MSBC80, Gil80, GDS83, SC81b]. Station [Sul80, lWJ82]. statistical [SH89]. statistics [AH87, GBF86]. status [FW82, HS84b, PLN84, Pra90, Söd87, Som89b, VJOR88]. steady [DP81, KM83, RM86b, SAK89, Som85]. Stenosemella [RE81]. Sterols [LHH80]. stimulation [GK88]. stimuli [Dod88]. stimulus [D’A80]. stocked [GR84]. stoichiometric [ZST89]. Stokes [KA89]. Stommel [Wel88]. stoneflies [AFM87]. Stony [Ano80e, KDR89, Sel81]. storage [MMW86]. stored [BS86]. STPM [Fen80b]. Strait [MPsCS80]. strata [SK82]. strategies [Lun84, Str83]. strategy [NJF84]. Stratification [HH80, DS87, FHHB80, KK80, Kli88, Ste85, Ven84]. stratified [BBG89, Brz88, CB87, HGH82, KOTeI86, Mon85, Nür88, Sta82, WC87]. stratigraphy [Haw80, Zül81]. Stream [O’D85, Bri82, CSW^{+}80, HDLP85, HK88b, KB82, Kir89, Klo85, MMG86, Mul81, Mur84a, Pet86b, PHCK83, SC88, WH80, Wil80a, GBF86, Han87, HLS87, New86, YAL^{+}81]. Stream-dwelling [O’D85]. streambed [KB89a]. streams [Mor84, Sel81, WBG^{+}87, Rzó81b]. stress [HM86, McK81, MB80, Sta82]. Strickler [Lam87]. Stromatolites [Shi85]. Stroudsburg [Man83]. structural [HF89, McL86]. structure [Alt89, BCL85, DdNK82, DMdK84, DHV^{+}87, ER88, EJ86, Fry88, Gil88, Hrb82, JRC83, Leh88a, Leh88c, Lem87, LS81, MDM80, Óla80, Pac86, Pat88, PHI84, SA81, TD87, THCP81, Wei80, Pim83]. structures [GK89]. Stud [Lon86]. Student [Ano88e, Ano88m, Ano88n, Ano89n]. Studer [Fre88]. studied [JERC85, MF86]. Studies [Bur80, Epp82, Fen84, NL87, Ril82, WK80, Woo80, Ami88, ADO85, BW89, BH88a, BB86, BB84b, Che80a, HH86, HI88, HA85, Høj89, HFA80, HW86, JZF80, Joh89, JRC83, JD88, LCPL88, OGMG88, QHHLS89, RNVB81, RJBC83, RF88, SGMB87, WBG^{+}87, lWR89]. Study [Thi83, Alt89, AC82, BS80b, Bow84, Bur84, Car89, CK79, Den85, DHV^{+}87, Elm83, Fra87, GG81, HF82, LFRT86, Lon84, McK81, Pri89, QEQD86, TD87, WD87, WBV80, YJ84]. studying [CW85, NMM83]. Stumm [Cam87, Mor88a]. Stuttgart [Edm81, Edm82b]. subalpine [AGG82, GJP89]. subantarctic [Fre88, Ohm87]. subaquatic [Blo81]. subarctic [DW87, WKLK89]. sublethal [Man80]. sublittoral [DTBB87, HP89]. submarine [KA89]. submerged [BB84b, CK82, DK86, LS82a]. submersed [KLJ86, LKS89, Ste88b, TS82]. submission [Jum88b]. suboxic [GB82]. subsampling [GHW84, SMR84]. subsequent [Pet86c]. subsidence [HDP83]. substance [BMP81, DCV87]. substances [CS81, CV82, EHD^{+}86, HD86, MTW^{+}88, TW83]. substoichiometric [HS82]. substrata [KB89a]. substrate [But86, CK79, FS84, GCL89, GCD87, GG81, Sel81, SB89a, Wil80a]. substrates [Chr84, Nic89]. substratum [Bor87]. Subsurface [HF82, Bur84]. subtidal [BW89, NJF84, RF88]. subtropical [BSON84, Elm83, FF85, LDR87, SB89c]. success [GM88]. succession [Eva88, HJP87, Som88]. suitable [lWJ82]. Sulfate [HE88, HKT89, HG83, Kin83, TAJ89, Chr84, CW83, HM84, HDK84, IB82a, IJ85, Kin88, KM89, KB80b, RKF86, SO87, WB84]. sulfate-methane [IJ85]. sulfates [UB89]. sulfide [DH80, HGH82, MPF88, SAJB86, TML^{+}88]. sulfoxide [FH84]. Sulfur [DM85, PH87, AHRF89, Fra87, Fry86, GMPA^{+}85, HY89, KS82b, Kin88, LGV85, PB81, RKF86, VAI85, WLG85, vGMMG85]. sulfuric [RKS^{+}86]. Summer [CM81, LSM81, CP83, FHHB80, FVH^{+}84, Han87, IP84]. sun [Gor89b]. sunlight [MH82]. superba [HHOC89]. Superior [JSS81, JEE82]. Suppl [Ban84a]. supplies [LSM81]. supply [Che80a, Gro89a, OJR^{+}86, Smi87, SH88, TC85]. supplying [FBG89]. Suppression [Gil88, Van86]. Surface [Høj89, MB89b, Sir87, VW83, ACC86, BF87, BS89, BA87, BHK^{+}82, BD81, Car83, CL81, CRFB80, CV82, Csa80, DP81, DP86, DH85a, FS83, Ger83a, Ham82, HACF85, HS88, Hun80, HL82, KSH88, LM86, Lew83a, LL82, LL85b, Mor88a, OK87, PTB83, PGB88, PP85, Söd87, SRK86, SG84, Syz82, VW80, VN87, WL83, WLR88, Wot82, dAL87, vdBD84]. surface-active [CV82, Ger83a, Hun80]. surfaces [DM83, Mob89, ZoB80]. surficial [HCC88b, JBvdB89]. surrounding [Lan82]. Surveys [Pai80]. Survival [GR84, Pet86c, Lun84]. susceptibility [DF82b, HL85]. suspended [Cuk87, Han81, HL82, Jon87, Kir83, LL85b, NL87]. suspension [DeM89, Fen80c, FBG89, PB87, RFS84, SC88, WB86b]. suspension-feeding [DeM89, Fen80c, RFS84, WB86b]. suspensions [Ban88]. Sustaining [Ano81c, Ano82b, Ano83b, Ano84b, Ano86b, Ano89c, Koo85]. sustenance [BK89]. swamp [Mul81, BMH85, FF85, Pom80, Wet84]. swamp-stream [Mul81]. swarming [HGM80]. swarms [BWG83]. Swedish [Båm85, Pet86b]. Swimming [Pri89, BMS83, FCO87, MB86, PGO82, RS89, ZK80]. Swiss [BU86]. Switzerland [IJ84, LMS87]. Sylt [BHK^{+}82]. sylvestris [Men81]. symbioses [DC88b]. Symbiotic [JERC85, BQ86, Dou87, Dou88, MM82b]. Symp [Dav86, Edm82d, Hor83]. symposium [Hrb82]. Synechococcus [CI88, CW84, GMFLJ85, KGGW88, SS88b, WHM^{+}85]. Synthesis [Man83, ADF^{+}82, CJTL83, COL84b, KH88a, WK86b]. system [Bot85, BSC80, CL80, Den82b, DWO86, EHD^{+}86, GSB89, JT84, LSK^{+}83, Mac86, OGMG88, Ami88, Bur80]. systems [Fen84, Gar80, Loe81, LH87, VW83, WK86a, Wri89, Bro83].

T [Ano88h, Ano88j, Hal82a, Har89b, Hut88, Jum84, Lon84, MS87, Smi86b, Ste84, Thi83]. T. [Gie82]. table [AH87, HT85a, HBT87, Win81]. tables [ZEY89]. Tahoe [Gol88, VG80, APR80, ADP^{+}84, BWG83, Thr81]. Tanganyika [Rud80, Hab85, HK81a, HF81]. tanks [HT85b]. Tarn [Haw80]. Tasek [Wet84]. tatricus [GR84]. Taub [Edm87]. taxa [SMR84]. taxon [SMR84]. technique [BS85b, FO82, FS84, Har87, HHH83, HDW85, Jef81, JD86b, KN81, Kep85, SMR84, SRW84, SH87a, TC84, YHM^{+}83, ZK80, vdLAH^{+}84]. Techniques [CMB87, SGSS84, VW80]. Teil [Edm82b]. telemetry [Den82b]. Temora [WC86]. temperate [DWO86, Har83b, Jon82, MP89, RKS^{+}86, Sta80, VL87, WLB84, WL85a]. temperate-region [DWO86]. Temperature [FCO87, Lew83b, MDM80, Óla80, RG81, SB89b, Str83, CM82, Cor89, GSB89, Ger82, HK80, HP87, Hun83b, JS83, KB89a, Kel89, KC81, Kri83, OP83, PS84, TB83, TML81, WC81, ZD89]. Temperature-length [Str83]. temperatures [Fei82]. Temporal [BSON84, Car83, Neh87, WMM83, CM83, EHE^{+}83, HK81a, Lew87, Lin87, MRE84, Yen85]. temporally [McC86]. Tennessee [Ano81f]. Tenore [Jum82]. term [Bur84, CMR89, CRGC87, CW84, Dug81, Gil89, Hak84, HLS87, Pat89, RKF86, SP89b, Ste85, Two83, WHFT89, WGM82, Joh89]. terms [Le 82, MBT80]. test [BB89a, CR80, LP81, Ols87, TML81]. testable [Mor88b]. tested [WB84]. testing [New86]. Tests [LM86, PHCK83]. testudo [Ste88a]. Tevesz [Now84]. Texas [Ano81e]. Texel [Dav86]. text [Edm81]. Th [CB85, CB87, FTKB87, TCN^{+}82]. Thalassiosira [AM82, HM86, LB80, PH88b, RM81, SH83]. Their [CSHO89, O’D85, RW89, BKB85, BWG82, CM89, CK79, CK80, Fra87, GG81, GM88, Har84, MG83, MV88, Mon85, PU82, RNVB81, Rie84, SB89c, VP85]. theoretical [Dri81, Smi82c]. theory [DeM89, Kim87, TML81, Zan89, New86]. Thermal [PH88a, SC81a]. thermally [KOTeI86, Nür88, Rah85]. thermocline [QBHS80]. thermodynamic [CM86]. theses [Ano88j]. thiebautii [LGM80, SNM^{+}87]. thin [Jef81]. thin-layer [Jef81]. thiols [MV88]. Thompson [Bot85]. thorns [Ols87]. Thornton [Hec84]. those [WL83]. thousand [ZEY89]. thousand-year-old [ZEY89]. three [Fry86, Hun81, OKG81, PB80b, SB89c]. Thurman [Hen87]. Thymidine [BA87, CDG88, DH85b, RBNF87, RW89]. Thysanoessa [McC86]. Tidal [Boo80, Pet88, CMK89, Duc82, GL81, HHS80, HA85, MPsCS80, PB87, VTV^{+}80]. tidal-powered [HHS80]. Tidally [DP89b, MBT80]. tide [CS85, TS81b, Unc81, ZEY89]. tide-prediction [ZEY89]. tides [CB86, VCG^{+}87]. Tien [Hyn81]. Tilman [Pim83]. timber [RD89]. Time [DG83, McC86, Alt89, AM83, BC80, CWD^{+}86, GPHI83, Har83b, Iba81, JZF80, KLC81, MH86, MLG87, San84, SE82, Sta80, VN87, WC81]. time-course [Har83b]. time-lapse [BC80]. Time-series [McC86, Alt89, JZF80]. times [HACF85]. tintinnid [RE81]. tintinnids [Ver85]. Tissue [MS85b]. Titicaca [VWVR84]. titles [Ano80c, Ano81d, Ano82c]. titration [lWJ82]. Tokyo [HFFS85]. tolerance [SC81a]. toluene [BR86]. Tomales [SHDV89]. Tomczak [Dug80]. Tonolli [Edm86b]. tonsa [DDSmV83, Rom84]. tonsus [Ohm87]. tool [NMM83]. Topics [Now84]. torrent [Hyn81]. Total [JEE82, AHRF89, AHI^{+}86, BF87, Bur86b, FDKM86, FDK^{+}86, HFL84, HS82, KW86, KS82b, KM84, RPC89, SS80a, SS80b]. toxic [NDS86]. toxicity [FM82, RM81, Rue83]. toxin [DC87]. toxins [Whi81]. Trace [Bra80, Dav86, Hun83a, Kep85, PH87, San84, San88, Mar83]. traced [IJ84]. tracer [Car89, CRFB80, CL80, LCP80, QBHS80, SW80, SC84]. tracers [Kin83]. tracing [GM84]. trails [Ano88k, HP89]. Trans [Man80]. transect [APR80]. transfer [CM89, Csa86, DS87, DP86, HHH83, SA86]. transformation [OAI^{+}85]. transformations [PSGD86]. Transient [Col82, Col80b, BSB^{+}87]. transition [IJ85, Lyl83]. transitions [EEMC88]. translational [KMK88]. translocation [Wil84]. transmission [KK80]. transmitted [SB80]. transparency [Gol88, Kli88, MB89a]. Transport [LHM84, TNJ84, CMK89, FS83, GM84, IBC^{+}83, KLB80, LCP80, LL82, MJN84, SBN^{+}83, SSK87, SH85, TS83, WBG^{+}87, WOS89, vdLAH^{+}84]. transportable [lWJ82]. transuranic [FBF83]. tranteri [KM87]. trap [JZF80, LMS87, LSB81]. trapped [DF82b]. trapping [Blo81, RG83]. traps [DWR82, DL85b, HTWR84]. Travel [Ano88e, Ano88n]. treatment [McK81]. Trench [HE88]. Trends [CW83]. tributaries [Pet81]. Trichodesmium [BB84a]. Tritiated [DH85b]. Tritium [HI88, QBHS80]. Trophic [BRH^{+}88, DC83, SA85, BC82, BH88a, BB86, CM89, FW82, HFA80, Ive88, MSBC80, OAI^{+}85, Pac86, Söd87, Sta85a, Sta85b, Swa83, Thr88, Whi89, Lam87]. tropical [Bay88, Bir88, Bri82, Bur80, BRBL83, DL85b, Fle89, GK83, Har83b, Hec84, Hec86, Le 82, Lew82b, LK89, RD89, She84, SG86, Two83, VWVR84, Woo80]. troughs [Bot85]. trout [Lue88]. true [Fra87]. tubes [Luc86]. Tucker [Tut80]. tundra [Man83, PHCK83]. tunicates [PD80]. turbid [MG80]. Turbidity [SM82b, LG88]. Turbulent [IJ84, Han84, Lem89, Csa80]. Turkana [BSB^{+}87]. turnover [Bor87, Nür88, SH84, SKW86]. twilight [SLL85]. Twin [LCPL88]. two [BMH85, Bor83, CM89, CVG89, DGN82, Edw87, GJP89, GBF86, GMPA^{+}85, HK88b, IL85, JK84, LW84, MCK89, POHMD84, PG81, QBHS80, SMSM86, SC81a, SNL84, SW89, TNEJ84, TS82, Wil80b]. two-component [MCK89]. two-flow [SW89]. two-stage [Bor83]. type [TC85, WBV80]. types [BH83, TH80, WC81, WHM^{+}85]. Typewriter [Ban84b, Csa80, Dav86, Dug80, Fen83, Hor83, Lis85, Lon85, Mac85, Mar83, MS84, Jum82, Rzó81b, Wet82]. typical [BBA84, HD84]. typicus [CS83, DG80].

U [Woo85, CB85, CB87, FTKB87]. U.K [Man80]. U.S. [YMBO87]. U.S.A. [HI88]. Uchupi [McM86]. Ulanowicz [Jum84, Ste84]. ultra [Gol88]. ultra-oligotrophic [Gol88]. ultraphytoplankton [GMFLJ85]. ultraplankton [MH85]. ultraviolet [Fle89, JY84, MS84]. unacidified [Sto87]. unattended [HK80]. uncertainties [LW84]. uncontaminated [McA86]. under-ice [KE80]. undergoing [MC82]. Underice [RI88]. underlying [SRK86]. understanding [Leh86]. Underwater [KBO84]. undulata [DH88]. UNESCO [Jum84, Ste84]. unicellular [Ban82, CW84, SH85]. unique [SC88]. United [Gil80, GDS83]. Univ [Ano86j, Epp82, Fen84, Hal82a, Hrb82, Jah83, Jum82, Lon86, Lun84, Mor80b, Pom80, Smi82a, Smi88, Woo85]. University [Bro80, Gar80, HG88, Pim83, Smi86b, Wel88, Ano82e, Ano88f, Ano89g]. Unknown [GK83]. unusual [TS83]. upland [Ban87]. upon [HM82]. Upper [BRBL83, DG83, KBB^{+}89, KM81, LCPL88, SMP^{+}89]. Uptake [PH88b, Wil84, AM82, Atk87, AS87, Bal85, BSS^{+}84, BMPD84, CD85, Cof89, Col80b, CMR89, CK84b, DC88b, DL83a, DW86b, FOB80, Gar81b, Gar84, GG84, GLMA82, GG82, HAvdL^{+}89, Har83b, HW88, JR85, KSH85, KH88a, KH86a, KW87, Law85, MWB82, MF89, NG83, Ols80a, PP85, RLG86, RM86a, RS81, RM81, STC88, SH88, VO89, WGM82, WM82, WKLK89, ZFFC88, ZKD87]. upwelled [OH87]. upwelling [And85, BMPD84, BSC80, DP89a, GZG80, HF84, HFL84, Kir89, KW87, MMP86, PA83, WD87, ZKD87, Dug80]. uranium [DSB86]. Urban [KR89]. urchin [Dug81, ER88]. urchins [Dug81]. urea [PH88b]. US/IBP [Man83]. Use [HA85, Lor80, Vos89, BK82, BS85b, DW86b, Fee79, HFA80, IL85, Kin83, LHH80, Mac86, Man88, MF86, Pat80, PF80, Rue83, SL87a, WM82, WD87, WLR88, ZBM82, Zül81, ZEY89, lWJ82]. used [BB84b, Car89, PH87, PG81, SC88]. using [Ban88, Bot85, CS88, DWR82, Doy80, Dri81, FO82, GWB89, GPS85, HFF87, HH86, HS82, OSML87, Rod87, SAJB86, SGSS84, TS85, TCN^{+}82, ZST89, ZPVM83]. Utilization [FS84, Rom84, WK86a, Bel80, BS80b, Bel83, Bur86b, DP89a, DB81, FP87, HP89, KOTeI86, MH89, MK84]. UV [BMP81, FH82]. UV-sensitive [FH82].

V [Ban84a, Ban84b, Bro82, Edm81, EA84, Fen82, Fen84, Hor83, Lon84, Lon85, Man83, Mil84b, Mur84b, Now84, Pom89, Rzó81b, Wet83, Wet84, Whi86, Csa80, Edm82a, Edm82d, Epp82, Shi85, Wet82]. vacuolation [vRS85]. Valencia [LL84a, LL85a, LW81, Lew83b, LL84b]. validation [KH86c]. validus [OBP87]. Valiela [Boo80]. Valley [WH80]. values [RJ84, SLP87]. valve [SWSC85]. valves [McB88]. Vancouver [Ano85d]. vanishing [GLMA85, Law84]. Variability [Hun83b, AM83, BSON84, Bra82, Car83, GL81, GSBP87, Gro89a, KSH85, MH82, PS84]. variable [BB89a, BF89, IP85, KMK88]. variables [FW82]. variance [DPF87, GHW84]. Variation [GC81, ER88, Hai80, Hai88, KTM83, Ker88, LL84a, Lin87, PGB88, Rin81, TML^{+}88]. Variations [KLC81, SLP87, SWSC85, ARP82, BC80, CCW87, CP83, DL85b, GPHI83, JKMN87, Lew87, ML82, MRE84, Neh87, TW83, VAI85]. various [CS83, Man88]. Vechten [SK82]. vector [KA89]. vegetated [WMSW89]. velocity [BR80, FCO87, KMK88, LKS89]. velum [CCF89]. Venezuela [HL87, LL84a, LL85a, LW81, Lew83b, LL84b, LSLW86, SL89]. venir [Fow80]. vent [Smi85, VFB87]. ventfield [BT84]. ventricosa [RE81]. verification [OØ85]. verified [Yen85]. Verlag [Bur80, Dug80, Edm82b, McM86, Pic82, Ser80, WK80]. vernal [KB89a]. versus [Nür87, Sch80, Smi84]. Vertical [AATC82, Brz88, CA87, GPHI83, GMPA^{+}85, HP87, Her83, Lew82b, MH80, Mul89, QBHS80, Sch85, SSK87, SKW86, Wei84, Alt89, BF89, CM82, DFW89, DG83, FCS84, Ger82, Her82, IP84, JJ87, LaZ80, MG83, OP83, PHI84, SLL85, SG86, Sul80, Unc81, Ven84, WC87, vRS85, CCW87]. vesicle [OW84]. vessels [HT85b]. VI [DSB86, Rzó81a, Tut80]. via [DMM82, IB82a, WJ89]. viable [McA86, TD83]. vicinity [HFFS85]. view [MB89b, Wel88]. viii [Man80]. Villefranche [Iba81]. Vincent [Pae89]. Vineyard [Fei82].

Virginia [Pom80, Duc82]. viscosities [CMB87]. visibility [Zar81]. visible [BMP81]. Visual [Pre86, RS86a, YTW84]. Vitamin [PHCT89]. vivo [FN85, HM87, MH87, SG84]. Volatile [GZG80, Sau80, NG87]. Voltammetric [RNVB81]. Volume [Ano80c, Ano81d, Ano82c, Ano83e, Ano84f, Ano85c, Ano86i, Ano87f, Ano88g, Ano89h, PS89, RS88]. volumes [Ban82]. Volumetric [OH87]. Volvox [SG86]. Vortex [SC88]. vs [TLH89]. vulnerability [SRW81].

W [Ano86j, Ban84b, Bro82, Bur80, Cam87, Cat85, Dav86, Edm81, Edm82a, Edm82c, Edm82d, Fry80, Gar80, Gil80, Har89a, HG88, Hec84, Hec86, Hrb82, Hyn81, Jum82, Mac85, Mac80, Mil84b, Mor88a, Mur84b, New86, Pae89, Pai84b, Pom80, Shi85, Wet83, Wet84, Woo80]. W. [Ano88h, Ano88j, Hut88]. Wadden [Ban84b]. Wahl [Mol89]. Wales [Ban87]. Ward [Rzó81b]. Warm [Woo85, CRGC87, DH85a, DH85b, GBF86, Han87, HLS87, MN86]. warm-core [CRGC87, Han87, HLS87, MN86]. Warren [Pon82]. Washington [CR83, Lis85, Abe88, BBS80, DP89a, EL81, EL82, HTE84, IL85, IA85, JM84, JM85, KM84, KMD^{+}88, LE83, Leh88c, MTW^{+}88, Mur81b, QEQD86, RDE82, Run85]. Water [HH81, Kir85, McL89, PS84, And85, AB80, BZH^{+}85, BMH^{+}87, BH88a, Ber89, BB86, BRBL83, CBH83, Car84, CWD^{+}86, CB86, Chr89, Chu85, DC88a, Fle89, FP81, GCLC87, Gel86, Gor89a, GM84, HFF87, HE88, HHS80, HCC88b, HH85, Hob88, Høj89, HT85b, HS88, IBC^{+}83, Jah88, JW84, KB89a, KE80, Kir84, KBO84, KM89, Kon89, Kor80, LFRT86, Lew83a, LBB88, MMD^{+}80, Mor88a, Mul81, Nag88, NSL86, Óla80, Ore83b, Pai86, Pse88, RGS89, RNVB81, RS86c, SP89b, SSGB89, SG84, TW83, THCP81, TB84, VL89, WS81, Win81, WLR88, vdBD84, vdLAH^{+}84, BW86, Gil80]. water-quality [Høj89]. water-table [Win81]. waterborne [WBS83]. waters [ACC86, BS82, Ban87, BH83, Brz88, Bur86b, CP86, CS81, CL81, CM86, CSAO86, Cof89, DM85, Dav86, DH85a, EHE^{+}83, Fen84, FNL85, FVM82, FS83, FP82, Gor89b, Gri88, HHOC89, HK86, HL82, INA89, KSH88, KS82a, KS82b, KH86a, Kir89, KJP89, Lan82, LRH^{+}84, LL81, MV88, Mil84a, Mil86, Mob89, Neh87, NSL86, OH87, Ols80a, PCP87, Pae88, Pre86, PS81, PP85, PS89, RDW^{+}88, RPC89, San84, San88, Sau80, SNL84, SC81b, SPBB89, SS80a, SS80b, SB83, TML^{+}88, TB80, VL87, VN87, WLB84, Wei80, Wof83, dAL87, Bur86a, Den82a, Mur83, Pic85, Whi86, Bur88, Hen87]. Watershed [Pse88, Cor81, EH88, RGS89]. waterways [BB81]. Watson [RF89]. Wave [Csa87, Den85, Csa80]. waveguides [Cha85]. waves [Den82b, Lem87, WL83, WC87]. Wax [CVG89]. weak [Som89a]. web [Fry88, Leh88c]. webs [BLKH88, SS88a, TTF88, WKLK89]. weight [RM86b, SW80]. weinecki [Fre88]. weissflogii [AM82, HM86]. well [WLB84]. Wellington [Ano86j]. West [Kli88, LHM84, Båm85, CS85, ER88, VCG^{+}87]. Westemport [KM87]. Western [SA83, BS89, BBS80, KSH88, KCB89, SSS87]. wet [WLG85]. wetland [BMH85, CMK89, TM84, Wet82, WLG85]. wetlands [HB81, LBB88, RPL89]. Wetzel [Cat85]. WH7803 [CW84, KGGW88]. while [SMR84]. Whitfield [Dav83]. Whole [LCK89, BMH^{+}87, BKF^{+}87, Fee79, Pat80]. whole-core [BMH^{+}87]. Whole-lake [LCK89, BKF^{+}87, Fee79, Pat80]. Wiley [Cam87, Cha85, Dav83, Gil80, Ill80, Lec85, Lon84, Mil82, Mor88a, Pic85, Pla86, Pra80, Whi86, ZoB80]. Wiley-Interscience [Whi86]. Williams [Duc87, Woo80]. Wind [DTBB87, HMS86, Mon85, Sta82, IP85, Mob89]. Wind-forced [Mon85]. Wind-induced [HMS86]. wind-roughened [Mob89]. Windermere [HTWR84, RMB82]. window [Jan81]. Windpower [Sta80]. Winkler [Ami88, OGMG88, lWJ82]. Winners [Ano88n]. Winter [Ven84, PH88a, Smi81a, SM82b, WRS80]. Wisconsin [Ano82h, Cap82, FB80, HH80, LKS89, Sta85a, Sta85b]. within [CB86, Chu85, CB85, JKMN87, Lam87, Sel81]. Wolderwijd [ZBM82]. Wolff [Ban84b]. woodland [CSW^{+}80]. Woods [Pic82]. Workshop [Wet83]. world [Koh87, Pom89]. wrightii [MK84]. Wunsch [Pon82].

x [Bro80, Dug80, Fry80, Kru80, Ser80, MD87]. xi [Hyn81, Rzó81b, ZoB80]. xii [Gar80, Pom80]. xiii [Csa80, Fen80b, Mac80]. xiv [Den82a, Fow80, Mor80b, Pra80, WK80]. xv [Gil80]. xvii [Fen80a]. xviii [Woo80].

Yale [Hal82a]. year [ZEY89]. yellow [BMP81, DCV87, MMTL88]. yield [KR88, KOTeI86, LUP88, MLG87]. York [Bro80, Bur80, Cam87, Cha85, Csa80, DG80, Den82a, Dug80, Edm84b, Epp82, Fen80b, Fen82, Fre84, Fry86, Gar81b, Gil80, HN81, Hor83, Ill80, Koh82, Kru80, Lon84, Lon85, Lun84, Mar83, McM86, MS84, Mor80b, New86, Now84, Pic82, Pra80, Ril82, Rzó81b, Ser80, Shi85, Smi82a, Smi86b, Smi88, TCN^{+}82, Wan82, WK80, Wet82, Whi86, Woo85, ZoB80, Ano80e, CD85, Chu85, Duc82, THCP81]. young [SVG89].

Zaret [Hal82a]. Zealand [PSGD86]. Zeaxanthin [KGGW88, GMFLJ85]. zeaxanthin-dominant [GMFLJ85]. Zehnder [Ben89, Bez89]. zinc [BSG83, Bru89a, KDC86, RM81, vdBD84]. Zn [RNVB81]. Zone [Mor88b, AN81, BB84b, Chu85, CB87, Cor81, DTBB87, GPS85, HK81a, Lem89, VHJ^{+}89]. Zoogeographic [SCR84]. Zoogeography [Mil82]. zooplankter [Hal82b]. zooplankters [GS81]. Zooplanktivorous [BF89, TM84, WL85b]. Zooplankton [CRGC87, Edm82b, EEMC88, Fie82, GL81, HK81b, Le 82, SL89, Smi82d, WL85a, BS80a, BCL85, BS86, BBP82, BB89b, CVG89, CDRW86, CPT86, Dod88, DPF87, DL85a, DdNK82, DTS^{+}89, FP87, GSBP87, Gil88, GWB89, Håk87, Hrb82, IH81, Jan81, KN81, LFRT86, LCK89, Leh88a, LO83, Lyn83, MTL84, MRE84, NL89, OØ85, OJR^{+}86, OP83, PO81, Pac86, PD84, PG81, RS89, Ree81, TS81a, Tay88, Thr88, TM84, Two83, Van86, Whi81, Bur80]. Zooplankton-mediated [EEMC88]. zooxanthellae [DC88b, MMM81, TD83]. zooxanthellae-invertebrate [DC88b]. Zostera [Bor87]. Zulstra [Pom89].

Zurich [HMS86, SSK87, Sch85].

References

Anderson:1982:VHD

[AATC82] Donald M. Anderson, David G. Aubrey, Mary A. Tyler, and D. Wayne Coats. Vertical and horizontal distributions of dinoflagellate cysts in sediments. Limnology and Oceanography, 26(5): 757–765, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Atlas:1980:DMS

[AB80] David Atlas and T. T. Bannister. Dependence of mean spectral extinction coefficient of phytoplankton on depth, water color, and species. Limnology and Oceanography, 25(1):157–159, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Abella:1988:EMM

[Abe88] Sally E. B. Abella. The effect of the Mt. Mazama ashfall on the planktonic diatom community of Lake Washington. Limnology and Oceanography, 33(6):1376–1385, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Anderson:1982:FDS

[AC82] S. Miles Anderson and A. C. Charters. A fluid dynamics study of seawater flow through Gelidium nudifrons. Limnology and Oceanography, 27(3):399–412, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Alldredge:1986:PHB

[ACC86] Alice L. Alldredge, Jonathan J. Cole, and David A. Caron. Production of heterotrophic bacteria inhabiting macroscopic organic aggregates (marine snow) from surface waters. Limnology and Oceanography, 31(1):68–78, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Auclair:1982:HFE

[ADF^{+}82] J. C. Auclair, S. Demers, M. Fréchette, L. Legendre, and C. L. Trump. High frequency endogenous periodicities of chlorophyll synthesis in estuarine phytoplankton. Limnology and Oceanography, 27(2):348–352, March 1982. CODEN LIOCAH. ISSN 0024-

Agusti:1987:ACS

[ADK87] Susana Agusti, Carlos M. Duarte, and Jacob Kalff. Algal cell size and the maximum density and biomass of phytoplankton. Limnology and Oceanography, 32(4):983–986, July 1987. CODEN LI-OCAH. ISSN 0024-3590.

Appleby:1985:MSE

[ADO85] P. G. Appleby, J. A. Dearing, and F. Oldfield. Magnetic studies of erosion in a Scottish lake catchment. 1. Core chronology and correlation. Limnology and Oceanography, 30(6):1144–1153, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Abbott:1984:MDD

[ADP^{+}84] Mark R. Abbott, Kenneth L. Denman, Thomas M. Powell, Peter J. Richerson, Robert C. Richards, and Charles R. Goldman. Mixing and the dynamics of the deep chlorophyll maximum in Lake Tahoe. Limnology and Oceanography, 29(4):862–878, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Andersen:1985:BMF

[AF85] P. Andersen and T. Fenchel. Bacterivory by microheterotrophic flagellates in seawater samples. Limnology and Oceanography, 30 (1):198–202, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Allan:1986:DEA

[AFM86] J. D. Allan, A. S. Flecker, and N. L. McClintock. Diel epibenthic activity of mayfly nymphs, and its nonconcordance with behavioral drift. Limnology and Oceanography, 31(5):1057–1065, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Allan:1987:PSS

[AFM87] J. David Allan, Alexander S. Flecker, and Nancy L. McClintock. Prey size selection by carnivorous stoneflies. Limnology and Oceanography, 32(4):864–872, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Alldredge:1988:SSB

[AG88] Alice L. Alldredge and Chris Gotschalk. In situ settling behavior of marine snow. Limnology and Oceanography, 33(3):339–351, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Axler:1982:INA

[AGG82] Richard P. Axler, Richard M. Gersberg, and Charles R. Goldman. Inorganic nitrogen assimilation in a subalpine lake. Limnology and Oceanography, 27(1):53–65, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Aksnes:1987:OLT

[AH87] Dag L. Aksnes and Tore J. Høisceter. Obtaining life table data from stage-frequency distributional statistics. Limnology and Oceanography, 32(2):514–517, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1986:BRM

[AHI^{+}86] Leif G. Anderson, Per O. J. Hall, Ake Iverfeldt, Michiel M. Rutgers van der Loejf, Bjørn Sundby, and Stig F. G. Westerlund. Benthic respiration measured by total carbonate production. Limnology and Oceanography, 31(2):319–329, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Amaral:1989:LTS

[AHRF89] J. A. Amaral, R. H. Hesslein, J. W. M. Rudd, and D. E. Fox. Loss of total sulfur and changes in sulfur isotopic ratios due to drying of lacustrine sediments. Limnology and Oceanography, 34(7):1351– 1358, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Agusti:1989:IGC

[AK89] Susana Agusti and Jacob Kalff. The influence of growth conditions on the size dependence of maximal algal density and biomass. Limnology and Oceanography, 34(6):1104–1108, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Alldredge:1981:IAG

[All81] Alice L. Alldredge. The impact of appendicularian grazing on natural food concentrations in situ. Limnology and Oceanography, 26(2):247–257, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1985:SCD

[ALRP85] Donald M. Anderson, John J. Lively, Ellen M. Reardon, and Carl A. Price. Sinking characteristics of dinoflagellate cysts. Limnology and Oceanography, 30(5):1000–1009, September 1985. CO-DEN LIOCAH. ISSN 0024-3590.

Altabet:1989:TSS

[Alt89] Mark A. Altabet. A time-series study of the vertical structure of nitrogen and particle dynamics in the Sargasso Sea. Limnology and Oceanography, 34(7):1185–1201, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1982:IAI

[AM82] Michael A. Anderson and Francois M. M. Morel. The influence of aqueous iron chemistry on the uptake of iron by the coastal diatom Thalassiosira weissflogii. Limnology and Oceanography, 27(5):789–813, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Andrews:1983:STV

[AM83] John C. Andrews and Hans Müller. Space-time variability of nutrients in a lagoonal patch reef. Limnology and Oceanography, 28 (2):215–227, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Aminot:1988:PAD

[Ami88] Alain Aminot. Precision and accuracy of dissolved oxygen measurements. A comment on the paper by Oudot et al. 1988: Precise shipboard determination of dissolved oxygen (Winkler procedure) for productivity studies with a commercial system. Limnology and Oceanography, 33(6 (part 2)):1646–1648, November 1988. CODEN LIOCAH. ISSN 0024-3590. See [OGMG88] and reply [Oud89].

Arts:1981:IAD

[AMP81] Michael T. Arts, Edward J. Maly, and Maria Pasitschniak. The influence of Acilius (Dytiscidae) predation on Daphnia in a small pond. Limnology and Oceanography, 26(6):1172–1175, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Alexander:1981:OEB

[AN81] V. Alexander and H. J. Niebauer. Oceanography of the eastern Bering Sea ice-edge zone in spring. Limnology and Oceanography, 26(6):1111–1125, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Andreae:1980:DMF

[And80] Meinart O. Andreae. Dimethylsulfoxide in marine and freshwaters. Limnology and Oceanography, 25(6):1054–1063, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Andreae:1985:DWC

[And85] Meinrat O. Andreae. Dimethylsulfide in the water column and the sediment porewaters of the Peru upwelling area. Limnology and Oceanography, 30(6):1208–1218, November 1985. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1980:ASL

[Ano80a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement 1 January–31 December 1978. Limnology and Oceanography, 25(1):195–196, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1980:A

[Ano80b] Anonymous. Announcements. Limnology and Oceanography, 49 (6):582, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1980:ATL

[Ano80c] Anonymous. Authors and titles in Limnology and Oceanography volume 25, 1980. Limnology and Oceanography, 25(6):1154–1159, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1980:BR

[Ano80d] Anonymous. Books received. Limnology and Oceanography, 25(2): 387–388, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1980:PAS

[Ano80e] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. 42nd Annual Meeting 20 June 1979 Stony Brook, New York. Limnology and Oceanography, 25(1):193–194, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1981:ASLa

[Ano81a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 26(1):198– 199, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1981:ASLb

[Ano81b] Anonymous. American Society of Limnology and Oceanography, Inc. publication announcement. Limnology and Oceanography, 26 (2):399, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1981:ASLc

[Ano81c] Anonymous. The American Society of Limnology and Oceanography, Inc. sustaining members 1981. Limnology and Oceanography, 26(6):1183, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1981:ATL

[Ano81d] Anonymous. Authors and titles in Limnology and Oceanography volume 26, 1981. Limnology and Oceanography, 26(6):1184–1188, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1981:OSA

[Ano81e] Anonymous. Ocean sciences: AGU/ASLO Joint Meeting San Antonio, Texas 16–19 February 1982. Limnology and Oceanography, 27(4):999, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1981:PAS

[Ano81f] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. 43rd Annual Meeting 17 June 1980 Knoxville, Tennessee. Limnology and Oceanography, 26(3):598– 599, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:ASLa

[Ano82a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement 1 January–31 December 1980. Limnology and Oceanography, 27(1):193–194, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:ASLb

[Ano82b] Anonymous. The American Society of Limnology and Oceanography, Inc. sustaining members 1982. Limnology and Oceanography, 27(6):1176, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:ATL

[Ano82c] Anonymous. Authors and titles in Limnology and Oceanography volume 27, 1982. Limnology and Oceanography, 27(6):1177, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:BR

[Ano82d] Anonymous. Books received. Limnology and Oceanography, 27(1): 199–200, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:DUP

[Ano82e] Anonymous. Dalhousie University publication announcement. Limnology and Oceanography, 27(6):1112, November 1982. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:E

[Ano82f] Anonymous. Errata. Limnology and Oceanography, 27(4):698, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:GEH

[Ano82g] Anonymous. G. Evelyn Hutchinson Medal announcement and call for nominations. Limnology and Oceanography, 27(6):1175, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:PAS

[Ano82h] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. 44th Annual Meeting 16 June 1981 Milwaukee, Wisconsin. Limnology and Oceanography, 27(1):191–192, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1982:PA

[Ano82i] Anonymous. Publication announcement. Limnology and Oceanography, 27(3):593–594, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1983:ASLa

[Ano83a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 28(1):197, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1983:ASLb

[Ano83b] Anonymous. The American Society of Limnology and Oceanography, Inc. sustaining members 1983. Limnology and Oceanography, 28(5):1046–1047, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1983:Aa

[Ano83c] Anonymous. Announcements. Limnology and Oceanography, 28 (4):800, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1983:Ab

[Ano83d] Anonymous. Announcements. Limnology and Oceanography, 28 (5):1040, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1983:ALO

[Ano83e] Anonymous. Authors in Limnology and Oceanography volume 28, 1983. Limnology and Oceanography, 28(6):1281–1283, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1983:PAS

[Ano83f] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. 45th Annual Meeting 15 June 1982 Raleigh, North Carolina. Limnology and Oceanography, 28(1): 194–196, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:ASLa

[Ano84a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 29(1):219– 220, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:ASLb

[Ano84b] Anonymous. The American Society of Limnology and Oceanography, Inc. sustaining member 1984. Limnology and Oceanography, 29(6):1342–1346, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:Aa

[Ano84c] Anonymous. Announcement. Limnology and Oceanography, 29 (4):902, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:Ab

[Ano84d] Anonymous. Announcement. Limnology and Oceanography, 29 (5):1148, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:AGE

[Ano84e] Anonymous. Announcement: G. Evelyn Hutchinson Medal. Limnology and Oceanography, 29(6):1354, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:ALO

[Ano84f] Anonymous. Authors in Limnology and Oceanography volume 29, 1984. Limnology and Oceanography, 29(6):1355–1358, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:BR

[Ano84g] Anonymous. Books received. Limnology and Oceanography, 29(6): 1352–1354, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1984:PAS

[Ano84h] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. 46th Annual Meeting 15 June 1983 St.

John’s, Newfoundland. Limnology and Oceanography, 29(1):217– 218, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1985:ASL

[Ano85a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement 1 January–31 December 1983. Limnology and Oceanography, 30(1):233–234, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1985:A

[Ano85b] Anonymous. Announcements. Limnology and Oceanography, 30 (4):918, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1985:ALO

[Ano85c] Anonymous. Authors in Limnology and Oceanography volume 30, 1985. Limnology and Oceanography, 30(6):1351–1354, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1985:BRP

[Ano85d] Anonymous. Book review: Proceedings of the American Society of Limnology and Oceanography, Inc. 47th Annual Meeting 12 June 1984 Vancouver, British Columbia. Limnology and Oceanography, 30(1):231–232, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1985:CNGa

[Ano85e] Anonymous. Call for nominations: G. Evelyn Hutchinson Medal. Limnology and Oceanography, 30(5):1130, September 1985. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1985:CNGb

[Ano85f] Anonymous. Call for nominations: G. Evelyn Hutchinson Medal. Limnology and Oceanography, 30(6):1297, November 1985. CO-DEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:ASLa

[Ano86a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 31(1):229– 230, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:ASLb

[Ano86b] Anonymous. The American Society of Limnology and Oceanography, Inc. sustaining member 1986. Limnology and Oceanography, 31(5):1167, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:Aa

[Ano86c] Anonymous. Announcement. Limnology and Oceanography, 31 (1):228, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:Ab

[Ano86d] Anonymous. Announcement. Limnology and Oceanography, 31 (2):413, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:Ac

[Ano86e] Anonymous. Announcement. Limnology and Oceanography, 31 (3):672, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:Ad

[Ano86f] Anonymous. Announcement. Limnology and Oceanography, 31 (4):906, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:Ae

[Ano86g] Anonymous. Announcement. Limnology and Oceanography, 31 (6):1383, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:AGE

[Ano86h] Anonymous. Announcement: G. Evelyn Hutchinson Medal. Limnology and Oceanography, 31(5):1142, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:ALO

[Ano86i] Anonymous. Authors in Limnology and Oceanography volume 31, 1986. Limnology and Oceanography, 31(6):1401–1402, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:BRG

[Ano86j] Anonymous. Book review: Glynn, P. W., and G. M. Wellington. 1984. Corals and coral reefs of the Galapagos Islands. Univ. California Press, Berkeley. 332 p. $45.00. Limnology and Oceanography, 31(1):235, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:N

[Ano86k] Anonymous. Notice. Limnology and Oceanography, 31(3):627–636, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1986:PAS

[Ano86l] Anonymous. Proceedings of the American Society of Limnology and Oceanography, Inc. Limnology and Oceanography, 31(1):231– 232, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:ASL

[Ano87a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 32(1):273– 274, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:Ab

[Ano87b] Anonymous. Announcement. Limnology and Oceanography, 32 (4):1010, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:Ac

[Ano87c] Anonymous. Announcement. Limnology and Oceanography, 32 (4):1014, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:Ad

[Ano87d] Anonymous. Announcement. Limnology and Oceanography, 32 (5):1180, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:Aa

[Ano87e] Anonymous. Announcements. Limnology and Oceanography, 32 (3):780, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:AV

[Ano87f] Anonymous. Authors in volume 32, 1987. Limnology and Oceanography, 32(6):1384–1387, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:BR

[Ano87g] Anonymous. Books received. Limnology and Oceanography, 32(6): 1382–1383, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1987:PAS

[Ano87h] Anonymous. Proceedings of the American Society of Limnology Oceanography, Inc. Limnology and Oceanography, 32(1):271–272, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:ASL

[Ano88a] Anonymous. American Society of Limnology and Oceanography, Inc. financial statement. Limnology and Oceanography, 33(1):155– 156, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:Aa

[Ano88b] Anonymous. Announcement. Limnology and Oceanography, 33 (4):648, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:Ab

[Ano88c] Anonymous. Announcement. Limnology and Oceanography, 33 (5):1228, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:AGE

[Ano88d] Anonymous. Announcement: G. Evelyn Hutchinson Medal. Limnology and Oceanography, 33(2):312, March 1988. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1988:AST

[Ano88e] Anonymous. ASLO-’89 student travel awards. Limnology and Oceanography, 33(6 (part 2)):1652, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1988:AUA

[Ano88f] Anonymous. ASLO-’89 University of Alaska, fairbanks. Limnology and Oceanography, 33(6 (part 2)):1652, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:ALO

[Ano88g] Anonymous. Authors in Limnology and Oceanography, volume 33, 1988. Limnology and Oceanography, 33(6 (part 2)):1656, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:BWE

[Ano88h] Anonymous. Bibliography of W. T. Edmondson. Limnology and Oceanography, 33(6):1241–1243, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1988:CP

[Ano88i] Anonymous. Call for papers. Limnology and Oceanography, 33(6 (part 2)):1655, November 1988. CODEN LIOCAH. ISSN 0024-

Anonymous:1988:DTC

[Ano88j] Anonymous. Dissertations and theses completed under the direction of W. T. Edmondson. Limnology and Oceanography, 33(6): 1243–1244, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:FMF

[Ano88k] Anonymous. A fast multichannel fluorometer for investigating aquatic chemoreception and odor trails. Limnology and Oceanography, 33(6 (part 2)):1586–1595, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1988:PA

[Ano88l] Anonymous. Publication announcement. Limnology and Oceanography, 33(2):310, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:SAA

[Ano88m] Anonymous. Student awards for the 1968 annual meeting. Limnology and Oceanography, 33(6 (part 2)):1654, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1988:STA

[Ano88n] Anonymous. Student travel award winners. Limnology and Oceanography, 33(3):487, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:ASLc

[Ano89a] Anonymous. American society of limnology and oceanography executive secretary. Limnology and Oceanography, 34(7):1366, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:ASLa

[Ano89b] Anonymous. American Society of Limnology and Oceanography. Inc. financial statement. Limnology and Oceanography, 34(1):265– 266, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:ASLb

[Ano89c] Anonymous. American society of limnology and oceanography, sustaining member. Limnology and Oceanography, 34(6):1148, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:Aa

[Ano89d] Anonymous. Announcement. Limnology and Oceanography, 34 (4):789, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:Ac

[Ano89e] Anonymous. Announcement. Limnology and Oceanography, 34 (8):1766, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:Ab

[Ano89f] Anonymous. Announcements. Limnology and Oceanography, 34 (5):971, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:AUA

[Ano89g] Anonymous. ASLO-’89 University of Alaska, fairbanks. Limnology and Oceanography, 34(1):264, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:ALO

[Ano89h] Anonymous. Authors in Limnology and Oceanography volume 34, 1989. Limnology and Oceanography, 34(8):1762–1765, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:BRA

[Ano89i] Anonymous. Book review: Autotrophic bacteria (H. G. Schlegel and B. Bowien). Limnology and Oceanography, 34(6):1150–1152, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:GEH

[Ano89j] Anonymous. G. Evelyn Hutchinson Medal. Limnology and Oceanography, 34(7):1385–1388, November 1989. CODEN LI-OCAH. ISSN 0024-3590.

Anonymous:1989:MPK

[Ano89k] Anonymous. In memoriam — Peter Kilham. Limnology and Oceanography, 34(5):967, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:PSO

[Ano89l] Anonymous. Prepublication special offer. Limnology and Oceanography, 34(7):1384, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:PAS

[Ano89m] Anonymous. Proceedings of the American society of limnology and oceanography. Limnology and Oceanography, 34(1):267, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Anonymous:1989:SAA

[Ano89n] Anonymous. Student awards for the 1989 annual meeting. Limnology and Oceanography, 34(7):1388, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Abbott:1980:ETD

[APR80] Mark R. Abbott, Thomas M. Powell, and Peter J. Richerson. The effects of transect direction on observed spatial patterns of chlorophyll in Lake Tahoe. Limnology and Oceanography, 49(6):534–537, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1988:DMM

[AR88] N. J. Anderson and B. Rippey. Diagenesis of magnetic minerals in the recent sediments of a eutrophic lake. Limnology and Oceanography, 33(6 (part 2)):1476–1492, November 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Abbott:1982:SRP

[ARP82] Mark R. Abbott, Peter J. Richerson, and Thomas M. Powell. In situ response of phytoplankton fluorescence to rapid variations in light. Limnology and Oceanography, 27(2):218–225, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:1983:CPR

[AS83] M. J. Atkinson and S. V. Smith. C : N : P ratios of benthic marine plants. Limnology and Oceanography, 28(3):568–574, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:1987:SUB

[AS87] M. J. Atkinson and D. F. Smith. Slow uptake of ^{32}P over a barrier reef flat. Limnology and Oceanography, 32(2):436–441, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:1981:CEA

[ASS81] M. Atkinson, S. V. Smith, and E. D. Stroup. Circulation in Enewetak Atoll lagoon. Limnology and Oceanography, 26(6):1074–1083, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Anderson:1987:EDA

[ATA87] Donald M. Anderson, Craig D. Taylor, and E. Virginia Armbrust. The effects of darkness and anaerobiosis on dinoflagellate cyst germination. Limnology and Oceanography, 32(2):340–351, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:1987:RPU

[Atk87] M. J. Atkinson. Rates of phosphate uptake by coral reef flat communities. Limnology and Oceanography, 32(2):426–435, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Atkinson:1988:FRO

[Atk88] M. J. Atkinson. Fast-response oxygen sensor for a free-fall CTD. Limnology and Oceanography, 33(1):141–145, January 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Avnimelech:1983:PCC

[Avn83] Yoram Avnimelech. Phosphorus and calcium carbonate solubilities in Lake Kinneret. Limnology and Oceanography, 28(4):640– 645, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Bell:1987:TIM

[BA87] Russell T. Bell and Ingemar Ahlgren. Thymidine incorporation and microbial respiration in the surface sediment of a hypereutrophic lake. Limnology and Oceanography, 32(2):476–482, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Balzer:1984:OMD

[Bal84] Wolfgang Balzer. Organic matter degradation and biogenic element cycling in a nearshore sediment (Kiel Bight). Limnology and Oceanography, 29(6):1231–1246, November 1984. CODEN LI-OCAH. ISSN 0024-3590.

Balch:1985:LEL

[Bal85] William M. Balch. Lack of an effect of light on methylamine uptake by phytoplankton. Limnology and Oceanography, 30(3):665–674, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Baamstedt:1985:SER

[Båm85] Ulf Båmstedt. Seasonal excretion rates of macrozooplankton from the Swedish west coast. Limnology and Oceanography, 30(3):607– 617, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Banse:1982:CVM

[Ban82] Karl Banse. Cell volumes, maximal growth rates of unicellular algae and ciliates, and the role of ciliates in the marine pelagial. Limnology and Oceanography, 27(6):1059–1071, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Banse:1984:BRB

[Ban84a] Karl Banse. Book review: Bougis, P. [Ed.] 1982. Marine pelagic Protozoa and microzooplankton ecology. Ann. Inst. Oceanogr., Paris, V. 58 (Suppl.). Masson, Paris. 352 p. Limnology and Oceanography, 29(2):445–446, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Banse:1984:BRW

[Ban84b] Karl Banse. Book review: Wolff, W. J. [Ed.] 1983. Ecology of the Wadden Sea. A. A. Balkema, Rotterdam. 3 V. + appendix maps. [Typewriter composition] 2,077 p. $85.00. Limnology and Oceanography, 29(6):1348–1349, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Banens:1987:GCU

[Ban87] R. J. Banens. The geochemical character of upland waters of northeast New South Wales. Limnology and Oceanography, 32(6): 1291–1306, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Bannister:1988:EAC

[Ban88] T. T. Bannister. Estimation of absorption coefficients of scattering suspensions using opal glass. Limnology and Oceanography, 33(4): 607–615, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Bayley:1988:AEW

[Bay88] Peter B. Bayley. Accounting for effort when comparing tropical fisheries in lakes, river-floodplains, and lagoons. Limnology and Oceanography, 33(4 (part 2)):963–972, July 1988. CODEN LI-OCAH. ISSN 0024-3590.

Boto:1981:DOP

[BB81] Kevin G. Boto and John S. Bunt. Dissolved oxygen and pH relationships in northern Australian mangrove waterways. Limnology and Oceanography, 26(6):1176–1178, November 1981. CODEN LI-OCAH. ISSN 0024-3590.

Borstad:1984:NPO

[BB84a] Gary A. Borstad and Lorraine E. Borstad. Needlelike projectiles observed issuing from lysing Trichodesmium (Cyanophyta) cells. Limnology and Oceanography, 29(5):1097–1099, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Bottomley:1984:SPS

[BB84b] E. Z. Bottomley and I. L. Bayly. A sediment porewater sampler used in root zone studies of the submerged macrophyte, Myriophyllum spicatum. Limnology and Oceanography, 29(3):671–673, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Binford:1986:DOS

[BB86] Michael W. Binford and Mark Brenner. Dilution of ^{210}Pb by organic sedimentation in lakes of different trophic states, and application to studies of sediment-water interactions. Limnology and Oceanography, 31(3):584–595, May 1986. CODEN LIOCAH. ISSN 0024-3590. See comment [BH88a].

Binford:1988:RCB

[BB88] Michael W. Binford and Mark Brenner. Reply to comment by Benoit and Hemond. Limnology and Oceanography, 33(2):304–310, March 1988. CODEN LIOCAH. ISSN 0024-3590. See [BH88a].

Bennett:1989:ETE

[BB89a] William N. Bennett and Martin E. Boraas. An experimental test of the egg-ratio method with instantaneous birth rate as an independent variable. Limnology and Oceanography, 34(6):1120–1125, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Bloesch:1989:CPZ

[BB89b] Jürg Bloesch and Hans-Rudolf Bürgi. Changes in phytoplankton and zooplankton biomass and composition reflected by sedimentation. Limnology and Oceanography, 34(6):1048–1061, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Bowen:1984:MDD

[BBA84] Stephen H. Bowen, Argentino A. Bonetto, and Molly O. Ahlgren. Microorganisms and detritus in the diet of a typical neotropical riverine detritivore, Prochilodus platensis (Pisces: Prochilodontidae). Limnology and Oceanography, 29(5):1120–1122, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Bloem:1989:BAP

[BBG89] Jaap Bloem and Marie-José B. Bär-Gilissen. Bacterial activity and protozoan grazing potential in a stratified lake. Limnology and Oceanography, 34(2):297–309, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Baird:1985:APM

[BBH85] Fay Baird, Donald C. Buso, and James W. Hornbeck. Access pipes for multiple sampling under ice. Limnology and Oceanography, 30 (5):1129–1130, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Blazka:1982:OCA

[BBP82] Pavel Blažka, Zdeněk Brandi, and Lidmila Procházková. Oxygen consumption and ammonia and phosphate excretion in pond zooplankton. Limnology and Oceanography, 27(2):294–303, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Birch:1980:RSR

[BBS80] Peter B. Birch, Robert S. Barnes, and Dimitris E. Spyridakis. Recent sedimentation and its relationship with primary productivity in four western Washington lakes. Limnology and Oceanography, 25(2):240–247, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Barnes:1980:DSV

[BC80] D. J. Barnes and C. J. Crossland. Diurnal and seasonal variations in the growth of a staghorn coral measured by time-lapse photography. Limnology and Oceanography, 25(6):1113–1117, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Beaver:1982:TRC

[BC82] John R. Beaver and Thomas L. Crisman. The trophic response of ciliated protozoans in freshwater lakes. Limnology and Oceanography, 27(2):246–253, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Bohle-Carbonell:1986:CLG

[BC86] M. Bohle-Carbonell. Currents in Lake Geneva. Limnology and Oceanography, 31(6):1255–1266, November 1986. CODEN LI-OCAH. ISSN 0024-3590.

Bergquist:1985:SPS

[BCL85] Ann M. Bergquist, Stephen R. Carpenter, and John C. Latino. Shifts in phytoplankton size structure and community composition during grazing by contrasting zooplankton assemblages. Limnology and Oceanography, 30(5):1037–1045, September 1985. CO-DEN LIOCAH. ISSN 0024-3590.

Bethoux:1986:BFA

[BCM86] J. Pierre Béthoux and Gérard Copin-Montégut. Biological fixation of atmospheric nitrogen in the Mediterranean Sea. Limnology and Oceanography, 31(6):1353–1358, November 1986. CODEN LI-OCAH. ISSN 0024-3590.

Baier:1989:FCO

[BCY89] Rodger W. Baier, Terry L. Cucci, and Clarice M. Yentsch. Flow cytometry for observing the distribution of adsorption sites on particles. Limnology and Oceanography, 34(5):947–952, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Brush:1981:SDP

[BD81] Grace S. Brush and Ruth S. DeFries. Spatial distributions of pollen in surface sediments of the Potomac estuary. Limnology and Oceanography, 26(2):295–309, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Bell:1980:BUAa

[Bel80] Wayne H. Bell. Bacterial utilization of algal extracellular products. 1. The kinetic approach. Limnology and Oceanography, 25(6): 1007–1020, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Bell:1983:BUA

[Bel83] Wayne H. Bell. Bacterial utilization of algal extracellular products. 3. The specificity of algal-bacterial interaction. Limnology and Oceanography, 28(6):1131–1143, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Benner:1989:BRB

[Ben89] Ronald Benner. Book review: Biology of anaerobic microorganisms (J. B. A. Zehnder [ed.]). Limnology and Oceanography, 34 (3):647, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Bernhard:1989:DBF

[Ber89] Joan M. Bernhard. The distribution of benthic Foraminifera with respect to oxygen concentration and organic carbon levels in shallow-water Antarctic sediments. Limnology and Oceanography, 34(6):1131–1141, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Bezdek:1989:BRB

[Bez89] Kristin T. Bezdek. Book review: Biology of anaerobic microorganisms (J. B. A. Zehnder [ed.]). Limnology and Oceanography, 34(3):648, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Battarbee:1984:ICD

[BF84] R. W. Battarbee and R. J. Flower. The inwash of catchment diatoms as a source of error in the sediment-based reconstruction of pH in an acid lake. Limnology and Oceanography, 29(6):1325– 1329, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Baker:1987:RBP

[BF87] Karen S. Baker and Robert Frouin. Relation between photosynthetically available radiation and total insolation at the ocean surface under clear skies. Limnology and Oceanography, 32(6):1370– 1377, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Bollens:1989:ZFV

[BF89] Stephen M. Bollens and Bruce W. Frost. Zooplanktivorous fish and variable diel vertical migration in the marine planktonic copepod Calanus pacificus. Limnology and Oceanography, 34(6):1072–1083, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Brabrand:1984:CID

[BFKN84] Age Brabrand, Bjørn Faafeng, Torsten Källqvist, and Jens Petter Nilssen. Can iron defecation from fish influence phytoplankton production and biomass in eutrophic lakes? Limnology and Oceanography, 29(6):1330–1334, November 1984. CODEN LI-OCAH. ISSN 0024-3590.

Bogdan:1982:SPF

[BG82] Kenneth G. Bogdan and John J. Gilbert. Seasonal patterns of feeding by natural populations of Keratella, Polyarthra, and Bosmina: Clearance rates, selectivities, and contributions to community grazing. Limnology and Oceanography, 27(5):918–934, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Burns:1986:DOM

[BG86a] Carolyn W. Burns and John J. Gilbert. Direct observations of the mechanisms of interference between Daphnia and Keratella cochlearis. Limnology and Oceanography, 31(4):859–866, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Burns:1986:EDS

[BG86b] Carolyn W. Burns and John J. Gilbert. Effects of daphnid size and density on interference between Daphnia and Keratella cochlearis. Limnology and Oceanography, 31(4):848–858, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Bender:1987:CFM

[BGJ^{+}87] Michael Bender, Karen Grande, Kenneth Johnson, John Marra, Peter J. LeB. Williams, John Sieburth, Michael Pilson, Chris Langdon, Gary Hitchcock, Joseph Orchardo, Carleton Hunt, Percy Donaghay, and Kristina Heinemann. A comparison of four methods for determining planktonic community production. Limnology and Oceanography, 32(5):1085–1098, September 1987. CO-DEN LIOCAH. ISSN 0024-3590.

Blackburn:1983:NCD

[BH83] T. H. Blackburn and K. Henriksen. Nitrogen cycling in different types of sediments from Danish waters. Limnology and Oceanography, 28(3):477–493, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Boavida:1984:PRD

[BH84] M. J. Boavida and Robert T. Heath. Are the phosphatases released by Daphnia magna components of its food? Limnology and Oceanography, 29(3):641–645, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Benoit:1988:CDO

[BH88a] Gaboury Benoit and Harold F. Hemond. Comment on “Dilution of ^{210}Pb by organic sedimentation in lakes of different trophic states, and application to studies of sediment-water interaction”‘ (Binford and Brenner). Limnology and Oceanography, 33(2):299– 304, March 1988. CODEN LIOCAH. ISSN 0024-3590. See reply [BB88].

Benoit:1988:IMM

[BH88b] Gaboury Benoit and Harold F. Hemond. Improved methods for the measurement of ^{210}Po. ^{210}Pb, and ^{226}Ra. Limnology and Oceanography, 33(6 (part 2)):1618–1622, November 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Brockmann:1982:ASF

[BHK^{+}82] Uwe H. Brockmann, Heinrich Huhnerfuss, Gerhard Kattner, Hans-Christoph Broecker, and Gunter Hentzschel. Artificial surface films in the sea area near Sylt. Limnology and Oceanography, 27(6):1050–1058, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Birkeland:1988:BRE

[Bir88] Charles Birkeland. Book review: Ecology of tropical oceans (A. R. Longhurst and D. Pauly). Limnology and Oceanography, 33(5): 1229–1230, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Bjorrisen:1988:PEO

[Bjø88] Peter Koefoed Bjørrisen. Phytoplankton exudation of organic matter: Why do healthy cells do it? Limnology and Oceanography, 33 (1):151–154, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Benson:1980:CIF

[BK80] Bruce B. Benson and Daniel Krause, Jr. The concentration and isotopic fractionation of gases dissolved in freshwater in equilibrium with the atmosphere. 1. Oxygen. Limnology and Oceanography, 25(4):662–671, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Battarbee:1982:UEC

[BK82] R. W. Battarbee and M. J. Kneen. The use of electronically counted microspheres in absolute diatom analysis. Limnology and Oceanography, 27(1):184–188, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Benson:1984:CIF

[BK84] Bruce B. Benson and Daniel Krause, Jr. The concentration and isotopic fractionation of oxygen dissolved in freshwater and seawater in equilibrium with the atmosphere. Limnology and Oceanography, 29(3):620–632, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Bird:1987:APR

[BK87] David F. Bird and Jacob Kalff. Algal phagotrophy: Regulating factors and importance relative to photosynthesis in Dinobryon (Chrysophyceae). Limnology and Oceanography, 32(2):277–284, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Bird:1989:PSM

[BK89] D. F. Bird and J. Kalff. Phagotrophic sustenance of a metalimnetic phytoplankton peak. Limnology and Oceanography, 34(1):155– 162, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Bidigare:1985:RDC

[BKB85] R. R. Bidigare, M. C. Kennicutt II, and J. M. Brooks. Rapid determination of chlorophylls and their degradation products by highperformance liquid chromatography. Limnology and Oceanography, 30(2):432–435, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Bower:1987:SMP

[BKF^{+}87] Peter M. Bower, Carol A. Kelly, Everett J. Fee, John A. Shearer, Doug R. DeClercq, and David W. Schindler. Simultaneous measurement of primary production by whole-lake and bottle radiocarbon additions. Limnology and Oceanography, 32(2):299–312, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Benner:1988:CCE

[BLKH88] Ronald Benner, Jan Lay, Elizabeth K’nees, and Robert E. Hodson. Carbon conversion efficiency for bacterial growth on lignocellulose: Implications for detritus-based food webs. Limnology and Oceanography, 33(6 (part 2)):1514–1526, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Blomqvist:1981:STS

[Blo81] Sven Blomqvist. Sediment trapping — a subaquatic in situ experiment. Limnology and Oceanography, 26(3):585–590, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Boatman:1982:MEN

[BM82] Charles D. Boatman and James W. Murray. Modeling exchangeable NH_{4}^{+} adsorption in marine sediments: Process and controls of adsorption. Limnology and Oceanography, 27(1):99–110, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Benner:1985:EPP

[BMH85] Ronald Benner, Mary Ann Moran, and Robert E. Hodson. Effects of pH and plant source on lignocellulose biodegradation rates in two wetland ecosystems, the Okefenokee Swamp and a Georgia salt marsh. Limnology and Oceanography, 30(3):489–499, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Benner:1986:BCL

[BMH86] Ronald Benner, Mary Ann Moran, and Robert E. Hodson. Biogeochemical cycling of lignocellulosic carbon in marine and freshwater ecosystems: Relative contributions of procaryotes and eucaryotes.

Limnology and Oceanography, 31(1):89–100, January 1986. CO-DEN LIOCAH. ISSN 0024-3590.

Bender:1987:WCS

[BMH^{+}87] Michael Bender, William Martin, Jennifer Hess, Fred Sayles, Lary Ball, and Claude Lambert. A whole-core squeezer for interfacial pore-water sampling. Limnology and Oceanography, 32(6):1214– 1225, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Bricaud:1981:ADO

[BMP81] Annick Bricaud, Andre Morel, and Louis Prieur. Absorption by dissolved organic matter of the sea (yellow substance) in the UV and visible domains. Limnology and Oceanography, 26(1):43–53, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Bricaud:1983:OEF

[BMP83] Annick Bricaud, Andre Morel, and Louis Prieur. Optical efficiency factors of some phytoplankters. Limnology and Oceanography, 28 (5):816–832, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Blasco:1984:RBN

[BMPD84] Dolors Blasco, J. J. Maclsaac, T. T. Packard, and R. C. Dugdale. Relationship between nitrate reductase and nitrate uptake in phytoplankton in the Peru upwelling region. Limnology and Oceanography, 29(2):275–286, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Buskey:1983:EDB

[BMS83] Edward Buskey, Lesley Mills, and Elijah Swift. The effects of dinoflagellate bioluminescence on the swimming behavior of a marine copepod. Limnology and Oceanography, 28(3):575–579, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Bauer:1988:MBH

[BMS^{+}88] James E. Bauer, Paul A. Montagna, Robert B. Spies, Maria C. Prieto, and Dane Hardin. Mierobial biogeochemistry and heterotrophy in sediments of a marine hydrocarbon seep. Limnology and Oceanography, 33(6 (part 2)):1493–1513, November 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Boehlert:1989:BRM

[Boe89] George W. Boehlert. Book review: Marine populations: an essay on population regulation and speciation (M. Sinclair). Limnology and Oceanography, 34(5):968–969, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Boon:1980:CNP

[Boo80] John D. Boon III. Comment on “Nutrient and particulate fluxes in a salt marsh ecosystem: Tidal exchanges and inputs by precipitation and groundwater” (Valiela et al.). Limnology and Oceanography, 25(1):182–183, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Boraas:1983:PDF

[Bor83] Martin E. Boraas. Population dynamics of food-limited rotifers in two-stage chemostat culture. Limnology and Oceanography, 28(3): 546–563, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Borum:1987:DEE

[Bor87] Jens Borum. Dynamics of epiphyton on eelgrass (Zostera marina L.) leaves: Relative roles of algal growth, herbivory, and substratum turnover. Limnology and Oceanography, 32(4):986–992, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Bothwell:1985:PLL

[Bot85] Max L. Bothwell. Phosphorus limitation of lotic periphyton growth rates: an intersite comparison using continuous-flow troughs (Thompson River system, British Columbia). Limnology and Oceanography, 30(3):527–542, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Bowden:1984:NID

[Bow84] William B. Bowden. A nitrogen-15 isotope dilution study of ammonium production and consumption in a marsh sediment. Limnology and Oceanography, 29(5):1004–1015, September 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Bradford:1987:RBC

[BP87] Marie E. Bradford and Robert Henry Peters. The relationship between chemically analyzed phosphorus fractions and bioavailable phosphorus. Limnology and Oceanography, 32(5):1124–1137, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Bird:1986:ISD

[BQ86] A. E. Bird and R. J. Quinn. Isolation of symbiotic dinoflagellates by centrifugal elutriation. Limnology and Oceanography, 31(1): 225–228, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Burns:1980:SMS

[BR80] N. M. Burns and F. Rosa. In situ measurement of the settling velocity of organic carbon particles and 10 species of phytoplankton. Limnology and Oceanography, 25(5):855–864, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Button:1986:DHM

[BR86] D. K. Button and B. R. Robertson. Dissolved hydrocarbon metabolism: The concentration-dependent kinetics of toluene oxidation in some North American estuaries. Limnology and Oceanography, 31(1):101–111, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Bell:1989:AID

[BR89] Russell T. Bell and Bo Riemann. Adenine incorporation into DNA as a measure of microbial production in freshwaters. Limnology and Oceanography, 34(2):435–444, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Brass:1980:TEA

[Bra80] G. W. Brass. Trace elements in acantharian skeletons. Limnology and Oceanography, 25(1):146–149, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Brand:1982:GVS

[Bra82] Larry E. Brand. Genetic variability and spatial patterns of genetic differentiation in the reproductive rates of the marine coccolithophores Emiliania huxleyi and Gephyrocapsa oceanica. Limnology and Oceanography, 27(2):236–245, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Burke:1983:UWC

[BRBL83] Roger A. Burke, Jr., David F. Reid, James M. Brooks, and Dennis M. Lavoie. Upper water column methane geochemistry in the eastern tropical North Pacific. Limnology and Oceanography, 28 (1):19–32, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Bjornsen:1988:TIB

[BRH^{+}88] Peter Koefoed Bjørnsen, Bo Riemann, Steen Jesper Horsted, Torkel Gissel Nielsen, and Jan Pock-Sten. Trophic interactions between heterotrophic nanoflagellates and bacterioplankton in manipulated seawater enclosures. Limnology and Oceanography, 33 (3):409–420, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Bright:1982:SBP

[Bri82] Greg R. Bright. Secondary benthic production in a tropical island stream. Limnology and Oceanography, 27(3):472–480, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Brock:1980:BRS

[Bro80] Thomas D. Brock. Book review: Sieburth, J. McN. 1979. Sea microbes. Oxford University Press, New York, x + 491 p. $49.95. Limnology and Oceanography, 25(1):199–200, January 1980. CO-DEN LIOCAH. ISSN 0024-3590.

Brock:1982:BRD

[Bro82] Thomas D. Brock. Book review: Droop, M. R., and H. W. Jannasch [Eds.] 1980. Advances in aquatic microbiology. V. 2. Academic Press Inc. (London) Ltd. 356 p. $46.00. Limnology and Oceanography, 27(1):198, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Brock:1983:BRS

[Bro83] Thomas D. Brock. Book review: Seki, H. 1982. Organic materials in aquatic systems. CRC Press, Boca Raton, Florida. 201 p. $64.00. Limnology and Oceanography, 28(5):1041, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Bruland:1989:CZN

[Bru89a] Kenneth W. Bruland. Complexation of zinc by natural organic ligands in the central North Pacific. Limnology and Oceanography, 34(2):269–285, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Brush:1989:RPE

[Bru89b] Grace S. Brush. Rates and patterns of estuarine sediment accumulation. Limnology and Oceanography, 34(7):1235–1246, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Brzezinski:1988:VDA

[Brz88] Mark A. Brzezinski. Vertical distribution of ammonium in stratified oligotrophic waters. Limnology and Oceanography, 33(5): 1176–1182, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Bdmstedt:1980:RCZ

[BS80a] Ulf Bdmstedt and Hein Rune Skjoldal. RNA concentration of zooplankton: Relationship with size and growth. Limnology and Oceanography, 25(2):304–316, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Bell:1980:BUAb

[BS80b] Wayne H. Bell and Egil Sakshaug. Bacterial utilization of algal extracellular products. 2. A kinetic study of natural populations. Limnology and Oceanography, 25(6):1021–1033, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Baker:1982:BOC

[BS82] Karen S. Baker and Raymond C. Smith. Bio-optical classification and model of natural waters. 2. Limnology and Oceanography, 27 (3):500–509, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Bleiwas:1985:CLS

[BS85a] Agnes H. Bleiwas and Pamela M. Stokes. Collection of large and small food particles by Bosmina. Limnology and Oceanography, 30(5):1090–1092, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Braunwarth:1985:ASG

[BS85b] C. Braunwarth and U. Sommer. Analyses of the in situ growth rates of cryptophyceae by use of the mitotic index technique. Limnology and Oceanography, 30(4):893–897, July 1985. CODEN LI-OCAH. ISSN 0024-3590.

Bjorkman:1986:MLO

[BS86] Bjorn Bjorkman and Joseph Shapiro. Measurement of the lipid ovary index on stored zooplankton samples. Limnology and Oceanography, 31(5):1138–1139, September 1986. CODEN LI-OCAH. ISSN 0024-3590.

Batchelder:1989:ENS

[BS89] Harold P. Batchelder and Elijah Swift. Estimated near-surface mesoplanktonic bioluminescence in the western North Atlantic during July 1986. Limnology and Oceanography, 34(1):113–128, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Barton:1987:CBT

[BSB^{+}87] C. E. Barton, D. K. Solomon, J. R. Bowman, T. E. Cerling, and M. D. Sayer. Chloride budgets in transient lakes: Lakes baringo, naivasha, and turkana. Limnology and Oceanography, 32(3):745– 751, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Boyd:1980:GPC

[BSC80] Carl M. Boyd, Sharon L. Smith, and Timothy J. Cowles. Grazing patterns of copepods in the upwelling system off Peru. Limnology and Oceanography, 25(4):583–596, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Butler:1981:IPI

[BSF81] Edward C. V. Butler, J. David Smith, and Nicholas S. Fisher. Influence of phytoplankton on iodine speciation in seawater. Limnology and Oceanography, 26(2):382–386, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Brand:1983:LMP

[BSG83] Larry E. Brand, William G. Sunda, and Robert R. L. Guillard. Limitation of marine phytoplankton reproductive rates by zinc, manganese, and iron. Limnology and Oceanography, 28(6):1182– 1198, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Baker:1983:CDF

[BSN83] Karen S. Baker, Raymond C. Smith, and James R. Nelson. Chlorophyll determinations with filter fluorometer: Lamp/filter combination can minimize error. Limnology and Oceanography, 28(5): 1037–1040, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Bienfang:1984:TSV

[BSON84] P. K. Bienfang, J. P. Szyper, M. Y. Okamoto, and E. K. Noda. Temporal and spatial variability of phytoplankton in a subtropical ecosystem. Limnology and Oceanography, 29(3):527–539, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Berman:1984:CAN

[BSS^{+}84] T. Berman, B. F. Sherr, E. Sherr, D. Wynne, and J. J. McCarthy. The characteristics of ammonium and nitrate uptake by phytoplankton in Lake Kinneret. Limnology and Oceanography, 29(2): 287–297, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Blanchard:1981:BSB

[BSW81] Duncan C. Blanchard, Lawrence D. Syzdek, and Martin E. Weber. Bubble scavenging of bacteria in freshwater quickly produces bacterial enrichment in airborne jet drops. Limnology and Oceanography, 27(4):961–964, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Bennett:1984:RAB

[BT84] J. T. Bennett and K. K. Turekian. Radiometric ages of brachyuran crabs from the Galapagos spreading-center hydrothermal ventfield. Limnology and Oceanography, 29(5):1088–1091, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Bloesch:1986:HSD

[BU86] Jurg Bloesch and Urs Uehlinger. Horizontal sedimentation differences in a eutrophic Swiss lake. Limnology and Oceanography, 31 (5):1094–1109, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Burgis:1980:BRL

[Bur80] Mary J. Burgis. Book review: Lewis, W. M., Jr. 1979. Zooplankton community analysis: Studies on a tropical system. Springer-Verlag, New York, ix + 163 p. Limnology and Oceanography, 25 (5):966, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Burney:1984:SFP

[Bur84] Curtis Burney. A subsurface flexible plastic enclosure for the in situ study of short term microbiological and chemical dynamics. Limnology and Oceanography, 29(5):1140–1144, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Burgis:1986:BRD

[Bur86a] Mary Burgis. Book review: Downing, J. A., and F. H. Rigler [Eds.]. 1984. A manual on methods for the assessment of secondary production in fresh waters, IBP Handbook No. 17, 2nd ed. Blackwell Sci. Publ., Oxford. 500 p. $55.00. Limnology and Oceanography, 31(3):662–667, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Burney:1986:BUT

[Bur86b] Curtis M. Burney. Bacterial utilization of total in situ dissolved carbohydrate in offshore waters. Limnology and Oceanography, 31 (2):427–431, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Burgess:1988:BRB

[Bur88] Robert L. Burgess. Book review: Breaking new waters (A. L. Beckel). Limnology and Oceanography, 33(6 (part 2)):1650–1651, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Butler:1982:PDS

[But82] Malcolm G. Butler. Production dynamics of some Arctic Chironomus larvae. Limnology and Oceanography, 26(5):728–736, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Button:1986:AOS

[But86] D. K. Button. Affinity of organisms for substrate. Limnology and Oceanography, 31(2):453–456, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Bucklin:1986:GHE

[BW86] Ann Bucklin and Peter H. Wiebe. Genetic heterogeneity in euphausiid populations: Euphausia krohnii and Nematoscelis megalops in North Atlantic Slope Water. Limnology and Oceanography, 31(6):1346–1352, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Barnett:1989:CLS

[BW89] P. R. O. Barnett and J. Watson. A comment on “Laboratory studies on core sampling with application to subtidal meiobenthos collection” (Rutledge and Fleeger). Limnology and Oceanography, 34(6):1142–1143, September 1989. CODEN LIOCAH. ISSN 0024-3590. See reply [RF89].

Boon:1982:PMS

[BWG82] Jaap J. Boon, Robert G. Wetzel, and Gordon L. Godshalk. Pyrolysis mass spectrometry of some scirpus species and their decomposition products. Limnology and Oceanography, 27(5):839–848, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Byron:1983:OCS

[BWG83] E. R. Byron, P. T. Whitman, and C. R. Goldman. Observations of copepod swarms in Lake Tahoe. Limnology and Oceanography, 28(2):378–382, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Bryant:1980:SCS

[BWJ80] R. Bryant, D. J. A. Williams, and A. E. James. A sampler for cohesive sediment in the benthic boundary layer. Limnology and Oceanography, 49(6):572–576, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Berman:1985:SDD

[BWS^{+}85] T. Berman, P. D. Walline, A. Schneller, J. Rothenberg, and D. W. Townsend. Secchi disk depth record: a claim for the eastern Mediterranean. Limnology and Oceanography, 30(2):447–448, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Bayley:1985:EMN

[BZH^{+}85] Suzanne E. Bayley, John Zoltek, Jr., Albert J. Hermann, Thomas J. Dolan, and Louis Tortora. Experimental manipulation of nutrients and water in a freshwater marsh: Effects on biomass, decomposition, and nutrient accumulation. Limnology and Oceanography, 30(3):500–512, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Colman:1983:HDS

[CA83] John A. Colman and David E. Armstrong. Horizontal diffusivity in a small, ice-covered lake. Limnology and Oceanography, 28(5): 1020–1026, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Colman:1987:VED

[CA87] John A. Colman and David E. Armstrong. Vertical eddy diffusivity determined with ^{222}Rn in the benthic boundary layer of ice-covered lakes. Limnology and Oceanography, 32(3):577–590, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Campbell:1989:PAR

[CA89] Janet W. Campbell and Thorkild Aarup. Photosynthetically available radiation at high latitudes. Limnology and Oceanography, 34 (8):1490–1499, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Campbell:1987:BRS

[Cam87] P. G. C. Campbell. Book review: Stumm, W. [ed.]. 1985. Chemical processes in lakes. John Wiley and Sons, Inc., New York. 435 p. $59.95. Limnology and Oceanography, 32(5):1177–1178, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Capelli:1982:DNW

[Cap82] Gregory M. Capelli. Displacement of northern Wisconsin crayfish by Orconectes rusticus (Girard). Limnology and Oceanography, 26 (5):741–745, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Carlson:1980:MCC

[Car80] Robert E. Carlson. More complications in the chlorophyll-secchi disk relationship. Limnology and Oceanography, 25(2):379–382, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Carlson:1983:DOM

[Car83] David J. Carlson. Dissolved organic materials in surface microlayers: Temporal and spatial variability and relation to sea state. Limnology and Oceanography, 28(3):415–431, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1984:IWS

[Car84] R. Carignan. Interstitial water sampling by dialysis: Methodological notes. Limnology and Oceanography, 29(3):667–670, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1985:RBP

[Car85] Robert C. Carpenter. Relationships between primary production and ii radiance in coral reef algal communities. Limnology and Oceanography, 30(4):784–793, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Carey:1989:FDT

[Car89] Drew A. Carey. Fluorometric detection of tracer particles used to study animal-particle dynamics. Limnology and Oceanography, 34 (3):630–635, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Conrad:1983:PCC

[CAS83] Ralf Conrad, Michel Aragno, and Wolfgang Seiler. Production and consumption of carbon monoxide in a eutrophic lake. Limnology and Oceanography, 28(1):42–49, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Cattaneo:1983:GE

[Cat83] A. Cattaneo. Grazing on epiphytes. Limnology and Oceanography, 28(1):124–132, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Cattaneo:1985:BRW

[Cat85] Antonella Cattaneo. Book review: Wetzel, R. [Ed.] 1983. Periphyton of freshwater ecosystems. Develop. Hydrobiol. 17. Dr. W. Junk BY Publ., The Hague. 346 p. $87.50. Limnology and Oceanography, 30(4):917, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Cutter:1984:MBS

[CB84] Gregory A. Cutter and Kenneth W. Bruland. The marine biogeochemistry of selenium: a re-evaluation. Limnology and Oceanography, 29(6):1179–1192, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Coale:1985:DWC

[CB85] Kenneth H. Coale and Kenneth W. Bruland. ^{234}Th : ^{238}U disequilibria within the California Current. Limnology and Oceanography, 30(1):22–33, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Carr:1986:EIT

[CB86] Alan P. Carr and Michael W. L. Blackley. The effects and implication of tides and rainfall on the circulation of water within salt marsh sediments. Limnology and Oceanography, 31(2):266–276, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Coale:1987:OSE

[CB87] Kenneth H. Coale and Kenneth W. Bruland. Oceanic stratified euphotic zone as elucidated by ^{234}Th : ^{238}U disequilibria. Limnology and Oceanography, 32(1):189–200, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Coale:1988:CCN

[CB88] Kenneth H. Coale and Kenneth W. Bruland. Copper complexation in the Northeast Pacific. Limnology and Oceanography, 33(5): 1084–1101, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Canfield:1983:FSH

[CBH83] Daniel E. Canfield, Jr., Roger W. Bachmann, and Mark V. Hoyer. Freeze-out of salts in hard-water lakes. Limnology and Oceanography, 28(5):970–977, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Carlson:1984:REF

[CC84] David J. Carlson and Mary L. Carlson. Reassessment of exudation by fucoid macroalgae. Limnology and Oceanography, 29(5):1077– 1087, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Conway:1989:REB

[CCF89] Noellette Conway, Judith McDowell Capuzzo, and Brian Fry. The role of endosymbiotic bacteria in the nutrition of Solemya velum: Evidence from a stable isotope analysis of endosymbionts and host.

Limnology and Oceanography, 34(1):249–255, January 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Cloern:1983:APM

[CCO83] James E. Cloern, Brian E. Cole, and Ronald S. Oremland. Autotrophic processes in meromictic Big Soda Lake, Nevada. Limnology and Oceanography, 28(6):1049–1061, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Cole:1989:DOC

[CCS^{+}89] Jonathan J. Cole, Nina F. Caraco, David L. Strayer, Clifford Ochs, and Scott Nolan. A detailed organic carbon budget as an ecosystem-level calibration of bacterial respiration in an oligotrophic lake during midsummer. Limnology and Oceanography, 34(2):286–296, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Cloern:1987:BSL

[CCW87] James E. Cloern, Brian E. Cole, and Sally M. Wienke. Big Soda Lake (Nevada). 4. Vertical fluxes of particulate matter: Seasonality and variations across the chemocline. Limnology and Oceanography, 32(4):815–824, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1985:NNU

[CD85] Edward J. Carpenter and Susan Dunham. Nitrogenous nutrient uptake, primary production, and species composition of phytoplankton in the Carmans River estuary, Long Island, New York. Limnology and Oceanography, 30(3):513–526, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Carman:1988:CTC

[CDG88] Kevin R. Carman, Fred C. Dobbs, and James B. Guckert. Consequences of thymidine catabolism for estimates of bacterial production: an example from a coastal marine sediment. Limnology and Oceanography, 33(6 (part 2)):1595–1606, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Cohen:1984:EAC

[CDPC84] Ronald R. H. Cohen, Paul V. Dresler, Elizabeth J. P. Phillips, and Robert L. Cory. The effect of the Asiatic clam, Corbicula fluminea, on phytoplankton of the Potomac River, Maryland. Limnology and Oceanography, 29(1):170–180, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Conover:1986:PLC

[CDRW86] R. J. Conover, Ravi Durvasula, Suzanne Roy, and Rong Wang. Probable loss of chlorophyll-derived pigments during passage through the gut of zooplankton, and some of the consequences. Limnology and Oceanography, 31(4):878–886, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Carpenter:1986:CPD

[CEE86] Stephen R. Carpenter, Monica M. Elser, and James J. Elser. Chlorophyll production, degradation, and sedimentation: Implications for paleolimnology. Limnology and Oceanography, 31(1): 112–124, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Carignan:1981:PMP

[CF81] R. Carignan and R. J. Flett. Postdepositional mobility of phosphorus in lake sediments. Limnology and Oceanography, 26(2): 361–366, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Coffin:1989:CIC

[CFPW89] Richard B. Coffin, Brian Fry, Bruce J. Peterson, and Richard T. Wright. Carbon isotopic compositions of estuarine bacteria. Limnology and Oceanography, 34(7):1305–1310, November 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Cooney:1980:RBE

[CG80] John D. Cooney and Carl W. Gehrs. The relationship between egg size and naupliar size in the calanoid copepod Diaptomus clavipes Schacht. Limnology and Oceanography, 49(6):549–552, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Childress:1989:BBD

[CGCG89] James J. Childress, David L. Gluck, Robert S. Carney, and Marcia M. Gowing. Benthopelagic biomass distribution and oxygen consumption in a deep-sea benthic boundary layer dominated by gelatinous organisms. Limnology and Oceanography, 34(5):913– 930, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Carleton:1987:DOD

[CH87] J. H. Carleton and W. M. Hamner. A diver-operated device for the capture of mobile epibenthic organisms. Limnology and Oceanography, 32(2):503–510, March 1987. CODEN LIOCAH. ISSN 0024-

Chapman:1985:BRB

[Cha85] Ross Chapman. Book review: Boyles, C. A. 1984. Acoustic waveguides. Applications to oceanic science. John Wiley & Sons, Inc., New York. 321 p. $46.95. Limnology and Oceanography, 30(1): 237–238, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Checkley:1980:EPM

[Che80a] David M. Checkley, Jr. The egg production of a marine planktonic copepod in relation to its food supply: Laboratory studies. Limnology and Oceanography, 25(3):430–446, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Checkley:1980:FLE

[Che80b] David M. Checkley, Jr. Food limitation of egg production by a marine, planktonic copepod in the sea off southern California. Limnology and Oceanography, 25(6):991–998, November 1980. CO-DEN LIOCAH. ISSN 0024-3590.

Christensen:1984:DSO

[Chr84] Dorte Christensen. Determination of substrates oxidized by sulfate reduction in intact cores of marine sediments. Limnology and Oceanography, 29(1):189–191, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Chrost:1989:CSG

[Chr89] Ryszard J. Chróst. Characterization and significance of βglucosidase activity in lake water. Limnology and Oceanography, 34(4):660–672, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Churchill:1985:IOS

[Chu85] J. H. Churchill. Intrusions of outer shelf and slope water within the nearshore zone off Long Island, New York. Limnology and Oceanography, 30(5):972–986, September 1985. CODEN LI-OCAH. ISSN 0024-3590.

Campbell:1988:ISS

[CI88] Lisa Campbell and Rodolfo Iturriaga. Identification of Synechococcus spp. in the Sargasso Sea by immunofluorescence and fluorescence excitation spectroscopy performed on individual cells. Limnology and Oceanography, 33(5):1196–1201, September 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Cuhel:1983:MGM

[CJTL83] Russell L. Cuhel, Holger W. Jannasch, Craig D. Taylor, and David R. S. Lean. Microbial growth and macromolecular synthesis in the northwestern Atlantic Ocean. Limnology and Oceanography, 28(1): 1–18, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Cattaneo:1979:PPA

[CK79] A. Cattaneo and J. Kalff. Primary production of algae growing on natural and artificial aquatic plants: a study of interactions between epiphytes and their substrate. Limnology and Oceanography, 24(6):1031–1037, November 1979. CODEN LIOCAH. ISSN 0024-3590. See comment [GG81].

Cattaneo:1980:RCA

[CK80] A. Cattaneo and J. Kalff. The relative contribution of aquatic macrophytes and their epiphytes to the production of macrophyte beds. Limnology and Oceanography, 25(2):280–289, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Cattaneo:1981:RCG

[CK81] A. Cattaneo and J. Kalff. Reply to comment by Gough and Gough. Limnology and Oceanography, 27(4):988–989, September 1981. CODEN LIOCAH. ISSN 0024-3590. See [GG81].

Carignan:1982:PRS

[CK82] R. Carignan and J. Kalff. Phosphorus release by submerged macrophytes: Significance to epiphyton and phytoplankton. Limnology and Oceanography, 27(3):419–427, May 1982. CODEN LI-OCAH. ISSN 0024-3590.

Currie:1984:CAF

[CK84a] David J. Currie and Jaap Kalff. A comparison of the abilities of freshwater algae and bacteria to acquire and retain phosphorus. Limnology and Oceanography, 29(2):298–310, March 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Currie:1984:RIB

[CK84b] David J. Currie and Jaap Kalff. The relative importance of bacterioplankton and phytoplankton in phosphorus uptake in freshwater. Limnology and Oceanography, 29(2):311–321, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Capone:1988:CMD

[CK88] Douglas G. Capone and Ronald P. Kiene. Comparison of microbial dynamics in marine and freshwater sediments: Contrasts in anaerobic carbon catabolism. Limnology and Oceanography, 33(4 (part 2)):725–749, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Cook:1986:MHI

[CKST86] R. B. Cook, C. A. Kelly, D. W. Schindler, and M. A. Turner. Mechanisms of hydrogen ion neutralization in an experimentally acidified lake. Limnology and Oceanography, 31(1):134–148, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Conley:1989:DSC

[CKT89] Daniel J. Conley, Susan S. Kilham, and Edward Theriot. Differences in silica content between marine and freshwater diatoms. Limnology and Oceanography, 34(1):205–212, January 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Copping:1980:CBM

[CL80] Andrea E. Copping and Carl J. Lorenzen. Carbon budget of a marine phytoplankton-herbivore system with carbon-14 as a tracer. Limnology and Oceanography, 25(5):873–882, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Chapman:1981:SSM

[CL81] Piers Chapman and Peter S. Liss. The sea surface microlayer: Measurements of dissolved iodine species and nutrients in coastal waters. Limnology and Oceanography, 26(2):387–390, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Cole:1982:PPD

[CLS82] Jonathan J. Cole, Gene E. Likens, and David L. Strayer. Photosynthetically produced dissolved organic carbon: an important carbon source for planktonic bacteria. Limnology and Oceanography, 27(6):1080–1090, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Cranston:1980:CSC

[CM80] R. E. Cranston and J. W. Murray. Chromium species in the Columbia River and estuary. Limnology and Oceanography, 25(6): 1104–1112, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Cuker:1981:SPF

[CM81] Benjamin E. Cuker and Samuel C. Mozley. Summer population fluctuations, feeding, and growth of Hydra in an Arctic lake. Limnology and Oceanography, 26(4):697–708, July 1981. CODEN LI-OCAH. ISSN 0024-3590.

Calaban:1982:ETD

[CM82] M. J. Calaban and J. C. Makarewicz. The effect of temperature and density on the amplitude of vertical migration of Daphnia magna. Limnology and Oceanography, 27(2):262–271, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Crill:1983:STF

[CM83] Patrick M. Crill and Christopher S. Martens. Spatial and temporal fluctuations of methane production in anoxic coastal marine sediments. Limnology and Oceanography, 28(6):1117–1130, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Chen:1986:TPN

[CM86] Chen-Tung A. Chen and Frank J. Millero. thermodynamic properties for natural waters covering only the limnological range. Limnology and Oceanography, 31(3):657–662, May 1986. CODEN LI-OCAH. ISSN 0024-3590.

Carlough:1989:PTS

[CM89] Leslie A. Carlough and Judy L. Meyer. Protozoans in two southeastern blackwater rivers and their importance to trophic transfer. Limnology and Oceanography, 34(1):163–177, January 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Carlson:1987:TFD

[CMB87] David J. Carlson, Lynn E. Morrill, and Jennifer E. Brophy. Techniques of fluorescence depolarization for measuring seawater viscosities. Limnology and Oceanography, 32(6):1377–1381, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Chanton:1989:GTM

[CMK89] Jeffrey P. Chanton, Christopher S. Martens, and Cheryl A. Kelley. Gas transport from methane-saturated, tidal freshwater and wetland sediments. Limnology and Oceanography, 34(5):807–819, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Chapnick:1982:MMM

[CMN82] Susan D. Chapnick, Willard S. Moore, and Kenneth H. Nealson. Microbially mediated manganese oxidation in a freshwater lake. Limnology and Oceanography, 27(6):1004–1014, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Collos:1989:HLT

[CMR89] Y. Collos, S. Y. Maestrini, and J.-M. Robert. High long-term nitrate uptake by oyster-pond microalgae in the presence of high ammonium concentrations. Limnology and Oceanography, 34(5): 957–964, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Coffin:1989:BUD

[Cof89] Richard B. Coffin. Bacterial uptake of dissolved free and combined amino acids in estuarine waters. Limnology and Oceanography, 34 (3):531–542, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Coleman:1980:EDB

[Col80a] Annette W. Coleman. Enhanced detection of bacteria in natural environments by fluorochrome staining of DNA. Limnology and Oceanography, 25(5):948–951, September 1980. CODEN LI-OCAH. ISSN 0024-3590.

Collos:1980:TSN

[Col80b] Yves Collos. Transient situations in nitrate assimilation by marine diatoms. 1. Changes in uptake parameters during nitrogen starvation. Limnology and Oceanography, 25(6):1075–1081, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Colios:1982:TSN

[Col82] Yves Colios. Transient situations in nitrate assimilation by marine diatoms. 2. Changes in nitrate and nitrite following a nitrate perturbation. Limnology and Oceanography, 27(3):528–535, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Collienne:1983:PIE

[Col83] R. H. Collienne. Photoreduction of iron in the epilimnion of acidic lakes. Limnology and Oceanography, 28(1):83–100, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Cole:1984:EAP

[Col84a] Brian E. Cole. Evaluation of an ATP photometer for measuring 14C activity. Limnology and Oceanography, 29(3):663–666, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Cuhel:1984:NSP

[COL84b] Russell L. Cuhel, Peter B. Ortner, and David R. S. Lean. Night synthesis of protein by algae. Limnology and Oceanography, 29(4): 731–744, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Collier:1985:MNP

[Col85] Robert W. Collier. Molybdenum in the Northeast Pacific Ocean. Limnology and Oceanography, 30(6):1351–1354, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Correll:1981:NMB

[Cor81] David L. Correll. Nutrient mass balances for the watershed, headwaters intertidal zone, and basin of the Rhode River Estuary. Limnology and Oceanography, 26(6):1142–1149, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Cornett:1989:PCH

[Cor89] R. J. Cornett. Predicting changes in hypolimnetic oxygen concentrations with phosphorus retention, temperature, and morphometry. Limnology and Oceanography, 34(7):1359–1366, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Cote:1983:DDV

[CP83] Brenda Côté and Trevor Platt. Day-to-day variations in the spring-summer photosynthetic parameters of coastal marine phytoplankton. Limnology and Oceanography, 28(2):320–344, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Caraco:1986:MBC

[CP86] N. Caraco and A. H. Puccoon. The measurement of bacterial chlorophyll and algal chlorophyll-a in natural waters. Limnology and Oceanography, 31(4):889–893, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Caraco:1989:RCK

[CP89] N. F. Caraco and A. H. Puccoon. A reply to the comment by Knowlton et al. Limnology and Oceanography, 34(7):1383–1384, November 1989. CODEN LIOCAH. ISSN 0024-3590. See [KJP89].

Cryer:1986:RIB

[CPT86] Martin Cryer, Graeme Peirson, and Colin R. Townsend. Reciprocal interactions between roach, Rutilus rutilus, and zooplankton in a small lake: Prey dynamics and fish growth and recruitment.

Limnology and Oceanography, 31(5):1022–1038, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Cornett:1980:AHO

[CR80] R. J. Cornett and F. H. Rigler. The areal hypolimnetic oxygen deficit: an empirical test of the model. Limnology and Oceanography, 25(4):672–679, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Chapra:1983:CEC

[CR83] Steven C. Chapra and Kenneth H. Reckhow. Comment on “The effect of changes in the nutrient income on the condition of Lake Washington” (Edmondson and Lehman). Limnology and Oceanography, 28(4):792–795, July 1983. CODEN LIOCAH. ISSN 0024-3590. See [EL81].

Cook:1980:DLS

[CRFB80] David O. Cook, James W. Richardson, Timothy L. Flynn, and Kenneth J. Berger. Design of a Lagrangian surface current tracer. Limnology and Oceanography, 25(6):1123–1127, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Cowles:1987:STC

[CRGC87] T. J. Cowles, M. R. Roman, A. L. Gauzens, and N. J. Copley. Short-term changes in the biology of a warm-core ring: Zooplankton biomass and grazing. Limnology and Oceanography, 32(3): 653–664, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Crumpton:1987:SRM

[Cru87] W. G. Crumpton. A simple and reliable method for making permanent mounts of phytoplankton for light and fluorescence microscopy. Limnology and Oceanography, 32(5):1154–1159, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Carlson:1981:DHS

[CS81] Robert E. Carlson and Joseph Shapiro. Dissolved humic substances: a major source of error in fluorometric analyses involving lake waters. Limnology and Oceanography, 26(4):785–790, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Cowles:1983:CFA

[CS83] Timothy J. Cowles and J. Rudi Strickier. Characterization of feeding activity patterns in the planktonic copepod Centropages typicus Krøyer under various food conditions. Limnology and Oceanography, 28(1):106–115, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Carder:1985:RSR

[CS85] K. L. Carder and R. G. Steward. A remote-sensing reflectance model of a red-tide dinoflagellate off west Florida. Limnology and Oceanography, 30(2):286–298, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Charles:1988:NMU

[CS88] Donald F. Charles and John P. Smol. New methods for using diatoms and chrysophytes to infer past pH of low-alkalinity lakes. Limnology and Oceanography, 33(6 (part 2)):1451–1462, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Csanadi:1980:BRF

[Csa80] G. T. Csanadi. Book review: Favre, A., and K. Hasselmann [Eds.]. 1978. Turbulent fluxes through the sea surface, wave dynamics, and prediction. NATO Conf. Ser. 5. Air-sea interaction, v. 1. Plenum Press, New York, xiii + 677 p. $49.50. (Typewriter composition). Limnology and Oceanography, 25(1):198–199, January 1980. CO-DEN LIOCAH. ISSN 0024-3590.

Csanady:1986:MTS

[Csa86] G. T. Csanady. Mass transfer to and from small particles in the sea. Limnology and Oceanography, 31(2):237–248, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Csanady:1987:WIF

[Csa87] G. T. Csanady. Wave interactions and fluid flows. Limnology and Oceanography, 32(5):1177, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Chrost:1986:MDE

[CSAO86] Ryszard J. Chrost, Waldemar Siuda, Dieter Albrecht, and Jurgen Overbeck. A method for determining enzymatically hydrolyzable phosphate (EHP) in natural waters. Limnology and Oceanography, 31(3):662–667, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Cifuentes:1988:SCN

[CSF88] L. A. Cifuentes, J. H. Sharp, and Marilyn L. Fogel. Stable carbon and nitrogen isotope biogeochemistry in the Delaware estuary. Limnology and Oceanography, 33(5):1102–1115, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Carder:1989:MHF

[CSHO89] Kendall L. Carder, Robert G. Steward, George R. Harvey, and Peter B. Ortner. Marine humic and fulvic acids: Their effects on remote sensing of ocean chlorophyll. Limnology and Oceanography, 34(1):68–81, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Carder:1986:RBC

[CSPV86] K. L. Carder, R. G. Steward, J. H. Paul, and G. A. Vargo. Relationships between chlorophyll and ocean color constituents as they affect remote-sensing reflectance models. Limnology and Oceanography, 31(2):403–412, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Cummins:1980:PCN

[CSW^{+}80] Kenneth W. Cummins, George L. Spengler, G. Milton Ward, Robert M. Speaker, Roger W. Ovink, David C. Mahan, and Rosanna L. Mattingly. Processing of confined and naturally entrained leaf litter in a woodland stream ecosystem. Limnology and Oceanography, 25(5):952–957, September 1980. CODEN LI-OCAH. ISSN 0024-3590.

Cuker:1983:GNI

[Cuk83] Benjamin E. Cuker. Grazing and nutrient interactions in controlling the activity and composition of the epilithic algal community of an Arctic lake. Limnology and Oceanography, 28(1):133–141, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Cuker:1987:FEI

[Cuk87] Benjamin E. Cuker. Field experiment on the influences of suspended clay and P on the plankton of a small lake. Limnology and Oceanography, 32(4):840–847, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Culberson:1980:CSP

[Cul80] Charles H. Culberson. Calculation of the in situ pH of seawater. Limnology and Oceanography, 25(1):150–152, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Currie:1984:MNP

[Cur84] David J. Currie. Microscale nutrient patches: Do they matter to the phytoplankton? Limnology and Oceanography, 29(1):211–214, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Cosovic:1982:AAP

[CV82] Božena Cosović and Vjeročka Vojvodić. The application of ac polarography to the determination of surface-active substances in seawater. Limnology and Oceanography, 27(2):361–369, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Cavaletto:1989:WET

[CVG89] Joann F. Cavaletto, Henry A. Vanderploeg, and Wayne S. Gardner. Wax esters in two species of freshwater zooplankton. Limnology and Oceanography, 34(4):785–789, June 1989. CODEN LI-OCAH. ISSN 0024-3590.

Crumpton:1981:MPP

[CW81] William G. Crumpton and Robert G. Wetzel. A method for preparing permanent mounts of phytoplankton for critical microscopy and cell counting. Limnology and Oceanography, 27(4): 976–980, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Clair:1983:TPC

[CW83] Thomas A. Clair and Paul H. Whitfield. Trends in pH, calcium, and sulfate of rivers in Atlantic Canada. Limnology and Oceanography, 28(1):160–165, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Cuhel:1984:BCS

[CW84] Russell L. Cuhel and John B. Waterbury. Biochemical composition and short term nutrient incorporation patterns in a unicellular marine cyanobacterium, Synechococcus (WH7803). Limnology and Oceanography, 29(2):370–374, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Carlton:1985:BCS

[CW85] Richard G. Carlton and Robert G. Wetzel. A box corer for studying metabolism of epipelic microorganisms in sediment under in situ conditions. Limnology and Oceanography, 30(2):422–426, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Carlton:1988:PFL

[CW88] Richard G. Carlton and Robert G. Wetzel. Phosphorus flux from lake sediments: Effect of epipelic algal oxygen production. Limnology and Oceanography, 33(4):562–570, July 1988. CODEN LI-OCAH. ISSN 0024-3590.

Carmack:1986:MIC

[CWD^{+}86] Eddy C. Carmack, Ronald C. Wiegand, Ralph J. Daley, Colin B. J. Gray, Steve Jasper, and Christopher H. Pharo. Mechanisms influencing the circulation and distribution of water mass in a medium residence-time lake. Limnology and Oceanography, 31(2): 249–265, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Corredor:1988:NRC

[CWV^{+}88] Jorge E. Corredor, Clive R. Wilkinson, Vance P. Vicente, Julio M. Morell, and Ernesto Otero. Nitrate release by Caribbean reef sponges. Limnology and Oceanography, 33(1):114–120, January 1988. CODEN LIOCAH. ISSN 0024-3590.

DAbramo:1980:IRD

[D’A80] Louis R. D’Abramo. Ingestion rate decrease as the stimulus for sexuality in populations of Moina macrocopa. Limnology and Oceanography, 25(3):422–429, May 1980. CODEN LIOCAH. ISSN 0024-3590.

deAngelis:1987:MSW

[dAL87] Marie A. de Angelis and Marvin D. Lilley. Methane in surface waters of Oregon estuaries and rivers. Limnology and Oceanography, 32(3):716–722, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Davison:1983:BRW

[Dav83] W. Davison. Book review: Whitfield, M., and D. Jagner [Eds.]. 1981. Marine electrochemistry. A practical introduction. John Wiley and Sons, Somerset, N.J. 529 p. $71.95. Limnology and Oceanography, 28(5):1046–1047, September 1983. CODEN LI-OCAH. ISSN 0024-3590.

Davison:1986:BRK

[Dav86] W. Davison. Book review: Kramer, C. J. M., and J. C. Duinker [Eds.]. 1984. Complexation of trace metals in natural waters. Developments in Biogeochemistry. Proc. Int. Symp., Texel, 1983. Martinus Nijhoff/Dr. W. Junk, Publishers, Dordrecht, The Netherlands. 472 p. (Typewriter composition.) $67.50. Limnology and Oceanography, 31(3):672, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Dubinsky:1981:LUP

[DB81] Zvy Dubinsky and Thomas Berman. Light utilization by phytoplankton in Lake Kinneret (Israel). Limnology and Oceanography, 26(4):660–670, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1983:TRF

[DC83] Stanley I. Dodson and Scott D. Cooper. Trophic relationships of the freshwater jellyfish Craspedacusta sowerbyi Lankester 1880. Limnology and Oceanography, 28(2):345–351, March 1983. CO-DEN LIOCAH. ISSN 0024-3590.

Dumont:1987:FPM

[DC87] Henri J. Dumont and Inez Carets. Flatworm predator (Mesostoma cf. lingua) releases a toxin to catch planktonic prey (Daphnia magna). Limnology and Oceanography, 32(3):699–702, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Davies-Colley:1988:MWC

[DC88a] Robert J. Davies-Colley. Measuring water clarity with a black disk. Limnology and Oceanography, 33(4):616–623, July 1988. CODEN LIOCAH. ISSN 0024-3590.

DElia:1988:MUZ

[DC88b] Christopher F. D’Elia and Clayton B. Cook. Methylamine uptake by zooxanthellae-invertebrate symbioses: Insights into host ammonium environment and nutrition. Limnology and Oceanography, 33(5):1153–1165, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Davies-Colley:1987:ALY

[DCV87] R. J. Davies-Colley and W. N. Vant. Absorption of light by yellow substance in freshwater lakes. Limnology and Oceanography, 32(2): 416–425, March 1987. CODEN LIOCAH. ISSN 0024-3590.

DeMeester:1989:PDI

[DD89] Luc De Meester and Henri J. Dumont. Phototaxis in Daphnia: Interaction of hunger and genotype. Limnology and Oceanography, 34(7):1322–1325, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Drenner:1982:SIF

[DdNK82] Ray W. Drenner, Frank de Noyelles, Jr., and Dean Kettle. Selective impact of filter-feeding gizzard shad on zooplankton community structure. Limnology and Oceanography, 27(5):965–968, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Durbin:1983:FLP

[DDSmV83] Edward G. Durbin, Ann G. Durbin, Thomas J. S-mayda, and Peter G. Verity. Food limitation of production by adult Acartia tonsa in Narragansett Bay, Rhode Island. Limnology and Oceanography, 28(6):1199–1213, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Deevey:1988:EDL

[Dee88] Edward S. Deevey, Jr. Estimation of downward leakage from Florida lakes. Limnology and Oceanography, 33(6):1308–1320, November 1988. CODEN LIOCAH. ISSN 0024-3590.

DeMott:1982:FSR

[DeM82] William R. DeMott. Feeding selectivities and relative ingestion rates of Daphnia and Bosmina. Limnology and Oceanography, 27 (3):518–527, May 1982. CODEN LIOCAH. ISSN 0024-3590.

DeMott:1988:DBA

[DeM88] William R. DeMott. Discrimination between algae and artificial particles by freshwater and marine copepods. Limnology and Oceanography, 33(3):397–408, May 1988. CODEN LIOCAH. ISSN 0024-3590.

DeMott:1989:OFT

[DeM89] William R. DeMott. Optimal foraging theory as a predictor of chemically mediated food selection by suspension-feeding copepods. Limnology and Oceanography, 34(1):140–154, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Denman:1982:BRF

[Den82a] Kenneth L. Denman. Book review: Fischer, H. B., E. J. List, R. C. Y. Koh, J. Imberger, and N. A. Brooks. 1979. Mixing in inland and coastal waters. Academic Press, Inc., New York, xiv + 483 p. $35.00. Limnology and Oceanography, 27(3):593–594, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Denny:1982:FIO

[Den82b] Mark W. Denny. Forces on intertidal organisms due to breaking ocean waves: Design and application of a telemetry system. Limnology and Oceanography, 27(1):178–183, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Denny:1983:SDR

[Den83] Mark W. Denny. A simple device for recording the maximum force exerted on intertidal organisms. Limnology and Oceanography, 28 (6):1269–1274, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Denny:1985:WFI

[Den85] Mark W. Denny. Wave forces on intertidal organisms: a case study. Limnology and Oceanography, 30(6):1171–1187, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Devol:1983:MOR

[Dev83] Allan H. Devol. Methane oxidation rates in the anaerobic sediments of Saanich Inlet. Limnology and Oceanography, 28(4):738– 742, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Davis:1982:SFM

[DF82a] Margaret Bryan Davis and Mary S. (Jesse) Ford. Sediment focusing in Mirror Lake, New Hampshire. Limnology and Oceanography, 27(1):137–150, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Dearing:1982:MSS

[DF82b] J. A. Dearing and R. J. Flower. The magnetic susceptibility of sedimenting material trapped in Lough Neagh, Northern Ireland, and its erosional significance. Limnology and Oceanography, 27(5): 969–975, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Dagg:1989:CDM

[DFW89] M. J. Dagg, B. W. Frost, and W. E. Walser, Jr. Copepod diel migration, feeding, and the vertical flux of pheopigments. Limnology and Oceanography, 34(6):1062–1071, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Dagg:1980:NFR

[DG80] M. J. Dagg and D. W. Grill. Natural feeding rates of Centropages typicus females in the New York Bight. Limnology and Oceanography, 25(4):597–609, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Denman:1983:TSS

[DG83] K. L. Denman and A. E. Gargett. Time and space scales of vertical mixing and advection of phytoplankton in the upper ocean. Limnology and Oceanography, 28(5):801–815, September 1983. CO-DEN LIOCAH. ISSN 0024-3590.

Davis:1982:LAE

[DGN82] Andrew O. Davis, James N. Galloway, and D. Kirk Nordstrom. Lake acidification: Its effect on lead in the sediment of two Adirondack lakes. Limnology and Oceanography, 27(1):163–167, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Davison:1980:DSF

[DH80] W. Davison and S. I. Heaney. Determination of the solubility of ferrous sulfide in a seasonally anoxic marine basin. Limnology and Oceanography, 25(1):153–156, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1985:GHB

[DH85a] Hugh W. Ducklow and Suzanne M. Hill. The growth of heterotrophic bacteria in the surface waters of warm core rings. Limnology and Oceanography, 30(2):239–259, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1985:TTI

[DH85b] Hugh W. Ducklow and Suzanne M. Hill. Tritiated thymidine incorporation and the growth of heterotrophic bacteria in warm core rings. Limnology and Oceanography, 30(2):260–272, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1988:IPE

[DH88] Stanley I. Dodson and John E. Havel. Indirect prey effects: Some morphological and life history responses of Daphnia pulex exposed to Notonecta undulata. Limnology and Oceanography, 33(6):1274– 1285, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Drenner:1987:ESS

[DHV^{+}87] Ray W. Drenner, K. David Hambright, Gary L. Vinyard, Moshe Gophen, and Utsa Pollingher. Experimental study of size-selective phytoplankton grazing by a filter-feeding cichlid and the cichlid’s effects on plankton community structure. Limnology and Oceanography, 32(5):1138–1144, September 1987. CODEN LIOCAH. ISSN 0024-3590.

deJonge:1989:PRC

[dJV89] V. N. de Jonge and L. A. Villerius. Possible role of carbonate dissolution in estuarine phosphate dynamics. Limnology and Oceanography, 34(2):332–340, March 1989. CODEN LIOCAH. ISSN 0024-

Duarte:1986:LSP

[DK86] Carlos M. Duarte and Jacob Kalff. Littoral slope as a predictor of the maximum biomass of submerged macrophyte communities. Limnology and Oceanography, 31(5):1072–1080, September 1986. CODEN LIOCAH. ISSN 0024-3590.

DeFrenza:1986:RSS

[DKMvL86] Joe DeFrenza, Robert J. Kirner, Edward J. Maly, and Henry C. van Leeuwen. The relationships of sex size ratio and season to mating intensity in some calanoid copepods. Limnology and Oceanography, 31(3):491–496, May 1986. CODEN LIOCAH. ISSN 0024-3590.

DiTullio:1983:EPU

[DL83a] Giacomo R. DiTullio and Edward A. Laws. Estimates of phytoplankton N uptake based on ^{14}CO_{2} incorporation into protein. Limnology and Oceanography, 28(1):177–185, January 1983. CO-DEN LIOCAH. ISSN 0024-3590.

Dodge:1983:ECH

[DL83b] Richard E. Dodge and Judith C. Lang. Environmental correlates of hermatypic coral (Montastrea annularis) growth on the East Flower Gardens Bank, northwest Gulf of Mexico. Limnology and Oceanography, 28(2):228–240, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Downs:1985:CPR

[DL85a] Jan Newton Downs and Carl J. Lorenzen. Carbon: pheopigment ratios of zooplankton fecal pellets as an index of herbivorous feeding. Limnology and Oceanography, 30(5):1024–1036, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Dymond:1985:FCB

[DL85b] Jack Dymond and Mitchell Lyle. Flux comparisons between sediments and sediment traps in the eastern tropical Pacific: Implications for atmospheric CO_{2} variations during the Pleistocene. Limnology and Oceanography, 30(4):699–712, July 1985. CODEN LIOCAH. ISSN 0024-3590.

DeFlaun:1983:RBB

[DM83] Mary F. DeFlaun and Lawrence M. Mayer. Relationships between bacteria and grain surfaces in intertidal sediments. Limnology and Oceanography, 28(5):873–881, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Davenport:1984:QGD

[DM84] James B. Davenport and Bassett Maguire, Jr. Quantitative grain density autoradiography and the intraspecific distribution of primary productivity in phytoplankton. Limnology and Oceanography, 29(2):410–416, March 1984. CODEN LIOCAH. ISSN 0024-3590.

David:1985:SCC

[DM85] Mark B. David and Myron J. Mitchell. Sulfur constituents and cycling in waters, seston, and sediments of an oligotrophic lake. Limnology and Oceanography, 30(6):1196–1207, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Drenner:1984:SPI

[DMdK84] Ray W. Drenner, John R. Mummert, Frank deNoyelles, Jr., and Dean Kettle. Selective particle ingestion by a filter-feeding fish and its impact on phytoplankton community structure. Limnology and Oceanography, 29(5):941–948, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Durako:1982:PMR

[DMM82] Michael J. Durako, Richard A. Medlyn, and Mark D. Moffler. Particulate matter resuspension via metabolically produced gas bubbles from benthic estuarine microalgae communities. Limnology and Oceanography, 26(5):752–756, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Dodson:1988:ERC

[Dod88] Stanley Dodson. The ecological role of chemical stimuli for the zooplankton: Predator-avoidance behavior in Daphnia. Limnology and Oceanography, 33(6 (part 2)):1431–1439, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Douglas:1987:IHC

[Dou87] A. E. Douglas. The influence of host contamination on maltose release by symbiotic Chlorella. Limnology and Oceanography, 32 (6):1363–1365, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Douglas:1988:IHC

[Dou88] A. E. Douglas. The influence of host contamination on maltose release by symbiotic Chlorella. Limnology and Oceanography, 33 (2):295–297, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Doyle:1980:LUI

[Doy80] Roger W. Doyle. On the logic of using individual animals. Limnology and Oceanography, 25(6):1145–1147, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Downing:1980:EBS

[DP80] John A. Downing and Robert H. Peters. The effect of body size and food concentration on the in situ filtering rate of Sida crystallina. Limnology and Oceanography, 25(5):883–895, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Dragcevic:1981:PSA

[DP81] Dj. Dragcevic and V. Pravdic. Properties of the seawater–air interface. 2. Rates of surface film formation under steady state conditions. Limnology and Oceanography, 26(3):492–499, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Dragcevic:1986:PSA

[DP86] Dj. Dragčević and V. Pravdić. Properties of the seawater–air interface. 3. The electrification and energy transfer through surface films. Limnology and Oceanography, 31(3):525–532, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Dortch:1989:BIU

[DP89a] Quay Dortch and James R. Postel. Biochemical indicators of N utilization by phytoplankton during upwelling off the Washington coast. Limnology and Oceanography, 34(4):758–773, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Dustan:1989:TIE

[DP89b] Phillip Dustan and James L. Pinckney, Jr. Tidally induced estuarine phytoplankton patchiness. Limnology and Oceanography, 34 (2):410–419, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Downing:1987:EIV

[DPF87] John A. Downing, Martin Pérusse, and Yves Frenette. Effect of interreplicate variance on zooplankton sampling design and data analysis. Limnology and Oceanography, 32(3):673–679, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Devol:1987:RGE

[DQRM87] Allan H. Devol, Paul D. Quay, Jeffrey E. Richey, and Luiz A. Martinelli. The role of gas exchange in the inorganic carbon, oxygen, and ^{222}Rn budgets of the Amazon River. Limnology and Oceanography, 32(1):235–248, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Downing:1988:SPL

[DR88] John A. Downing and Laurel C. Rath. Spatial patchiness in the lacustrine sedimentary environment. Limnology and Oceanography, 33(3):447–458, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Dring:1981:CAP

[Dri81] M. J. Dring. Chromatic adaptation of photosynthesis in benthic marine algae: an examination of its ecological significance using a theoretical model. Limnology and Oceanography, 26(2):271–284, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Dirks:1987:RTS

[DS87] R. W. J. Dirks and D. Spitzer. On the radiative transfer in the sea, including fluorescence and stratification effects. Limnology and Oceanography, 32(4):942–953, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Djogic:1986:CUV

[DSB86] Renata Djogić, Laszlo Sipos, and Marko Branica. Characterization of uranium(VI) in seawater. Limnology and Oceanography, 31(5): 1122–1131, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Davies:1982:GAI

[DSH82] Anthony G. Davies, Jillian A. Sleep, and D. S. Harbour. Germanic acid inhibition of carbon fixation in natural phytoplankton assemblages. Limnology and Oceanography, 27(2):357–361, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Demers:1987:RSS

[DTBB87] Serge Demers, Jean-Claude Therriault, Edwin Bourget, and Abdourrahmane Bah. Resuspension in the shallow sublittoral zone of a macrotidal estuarine environment: Wind influence. Limnology and Oceanography, 32(2):327–339, March 1987. CODEN LI-OCAH. ISSN 0024-3590.

Drenner:1989:IPF

[DTS^{+}89] Ray W. Drenner, Stephen T. Threlkeld, J. Durward Smith, John R. Mummert, and Patricia A. Cantrell. Interdependence of phosphorus, fish, and site effects on phytoplankton biomass and zooplankton. Limnology and Oceanography, 34(7):1315–1321, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1982:CBN

[Duc82] Hugh W. Ducklow. Chesapeake Bay nutrient and plankton dynamics. 1. Bacterial biomass and production during spring tidal destratification in the York River, Virginia, estuary. Limnology and Oceanography, 27(4):651–659, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Ducklow:1987:BRH

[Duc87] Hugh W. Ducklow. Book review: Heterotrophic activity in the sea (J. H. Hobbie and P. J. LeB. Williams [eds.]). Limnology and Oceanography, 32(2):522–523, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1980:BRB

[Dug80] Richard C. Dugdale. Book review: Boje, R., and M. Tomczak [Eds.] 1978. Upwelling ecosystems. Springer-Verlag, Berlin, Heidelberg, New York, x + 303 p. $27.00. (Typewriter composition). Limnology and Oceanography, 49(6):578, May 1980. CODEN LI-OCAH. ISSN 0024-3590.

Duggins:1981:SUK

[Dug81] David O. Duggins. Sea urchins and kelp: The effects of short term changes in urchin diet. Limnology and Oceanography, 26(2): 391–394, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Dagg:1986:EFC

[DW86a] Michael J. Dagg and W. Edward Walser, Jr. The effect of food concentration on fecal pellet size in marine copepods. Limnology and Oceanography, 31(5):1066–1071, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Dugdale:1986:UMN

[DW86b] R. C. Dugdale and F. P. Wilkerson. The use of ^{15}N to measure nitrogen uptake in eutrophic oceans; experimental considerations. Limnology and Oceanography, 31(4):673–689, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Dagg:1987:IGP

[DW87] Michael J. Dagg and W. Edward Walser, Jr. Ingestion, gut passage, and egestion by the copepod Neocalanus plumchrus in the laboratory and in the subarctic Pacific Ocean12. Limnology and Oceanography, 32(1):178–188, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Dustin:1986:LLM

[DWO86] Nina M. Dustin, Bruce H. Wilkinson, and Robert M. Owen. Littlefield Lake, Michigan: Carbonate budget of Holocene sedimentation in a temperate-region lacustrine system. Limnology and Oceanography, 31(6):1301–1311, November 1986. CODEN LI-OCAH. ISSN 0024-3590.

Davison:1982:DIM

[DWR82] W. Davison, C. Woof, and E. Rigg. The dynamics of iron and manganese in a seasonally anoxic lake; direct measurement of fluxes using sediment traps. Limnology and Oceanography, 27(6):987– 1003, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Edwards:1984:BRR

[EA84] John S. Edwards and Mark Andersen. Book review: Resh, V. H., and D. M. Rosenberg [Eds.] 1984. The ecology of aquatic insects. Praeger, N.Y. 1984. 625 p. $35.00. Limnology and Oceanography, 29(6):1350–1351, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Ebert:1981:EMG

[Ebe81] Thomas A. Ebert. Estimating mortality from growth parameters and a size distribution when recruitment is periodic. Limnology and Oceanography, 26(4):764–769, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Entsch:1983:PNC

[EBSW83] Barrie Entsch, Kevin G. Boto, Robin G. Sim, and John T. Wellington. Phosphorus and nitrogen in coral reef sediments. Limnology and Oceanography, 28(3):465–476, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Eckman:1983:HPA

[Eck83] James E. Eckman. Hydrodynamic processes affecting benthic recruitment. Limnology and Oceanography, 28(2):241–257, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Estep:1986:HIO

[EDKS86] Kenneth W. Estep, Paul G. Davis, Maureen D. Keller, and John McN. Sieburth. How important are oceanic algal nanoflagellates in bacterivory? Limnology and Oceanography, 31(3):646–650, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1980:SDC

[Edm80a] W. T. Edmondson. Secchi disk and chlorophyll. Limnology and Oceanography, 25(2):378–379, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1980:ECR

[Edm80b] Yvette H. Edmondson. Editorial comment: Rate of processing of manuscripts for publication. Limnology and Oceanography, 25(6): 1142, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1981:BRK

[Edm81] W. T. Edmondson. Book review: Koste, W. 1978. Rotatoria. Die Radertiere Mittel-europas, 2nd ed. Gebruder Borntraeger, Berlin and Stuttgart. V. 1, text, 673 p.; V. 2, plates, 476 p. DM238. Limnology and Oceanography, 26(2):400, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1982:BRD

[Edm82a] W. T. Edmondson. Book review: Dokulil, M., H. Metz, and D. Jewson [Eds.] 1980. Shallow lakes. Developments in Hydrobiology, v. 3. Dr. W. Junk Publ., The Hague. 218 p. $59.50. Limnology and Oceanography, 27(2):396, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1982:BRK

[Edm82b] W. T. Edmondson. Book review: Kiefer, F., and G. Fryer. 1978. Das Zooplankton der Binnengewässer. Teil 2. E. Schweizerbartsche Verlag, Stuttgart. 380 p. DM108. Limnology and Oceanography, 27(1):196, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1982:BRL

[Edm82c] W. T. Edmondson. Book review: Löffler, H. [Ed.] 1980. Neusiedlersee. The limnology of a shallow lake in Central Europe. Monogr. Biol. 37. Dr. W. Junk Publ., The Hague. 559 p. $102.65. Limnology and Oceanography, 27(1):198–199, January 1982. CO-DEN LIOCAH. ISSN 0024-3590.

Edmondson:1982:BRR

[Edm82d] W. T. Edmondson. Book review: Rotatoria. 1980. Proc. 2nd Int. Rotifer Symp., Gent, 1979. Developments in Hydrobiology, v. 1. H. J. Dumont and J. Green, Eds. Dr. W. Junk Publ., The Hague. [Reprinted from Hydrobiologia 73(1–3). 1980.] 268 p. $79.00. Limnology and Oceanography, 27(2):397–398, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1982:MFH

[Edm82e] Yvette H. Edmondson. In memoriam: [F. H. Rigler — 1928–1982 and R. Y. Stanier — 1916–1982]. Limnology and Oceanography, 27(5):986, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1984:BRM

[Edm84a] W. T. Edmondson. Book review: Margalef, R. 1983. Limnología. Ediciones Omega, S.A., Barcelona. 1010 p. Limnology and Oceanography, 29(6):1349, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1984:BRP

[Edm84b] W. T. Edmondson. Book review: Pourriot, R., J. Capblancq, P. Champ, and J. A. Meyer. 1982. Écologie du plancton des eaux continentales. Collection d’Ecologie 16, Masson, Paris and New York. 198 p. (Softcover). Limnology and Oceanography, 29(6): 1349–1350, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1984:ECS

[Edm84c] Yvette H. Edmondson. Editorial comment: Scope of the journal. Limnology and Oceanography, 29(2):440–441, March 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Edmondson:1984:ECN

[Edm84d] Yvette H. Edmondson. Editorial comment: The natural history of a manuscript. Limnology and Oceanography, 29(5):1145–1148, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1986:ECF

[Edm86a] Yvette H. Edmondson. Editorial comment: Final report. Limnology and Oceanography, 31(6):1401–1402, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1986:MLT

[Edm86b] Yvette H. Edmondson. In memoriam: [Livia Tonolli — 1909– 1985]. Limnology and Oceanography, 31(1):234, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1987:BRL

[Edm87] W. T. Edmondson. Book review: Lakes and reservoirs (F. B. Taub [ed.]). Limnology and Oceanography, 32(3):776–777, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Edwards:1987:SBF

[Edw87] R. T. Edwards. Sestonic bacteria as a food source for filtering invertebrates in two southeastern blackwater rivers. Limnology and Oceanography, 32(1):221–234, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Elser:1988:ZMT

[EEMC88] James J. Elser, Monica M. Elser, Neil A. MacKay, and Stephen R. Carpenter. Zooplankton-mediated transitions between n- and plimited algal growth. Limnology and Oceanography, 33(1):1–14, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Eshleman:1988:AMI

[EH88] Keith N. Eshleman and Harold F. Hemond. Alkalinity and major ion budgets for a Massachusetts reservoir and watershed. Limnology and Oceanography, 33(2):174–185, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Ertel:1986:DHS

[EHD^{+}86] John R. Ertel, John I. Hedges, Allan H. Devol, Jeffrey E. Richey, and Maria de Nazaré Góes Ribeiro. Dissolved humic substances of the Amazon River system. Limnology and Oceanography, 31(4): 739–754, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Exton:1983:SDT

[EHE^{+}83] R. J. Exton, W. M. Houghton, W. Esaias, L. W. Haas, and D. Hayward. Spectral differences and temporal stability of phycoerythrin fluorescence in estuarine and coastal waters due to the domination of labile cryptophytes and stabile cyanobacteria. Limnology and Oceanography, 28(6):1225–1231, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Evans:1986:RSD

[EJ86] Marlene S. Evans and David J. Jude. Recent shifts in Daphnia community structure in southeastern Lake Michigan: a comparison of the inshore and offshore regions. Limnology and Oceanography, 31(1):56–67, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Edmondson:1981:ECN

[EL81] W. T. Edmondson and John T. Lehman. The effect of changes in the nutrient income on the condition of Lake Washington. Limnology and Oceanography, 26(1):1–29, January 1981. CODEN LI-OCAH. ISSN 0024-3590. See comment [CR83].

Edmondson:1982:DLW

[EL82] W. T. Edmondson and Ami H. Litt. Daphnia in Lake Washington. Limnology and Oceanography, 27(2):272–293, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Elmore:1983:FID

[Elm83] J. L. Elmore. Factors influencing Diaptomus distributions: an experimental study in subtropical Florida. Limnology and Oceanography, 28(3):522–532, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Eppley:1982:BRM

[Epp82] Richard W. Eppley. Book review: Morris, I. [Ed.] 1981. The physiological ecology of phytoplankton. Studies in ecology, v. 7. Univ. California Press, Berkeley and New York. 625 p. $74.50. Limnology and Oceanography, 27(3):592–593, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Ebert:1988:LVS

[ER88] Thomas A. Ebert and Michael P. Russell. Latitudinal variation in size structure of the west coast purple sea urchin: a correlation with headlands. Limnology and Oceanography, 33(2):286–294, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Ellis:1982:CPC

[ES82] Bonnie K. Ellis and Jack A. Stanford. Comparative photoheterotrophy, chemoheterotrophy, and photolithotrophy in a eutrophic reservoir and an oligotrophic lake. Limnology and Oceanography, 27(3):440–454, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Einarsson:1983:SAL

[ES83] Stefan Einarsson and Unnsteinn Stefánsson. The sources of alkalinity in Lake Miklavatn, north Iceland. Limnology and Oceanography, 28(1):50–57, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Evans:1980:PAL

[ET80] Geoffrey T. Evans and F. J. R. Taylor. Phytoplankton accumulation in Langmuir cells. Limnology and Oceanography, 25(5):840– 845, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Evans:1988:FDS

[Eva88] Geoffrey T. Evans. A framework for discussing seasonal succession and coexistence of phytoplankton species. Limnology and Oceanography, 33(5):1027–1036, September 1988. CODEN LI-OCAH. ISSN 0024-3590.

Farrell:1988:BRB

[Far88] Terence M. Farrell. Book review: Barnacle biology (A. E. Southward [ed.]). Limnology and Oceanography, 33(6 (part 2)):1651, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Fallon:1980:PBG

[FB80] Robert D. Fallon and Thomas D. Brock. Planktonic blue-green algae: Production, sedimentation, and decomposition in Lake Mendota, Wisconsin. Limnology and Oceanography, 25(1):72–88, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Fraleigh:1988:MPC

[FB88] Peter C. Fraleigh and Jeffrey C. Burnham. Myxococcal predation on cyanobacterial populations: Nutrient effects. Limnology and Oceanography, 33(3):476–483, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Fisher:1983:IMP

[FBF83] Nicholas S. Fisher, Poul Bjerregaard, and Scott W. Fowler. Interactions of marine plankton with transuranic elements. 1. Biokinetics of neptunium, plutonium, americium, and californium in phytoplankton. Limnology and Oceanography, 28(3):432–447, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Frechette:1989:IBL

[FBG89] Marcel Fréchette, Cheryl Ann Butman, and W. Rockwell Geyer. The importance of boundary-layer flows in supplying phytoplankton to the benthic suspension feeder, Mytilus edulis L. Limnology and Oceanography, 34(1):19–36, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Findlay:1986:BGM

[FCC^{+}86] Stuart Findlay, Les Carlough, M. Tad Crocker, H. Kay Gill, Judy L. Meyer, and Phillip J. Smith. Bacterial growth on macrophyte leachate and fate of bacterial production. Limnology and Oceanography, 31(6):1335–1341, November 1986. CODEN LI-OCAH. ISSN 0024-3590.

Farmer:1987:TVM

[FCO87] David D. Farmer, Gregory B. Crawford, and Thomas R. Osborn. Temperature and velocity microstructure caused by swimming fish. Limnology and Oceanography, 32(4):978–983, July 1987. CO-DEN LIOCAH. ISSN 0024-3590.

Forward:1984:CDV

[FCS84] Richard B. Forward, Jr., Thomas W. Cronin, and Donald E. Stearns. Control of diel vertical migration: Photoresponses of a larval crustacean. Limnology and Oceanography, 29(1):146–154, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Fuhrman:1986:DAI

[FDK^{+}86] Jed A. Fuhrman, Hugh W. Ducklow, David L. Kirchman, John Hudak, George B. McManus, and Jonathan Kramer. Does adenine incorporation into nucleic acids measure total microbial production? Limnology and Oceanography, 31(3):627–636, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Fuhrman:1986:ATM

[FDKM86] Jed A. Fuhrman, Hugh W. Ducklow, David L. Kirchman, and George B. McManus. Adenine and total microbial production: a reply. Limnology and Oceanography, 31(6):1395–1400, November 1986. CODEN LIOCAH. ISSN 0024-3590. See [KW86].

Forsberg:1988:FCN

[FDR^{+}88] Bruce R. Forsberg, Allan H. Devol, Jeffrey E. Richey, Luiz A. Martinelli, and Humberto dos Santos. Factors controlling nutrient concentrations in Amazon floodplain lakes. Limnology and Oceanography, 33(1):41–56, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Falkowski:1985:GIR

[FDW85] Paul G. Falkowski, Zvy Dubinsky, and Kevin Wyman. Growthirradiance relationships in phytoplankton. Limnology and Oceanography, 30(2):311–321, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Fontaine:1981:MFL

[FE81] Thomas D. Fontaine III and Katherine Carter Ewel. Metabolism of a Florida lake ecosystem. Limnology and Oceanography, 26(4): 754–763, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Fee:1979:RBL

[Fee79] Everett J. Fee. A relation between lake morphometry and primary productivity and its use in interpreting whole-lake eutrophication experiments. Limnology and Oceanography, 24(3):401–416, May 1979. CODEN LIOCAH. ISSN 0024-3590. See comment [Pat80].

Fee:1980:RCP

[Fee80] Everett Fee. Reply to comments by Patalas and Schindler. Limnology and Oceanography, 25(6):1152–1153, November 1980. CODEN LIOCAH. ISSN 0024-3590. See [Pat80, Sch80].

Feigenbaum:1982:FCS

[Fei82] David Feigenbaum. Feeding by the chaetognath, Sagitta elegans, at low temperatures in Vineyard Sound, Massachusetts. Limnology and Oceanography, 26(5):699–706, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Fenchel:1980:BRJ

[Fen80a] Tom Fenchel. Book review: Jefferies, R. L., and A. J. Davy [Eds.] 1979. Ecological processes in coastal environments. Blackwell Sci. Publ., Oxford, xvii + 684 p. Limnology and Oceanography, 25(5): 968, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Fenchel:1980:BRP

[Fen80b] Tom Fenchel. Book review: Phillips, R. C., and C. P. McRoy [Eds.] 1980. Handbook of seagrass biology: an ecosystem perspective. Garland STPM Press, New York and London, xiii + 353 p. Limnology and Oceanography, 49(6):579–580, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Fenchel:1980:RBP

[Fen80c] Tom Fenchel. Relation between particle size selection and clearance in suspension-feeding ciliates. Limnology and Oceanography, 25(4):733–738, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Fenchel:1982:BRE

[Fen82] Tom Fenchel. Book review: Estuarine perspectives. 1980. Proceedings of the Fifth Biennial International Estuarine Research Conference held at Jekyll Island, Georgia, 1979. V. S. Kennedy, Ed. Academic Press, New York and London. 533 p. Limnology and Oceanography, 27(2):394–395, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Fenchel:1983:BRN

[Fen83] T. M. Fenchel. Book review: Nedwell, D. B., and C. M. Brown [Eds.]. 1982. Sediment microbiology. Spec. Publ. 7 Soc. Gen. Microbiol. Academic Press Inc. (London) Ltd. 234 p. £14.20. (Typewriter composition). Limnology and Oceanography, 28(5):1041– 1042, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Fenchel:1984:BRM

[Fen84] Tom Fenchel. Book review: Mann, K. H. 1982. Ecology of coastal waters — a systems approach. Studies in Ecology, V. 8. Univ. California Press, Berkeley. 322 p. $36.00 hardcover; $18.00 softcover. Limnology and Oceanography, 29(2):446–447, March 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Fenchel:1987:MEP

[Fen87] Tom Fenchel. Marine ecological processes. Limnology and Oceanography, 32(3):778, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Freeman:1985:PFS

[FF85] Byron J. Freeman and Mary C. Freeman. Production of fishes in a subtropical blackwater ecosystem: The Okefenokee Swamp. Limnology and Oceanography, 30(3):686–692, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Francko:1982:USC

[FH82] David A. Francko and Robert T. Heath. UV-sensitive complex phosphorus: Association with dissolved humic material and iron in a bog lake. Limnology and Oceanography, 27(3):564–569, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Filbin:1984:ELP

[FH84] Gerald J. Filbin and R. Anton Hough. Extraction of ^{14}Clabeled photosynthate from aquatic plants with dimethyl sulfoxide (DMSO). Limnology and Oceanography, 29(2):426–428, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Filbin:1985:PPP

[FH85] Gerald J. Filbin and R. Anton Hough. Photosynthesis, photorespiration, and productivity in Lemna minor L. Limnology and Oceanography, 30(2):322–334, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Fallon:1980:RMI

[FHHB80] R. D. Fallon, S. Harrits, R. S. Hanson, and T. D. Brock. The role of methane in internal carbon cycling in Lake Mendota during summer stratification. Limnology and Oceanography, 25(2):357– 360, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Fiedler:1982:ZAR

[Fie82] Paul C. Fiedler. Zooplankton avoidance and reduced grazing responses to Gymnodinium splendens (Dinophyceae). Limnology and Oceanography, 27(5):961–965, September 1982. CODEN LI-OCAH. ISSN 0024-3590.

Fischer:1980:MPA

[Fis80] Hugo B. Fischer. Mixing processes on the Atlantic continental shelf, Cape Cod to Cape Hatteras. Limnology and Oceanography, 25(1):114–125, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Fisher:1986:RMM

[Fis86] Nicholas S. Fisher. On the reactivity of metals for marine phytoplankton. Limnology and Oceanography, 31(2):443–449, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Fitzwater:1982:MCE

[FKM82] Steve E. Fitzwater, George A. Knauer, and John H. Martin. Metal contamination and its effect on primary production measurements. Limnology and Oceanography, 27(3):544–551, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Fleischmann:1989:MPU

[Fle89] Esther M. Fleischmann. The measurement and penetration of ultraviolet radiation into tropical marine water. Limnology and Oceanography, 34(8):1623–1629, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Flynn:1988:CPP

[Fly88] Kevin J. Flynn. The concept of “primary production” in aquatic ecology. Limnology and Oceanography, 33(5):1215–1216, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Foster:1982:RCT

[FM82] Patricia L. Foster and Francois M. M. Morel. Reversal of cadmium toxicity in a diatom: an interaction between cadmium activity and iron. Limnology and Oceanography, 26(5):745–752, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Faust:1985:VSA

[FN85] Maria A. Faust and Karl H. Norris. In vivo spectrophotometric analysis of photosynthetic pigments in natural populations of phytoplankton. Limnology and Oceanography, 30(6):1316–1322, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Fernandez:1985:RSA

[FNL85] J. A. Fernández, F. X. Niell, and J. Lucena. A rapid and sensitive automated determination of phosphate in natural waters. Limnology and Oceanography, 30(1):227–230, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Falkowski:1982:TEP

[FO82] Paul G. Falkowski and Thomas G. Owens. A technique for estimating phytoplankton division rates by using a DNA-binding fluorescent dye. Limnology and Oceanography, 26(5):776–782, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Fisher:1980:PUM

[FOB80] Nicholas S. Fisher, Brenda L. Olson, and Vaughan T. Bowens. Plutonium uptake by marine phytoplankton in culture. Limnology and Oceanography, 25(5):823–839, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Forsberg:1985:FPP

[For85] Bruce R. Forsberg. The fate of planktonic primary production. Limnology and Oceanography, 30(4):807–819, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Fowler:1980:BRA

[Fow80] Scott W. Fowler. Book review: Ancellix, J., P. Guegueniat, and P. Germain. 1979. Radioecologie marine (etude du de-venir des radionucleides rejetes en milieu marin et applications a la radioprotection). Éditions Eyrolles, Paris, xiv + 256 p. 125 F. Limnology and Oceanography, 25(2):386–387, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Frape:1981:CIW

[FP81] S. K. Frape and R. J. Patterson. Chemistry of interstitial water and bottom sediments as indicators of seepage patterns in Perch Lake, Chalk River, Ontario. Limnology and Oceanography, 26(3): 500–517, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Fletcher:1982:SDS

[FP82] W. K. Fletcher and D. Poison. A simple device for sampling shallow waters. Limnology and Oceanography, 27(1):188–190, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Fulton:1987:ECM

[FP87] Rolland S. Fulton III and Hans W. Paerl. Effects of colonial morphology on zooplankton utilization of algal resources during blue-green algal (Microcystis aeruginosa) blooms. Limnology and Oceanography, 32(3):634–644, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Francois:1987:SEC

[Fra87] R. Francois. A study of the extraction conditions of sedimentary humic acids to estimate their true in situ sulfur content. Limnology and Oceanography, 32(4):964–972, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Frankignoulle:1988:FMA

[Fra88] M. Frankignoulle. Field measurements of air-sea CO_{2} exchange. Limnology and Oceanography, 33(3):313–322, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Frey:1984:BRG

[Fre84] David G. Frey. Book review: Goldman, C. R., and A. J. Horne. 1983. Limnology. McGraw-Hill Book Co., New York. 464 p. $31.95. Limnology and Oceanography, 29(2):447, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Frey:1988:AWS

[Fre88] David G. Frey. Alona weinecki Studer on the subantarctic islands, not Alona rectangula Sars (Chydoridae, Cladocera). Limnology and Oceanography, 33(6):1386–1411, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Froelich:1980:AOC

[Fro80] P. N. Froelich. Analysis of organic carbon in marine sediments. Limnology and Oceanography, 49(6):564–572, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Froelich:1988:KCD

[Fro88] Philip N. Froelich. Kinetic control of dissolved phosphate in natural rivers and estuaries: a primer on the phosphate buffer mechanism. Limnology and Oceanography, 33(4 (part 2)):649–668, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Fryer:1980:BRA

[Fry80] Geoffrey Fryer. Book review: Allanson, B. R. [Ed.] 1979. Lake Sibaya. Monographs in Biology 36. Dr. W. Junk Publ., The Hague, x + 364 p. $68.45. Limnology and Oceanography, 25(5):965–966, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Fry:1986:SCS

[Fry86] Brian Fry. Sources of carbon and sulfur nutrition for consumers in three meromictic lakes of New York State. Limnology and Oceanography, 31(1):79–88, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Fry:1988:FWS

[Fry88] Brian Fry. Food web structure on Georges Bank from stable C, N, and S isotopic compositions. Limnology and Oceanography, 33(5): 1182–1190, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Fischer:1983:OTS

[FS83] Hugo B. Fischer and Robert D. Smith. Observations of transport to surface waters from a plunging inflow to Lake Mead. Limnology and Oceanography, 28(2):258–272, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Ferguson:1984:UAA

[FS84] Randolph L. Ferguson and William G. Sunda. Utilization of amino acids by planktonic marine bacteria: Importance of clean technique and low substrate additions. Limnology and Oceanography, 29(2):258–274, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Fox:1985:FCC

[FSW85] Lewis E. Fox, Shawn L. Sager, and Steven C. Wofsy. Factors controlling the concentrations of soluble phosphorus in the Mississippi estuary. Limnology and Oceanography, 30(4):826–832, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Fisher:1987:ATP

[FTKB87] N. S. Fisher, J.-L. Teyssié, S. Krishnaswami, and M. Baskaran. Accumulation of Th, Pb, U, and Ra in marine phytoplankton and its geochemical significance. Limnology and Oceanography, 32(1): 131–142, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Furnas:1987:EPP

[Fur87] Miles J. Furnas. Effects of prescreening on productivity of sizefractionated phytoplankton. Limnology and Oceanography, 32(2): 483–491, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Fry:1985:MMM

[FV85] E. S. Fry and K. J. Voss. Measurement of the Mueller matrix for phytoplankton. Limnology and Oceanography, 30(6):1322–1326, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Fournier:1984:PFC

[FVH^{+}84] R. O. Fournier, M. Van Det, N. B. Hargreaves, J. S. Wilson, T. A. Clair, and R. Ernst. Physical factors controlling summer distribution of chlorophyll-a off southwestern Nova Scotia. Limnology and Oceanography, 29(3):517–526, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Fernendez:1982:DCH

[FVM82] Hugo Fernendez, Felipe Vazquez, and Frank J. Millero. The density and composition of hypersaline waters of a Mexican lagoon. Limnology and Oceanography, 27(2):315–321, March 1982. CO-DEN LIOCAH. ISSN 0024-3590.

Francko:1982:ICA

[FW82] David A. Francko and Robert G. Wetzel. The isolation of cyclic adenosine 3’:5’-monophosphate (cAMP) from lakes of differing trophic status: Correlation with planktonic metabolic variables. Limnology and Oceanography, 27(1):27–38, January 1982. CO-DEN LIOCAH. ISSN 0024-3590.

Gordon:1981:SMP

[GA81] Arnold L. Gordon and Frank Aikman III. Salinity maximum in the pycnocline of the Middle Atlantic Bight. Limnology and Oceanography, 26(1):123–130, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Garside:1980:BRD

[Gar80] Chris Garside. Book review: Dame, R. F. [Ed.] 1979. Marshestuarine systems simulation. The Belle W. Baruch Library in Marine Science No. 8. University of South Carolina Press, Columbia, xii + 260 p. $27.50. Limnology and Oceanography, 49(6):578, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Garrett:1981:COP

[Gar81a] William D. Garrett. Comment on “Organic particle and aggregate formation resulting from the dissolution of bubbles in seawater” (Johnson and Cooke). Limnology and Oceanography, 27(4):989– 992, September 1981. CODEN LIOCAH. ISSN 0024-3590. See reply [JC81].

Garside:1981:NAU

[Gar81b] C. Garside. Nitrate and ammonia uptake in the apex of the New York Bight. Limnology and Oceanography, 26(4):731–739, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Garside:1984:AUK

[Gar84] C. Garside. Apparent ^{15}N uptake kinetics resulting from remineralization. Limnology and Oceanography, 29(1):204–210, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1980:ENS

[GB80] Joel C. Goldman and Peter G. Brewer. Effect of nitrogen source and growth rate on phytoplankton-mediated changes in alkalinity. Limnology and Oceanography, 25(2):352–357, March 1980. CO-DEN LIOCAH. ISSN 0024-3590.

Grant:1981:PIC

[GB81] J. W. G. Grant and I. A. E. Bayly. Predator induction of crests in morphs of the Daphnia carinata King complex. Limnology and Oceanography, 26(2):201–218, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Goloway:1982:DMI

[GB82] Frances Goloway and Michael Bender. Diagenetic models of interstitial nitrate profiles in deep sea suboxic sediments. Limnology and Oceanography, 27(4):624–638, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Gerritsen:1987:EMN

[GB87] Jeroen Gerritsen and Susan W. Bradley. Electrophoretic mobility of natural particles and cultured organisms in freshwaters. Limnology and Oceanography, 32(5):1049–1058, September 1987. CO-DEN LIOCAH. ISSN 0024-3590.

Gifford:1981:SDD

[GBB81] Dian J. Gifford, Richard N. Bohrer, and Carl M. Boyd. Spines on diatoms: Do copepods care? Limnology and Oceanography, 26(6): 1057–1061, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Gould:1986:MSA

[GBF86] R. W. Gould, Jr., E. R. Balmori, and G. A. Fryxell. Multivariate statistics applied to phytoplankton data from two Gulf Stream warm core rings. Limnology and Oceanography, 31(5):951–968, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Geddes:1981:VSS

[GC81] Michael C. Geddes and Gerald A. Cole. Variation in sexual size differentiation in North American diaptomids (Copepoda: Calanoida): Does variation in the degree of dimorphism have ecological significance? Limnology and Oceanography, 26(2):367–374, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Griffith:1987:CAF

[GCAN87] Peter C. Griffith, John D. Cubit, Walter H. Adey, and James N. Norris. Computer-automated flow respirometry: Metabolism measurements on a Caribbean reef flat and in a microcosm. Limnology and Oceanography, 32(2):442–451, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1987:RGG

[GCD87] Joel C. Goldman, David A. Caron, and Mark R. Dennett. Regulation of gross growth efficiency and ammonium regeneration in bacteria by substrate c: N ratio. Limnology and Oceanography, 32(6):1239–1252, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1989:ONM

[GCL89] Wayne S. Gardner, Joann F. Chandler, and Gwenyth A. Laird. Organic nitrogen mineralization and substrate limitation of bacteria in Lake Michigan. Limnology and Oceanography, 34(2):478–485, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1987:SFD

[GCLC87] Wayne S. Gardner, Joann F. Chandler, Gwenyth A. Laird, and Hunter J. Carrick. Sources and fate of dissolved free amino acids in epilimnetic Lake Michigan water. Limnology and Oceanography, 32(6):1353–1362, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Gorham:1983:CCL

[GDS83] Eville Gorham, Walter E. Dean, and Jon E. Sanger. The chemical composition of lakes in the north-central United States. Limnology and Oceanography, 28(2):287–301, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Geller:1986:CME

[Gel86] Annette Geller. Comparison of mechanisms enhancing biodegradability of refractory lake water constituents. Limnology and Oceanography, 31(4):755–764, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Gerritsen:1982:BRD

[Ger82] Jeroen Gerritsen. Behavioral response of Daphnia to rate of temperature change: Possible enhancement of vertical migration. Limnology and Oceanography, 27(2):254–261, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Gershey:1983:BAD

[Ger83a] Robert M. Gershey. A bubble adsorption device for the isolation of surface-active organic matter in seawater. Limnology and Oceanography, 28(2):395–400, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Gershey:1983:CSO

[Ger83b] Robert M. Gershey. Characterization of seawater organic matter carried by bubble-generated aerosols. Limnology and Oceanography, 28(2):309–319, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Geyer:1989:GCE

[Gey89] W. Rockwell Geyer. Gravity currents: In the environment and the laboratory. Limnology and Oceanography, 34(1):264, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1985:MLA

[GFCP85] Wayne S. Gardner, William A. Frez, Elizabeth A. Cichocki, and Christopher C. Parrish. Micromethod for lipids in aquatic invertebrates. Limnology and Oceanography, 30(5):1099–1105, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Gough:1981:CPP

[GG81] Stephen B. Gough and Larry P. Gough. Comment on “Primary production of algae growing on natural and artificial aquatic plants: a study of interactions between epiphytes and their substrate” (Cattaneo and Kalff). Limnology and Oceanography, 27(4): 987–988, September 1981. CODEN LIOCAH. ISSN 0024-3590. See reply [CK81].

Goldman:1982:CRA

[GG82] Joel C. Goldman and Patricia M. Glibert. Comparative rapid ammonium uptake by four species of marine phytoplankton. Limnology and Oceanography, 27(5):814–827, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Garside:1984:CMU

[GG84] C. Garside and P. M. Glibert. Computer modeling of ^{15}N uptake and remineralization experiments. Limnology and Oceanography, 29(1):199–203, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Giblin:1984:PED

[GH84] Anne E. Giblin and Robert W. Howarth. Porewater evidence for a dynamic sedimentary iron cycle in salt marshes. Limnology and Oceanography, 29(1):47–63, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Govoni:1989:SDL

[GHC89] John J. Govoni, Donald E. Hoss, and David R. Colby. The spatial distribution of larvalfishes about the Mississippi River plume. Limnology and Oceanography, 34(1):178–187, January 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Gallegos:1980:PLR

[GHK80] Charles L. Gallegos, George M. Hornberger, and Mahlon G. Kelly. Photosynthesis — light relationships of a mixed culture of phytoplankton in fluctuating light. Limnology and Oceanography, 25(6): 1082–1092, November 1980. CODEN LIOCAH. ISSN 0024-3590.

George:1984:SPS

[GHW84] D. G. George, M. A. Hurley, and B. Winstanley. A simple plankton splitter with i note on its reduced subsampling variance. Limnology and Oceanography, 29(2):429–433, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Gieskes:1982:PSS

[Gie82] J. M. Gieskes. The practical salinity scale 1978: a reply to comments by T. R. Parsons. Limnology and Oceanography, 27(2): 387–389, March 1982. CODEN LIOCAH. ISSN 0024-3590. See [Par82].

Giguere:1980:MEC

[Gig80] Louis A. Giguère. Metabolic expenditures in Chaoborus larvae. Limnology and Oceanography, 25(5):922–928, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1980:BRP

[Gil80] John J. Gilbert. Book review: Pennak, R. W. 1978. Fresh-water invertebrates of the United States, 2nd ed. John Wiley & Sons, New York, xv + 803 p. $27.50. Limnology and Oceanography, 25 (2):383, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1988:SRP

[Gil88] John J. Gilbert. Suppression of rotifer populations by Daphnia: a review of the evidence, the mechanisms, and the effects on zooplankton community structure. Limnology and Oceanography, 33 (6):1286–1303, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1989:EDI

[Gil89] John J. Gilbert. The effect of Daphnia interference on a natural rotifer and ciliate community: Short-term bottle experiments. Limnology and Oceanography, 34(3):606–617, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1989:IFP

[GJP89] Charles R. Goldman, Alan Jassby, and Thomas Powell. Interannual fluctuations in primary production: Meteorological forcing at two subalpine lakes. Limnology and Oceanography, 34(2):310–323, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Gieskes:1983:UCD

[GK83] Winfried W. Gieskes and Gijsbert W. Kraay. Unknown chlorophyll-a derivatives in the North Sea and the tropical Atlantic Ocean revealed by HPLC analysis. Limnology and Oceanography, 28(4):757–766, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1988:ERR

[GK88] John J. Gilbert and Kevin L. Kirk. Escape response of the rotifer Keratella: Description, stimulation, fluid dynamics, and ecological significance. Limnology and Oceanography, 33(6 (part 2)):1440– 1450, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Ganor:1989:GEH

[GK89] Jiwchar Ganor and Amitai Katz. The geochemical evolution of halite structures in hypersaline lakes: The Dead Sea, Israel. Limnology and Oceanography, 34(7):1214–1223, November 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Gagnon:1981:ZSV

[GL81] Marc Gagnon and Guy Lacroix. Zooplankton sample variability in a tidal estuary: an interpretative model. Limnology and Oceanography, 26(3):401–413, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1982:IDM

[GLMA82] Patricia M. Gilbert, Fredric Lipschultz, James J. McCarthy, and Mark A. Altabet. Isotope dilution models of uptake and remineralization of ammonium by marine plankton. Limnology and Oceanography, 27(4):639–650, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1985:MVI

[GLMA85] P. M. Gilbert, F. Lipschultz, J. J. McCarthy, and M. A. Altabet. Has the mystery of the vanishing ^{15}N in isotope dilution experiments been resolved? Limnology and Oceanography, 30(2): 444–447, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Gruebel:1984:RTS

[GM84] Karen A. Gruebel and Christopher S. Martens. Radon-222 tracing of sediment-water chemical transport in an estuarine sediment. Limnology and Oceanography, 29(3):587–597, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Gachter:1985:DSS

[GM85] René Gächter and Antonin Mares. Does settling seston release soluble reactive phosphorus in the hypolimnion of lakes? Limnology and Oceanography, 30(2):364–371, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Grad:1988:SSR

[GM88] Gabriella Grad and Edward J. Maly. Sex size ratios and their influence on mating success in a calanoid copepod. Limnology and Oceanography, 33(6 (part 2)):1629–1634, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Guillard:1985:SSL

[GMFLJ85] R. R. L. Guillard, L. S. Murphy, P. Foss, and S. Liaaen-Jensen. Synechococcus spp. as likely zeaxanthin-dominant ultraphytoplankton in the North Atlantic. Limnology and Oceanography, 30(2):412–414, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Gloss:1980:ACN

[GMK80] Steven P. Gloss, Lawrence M. Mayer, and David E. Kidd. Advective control of nutrient dynamics in the epilimnion of a large reservoir. Limnology and Oceanography, 25(2):219–228, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Gachter:1988:CBR

[GMM88] René Gächter, Joseph S. Meyer, and Antonin Mares. Contribution of bacteria to release and fixation of phosphorus in lake sediments. Limnology and Oceanography, 33(6 (part 2)):1542–1558, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Guerrero:1985:PSB

[GMPA^{+}85] R. Guerrero, E. Montesinos, C. Pedrós-Alió, I. Esteve, J. Mas, H. van Gemerden, P. A. G. Hofman, and J. F. Bakker. Phototrophic sulfur bacteria in two Spanish lakes: Vertical distribution and limiting factors. Limnology and Oceanography, 30(5): 919–931, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1987:NMD

[GNM87] Wayne S. Gardner, Thomas F. Nalepa, and John M. Malczyk. Nitrogen mineralization and denitrification in Lake Michigan sediments. Limnology and Oceanography, 32(6):1226–1238, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1986:PGR

[Gol86] Joel C. Goldman. On phytoplankton growth rates and particulate c: N: P ratios at low light. Limnology and Oceanography, 31(6): 1358–1363, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Goldman:1988:PPN

[Gol88] Charles R. Goldman. Primary productivity, nutrients, and transparency during the early onset of eutrophication in ultraoligotrophic Lake Tahoe, California–Nevada. Limnology and Oceanography, 33(6):1321–1333, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Gophen:1989:BRA

[Gop89] Moshe Gophen. Book review: Autotrophic bacteria (H. G. Schlegel and B. Bowien). Limnology and Oceanography, 34(6): 1149, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Gordon:1989:CLB

[Gor89a] Howard R. Gordon. Can the Lambert-beer law be applied to the diffuse attenuation coefficient of ocean water? Limnology and Oceanography, 34(8):1389–1409, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Gordon:1989:DDR

[Gor89b] Howard R. Gordon. Dependence of the diffuse reflectance of natural waters on the sun angle. Limnology and Oceanography, 34(8): 1484–1489, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Gallegos:1983:PPA

[GPHI83] Charles L. Gallegos, Trevor Platt, W. G. Harrison, and Brian Irwin. Photosynthetic parameters of Arctic marine phytoplankton: Vertical variations and time scales of adaptation. Limnology and Oceanography, 28(4):698–708, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Guan:1985:ACM

[GPS85] Fumin Guan, José Peláez, and Robert H. Stewart. The atmospheric correction and measurement of chlorophyll concentration using the coastal zone color scanner. Limnology and Oceanography, 30(2):273–285, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Gliwicz:1984:SCA

[GR84] Z. Maciej Gliwicz and Marc G. Rowan. Survival of Cyclops abyssorum tatricus (Copepoda, Crustacea) in alpine lakes stocked with planktivorous fish. Limnology and Oceanography, 29(6):1290– 1299, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Griffith:1988:HPR

[Gri88] Peter C. Griffith. A high-precision respirometer for measuring small rates of change in the oxygen concentration of natural waters. Limnology and Oceanography, 33(4):632–638, July 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Grover:1989:ESP

[Gro89a] James P. Grover. Effects of Si : P supply ratio, supply variability, and selective grazing in the plankton: an experiment with a natural algal and protistan assemblage. Limnology and Oceanography, 34(2):349–367, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Grover:1989:PDG

[Gro89b] James P. Grover. Phosphorus-dependent growth kinetics of 11 species of freshwater algae. Limnology and Oceanography, 34(2): 341–348, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Gardner:1981:KEN

[GS81] Wayne S. Gardner and Donald Scavia. Kinetic examination of nitrogen release by zooplankters. Limnology and Oceanography, 26(5):801–810, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Gilbert:1984:AIP

[GS84] John J. Gilbert and Richard S. Stemberger. Asplanchna-induced polymorphism in the rotifer Keratella slacki. Limnology and Oceanography, 29(6):1309–1316, November 1984. CODEN LI-OCAH. ISSN 0024-3590.

Gilbert:1985:CKP

[GS85] John J. Gilbert and Richard S. Slemberger. Control of Keratella populations by interference competition from Daphnia. Limnology and Oceanography, 30(1):180–188, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Gliwicz:1986:FAD

[GS86] Maciej Z. Gliwicz and Agnieszka Sieniawska. Filtering activity of Daphnia in low concentrations of a pesticide. Limnology and Oceanography, 31(5):1132–1138, September 1986. CODEN LI-OCAH. ISSN 0024-3590.

Gavrieli:1989:SHF

[GSB89] Ittai Gavrieli, Avraham Starinsky, and Amos Bein. The solubility of halite as a function of temperature in the highly saline Dead Sea brine system. Limnology and Oceanography, 34(7):1224–1234, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Gerritsen:1987:IFV

[GSBP87] Jeroen Gerritsen, Robert W. Sanders, Susan W. Bradley, and Karen G. Porter. Individual feeding variability of protozoan and crustacean zooplankton analyzed with flow cytometry. Limnology and Oceanography, 32(3):691–699, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Greene:1989:AZS

[GWB89] Charles H. Greene, Peter H. Wiebe, and Janusz Burczynski. Analyzing zooplankton size distributions using high-frequency sound. Limnology and Oceanography, 34(1):129–139, January 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Gschwend:1980:VOC

[GZG80] Philip Gschwend, Oliver C. Zafiriou, and Robert B. Gagosian. Volatile organic compounds in seawater from the Peru upwelling region. Limnology and Oceanography, 25(6):1044–1053, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Holm:1981:RNL

[HA81] Nancy Peterson Holm and David E. Armstrong. Role of nutrient limitation and competition in controlling the populations of Asterionella formosa and Microcystis aeruginosa in semicontinuous culture. Limnology and Oceanography, 26(4):622–634, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Hollibaugh:1983:MDD

[HA83] J. T. Hollibaugh and F. Azam. Microbial degradation of dissolved proteins in seawater. Limnology and Oceanography, 28(6):1104– 1116, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Hess:1985:URC

[HA85] F. R. Hess and D. G. Aubrey. Use of radio-controlled miniature aircraft for drifter and dye current studies in a tidal inlet. Limnology and Oceanography, 30(2):426–431, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Haberyan:1985:REF

[Hab85] Kurt A. Haberyan. The role of eopepod fecal pellets in the deposition of diatoms in Lake Tanganyika. Limnology and Oceanography, 30(5):1010–1023, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Hardy:1985:SSM

[HACF85] J. T. Hardy, C. W. Apts, E. A. Crecelius, and G. W. Fellingham. The sea-surface microlayer: Fate and residence times of atmospheric metals. Limnology and Oceanography, 30(1):93–101, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1980:DVC

[Hai80] Nelson G. Hairston, Jr. On the diel variation of copepod pigmentation. Limnology and Oceanography, 25(4):742–747, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1988:IVS

[Hai88] Nelson G. Hairston, Jr. Interannual variation in seasonal predation: Its origin and ecological importance. Limnology and Oceanography, 33(6):1245–1253, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Hakanson:1984:LST

[Hak84] John L. Hakanson. The long and short term feeding condition in field-caught Calanus pacificus, as determined from the lipid content. Limnology and Oceanography, 29(4):794–804, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Haakanson:1987:FCC

[Håk87] John L. Håkanson. The feeding condition of Calanus pacificus and other zooplankton in relation to phytoplankton pigments in the California Current. Limnology and Oceanography, 32(4):881–894, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Hall:1982:BRZ

[Hal82a] D. J. Hall. Book review: Zaret, T. M. 1980. Predation and freshwater communities. Yale Univ. Press, New Haven, Connecticut. 187 p. $15.00. Limnology and Oceanography, 27(2):391–393, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Hall:1982:FIN

[Hal82b] T. M. Hall. Free ionie nickel accumulation and localization in the freshwater zooplankter, Daphnia magna. Limnology and Oceanography, 26(5):718–727, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Hammond:1981:AGS

[Ham81] L. S. Hammond. An analysis of grain size modification in biogenic carbonate sediments by deposit-feeding holothurians and echinoids (Echinodermata). Limnology and Oceanography, 26(5):898–906, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Hamblin:1982:FSO

[Ham82] P. F. Hamblin. On the free surface oscillations of Lake Ontario. Limnology and Oceanography, 27(6):1039–1049, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Hannan:1981:PLS

[Han81] Cheryl Ann Hannan. Polychaete larval settlement: Correspondence of patterns in suspended jar collectors and in the adjacent natural habitat in Monterey Bay, California. Limnology and Oceanography, 26(1):159–171, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Hannan:1984:PLA

[Han84] Cheryl Ann Hannan. Planktonic larvae 05 act like passive particles in turbulent near-bottom flows. Limnology and Oceanography, 29 (5):1108–1116, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Haney:1985:RCF

[Han85] James F. Haney. Regulation of cladoceran filtering rates in nature by body size, food concentration, and diel feeding patterns. Limnology and Oceanography, 30(2):397–411, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Haney:1987:EGS

[Han87] J. Christopher Haney. Effects of a Gulf Stream warm-core ring on summer seabird distribution in the northwest Atlantic Ocean. Limnology and Oceanography, 32(3):665–673, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Hansson:1988:ECI

[Han88] Lars-Anders Hansson. Effects of competitive interactions on the biomass development of planktonic and periphytic algae in lakes. Limnology and Oceanography, 33(1):121–128, January 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Harris:1983:DGC

[Har83a] J. R. W. Harris. The development and growth of Calanus copepodites. Limnology and Oceanography, 28(1):142–147, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1983:TCU

[Har83b] W. G. Harrison. The time-course of uptake of inorganic and organic nitrogen compounds by phytoplankton from the Eastern Canadian Arctic: a comparison with temperate and tropical populations. Limnology and Oceanography, 28(6):1231–1237, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1984:PJT

[Har84] Nancy M. Harrison. Predation on jellyfish and their associates by seabirds. Limnology and Oceanography, 29(6):1335–1337, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Harvey:1987:FST

[Har87] R. W. Harvey. A fluorochrome-staining technique for counting bacteria in saline, organically enriched, alkaline lakes. Limnology and Oceanography, 32(4):993–995, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Harbison:1989:BRP

[Har89a] G. R. Harbison. Book review: Pelagic snails: The biology of holoplanktonic gastropod mollusks (C. M. Lalli and R. W. Gilmer). Limnology and Oceanography, 34(5):969–970, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Hartwig:1989:BRC

[Har89b] Eric Hartwig. Book review: Congressional directory: Environment (J. T. Grupenhoff and B. Farley [eds.]). Limnology and Oceanography, 34(5):970, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Holm:1980:IAR

[HASI80] Thomas R. Holm, Marc A. Anderson, Robert R. Stanforth, and Dennis G. Iverson. The influence of adsorption on the rates of microbial degradation of arsenic species in sediments. Limnology and Oceanography, 25(1):23–30, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Havel:1985:CDP

[Hav85] John E. Havel. Cyclomorphosis of Daphnia pulex spined morphs. Limnology and Oceanography, 30(4):853–861, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Hall:1989:OUK

[HAvdL^{+}89] Per O. J. Hall, Leif G. Anderson, Michiel M. Rutgers van der Loeff, Bjørn Sundby, and Stig F. G. Westerlund. Oxygen uptake kinetics in the benthic boundary layer. Limnology and Oceanography, 34 (4):734–746, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Haworth:1980:CCP

[Haw80] Elizabeth Y. Haworth. Comparison of continuous phytoplankton records with the diatom stratigraphy in the recent sediments of blelham tarn. Limnology and Oceanography, 25(6):1093–1103, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Hayward:1981:MDD

[Hay81] Thomas L. Hayward. Mating and the depth distribution of an oceanic copepod. Limnology and Oceanography, 26(2):374–377, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Hay:1987:BRP

[Hay87] Mark E. Hay. Book review: The physiological ecology of seaweeds (C. S. Lobban, P. J. Harrison, and M. J. Duncan). Limnology and Oceanography, 32(5):1178–1179, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Hemond:1981:DMI

[HB81] H. F. Hemond and R. Burke. A device for the measurement of infiltration in intermittently flooded wetlands. Limnology and Oceanography, 26(4):795–800, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Herbst:1984:RBH

[HB84] G. N. Herbst and H. J. Bromley. Relationships between habitat stability, ionic composition, and the distribution of aquatic invertebrates in the desert regions of Israel. Limnology and Oceanography, 29(3):495–503, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Helder:1985:SCM

[HB85] Willem Helder and Joop F. Bakker. Shipboard comparison of micro- and minielectrodes for measuring oxygen distribution in marine sediments. Limnology and Oceanography, 30(5):1106–1109, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1987:PPM

[HBT87] Nelson G. Hairston, Jr., Moshe Braner, and Saran Twombly. Perspective on prospective methods for obtaining life table data. Limnology and Oceanography, 32(2):517–520, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Hedges:1988:FRO

[HCC88a] John I. Hedges, Wayne A. Clark, and Gregory L. Come. Fluxes and reactivities of organic matter in a coastal marine bay. Limnology and Oceanography, 33(5):1137–1152, September 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Hedges:1988:OMS

[HCC88b] John I. Hedges, Wayne A. Clark, and Gregory L. Come. Organic matter sources to the water column and surficial sediments of a marine bay. Limnology and Oceanography, 33(5):1116–1136, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Hedges:1986:CFP

[HCQ^{+}86] John I. Hedges, Wayne A. Clark, Paul D. Quay, Jeffrey E. Richey, Allan H. Devol, and M. Santos. Compositions and fluxes of particulate organic material in the Amazon River. Limnology and Oceanography, 31(4):717–738, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Havel:1984:CPT

[HD84] John E. Havel and Stanley I. Dodson. Chaoborus predation on typical and spined morphs of Daphnia pulex: Behavioral observations. Limnology and Oceanography, 29(3):487–494, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Henrichs:1986:DLO

[HD86] Susan M. Henrichs and Allen P. Doyle. Decomposition of ^{14}Clabeled organic substances in marine sediments. Limnology and Oceanography, 31(4):765–778, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Howes:1984:CFT

[HDK84] Brian L. Howes, John W. H. Dacey, and Gary M. King. Carbon flow through oxygen and sulfate reduction pathways in salt marsh sediments. Limnology and Oceanography, 29(5):1037–1051, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Hall:1985:PCB

[HDLP85] Ronald J. Hall, Charles T. Driscoll, Gene E. Likens, and J. Michael Pratt. Physical, chemical, and biological consequences of episodic aluminum additions to a stream. Limnology and Oceanography, 30 (1):212–220, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Hatton:1983:SAS

[HDP83] R. S. Hatton, R. D. DeLaune, and W. H. Patrick, Jr. Sedimentation, accretion, and subsidence in marshes of Barataria Basin, Louisiana. Limnology and Oceanography, 28(3):494–502, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Howes:1985:EST

[HDW85] Brian L. Howes, John W. H. Dacey, and Stuart G. Wakeham. Effects of sampling technique on measurements of porewater constituents in salt marsh sediments. Limnology and Oceanography, 30(1):221–227, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Hastings:1988:SRP

[HE88] David Hastings and Steven Emerson. Sulfate reduction in the presence of low oxygen levels in the water column of the Cariaco Trench. Limnology and Oceanography, 33(3):391–396, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1984:BRT

[Hec84] R. E. Hecky. Book review: Thornton, J. A. [Ed.] 1982. Lake McIlwaine: The eutrophication and recovery of a tropical African manmade lake. W. Junk, The Hague. 264 p. $49.50. Limnology and Oceanography, 29(2):444–445, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1986:BRC

[Hec86] R. E. Hecky. Book review: Carmouze, J. P., J. R. Durand, and C. Lévèque [Eds.]. 1983. Lake Chad: Ecology and productivity of a shallow tropical ecosystem. Monogr. Biol. 53. W. Junk Publ., The Hague. 575 p. $137.00. Limnology and Oceanography, 31(4):907, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Henrichs:1987:BRO

[Hen87] Susan M. Henrichs. Book review: Organic geochemistry of natural waters (E. M. Thurman [cd.]). Limnology and Oceanography, 32 (4):1015–1016, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Heron:1982:VFF

[Her82] A. C. Heron. A vertical free fall plankton net with no mouth obstructions. Limnology and Oceanography, 27(2):380–383, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Herman:1983:VDP

[Her83] Alex W. Herman. Vertical distribution patterns of copepods, chlorophyll, and production in northeastern Baffin Bay. Limnology and Oceanography, 28(4):709–719, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1981:PPR

[HF81] R. E. Hecky and E. J. Fee. Primary production and rates of algal growth in Lake Tanganyika. Limnology and Oceanography, 26(3): 532–547, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Hemond:1982:SFS

[HF82] Harold F. Hemond and Jayne L. Fifield. Subsurface flow in salt marsh peat: a model and field study. Limnology and Oceanography, 27(1):126–136, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Henrichs:1984:PURa

[HF84] Susan M. Henrichs and John W. Farrington. Peru upwelling region sediments near 15^{o}S. 1. Remineralization and accumulation of organic matter. Limnology and Oceanography, 29(1):1–19, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Hough:1988:IIC

[HF88] R. Anton Hough and Mark D. Fornwall. Interactions of inorganic carbon and light availability as controlling factors in aquatic macrophyte distribution and productivity. Limnology and Oceanography, 33(5):1202–1208, September 1988. CODEN LI-OCAH. ISSN 0024-3590.

Harvell:1989:CSD

[HF89] C. Drew Harvell and William Fenical. Chemical and structural defenses of Caribbean gorgonians (Pseudopterogorgia spp.): Intercolony localization of defense. Limnology and Oceanography, 34(2):382–389, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Hollibaugh:1980:RLN

[HFA80] J. T. Hollibaugh, J. A. Fuhrman, and F. Azam. Radioactively labeling of natural assemblages of bacterioplankton for use in trophic studies. Limnology and Oceanography, 25(1):172–181, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Hashimoto:1987:DPN

[HFF87] Shinya Hashimoto, Kitao Fujiwara, and Keiichiro Fuwa. Determination of phosphorus in natural water using hydride generation and gas chromatography. Limnology and Oceanography, 32(3): 729–735, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Hashimoto:1985:DCC

[HFFS85] Shinya Hashimoto, Kitao Fujiwara, Keiichiro Fuwa, and Toshiro Saino. Distribution and characteristics of carboxypeptidase activity in pond, river, and seawaters in the vicinity of Tokyo. Limnology and Oceanography, 30(3):631–645, May 1985. CODEN LI-OCAH. ISSN 0024-3590.

Henrichs:1984:PURb

[HFL84] Susan M. Henrichs, John W. Farrington, and Cindy Lee. Peru upwelling region sediments near 15^{o}S. 2. Dissolved free and total hydrolyzable amino acids. Limnology and Oceanography, 29(1): 20–34, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Howarth:1983:SRS

[HG83] Robert W. Howarth and Anne Giblin. Sulfate reduction in the salt marshes at Sapelo Island, Georgia. Limnology and Oceanography, 28(1):70–82, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Hebert:1985:CIS

[HG85a] Paul D. N. Hebert and Peter M. Grewe. Chaoborus-induced shifts in the morphology of Daphnia ambigua. Limnology and Oceanography, 30(6):1291–1297, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Horne:1985:NFO

[HG85b] A. J. Horne and D. L. Galat. Nitrogen fixation in an oligotrophic, saline desert lake: Pyramid Lake, Nevada. Limnology and Oceanography, 30(6):1229–1239, November 1985. CODEN LI-OCAH. ISSN 0024-3590.

Harvell:1988:BRK

[HG88] C. Drew Harvell and Charles H. Greene. Book review: Kerfoot, W. C., and A. Sih [eds.]. 1987. Predation: Direct and indirect impacts on aquatic communities. University Press of New England, Hanover. 386 p. $65.00. Limnology and Oceanography, 33(4):646– 647, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1982:PCB

[HGH82] W. M. Hamner, R. W. Gilmer, and P. P. Hamner. The physical, chemical, and biological characteristics of a stratified, saline, sulfide lake in Palau. Limnology and Oceanography, 27(5):896–909, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Hebert:1980:ISP

[HGM80] Paul D. N. Hebert, Allen G. Good, and Mona A. Mort. Induced swarming in the predatory copepod Heterocope septentrionalis. Limnology and Oceanography, 25(4):747–750, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Holm:1983:FAR

[HGS83] Nancy Peterson Holm, George G. Ganf, and Joseph Shapiro. Feeding and assimilation rates of Daphnia pulex fed Aphanizomenonflosaquae. Limnology and Oceanography, 28(4):677–687, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Harrits:1980:SAM

[HH80] Susan M. Harrits and Richard S. Hanson. Stratification of aerobic methane-oxidizing organisms in Lake Mendota, Madison, Wisconsin. Limnology and Oceanography, 25(3):412–421, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1981:EIM

[HH81] W. M. Hamner and I. R. Hauri. Effects of island mass: Water flow and plankton pattern around a reef in the Great Barrier Reef lagoon, Australia. Limnology and Oceanography, 26(6):1084–1102, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Herman:1985:SPH

[HH85] Peter M. J. Herman and Carlo Heip. Secondary production of the harpacticoid copepod Paronychocamptus nanus in a brackishwater habitat. Limnology and Oceanography, 30(5):1060–1066, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Heisey:1986:OAE

[HH86] Dennis M. Heisey and John M. Hoenig. Optimal allocation of effort in studies using the size-frequency method of estimating secondary production. Limnology and Oceanography, 31(1):211–215, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Hewes:1983:MRN

[HHH83] Christopher D. Hewes and Osmund Holm-Hansen. A method for recovering nanoplankton from filters for identification with the microscope: The filter-transfer-freeze (FTF) technique. Limnology and Oceanography, 28(2):389–394, March 1983. CODEN LI-OCAH. ISSN 0024-3590.

Hama:1987:DAA

[HHH87] Takeo Hama, Nobuhiko Handa, and Junko Hama. Determination of amino acid production rate of a marine phytoplankton population with ^{13}C and gas chromatography-mass spectrometry. Limnology and Oceanography, 32(5):1144–1153, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1989:FOO

[HHOC89] W. M. Hamner, P. P. Hamner, B. S. Obst, and J. H. Carleton. Field observations on the ontogeny of schooling of Euphausia superba furciliae and its relationship to ice in Antarctic waters. Limnology and Oceanography, 34(2):451–456, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Hayes:1980:TPW

[HHS80] D. W. Hayes, S. D. Harris, and R. S. Stoughton. A tidal-powered water sampler. Limnology and Oceanography, 25(4):761–764, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Howes:1981:ORP

[HHTV81] Brian L. Howes, Robert W. Howarth, John M. Teal, and Ivan Valiela. Oxidation-reduction potentials in a salt marsh: Spatial patterns and interactions with primary production. Limnology and Oceanography, 26(2):350–360, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Herczeg:1988:THS

[HI88] Andrew L. Herczeg and Dieter M. Imboden. Tritium hydrologic studies in four closed-basin lakes in the Great Basin, U.S.A. Limnology and Oceanography, 33(2):157–173, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Hilton:1985:CFP

[Hil85] J. Hilton. A conceptual framework for predicting the occurrence of sediment focusing and sediment redistribution in small lakes. Limnology and Oceanography, 30(6):1131–1143, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Hansson:1987:EFG

[HJP87] Lars-Anders Hansson, Lars Johansson, and Lennart Persson. Effects of fish grazing on nutrient release and succession of primary producers. Limnology and Oceanography, 32(3):723–729, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Hamblin:1980:EUC

[HK80] P. F. Hamblin and R. Kuehnel. An evaluation of an unattended current and temperature profiler for deep lakes. Limnology and Oceanography, 25(6):1128–1141, November 1980. CODEN LI-OCAH. ISSN 0024-3590.

Hecky:1981:PPE

[HK81a] R. E. Hecky and H. J. Kling. The phytoplankton and protozooplankton of the euphotic zone of Lake Tanganyika: Species composition, biomass, chlorophyll content, and spatio-temporal distribution. Limnology and Oceanography, 26(3):548–564, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Holtby:1981:ZFR

[HK81b] L. Blair Holtby and Roy Knoechel. Zooplankton filtering rates: Error due to loss of radioisotopic label in chemically preserved samples. Limnology and Oceanography, 26(4):774–780, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Hong:1986:RSI

[HK86] Huasheng Hong and Dana R. Kester. Redox state of iron in the offshore waters of Peru. Limnology and Oceanography, 31(3):512– 524, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Hecky:1988:NLP

[HK88a] R. E. Hecky and P. Kilham. Nutrient limitation of phytoplankton in freshwater and marine environments: a review of recent evidence on the effects of enrichment. Limnology and Oceanography, 33(4 (part 2)):796–822, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Hill:1988:CGE

[HK88b] Walter R. Hill and Allen W. Knight. Concurrent grazing effects of two stream insects on periphyton. Limnology and Oceanography, 33(1):15–26, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Hines:1989:SRO

[HKT89] Mark E. Hines, Stephen L. Knollmeyer, and Joyce B. Tugel. Sulfate reduction and other sedimentary biogeochemistry in a northern New England salt marsh. Limnology and Oceanography, 34 (3):578–590, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Hunter:1981:ACE

[HL81] Bonita L. Hunter and Edward A. Laws. ATP and chlorophylla as estimators of phytoplankton carbon biomass. Limnology and Oceanography, 26(5):944–956, September 1981. CODEN LI-OCAH. ISSN 0024-3590.

Hunter:1982:OMS

[HL82] K. A. Hunter and P. S. Liss. Organic matter and the surface charge of suspended particles in estuarine waters. Limnology and Oceanography, 27(2):322–335, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Hilton:1985:ERC

[HL85] J. Hilton and J. P. Lishman. The effect of redox changes on the magnetic susceptibility of sediments from a seasonally anoxic lake. Limnology and Oceanography, 30(4):907–909, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Hamilton:1987:CSC

[HL87] Stephen K. Hamilton and William M. Lewis, Jr. Causes of seasonality in the chemistry of a lake on the Orinoco River floodplain, Venezuela. Limnology and Oceanography, 32(6):1277–1290, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Humphries:1988:CBR

[HL88] Stella E. Humphries and Vincent D. Lyne. Cyanophyte blooms: The role of cell buoyancy. Limnology and Oceanography, 33(1): 79–91, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Hilton:1986:DPS

[HLA86] J. Hilton, J. P. Lishman, and P. V. Allen. The dominant processes of sediment distribution and focusing in a small, eutrophic, monomictic lake. Limnology and Oceanography, 31(1):125–133, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Hurlbert:1986:FFP

[HLM86] Stuart H. Hurlbert, Washington Loayza, and Tomas Moreno. Fishflamingo-plankton interactions in the Peruvian Andes. Limnology and Oceanography, 31(3):457–468, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Hitchcock:1987:STC

[HLS87] Gary L. Hitchcock, C. Langdon, and T. J. Smayda. Short-term changes in the biology of a Gulf Stream warm-core ring: Phytoplankton biomass and productivity. Limnology and Oceanography, 32(4):919–928, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Harbison:1980:FAC

[HM80] G. R. Harbison and V. L. McAlister. Fact and artifact in copepod feeding experiments. Limnology and Oceanography, 25(6):971–981, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Holmes:1982:POE

[HM82] Robert W. Holmes and Bruce E. Mahall. Preliminary observations on the effects of flooding and desiccation upon the net photosynthetic rates of high intertidal estuarine sediments. Limnology and Oceanography, 27(5):954–958, September 1982. CODEN LI-OCAH. ISSN 0024-3590.

Howarth:1984:PFM

[HM84] Robert W. Howarth and Susan Merkel. Pyrite formation and the measurement of sulfate reduction in salt marsh sediments. Limnology and Oceanography, 29(3):598–608, May 1984. CODEN LI-OCAH. ISSN 0024-3590.

Harrison:1986:RMD

[HM86] Gail I. Harrison and F. M. M. Morel. Response of the marine diatom Thalassiosira weissflogii to iron stress. Limnology and Oceanography, 31(5):989–997, September 1986. CODEN LI-OCAH. ISSN 0024-3590.

Haardt:1987:SVA

[HM87] Hauke Haardt and Helmut Maske. Specific in vivo absorption coefficient of chlorophyll-a at 675 nm. Limnology and Oceanography, 32(3):608–619, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Hudson:1989:DBE

[HM89] Robert J. M. Hudson and Francois M. M. Morel. Distinguishing between extra- and intracellular iron in marine phytoplankton. Limnology and Oceanography, 34(6):1113–1120, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Howarth:1988:NFFb

[HMC88] Robert W. Howarth, Roxanne Marino, and Jonathan J. Cole. Nitrogen fixation in freshwater, estuarine, and marine ecosystems. 2. Biogeochemical controls. Limnology and Oceanography, 33(4 (part 2)):688–701, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Harbison:1986:RSP

[HMG86] G. R. Harbison, V. L. McAlister, and R. W. Gilmer. The response of the salp, Pegea confoederata, to high levels of particulate material: Starvation in the midst of plenty. Limnology and Oceanography, 31(2):371–382, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Howarth:1988:NFFa

[HMLC88] Robert W. Howarth, Roxanne Marino, Judith Lane, and Jonathan J. Cole. Nitrogen fixation in freshwater, estuarine, and marine ecosystems. 1. Rates and importance. Limnology and Oceanography, 33(4 (part 2)):669–687, July 1988. CODEN LI-OCAH. ISSN 0024-3590.

Horn:1986:WII

[HMS86] W. Horn, C. H. Mortimer, and D. J. Schwab. Wind-induced internal seiches in Lake Zurich observed and modeled. Limnology and Oceanography, 31(6):1232–1254, November 1986. CODEN LI-OCAH. ISSN 0024-3590.

Han:1981:HOM

[HN81] Gregory Han and Terren Niedrauer. Hydrographic observations and mixing processes in the New York Bight, 1975–1977. Limnology and Oceanography, 26(6):1126–1141, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1985:A

[Hob85] John E. Hobbie. Announcement. Limnology and Oceanography, 30(1):230, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Hobbie:1988:CEP

[Hob88] John E. Hobbie. A comparison of the ecology of planktonic bacteria in fresh and salt water. Limnology and Oceanography, 33(4 (part 2)):750–764, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Hojerslev:1989:SWQ

[Høj89] Niels K. Højerslev. Surface water-quality studies in the interior marine environment of Denmark. Limnology and Oceanography, 34(8):1630–1639, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Horrigan:1981:PPU

[Hor81] Sarah G. Horrigan. Primary production under the Ross Ice Shelf, Antarctica. Limnology and Oceanography, 26(2):378–382, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Horne:1983:BRG

[Hor83] Alex J. Horne. Book review: Gower, J. F. [Ed.] 1981. Oceanography from space. Proc. COSPAR/SCOR/IUCRM Symp. 1980. Mar. Sci. V. 13. Plenum Publ. Corp., New York. 978 p. $95.00. (Typewriter composition). Limnology and Oceanography, 28(5): 1045–1046, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1987:VTG

[HP87] S. J. Harrison and A. P. Phizacklea. Vertical temperature gradients in muddy intertidal sediments in the Forth estuary, Scotland. Limnology and Oceanography, 32(4):954–963, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Herndl:1989:PMU

[HP89] Gerhard J. Herndl and Peter Peduzzi. Potential microbial utilization rates of sublittoral gastropod mucus trails. Limnology and Oceanography, 34(4):780–784, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Hrbacek:1982:BRE

[Hrb82] J. Hrbacek. Book review: Evolution and ecology of zooplankton communities. 1980. Proceedings of a symposium on the structure of zooplankton communities held at Dartmouth College, 20–25 August 1978. American Society of Limnology and Oceanography Special Symposium 3. W. C. Kerfoot, Ed. Univ. Press of New England, Hanover, New Hampshire. 794 p. $45.00. Limnology and Oceanography, 27(2):390–391, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Hanson:1980:GES

[HS80] Roger B. Hanson and James Snyder. Glucose exchanges in a salt marsh-estuary: Biological activity and chemical measurements.

Limnology and Oceanography, 25(4):633–642, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Horner:1982:MTI

[HS82] S. M. J. Horner and D. F. Smith. Measurement of total inorganic carbon in seawater by a substoichiometric assay using NaH^{14} CO_{3}. Limnology and Oceanography, 27(5):978–983, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Hedges:1984:CND

[HS84a] John I. Hedges and Jeffrey H. Stern. Carbon and nitrogen determinations of carbonate-containing solids. Limnology and Oceanography, 29(3):657–663, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Holm:1984:ELR

[HS84b] Nancy Peterson Holm and Joseph Shapiro. An examination of lipid reserves and the nutritional status of Daphnia pulex fed Aphanizomenon flos-aquae. Limnology and Oceanography, 29(5):1137– 1140, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1986:RSR

[HS86] William M. Hamner and David Schneider. Regularly spaced rows of medusae in the Bering Sea: Role of Langmuir circulation. Limnology and Oceanography, 31(1):171–176, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Horsch:1988:CCL

[HS88] G. M. Horsch and H. G. Stefan. Convective circulation in littoral water due to surface cooling. Limnology and Oceanography, 33(5): 1068–1083, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Hamner:1987:EOF

[HSMH87] W. M. Hamner, S. W. Strand, G. I. Matsumoto, and P. P. Hamner. Ethological observations on foraging behavior of the ctenophore Leucothea sp. in the open sea. Limnology and Oceanography, 32 (3):645–652, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1985:OLT

[HT85a] Nelson G. Hairston, Jr. and Saran Twombly. Obtaining life table data from cohort analyses: a critique of current methods. Limnology and Oceanography, 30(4):886–893, July 1985. CODEN LI-OCAH. ISSN 0024-3590.

Honjo:1985:GDO

[HT85b] Tsuneo Honjo and Kenji Tabata. Growth dynamics of Olisthodiscus luteus in outdoor tanks with flowing coastal water and in small vessels. Limnology and Oceanography, 30(3):653–664, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Hedges:1984:SDS

[HTE84] John I. Hedges, H. J. Turin, and John R. Ertel. Sources and distributions of sedimentary organic matter in the Columbia River drainage basin, Washington and Oregon. Limnology and Oceanography, 29(1):35–46, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Hamilton-Taylor:1984:DFM

[HTWR84] J. Hamilton-Taylor, M. Willis, and C. S. Reynolds. Depositional fluxes of metals and phytoplankton in Windermere as measured by sediment traps. Limnology and Oceanography, 29(4):695–710, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Hunter:1980:MPN

[Hun80] K. A. Hunter. Microelectrophoretic properties of natural surfaceactive organic matter in coastal seawater. Limnology and Oceanography, 25(5):807–822, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Huntley:1981:NNF

[Hun81] Mark Huntley. Nonselective, nonsaturated feeding by three calanid copepod species in the Labrador Sea. Limnology and Oceanography, 26(5):831–842, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Hunt:1983:IDM

[Hun83a] Carlton D. Hunt. Incorporation and deposition of Mn and other trace metals by flocculent organic matter in a controlled marine ecosystem. Limnology and Oceanography, 28(2):302–308, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Hunt:1983:VBM

[Hun83b] Carlton D. Hunt. Variability in the benthic Mn flux in coastal marine ecosystems resulting from temperature and primary production. Limnology and Oceanography, 28(5):913–923, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Hutchinson:1986:M

[Hut86] G. Evelyn Hutchinson. In memoriam. Limnology and Oceanography, 31(1):233, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Hutchinson:1988:WE

[Hut88] G. Evelyn Hutchinson. W. T. Edmondson. Limnology and Oceanography, 33(6):1231–1233, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Hay:1985:DCS

[HW85] Alex E. Hay and Rosemary Waters. A diatom chronology for sediments in a high accumulation rate environment: Rupert Inlet, British Columbia. Limnology and Oceanography, 30(4):898–906, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Huryn:1986:MOS

[HW86] Alexander D. Huryn and J. Bruce Wallace. A method for obtaining in situ growth rates of larval Chironomidae (Diptera) and its application to studies of secondary production. Limnology and Oceanography, 31(1):216–221, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Harrison:1988:INU

[HW88] W. G. Harrison and Laura J. E. Wood. Inorganic nitrogen uptake by marine picoplankton: Evidence for size partitioning. Limnology and Oceanography, 33(3):468–475, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Hairston:1983:CCS

[HWL83] Nelson G. Hairston, Jr., William E. Walton, and Kao T. Li. The causes and consequences of sex-specific mortality in a freshwater copepod. Limnology and Oceanography, 28(5):935–947, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Hsieh:1989:DMD

[HY89] Y. P. Hsieh and C. H. Yang. Diffusion methods for the determination ofreduced inorganic sulfur species in sediments. Limnology and Oceanography, 34(6):1126–1130, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1980:NCS

[Hyn80] H. B. N. Hynes. A name change in the secondary production business. Limnology and Oceanography, 25(4):778, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Hynes:1981:BRB

[Hyn81] H. B. N. Hynes. Book review: Brodsky, K. A. 1980. 0Mountain torrent of the Tien Shan. Monogr. Biol. 39. W. Junk, The Hague, xi + 311 p. $79.00. Limnology and Oceanography, 26(2):400, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Infante:1985:IDO

[IA85] Aida Infante and Sally E. B. Abella. Inhibition of Daphnia by Oscillatoria in Lake Washington. Limnology and Oceanography, 30(5):1046–1052, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Ingvorsen:1982:EFS

[IB82a] K. Ingvorsen and T. D. Brock. Electron flow via sulfate reduction and methanogenesis in the anaerobic hypolimnion of Lake Mendota. Limnology and Oceanography, 27(3):559–564, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Ivey:1982:EBC

[IB82b] G. N. Ivey and F. M. Boyce. Entrainment by bottom currents in Lake Erie. Limnology and Oceanography, 27(6):1029–1038, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Ibanez:1981:IDH

[Iba81] Frederic Ibanez. Immediate detection of heterogeneities in continuous multivariate, oceanographic recordings. Application to time series analysis of changes in the bay of Villefranche sur Mer. Limnology and Oceanography, 26(2):336–349, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Imberger:1983:IWM

[IBC^{+}83] J. Imberger, T. Berman, R. R. Christian, E. B. Sherr, D. E. Whitney, L. R. Pomeroy, R. G. Wiegert, and W. J. Wiebe. The influence of water motion on the distribution and transport of materials in a salt marsh estuary. Limnology and Oceanography, 28(2): 201–214, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Ittekkot:1982:MCA

[IDB82] Venugopalan Ittekkot, Egon T. Degens, and Uwe Brockmann. Monosaccharide composition of acid-hydrolyzable carbohydrates in particulate matter during a plankton bloom. Limnology and Oceanography, 26(5):770–776, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Ilamner:1981:LDH

[IH81] William M. Ilamner and Ivan R. Hauri. Long-distance horizontal migrations of zooplankton (scyphomedusae: Mastigias). Limnology and Oceanography, 26(3):414–423, May 1981. CODEN LI-OCAH. ISSN 0024-3590.

Imboden:1984:TMH

[IJ84] D. M. Imboden and Th. Joller. Turbulent mixing in the hypolimnion of Baldeggersee (Switzerland) traced by natural radon-222. Limnology and Oceanography, 29(4):831–844, July 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Iversen:1985:AMO

[IJ85] Niels Iversen and Bo Barker Jorgensen. Anaerobic methane oxidation rates at the sulfate-methane transition in marine sediments from Kattegat and Skagerrak (Denmark). Limnology and Oceanography, 30(5):944–955, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Infante:1985:DBT

[IL85] Aida Infante and Ami H. Litt. Differences between two species of Daphnia in the use of 10 species of algae in Lake Washington. Limnology and Oceanography, 30(5):1053–1059, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Illg:1980:BRG

[Ill80] Paul L. Illg. Book review: George, J. D., and J. J. George. 1979. Marine life: an illustrated encyclopedia of invertebrates in the sea. John Wiley & Sons, New York. 288 p. $39.95. Limnology and Oceanography, 49(6):577, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Imberger:1985:DML

[Imb85] J. Imberger. The diurnal mixed layer. Limnology and Oceanography, 30(4):737–770, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Iturriaga:1988:MAI

[IMK88] R. Iturriaga, B. G. Mitchell, and D. A. Kiefer. Microphotometric analysis of individual particle absorption spectra. Limnology and Oceanography, 33(1):128–135, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Iverson:1989:PDP

[INA89] R. L. Iverson, F. L. Nearhoof, and M. O. Andreae. Production of dimethylsulfonium propionate and dimethylsulfide by phytoplankton in estuarine and coastal waters. Limnology and Oceanography, 34(1):53–67, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Iversen:1987:BSL

[IOK87] Niels Iversen, Ronald S. Oremland, and Michael J. Klug. Big Soda Lake (Nevada). 3. Pelagic methanogenesis and anaerobic methane oxidation. Limnology and Oceanography, 32(4):804–814, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Ivey:1984:MVM

[IP84] G. N. Ivey and J. C. Patterson. A model of the vertical mixing in Lake Erie in summer. Limnology and Oceanography, 29(3):553– 563, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Imberger:1985:MLD

[IP85] Jörg Imberger and Geoffrey Parker. Mixed layer dynamics in a lake exposed to a spatially variable wind field. Limnology and Oceanography, 30(3):473–488, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Iturriaga:1989:MCP

[IS89] Rodolfo Iturriaga and David A. Siegel. Microphotometric characterization of phytoplankton and detrital absorption properties in the Sargasso Sea. Limnology and Oceanography, 34(8):1706–1726, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Iversen:1988:SPT

[Ive88] Torben Moth Iversen. Secondary production and trophic relationships in a spring invertebrate community. Limnology and Oceanography, 33(4):582–592, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Istvanovics:1986:CPN

[IVH^{+}86] Vera Istvánovics, Lajos Vörös, Sándor Herodek, László G. Tóth, and István Tátrai. Changes of phosphorus and nitrogen concentration and of phytoplankton in enriched lake enclosures. Limnology and Oceanography, 31(4):798–811, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Jackson:1987:SCR

[Jac87] George A. Jackson. Simulating chemosensory responses of marine microorganisms. Limnology and Oceanography, 32(6):1253–1266, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Jackson:1989:SBA

[Jac89] George A. Jackson. Simulation of bacterial attraction and adhesion to falling particles in an aquatic environment. Limnology and Oceanography, 34(3):514–530, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Jahnke:1983:BRB

[Jah83] R. A. Jahnke. Book review: Berner, R. A. 1980. Early diagenesis. Princeton Univ. Press, Princeton, N.J. 241 p. Softcover $9.50; hardcover $25.00. Limnology and Oceanography, 28(1):199, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Jahnke:1985:MMD

[Jah85] Richard Jahnke. A model of microenvironments in deep-sea sediments: Formation and effects on porewater profiles. Limnology and Oceanography, 30(5):966–971, September 1985. CODEN LI-OCAH. ISSN 0024-3590.

Jahnke:1988:SRI

[Jah88] Richard A. Jahnke. A simple, reliable, and inexpensive pore-water sampler. Limnology and Oceanography, 33(3):483–487, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Janssen:1981:SZJ

[Jan81] John Janssen. Searching for zooplankton just outside Snell’s window. Limnology and Oceanography, 26(6):1168–1171, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Javor:1983:PSC

[Jav83] Barbara J. Javor. Planktonic standing crop and nutrients in a saltern ecosystem. Limnology and Oceanography, 28(1):153–159, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1989:IBC

[JBvdB89] Richard K. Johnson, Bengt Boström, and Wouter van de Bund. Interactions between Chironomus plumosus (L.) and the microbial community in surficial sediments of a shallow, eutrophic lake. Limnology and Oceanography, 34(6):992–1003, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1980:OPA

[JC80] Bruce D. Johnson and Robert C. Cooke. Organic particle and aggregate formation resulting from the dissolution of bubbles in seawater. Limnology and Oceanography, 25(4):653–661, July 1980. CODEN LIOCAH. ISSN 0024-3590. See comment [Gar81a].

Johnson:1981:RCG

[JC81] Bruce D. Johnson and Robert C. Cooke. Reply to comment by Garrett. Limnology and Oceanography, 27(4):992–995, September 1981. CODEN LIOCAH. ISSN 0024-3590. See [Gar81a].

Jergensen:1986:SFO

[JD86a] Bo Barker Jergensen and David J. Des Marais. A simple fiber-optic microprobe for high resolution light measurements: Application in marine sediment. Limnology and Oceanography, 31(6):1376–1383, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Jergensen:1986:TAH

[JD86b] Bo Barker Jergensen and David J. Des Marais. A technique to assess the harmful effects of sampling and containment for determination of primary production. Limnology and Oceanography, 31 (6):1364–1373, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1988:OPB

[JD88] Bo Barker Jørgensen and David J. Des Marais. Optical properties of benthic photosynthetic communities: Fiber-optic studies of cyanobacterial mats. Limnology and Oceanography, 33(1):99–113, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1982:TOC

[JEE82] Thomas C. Johnson, James E. Evans, and Steven J. Eisenreich. Total organic carbon in Lake Superior sediments: Comparisons with hemipelagic and pelagic marine environments. Limnology and Oceanography, 27(3):481–491, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Jeffrey:1981:ITL

[Jef81] S. W. Jeffrey. An improved thin-layer chromatographic technique for marine phytoplankton pigments. Limnology and Oceanography, 26(1):191–197, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Jahnke:1982:MOR

[JEM82] Richard A. Jahnke, Steven R. Emerson, and James W. Murray. A model of oxygen reduction, denitrification, and organic matter mineralization in marine sediments. Limnology and Oceanography, 27(4):610–623, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1985:SPP

[JERC85] Bo Barker Jørgensen, Jonathan Erez, Peter Revsbech, and Yehuda Cohen. Symbiotic photosynthesis in a planktonic Foraminiferan, Globigerinoides sacculifer (Brady), studied with microelectrodes. Limnology and Oceanography, 30(6):1253–1267, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1982:DPS

[JGCS82] Bruce D. Johnson, Robert M. Gershey, Robert C. Cooke, and William H. Sutcliffe. A device for the production of small bubbles in seawater. Limnology and Oceanography, 27(2):369–373, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Johnsen:1987:EFL

[JJ87] Geir Helge Johnsen and Per Johan Jakobsen. The effect of food limitation on vertical migration in Daphnia longispina. Limnology and Oceanography, 32(4):873–880, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Jenkins:1984:CND

[JK84] Mark C. Jenkins and W. Michael Kemp. The coupling of nitrification and denitrification in two estuarine sediments. Limnology and Oceanography, 29(3):609–619, May 1984. CODEN LIOCAH. ISSN 0024-3590.

James:1987:SLV

[JKMN87] William F. James, Robert H. Kennedy, Robert H. Montgomery, and Joe Nix. Seasonal and longitudinal variations in apparent deposition rates within an Arkansas reservoir. Limnology and Oceanography, 32(5):1169–1176, September 1987. CODEN LI-OCAH. ISSN 0024-3590.

Jordan:1980:MPB

[JL80] Marilyn J. Jordan and Gene E. Likens. Measurement of planktonic bacterial production in an oligotrophic lake. Limnology and Oceanography, 25(4):719–732, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Jones:1984:NCC

[JM84] Carolyn J. Jones and James W. Murray. Nickel, cadmium, and copper in the northeast Pacific off the coast of Washington. Limnology and Oceanography, 29(4):711–720, July 1984. CODEN LI-OCAH. ISSN 0024-3590.

Jones:1985:GMN

[JM85] Carolyn J. Jones and James W. Murray. The geochemistry of manganese in the northeast Pacific Ocean off Washington. Limnology and Oceanography, 30(1):81–92, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1982:CDH

[Joh82] Kenneth S. Johnson. Carbon dioxide hydration and dehydration kinetics in seawater. Limnology and Oceanography, 27(5):849–855, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1989:BRL

[Joh89] Dale W. Johnson. Book review: Long-term studies in ecology: Approaches and alternatives (G. E. Likens [ed.]). Limnology and Oceanography, 34(4):791–792, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Jones:1982:NFT

[Jon82] Keith Jones. Nitrogen fixation in the temperate estuarine intertidal sediments of the River Lune. Limnology and Oceanography, 27(3):455–460, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Jonasz:1987:NSM

[Jon87] Miroslaw Jonasz. Nonsphericity of suspended marine particles and its influence on light scattering. Limnology and Oceanography, 32 (5):1059–1065, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1987:FAA

[Jør87] Niels O. G. Jørgensen. Free amino acids in lakes: Concentrations and assimilation rates in relation to phytoplankton and bacterial production. Limnology and Oceanography, 32(1):97–111, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1983:DNN

[JP83] Kenneth S. Johnson and Robert L. Petty. Determination of nitrate and nitrite in seawater by flow injection analysis. Limnology and Oceanography, 28(6):1260–1266, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Jorgensen:1985:DBL

[JR85] Bo Barker Jørgensen and Niels Peter Revsbech. Diffusive boundary layers and the oxygen uptake of sediments and detritus. Limnology and Oceanography, 30(1):111–122, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Jorgensen:1983:PSB

[JRC83] Bo Barker Jorgensen, Niels Peter Revsbech, and Yehuda Cohen. Photosynthesis and structure of benthic microbial mats: Microelectrode and SEM studies of four cyanobacterial communities. Limnology and Oceanography, 28(6):1075–1093, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Jewson:1981:LRS

[JRG81] D. H. Jewson, B. H. Rippey, and W. K. Gilmore. Loss rates from sedimentation, parasitism, and grazing during the growth, nutrient limitation, and dormancy of a diatom crop. Limnology and Oceanography, 26(6):1045–1056, November 1981. CODEN LI-OCAH. ISSN 0024-3590.

Jedrasik:1983:CDP

[JS83] J. Jedrasik and T. J. Simons. Comparison of different procedures for the simulation of temperature oscillations in Radunia Lake, Poland. Limnology and Oceanography, 28(5):1014–1020, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Jordan:1981:ROD

[JSS81] Thomas F. Jordan, Kirby R. Stortz, and Michael Sydor. Resonant oscillation in Duluth–Superior Harbor. Limnology and Oceanography, 26(1):186–190, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Jones:1984:DIS

[JT84] S. D. Jones and V. W. Truesdale. Dissolved iodine species in a British freshwater system. Limnology and Oceanography, 29(5): 1016–1024, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1982:BRM

[Jum82] Peter A. Jumars. Book review: Marine benthic dynamics. 1980. Belle W. Baruch Library in Marine Science No. 11. K. R. Tenore and B. C. Coull, Eds. Univ. South Carolina Press, Columbia. 451 p. (typewriter composition). $27.50. Limnology and Oceanography, 27(2):396, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1984:BRP

[Jum84] Peter A. Jumars. Book review: Platt, T., K. H. Mann, and R. E. Ulanowicz [Eds.]. 1981. Mathematical models in biological oceanography. Monogr. Oceanogr. Methodol. 7. UNESCO Press, Paris. 156 p. Limnology and Oceanography, 29(1):222–223, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1987:BRI

[Jum87a] Peter A. Jumars. Book review: The interpretation of ecological data (E. C. Pielou). Limnology and Oceanography, 32(2):523–524, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1987:ECE

[Jum87b] Peter A. Jumars. Editorial comment: The evolving natural history of a manuscript. Limnology and Oceanography, 32(4):1011–1014, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1988:ECC

[Jum88a] Peter A. Jumars. Editorial comment: Current concerns of readers. Limnology and Oceanography, 33(2):298–299, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Jumars:1988:ECS

[Jum88b] Peter A. Jumars. Editorial comment on submission of comments concerning printed comments. Limnology and Oceanography, 33 (5):1227–1228, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Jordan:1982:NBR

[JV82] Thomas E. Jordan and Ivan Valiela. A nitrogen budget of the ribbed mussel, Geukensia demissa, and its significance in nitrogen flow in a New England salt marsh. Limnology and Oceanography, 27(1):75–90, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Juttner:1984:EED

[JW84] Friedrich Jüttner and Karl Wurster. Evidence of ectocarpene and dictyopterenes A and C’ in the water of a freshwater lake. Limnology and Oceanography, 29(6):1322–1324, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Johnson:1985:SFP

[JW85] B. D. Johnson and P. J. Wangersky. Seawater filtration: Particle flow and impaction considerations. Limnology and Oceanography, 30(5):966–971, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Jokiel:1984:IUR

[JY84] P. L. Jokiel and R. H. York, Jr. Importance of ultraviolet radiation in photoinhibition of microalgal growth. Limnology and Oceanography, 29(1):192–198, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Jannasch:1980:SST

[JZF80] Hans W. Jannasch, Oliver C. Zafiriou, and John W. Farrington. A sequencing sediment trap for time-series studies of fragile particles. Limnology and Oceanography, 25(5):939–943, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Kattawar:1989:SVC

[KA89] George W. Kattawar and Charles N. Adams. Stokes vector calculations of the submarine light field in an atmosphere-ocean with scattering according to a Rayleigh phase matrix: Effect of interface refractive index on radiance and polarization. Limnology and Oceanography, 34(8):1453–1472, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Kantin:1983:CSA

[Kan83] Roger Kantin. Chemical speciation of antimony in marine algae. Limnology and Oceanography, 28(1):165–168, January 1983. CO-DEN LIOCAH. ISSN 0024-3590.

Krom:1980:APA

[KB80a] Michael D. Krom and Robert A. Berner. Adsorption of phosphate in anoxic marine sediments. Limnology and Oceanography, 25(5): 797–806, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Krom:1980:DCS

[KB80b] Michael D. Krom and Robert A. Berner. The diffusion coefficients of sulfate, ammonium, and phosphate ions in anoxic marine sediments. Limnology and Oceanography, 25(2):327–337, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Kaplan:1982:DFD

[KB82] Louis A. Kaplan and Thomas L. Bott. Diel fluctuations of DOC generated by algae in a Piedmont stream. Limnology and Oceanography, 27(6):1091–1100, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Kaplan:1989:DFB

[KB89a] Louis A. Kaplan and Thomas L. Bott. Diel fluctuations in bacterial activity on streambed substrata during vernal algal blooms: Effects of temperature, water chemistry, and habitat. Limnology and Oceanography, 34(4):718–733, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Karl:1989:MDD

[KB89b] David M. Karl and Megan D. Bailiff. The measurement and distribution of dissolved nucleic acids in aquatic environments. Limnology and Oceanography, 34(3):543–558, May 1989. CODEN LI-OCAH. ISSN 0024-3590.

Kennicutt:1989:USC

[KBB^{+}89] Mahlon C. Kennicutt II, James M. Brooks, Robert R. Bidigare, Susanne J. McDonald, Daniel L. Adkison, and Stephen A. Macko. An upper slope “cold” seep community: Northern California. Limnology and Oceanography, 34(3):635–640, May 1989. CODEN LI-OCAH. ISSN 0024-3590.

Kishino:1984:URE

[KBO84] Motoaki Kishino, Charles R. Booth, and Noboru Okami. Underwater radiant energy absorbed by phytoplankton, detritus, dissolved organic matter, and pure water. Limnology and Oceanography, 29(2):340–349, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Kelly:1981:CTI

[KC81] Carol A. Kelly and David P. Chynoweth. The contributions of temperature and of the input of organic matter in controlling rates of sediment methanogenesis. Limnology and Oceanography, 26(5): 891–897, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Kingsford:1985:FAD

[KC85] M. J. Kingsford and J. H. Choat. The fauna associated with drift algae captured with a plankton-mesh purse seine net. Limnology and Oceanography, 30(3):618–630, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1989:NFC

[KCB89] D. A. Kiefer, W. S. Chamberlin, and C. R. Booth. Natural fluorescence of chlorophyll a: Relationship to photosynthesis and chlorophyll concentration in the western South Pacific gyre. Limnology and Oceanography, 34(5):868–881, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Koop:1982:MFB

[KCN82] K. Koop, R. A. Carter, and R. C. Newell. Mannitol-fermenting bacteria as evidence for export from kelp beds. Limnology and Oceanography, 27(5):950–954, September 1982. CODEN LI-OCAH. ISSN 0024-3590.

Krueger:1981:EIP

[KD81] Donald A. Krueger and Stanley I. Dodson. Embryological induction and predation ecology in Daphnia pulex. Limnology and Oceanography, 26(2):219–223, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Kuwabara:1986:AGR

[KDC86] James S. Kuwabara, James A. Davis, and Cecily C. Y. Chang. Algal growth response to particle-bound orthophosphate and zinc. Limnology and Oceanography, 31(3):503–511, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Kleppel:1989:CPA

[KDR89] G. S. Kleppel, R. E. Dodge, and C. J. Reese. Changes in pigmentation associated with the bleaching of stony corals. Limnology and Oceanography, 34(7):1331–1335, November 1989. CODEN LIOCAH. ISSN 0024-3590.

King:1980:RSD

[KE80] Eric W. King and David A. Everitt. A remote sampling device for under-ice water, bottom biota, and sediments. Limnology and Oceanography, 25(5):935–938, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Kelly:1988:TIC

[Kel88] C. A. Kelly. Toward improving comparisons of alkalinity generation in lake basins. Limnology and Oceanography, 33(6 (part 2)): 1635–1637, November 1988. CODEN LIOCAH. ISSN 0024-3590. See reply [Sto88b].

Keller:1989:MET

[Kel89] Aimee A. Keller. Modeling the effects of temperature, light, and nutrients on primary productivity: an empirical and a mechanistic approach compared. Limnology and Oceanography, 34(1):82–95, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Kepkay:1985:KMM

[Kep85] Paul E. Kepkay. Kinetics of microbial manganese oxidation and trace metal binding in sediments: Results from an in situ dialysis technique. Limnology and Oceanography, 30(4):713–726, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Kerfoot:1988:DSI

[Ker88] W. Charles Kerfoot. Defensive spines: Inverse relationship between coefficients of variation and size. Limnology and Oceanography, 33(6):1412–1429, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Kana:1988:ZCS

[KGGW88] Todd M. Kana, Patricia M. Glibert, Ralf Goericke, and Nicholas A. Welschmeyer. Zeaxanthin and β-carotene in Synechococcus WH7803 respond differently to irradiance. Limnology and Oceanography, 33(6 (part 2)):1623–1626, November 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Krumgalz:1980:DCI

[KH80] Boris S. Krumgalz and Rella Holzer. On the determination of Ca^{2+} ion concentration in seawater. Limnology and Oceanography, 25 (2):367–370, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:1986:MRA

[KH86a] David L. Kirchman and Robert E. Hodson. Metabolic regulation of amino acid uptake in marine waters. Limnology and Oceanography, 31(2):339–350, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Knoechel:1986:CFR

[KH86b] Roy Knoechel and L. Blair Holtby. Cladoceran filtering rate: body length relationships for bacterial and large algal particles. Limnology and Oceanography, 31(1):195–199, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Knoechel:1986:CVB

[KH86c] Roy Knoechel and L. Blair Holtby. Construction and validation of a body-length-based model for the prediction of cladoceran community filtering rates. Limnology and Oceanography, 31(1):1–16, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Kanda:1988:AUS

[KH88a] Jota Kanda and Akihiko Hattori. Ammonium uptake and synthesis of cellular nitrogenous macromolecules in phytoplankton. Limnology and Oceanography, 33(6 (part 2)):1568–1579, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Kilham:1988:CEM

[KH88b] Peter Kilham and Robert E. Hecky. Comparative ecology of marine and freshwater phytoplankton. Limnology and Oceanography, 33(4 (part 2)):776–795, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Kilham:1982:APR

[Kil82] Peter Kilham. Acid precipitation: Its role in the alkalization of a lake in Michigan. Limnology and Oceanography, 27(5):856–867, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Kimmerer:1987:TSP

[Kim87] W. J. Kimmerer. The theory of secondary production calculations for continuously reproducing populations. Limnology and Oceanography, 32(1):1–13, January 1987. CODEN LIOCAH. ISSN 0024-3590.

King:1983:SRG

[Kin83] Gary M. King. Sulfate reduction in Georgia salt marsh soils: an evaluation of pyrite formation by use of ^{35}S and ^{55}Fe tracers. Limnology and Oceanography, 28(5):987–995, September 1983. CO-DEN LIOCAH. ISSN 0024-3590.

King:1988:PSR

[Kin88] Gary M. King. Patterns of sulfate reduction and the sulfur cycle in a South Carolina salt marsh. Limnology and Oceanography, 33 (3):376–390, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Kirchman:1983:PBA

[Kir83] David Kirchman. The production of bacteria attached to particles suspended in a freshwater pond. Limnology and Oceanography, 28 (5):858–872, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Kirk:1984:DRB

[Kir84] J. T. O. Kirk. Dependence of relationship between inherent and apparent optical properties of water on solar altitude. Limnology and Oceanography, 29(2):350–356, March 1984. CODEN LI-OCAH. ISSN 0024-3590.

Kirk:1985:WFP

[Kir85] Kevin L. Kirk. Water flows produced by Daphnia and Diaptomus: Implications for prey selection by mechanosensory predators. Limnology and Oceanography, 30(3):679–686, May 1985. CODEN LI-OCAH. ISSN 0024-3590.

Kirchman:1988:BRM

[Kir88] David Kirchman. Book review: Microbes in the sea (M. A. Sleigh [ed.]). Limnology and Oceanography, 33(6 (part 2)):1649–1650, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Kirk:1989:ULS

[Kir89] J. T. O. Kirk. The upwelling light stream in natural waters. Limnology and Oceanography, 34(8):1410–1425, December 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Knowlton:1989:CMB

[KJP89] Matthew F. Knowlton, John R. Jones, and Bruce D. Perkins. A comment on “The measurement of bacterial chlorophyll and algal chlorophyll-a in natural waters” (Caraco and Puccoon). Limnology and Oceanography, 34(7):1381–1382, November 1989. CODEN LIOCAH. ISSN 0024-3590. See reply [CP89].

Kitchell:1980:SSP

[KK80] Jennifer A. Kitchell and James F. Kitchell. Size-selective predation, light transmission, and oxygen stratification: Evidence from the recent sediments of manipulated lakes. Limnology and Oceanography, 25(3):389–402, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Kilham:1986:HRR

[KKH86] Peter Kilham, Susan S. Kilham, and Robert E. Hecky. Hypothesized resource relationships among African planktonic diatoms. Limnology and Oceanography, 31(6):1169–1181, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Kaplan:1980:PDO

[KLB80] Louis A. Kaplan, Richard A. Larson, and Thomas L. Bott. Patterns of dissolved organic carbon in transport. Limnology and Oceanography, 25(6):1034–1043, November 1980. CODEN LI-OCAH. ISSN 0024-3590.

Knowles:1981:NOC

[KLC81] Roger Knowles, David R. S. Lean, and Yiu Kwok Chan. Nitrous oxide concentrations in lakes: Variations with depth and time. Limnology and Oceanography, 26(5):855–866, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Kling:1988:CTD

[Kli88] George W. Kling. Comparative transparency, depth of mixing, and stability of stratification in lakes of Cameroon, West Africa. Limnology and Oceanography, 33(1):27–40, January 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Kemp:1986:CMM

[KLJ86] W. Michael Kemp, Marlon R. Lewis, and Thomas W. Jones. Comparison of methods for measuring production by the submersed macrophyte, Potamogeton perfoliatus L. Limnology and Oceanography, 31(6):1322–1334, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Klotz:1985:FCP

[Klo85] R. L. Klotz. Factors controlling phosphorus limitation in stream sediments. Limnology and Oceanography, 30(3):543–553, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Krueger:1980:CCI

[KM80] Charles C. Krueger and Frank B. Martin. Computation of confidence intervals for the size-frequency (Hynes) method of estimating secondary production. Limnology and Oceanography, 25(4): 773–777, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Knauer:1981:PPC

[KM81] George A. Knauer and John H. Martin. Primary production and carbon-nitrogen fluxes in the upper 1,500 m of the northeast Pacific. Limnology and Oceanography, 26(1):181–186, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1983:SSS

[KM83] Dale A. Kiefer and B. Greg Mitchell. A simple, steady state description of phytoplankton growth based on absorption cross section and quantum efficiency. Limnology and Oceanography, 28(4): 770–776, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Kuivila:1984:OMD

[KM84] Kathryn M. Kuivila and James W. Murray. Organic matter diagenesis in freshwater sediments: The alkalinity and total CO_{2} balance and methane production in the sediments of Lake Washington. Limnology and Oceanography, 29(6):1218–1230, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Kimmerer:1987:GMS

[KM87] W. J. Kimmerer and A. D. McKinnon. Growth, mortality, and secondary production of the copepod Acartia tranteri in Westemport Bay, Australia. Limnology and Oceanography, 32(1):14–28, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Klump:1989:SNR

[KM89] J. Val Klump and Christopher S. Martens. The seasonality of nutrient regeneration in an organic-rich coastal sediment: Kinetic modeling of changing pore-water nutrient and sulfate distributions. Limnology and Oceanography, 34(3):559–577, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Kuivila:1988:MCS

[KMD^{+}88] K. M. Kuivila, J. W. Murray, A. H. Devol, M. E. Lidstrom, and C. E. Reimers. Methane cycling in the sediments of Lake Washington. Limnology and Oceanography, 33(4):571–581, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Kamykowski:1988:OMC

[KMK88] Daniel Kamykowski, Steven A. McCollum, and Gary J. Kirkpatrick. Observations and a model concerning the translational velocity of a photosynthetic marine dinoflagellate under variable environmental conditions. Limnology and Oceanography, 33(1): 66–78, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Komar:1981:ASR

[KMSF81] Paul D. Komar, Alan P. Morse, Lawrence F. Small, and Scott W. Fowler. An analysis of sinking rates of natural copepod and euphausiid fecal pellets. Limnology and Oceanography, 26(1):172– 180, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Keen:1981:CIB

[KN81] Robert Keen and Raja Nassar. Confidence intervals for birth and death rates estimated with the egg-ratio technique for natural populations of zooplankton. Limnology and Oceanography, 26(1):131– 142, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1982:BRR

[Koh82] Alan J. Kohn. Book review: Rhoads, D. C., and R. A. Lutz [Eds.] 1980. Skeletal growth of aquatic organisms. Plenum Publ. Corp, New York. 750 p. $47.50. Limnology and Oceanography, 27(1): 197–198, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1984:BRB

[Koh84] Alan J. Kohn. Book review: Barnes, D. J. [Ed.] 1983. Perspectives on coral reefs. Australian Inst. Mar. Sci. Contrib. 200 AIMS. B. Clouston, Manuka, ACT 2603, Australia. 277 p. Limnology and Oceanography, 29(6):1348, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Kohn:1987:LCR

[Koh87] Alan J. Kohn. Living coral reefs of the world. Limnology and Oceanography, 32(4):1016, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Konopka:1982:PEM

[Kon82] Allan Konopka. Physiological ecology of a metalimnetic Oscillatoria rubescens population. Limnology and Oceanography, 27(6): 1154–1161, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Konopka:1989:MCH

[Kon89] Allan Konopka. Metalimnetic cyanobacteria in hard-water lakes: Buoyancy regulation and physiological state. Limnology and Oceanography, 34(7):1174–1184, November 1989. CODEN LI-OCAH. ISSN 0024-3590.

Koons:1985:ASL

[Koo85] C. B. Koons. The American Society of Limnology and Oceanography, Inc. sustaining member 1985. Limnology and Oceanography, 30(6):1351–1354, November 1985. CODEN LIOCAH. ISSN 0024-

Kortmann:1980:BAC

[Kor80] Robert W. Kortmann. Benthic and atmospheric contributions to the nutrient budgets of a soft-water lake. Limnology and Oceanography, 25(2):229–239, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Kishino:1986:LUE

[KOTeI86] Motoaki Kishino, Noboru Okami, Masayuki Takahashi, and Shun ei Ichimura. Light utilization efficiency and quantum yield of phytoplankton in a thermally stratified sea. Limnology and Oceanography, 31(3):557–566, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Kristiansen:1982:PLP

[KP82] S. Kristiansen and E. Paasche. Preparation of ^{15}N-labeled phytoplankton samples for optical emission spectrometry. Limnology and Oceanography, 27(2):373–375, March 1982. CODEN LI-OCAH. ISSN 0024-3590.

Kerr:1988:AFY

[KR88] S. R. Kerr and R. A. Ryder. The applicability of fish yield indices in freshwater and marine ecosystems. Limnology and Oceanography, 33(4 (part 2)):973–981, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Kelly:1989:RCU

[KR89] C. A. Kelly and J. W. M. Rudd. Reply to comment by Urban and Baker. Limnology and Oceanography, 34(6):1146–1148, September 1989. CODEN LIOCAH. ISSN 0024-3590. See [UB89].

Kelly:1982:PIB

[KRCS82] C. A. Kelly, John W. M. Rudd, R. B. Cook, and D. W. Schindler. The potential importance of bacterial processes in regulating rate of lake acidification. Limnology and Oceanography, 27(5):868–882, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Kelly:1984:ELA

[KRFS84] C. A. Kelly, J. W. M. Rudd, A. Furutani, and D. W. Schindler. Effects of lake acidification on rates of organic matter decomposition in sediments. Limnology and Oceanography, 29(4):687–694, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Kristiansen:1983:TON

[Kri83] S. Kristiansen. The temperature optimum of the nitrate reductase assay for marine phytoplankton. Limnology and Oceanography, 28 (4):776–780, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Krumbein:1980:BRN

[Kru80] W. E. Krumbein. Book review: Nissenbaum, A. [Ed.] 1980. Hypersaline brines and evaporilic environments. Developments in Sedimentology 28. Elsevier Sci. Publ. Co., Amsterdam and New York, x + 270 p. $48.75. Limnology and Oceanography, 25(5):967–968, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Krumbein:1982:BRE

[Kru82] W. E. Krumbein. Book review: Ehrlich, H. L. 1981. Geomicrobiology. Marcel Dekker, Inc. N.Y. 408 p. $29.75. Limnology and Oceanography, 27(5):984–985, September 1982. CODEN LI-OCAH. ISSN 0024-3590.

Koehl:1981:CFC

[KS81] M. A. R. Koehl and J. Rudi Strickier. Copepod feeding currents: Food capture at low Reynolds number. Limnology and Oceanography, 26(6):1062–1073, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Kiefer:1982:SAM

[KS82a] Dale A. Kiefer and Janice Beeler SooHoo. Spectral absorption by marine particles of coastal waters of Baja California. Limnology and Oceanography, 27(3):492–499, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Kijowski:1982:DTR

[KS82b] Wojciech Kijowski and Paul A. Steudler. Determination of total reduced sulfur in natural waters. Limnology and Oceanography, 27 (5):975–978, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Kanda:1985:NUN

[KSH85] Jota Kanda, Toshiro Saino, and Akihiko Hattori. Nitrogen uptake by natural populations of phytoplankton and primary production in the Pacific Ocean: Regional variability of uptake capacity. Limnology and Oceanography, 30(5):987–999, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Kanda:1988:NNP

[KSH88] Jota Kanda, Toshiro Saino, and Akihiko Hattori. Nitrogen nutrition and physiological state of natural populations of phytoplankton in surface waters of the western Pacific Ocean. Limnology and Oceanography, 33(6 (part 2)):1580–1585, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Kjerfve:1981:EMF

[KSP^{+}81] Bjorn Kjerfve, L. Harold Stevenson, Jeffrey A. Proehl, Thomas H. Chrzanowski, and Wiley M. Kitchens. Estimation of material fluxes in an estuarine cross section: a critical analysis of spatial measurement density and errors. Limnology and Oceanography, 26 (2):325–335, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Katz:1981:SGH

[KSTGB81] Amitai Katz, Avraham Starinsky, Nurit Taitel-Goldman, and Michael Beyth. Solubilities of gypsum and halite in the Dead Sea and in its mixtures with seawater. Limnology and Oceanography, 26(4):709–716, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Kelly:1983:LAV

[KTM83] Mahlon G. Kelly, Niels Thyssen, and Bjarne Moeslund. Light and the annual variation of oxygen- and carbon-based measurements of productivity in a macrophyte-dominated river. Limnology and Oceanography, 28(3):503–515, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Karl:1986:DAI

[KW86] David M. Karl and Christopher D. Winn. Does adenine incorporation into nucleic acids measure total microbial production?: a response to comments by Fuhrman et al. Limnology and Oceanography, 31(6):1384–1394, November 1986. CODEN LIOCAH. ISSN 0024-3590. See reply [FDKM86].

Kokkinakis:1987:NUP

[KW87] Steven A. Kokkinakis and Patricia A. Wheeler. Nitrogen uptake and phytoplankton growth in coastal upwelling regions. Limnology and Oceanography, 32(5):1112–1123, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Lampert:1987:BRT

[Lam87] Winfried Lampert. Book review: Trophic interactions within aquatic ecosystems (D. G. Meyers and J. R. Strickler [eds.]). Limnology and Oceanography, 32(3):778–779, May 1987. CODEN LI-OCAH. ISSN 0024-3590.

Landry:1980:DPC

[Lan80] M. R. Landry. Detection of prey by Calanus pacificus: Implications of the first antennae. Limnology and Oceanography, 49(6): 545–549, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Landers:1982:ENS

[Lan82] Dixon H. Landers. Effects of naturally senescing aquatic macrophytes on nutrient chemistry and chlorophyll-a of surrounding waters. Limnology and Oceanography, 27(3):428–439, May 1982. CO-DEN LIOCAH. ISSN 0024-3590.

Landry:1983:DMC

[Lan83] M. R. Landry. The development of marine calanoid copepods with comment on the isochronal rule. Limnology and Oceanography, 28 (4):614–624, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Laws:1984:IDM

[Law84] Edward Laws. Isotope dilution models and the mystery of the vanishing ^{15}N. Limnology and Oceanography, 29(2):379–386, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Laws:1985:AMN

[Law85] Edward A. Laws. Analytic models of NH_{4}^{+} uptake and regeneration experiments. Limnology and Oceanography, 30(6):1340–1350, November 1985. CODEN LIOCAH. ISSN 0024-3590.

LaZerte:1980:DHO

[LaZ80] Bruce D. LaZerte. The dominating higher order vertical modes of the internal seiche in a small lake. Limnology and Oceanography, 25(5):846–854, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Laws:1980:NLL

[LB80] Edward A. Laws and T. T. Bannister. Nutrient- and light-limited growth of Thalassiosira fluviatilis in continuous culture, with implications for phytoplankton growth in the ocean. Limnology and Oceanography, 25(3):457–473, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Lancelot:1984:AHB

[LB84] Christiane Lancelot and Gilles Billen. Activity of heterotrophic bacteria and its coupling to primary production during the spring phytoplankton bloom in the southern bight of the North Sea. Limnology and Oceanography, 29(4):721–730, July 1984. CODEN LI-OCAH. ISSN 0024-3590.

Lugo:1988:FWF

[LBB88] Ariel E. Lugo, Sandra Brown, and Mark M. Brinson. Forested wetlands in freshwater and salt-water environments. Limnology and Oceanography, 33(4 (part 2)):894–909, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Laws:1988:PPD

[LBZC88] Edward A. Laws, Paul K. Bienfang, David A. Ziemann, and Lytha D. Conquest. Phytoplankton population dynamics and the fate of production during the spring bloom in Auke Bay, Alaska. Limnology and Oceanography, 33(1):57–67, January 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Leavitt:1989:WLE

[LCK89] P. R. Leavitt, S. R. Carpenter, and J. F. Kitchell. Whole-lake experiments: The annual record of fossil pigments and zooplankton. Limnology and Oceanography, 34(4):700–717, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Lee:1980:GTS

[LCP80] David Robert Lee, John A. Cherry, and John F. Pickens. Groundwater transport of a salt tracer through a sandy lakebed. Limnology and Oceanography, 25(1):45–61, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Lai:1988:TCR

[LCPL88] D. Lai, Y. Chung, T. Platt, and T. Lee. Twin cosmogenic radiotracer studies of phosphorus recycling and chemical fluxes in the upper ocean. Limnology and Oceanography, 33(6 (part 2)): 1559–1567, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1986:SAC

[LD86] C. J. Lorenzen and J. Newton Downs. The specific absorption coefficients of chlorophyllide a and pheophorbide a in 90% acetone, and comments on the fluorometric determination of chlorophyll and pheopigments. Limnology and Oceanography, 31(2):449–452, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Laws:1987:HPG

[LDR87] E. A. Laws, G. R. DiTullio, and D. G. Redalje. High phytoplankton growth and production rates in the North Pacific subtropical gyre. Limnology and Oceanography, 32(4):905–918, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Legendre:1983:MPD

[LDYY83] Louis Legendre, Serge Demers, Clarice M. Yentsch, and Charles S. Yentsch. The ^{14}C method: Patterns of dark CO_{2} fixation and DCMU correction to replace the dark bottle. Limnology and Oceanography, 28(5):996–1003, September 1983. CODEN LI-OCAH. ISSN 0024-3590.

LeBorgne:1982:ZPE

[Le 82] Robert Le Borgne. Zooplankton production in the eastern tropical Atlantic Ocean: Net growth efficiency and P:B in terms of carbon, nitrogen, and phosphorus. Limnology and Oceanography, 27(4): 681–698, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1983:SPD

[LE83] John T. Lehman and W. T. Edmondson. The seasonality of phosphorus deposition in Lake Washington. Limnology and Oceanography, 28(4):796–800, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Lopez:1989:FDO

[LE89] Glenn Lopez and Ragnar Elmgren. Feeding depths and organic absorption for the deposit feeding benthic amphipods Pontoporera affinis and Pontoporeia femorata. Limnology and Oceanography, 34(6):982–991, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Leckie:1985:BRM

[Lec85] James O. Leckie. Book review: Morel, F. M. M. 1983. Principles of aquatic chemistry. John Wiley & Sons, Inc., Somerset, N.J. 446 p. $49.95. Limnology and Oceanography, 30(2):450, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Lee:1980:GSI

[Lee80] David Robert Lee. Groundwater-solute influx. Limnology and Oceanography, 25(1):183–186, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1980:RCN

[Leh80] John T. Lehman. Release and cycling of nutrients between planktonic algae and herbivores. Limnology and Oceanography, 25(4): 620–632, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1986:GUL

[Leh86] John T. Lehman. The goal of understanding in limnology. Limnology and Oceanography, 31(5):1160–1166, September 1986. CO-DEN LIOCAH. ISSN 0024-3590.

Lehman:1988:EPA

[Leh88a] John T. Lehman. Ecological principles affecting community structure and secondary production by zooplankton in marine and freshwater environments. Limnology and Oceanography, 33(4 (part 2)):931–945, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1988:GPE

[Leh88b] John T. Lehman. Good Professor Edmondson. Limnology and Oceanography, 33(6):1234–1240, November 1988. CODEN LI-OCAH. ISSN 0024-3590.

Lehman:1988:HML

[Leh88c] John T. Lehman. Hypolimnetic metabolism in Lake Washington: Relative effects of nutrient load and food web structure on lake productivity. Limnology and Oceanography, 33(6):1334–1347, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1989:BRC

[Leh89] John T. Lehman. Book review: Complex interactions in lake communities (S. R. Carpenter [ed.]). Limnology and Oceanography, 34 (2):487–489, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Lemmin:1987:SDI

[Lem87] U. Lemmin. The structure and dynamics of internal waves in baldeggersee. Limnology and Oceanography, 32(1):43–61, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Lemmin:1989:DHT

[Lem89] U. Lemmin. Dynamics of horizontal turbulent mixing in a nearshore zone of Lake Geneva. Limnology and Oceanography, 34(2):420–434, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1982:CPB

[Lew82a] William M. Lewis, Jr. Changes in pH and buffering capacity of lakes in the Colorado Rockies. Limnology and Oceanography, 27 (1):167–172, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1982:VED

[Lew82b] William M. Lewis, Jr. Vertical eddy diffusivities in a large tropical lake. Limnology and Oceanography, 27(1):161–163, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1983:CAM

[Lew83a] William M. Lewis, Jr. Collection of airborne materials by a water surface. Limnology and Oceanography, 28(6):1242–1246, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1983:THM

[Lew83b] William M. Lewis, Jr. Temperature, heat, and mixing in Lake Valencia, Venezuela. Limnology and Oceanography, 28(2):273–286, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1987:MGS

[Lew87] John B. Lewis. Measurements of groundwater seepage flux onto a coral reef: Spatial and temporal variations. Limnology and Oceanography, 32(5):1165–1169, September 1987. CODEN LI-OCAH. ISSN 0024-3590.

Lampert:1986:PCG

[LFRT86] Winfried Lampert, Walter Fleckner, Hakumat Rai, and Barbara E. Taylor. Phytoplankton control by grazing zooplankton: a study on the spring clear-water phase. Limnology and Oceanography, 31(3):478–490, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Lambert:1988:RRT

[LG88] André Lambert and Federico Giovanoli. Records of riverborne turbidity currents and indications of slope failures in the Rhone delta of Lake Geneva. Limnology and Oceanography, 33(3):458– 468, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Legendre:1989:NPE

[LG89] Louis Legendre and Michel Gosselin. New production and export of organic matter to the deep ocean: Consequences of some recent discoveries. Limnology and Oceanography, 34(7):1374–1380, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Li:1980:PCP

[LGM80] W. K. W. Li, H. E. Glover, and I. Morris. Physiology of carbon photoassimilation by Oscillatoria thiebautii in the Caribbean Sea.

Limnology and Oceanography, 25(3):447–456, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Luther:1985:PAS

[LGV85] George W. Luther III, Anne E. Giblin, and Richard Varsolona. Polarographic analysis of sulfur species in marine porewaters. Limnology and Oceanography, 30(4):727–736, July 1985. CODEN LI-OCAH. ISSN 0024-3590.

Larson:1980:OPD

[LH80] Richard A. Larson and John M. Hufnal, Jr. Oxidative polymerization of dissolved phenols by soluble and insoluble inorganic species. Limnology and Oceanography, 49(6):505–512, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Logan:1987:AMG

[LH87] Bruce E. Logan and James R. Hunt. Advantages to microbes of growth in permeable aggregates in marine systems. Limnology and Oceanography, 32(5):1034–1048, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Landry:1984:EFA

[LHF^{+}84] M. R. Landry, R. P. Hassett, V. Fagerness, J. Downs, and C. J. Lorenzen. Effect of food acclimation on assimilation efficiency of Calanus pacificus. Limnology and Oceanography, 29(2):361–364, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Lee:1980:SDS

[LHH80] Cindy Lee, Robert W. Howarth, and Brian L. Howes. Sterols in decomposing Spartina alterniflora and the use of ergosterol in estimating the contribution of fungi to detrital nitrogen. Limnology and Oceanography, 25(2):290–303, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Lesack:1984:TCN

[LHM84] Lance F. W. Lesack, Robert E. Hecky, and John M. Melack. Transport of carbon, nitrogen, phosphorus, and major solutes in the Gambia River, West Africa. Limnology and Oceanography, 29(4): 816–830, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Li:1983:CEE

[Li83] W. K. W. Li. Consideration of errors in estimating kinetic parameters based on Michaelis–Menten formalism in microbial ecology. Limnology and Oceanography, 28(1):185–190, January 1983. CO-DEN LIOCAH. ISSN 0024-3590.

Lemmin:1987:DBC

[LI87] U. Lemmin and D. M. Imboden. Dynamics of bottom currents in a small lake. Limnology and Oceanography, 32(1):62–75, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Lidstrom:1983:MCF

[Lid83] Mary E. Lidstrom. Methane consumption in framvaren, an anoxic marine fjord. Limnology and Oceanography, 28(6):1247–1251, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Lind:1987:STV

[Lin87] Owen T. Lind. Spatial and temporal variation in hypolimnetic oxygen deficits of a multidepression lake. Limnology and Oceanography, 32(3):740–744, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Liss:1985:BRN

[Lis85] P. S. Liss. Book review: National Research Council. 1983. Changing climate. Report of the carbon dioxide assessment committee. National Academy Press, Washington, D.C. 496 p. [Typewriter composition.] $29.50 (softcover). Limnology and Oceanography, 30 (1):236–237, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Liu:1989:ETP

[LK89] Kon-Kee Liu and Isaac R. Kaplan. The eastern tropical Pacific as a source of 15N-enriched nitrate in seawater off southern California. Limnology and Oceanography, 34(5):820–830, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Lodge:1989:PRB

[LKS89] David M. Lodge, David P. Krabbenhoft, and Robert G. Striegl. A positive relationship between groundwater velocity and submersed macrophyte biomass in sparkling Lake Wisconsin. Limnology and Oceanography, 34(1):235–239, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Lemon:1981:NOF

[LL81] Edgar Lemon and Donna Lemon. Nitrous oxide in fresh waters of the Great Lakes Basin. Limnology and Oceanography, 26(5): 867–879, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Lion:1982:ATC

[LL82] Leonard W. Lion and James O. Leckie. Accumulation and transport of Cd, Cu, and Pb in an estuarine salt marsh surface microlayer. Limnology and Oceanography, 27(1):111–125, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Levine:1984:DVN

[LL84a] Suzanne N. Levine and William M. Lewis, Jr. Diel variation of nitrogen fixation in Lake Valencia, Venezuela. Limnology and Oceanography, 29(4):887–893, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1984:LRN

[LL84b] William M. Lewis, Jr. and Suzanne N. Levine. The light response of nitrogen fixation in Lake Valencia, Venezuela. Limnology and Oceanography, 29(4):894–900, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Levine:1985:HHN

[LL85a] Suzanne N. Levine and William M. Lewis, Jr. The horizontal heterogeneity of nitrogen fixation in Lake Valencia, Venezuela. Limnology and Oceanography, 30(6):1240–1245, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Loder:1985:COC

[LL85b] T. C. Loder and P. S. Liss. Control by organic coatings of the surface charge of estuarine suspended particles. Limnology and Oceanography, 30(2):418–421, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Lonsdale:1989:EBL

[LL89] D. J. Lonsdale and J. S. Levinton. Energy budgets of latitudinally separated Scottolana canadensis (Copepoda: Harpacticoida). Limnology and Oceanography, 34(2):324–331, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Lapota:1986:BDI

[LLG86] David Lapota, Jon R. Losee, and Mark L. Geiger. Bioluminescence displays induced by pulsed light. Limnology and Oceanography, 31 (4):887–889, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Lemmin:1986:TEI

[LM86] U. Lemmin and C. H. Mortimer. Tests of an extension to internal seiches of Defant’s procedure for determination of surface seiche characteristics in real lakes. Limnology and Oceanography, 31(6): 1207–1231, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Lee:1987:CIF

[LMS87] Cindy Lee, Judith A. McKenzie, and Zurich Michael Sturm. Carbon isotope fractionation and changes in the flux and composition of particulate matter resulting from biological activity during a sediment trap experiment in Lake Greifen, Switzerland. Limnology and Oceanography, 32(1):83–96, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1986:NCP

[LN86] John T. Lehman and Trajče Naumoski. Net community production and hypolimnetic nutrient regeneration in a Michigan lake. Limnology and Oceanography, 31(4):788–797, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Luecke:1981:PFP

[LO81] Chris Luecke and W. John O’Brien. Phototoxicity and fish predation: Selective factors in color morphs in Heterocope. Limnology and Oceanography, 26(3):454–460, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Luecke:1983:EHP

[LO83] Chris Luecke and W. John O’Brien. The effect of Heterocope predation on zooplankton communities in Arctic ponds. Limnology and Oceanography, 28(2):367–377, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Loeb:1981:SMM

[Loe81] Stanford L. Loeb. An in situ method for measuring the primary productivity and standing crop of the epilithic periphyton community in lentic systems. Limnology and Oceanography, 26(2): 394–399, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Longhurst:1984:BRR

[Lon84] Alan Longhurst. Book review: Rowe, G. T. [Ed.] 1983. The sea (Ideas and observations on progress in the study of the seas). V. 8: Deep-sea biology. John Wiley and Sons, New York. 560 p. $69.95. Limnology and Oceanography, 29(6):1347–1348, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Longhurst:1985:BRR

[Lon85] Alan R. Longhurst. Book review: Rona, P. A., K. Boström, L. Laubier, and K. L. Smith, Jr. [Eds.] 1984. Hydrothermal processes at seafloor spreading centers. NATO Conf. Ser 4: Mar. Sci. V. 12.

Plenum Press, London and New York. 796 p. [Typewriter composition.] $110.00. Limnology and Oceanography, 30(1):235–236, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Longhurst:1986:BRD

[Lon86] Alan R. Longhurst. Book review: Deacon, G. 1985. The Antarctic Circumpolar Ocean. Stud. Polar Res. Cambridge Univ. Press, N.Y. 180 p. $24.95. Limnology and Oceanography, 31(4):908, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Lopez:1988:CEM

[Lop88] Glenn R. Lopez. Comparative ecology of the macrofauna of freshwater and marine muds. Limnology and Oceanography, 33(4 (part 2)):946–962, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1980:UCS

[Lor80] Marc W. Lorenzen. Use of chlorophyll–Secchi disk relationships. Limnology and Oceanography, 25(2):371–372, March 1980. CO-DEN LIOCAH. ISSN 0024-3590.

Lean:1981:PRL

[LP81] D. R. S. Lean and F. R. Pick. Photosynthetic response of lake plankton to nutrient enrichment: a test for nutrient limitation. Limnology and Oceanography, 26(6):1001–1019, November 1981. CODEN LIOCAH. ISSN 0024-3590.

LaFrance:1985:LCP

[LR85] Katherine LaFrance and Ernest Ruber. The life cycle and productivity of the amphipod Gammarus mucronatus on a northern Massachusetts salt marsh. Limnology and Oceanography, 30(5): 1067–1077, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Laws:1984:HPG

[LRH^{+}84] E. A. Laws, D. G. Redalje, L. W. Haas, P. K. Bienfang, R. W. Eppley, W. G. Harrison, D. M. Karl, and J. Marra. High phytoplankton growth and production rates in oligotrophic Hawaiian coastal waters. Limnology and Oceanography, 29(6):1161–1169, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Lynch:1981:PEP

[LS81] Michael Lynch and Joseph Shapiro. Predation, enrichment, and phytoplankton community structure. Limnology and Oceanography, 26(1):86–102, January 1981. CODEN LIOCAH. ISSN 0024-

LaZerte:1982:SCI

[LS82a] Bruce D. LaZerte and James E. Szalados. Stable carbon isotope ratio of submerged freshwater macrophytes. Limnology and Oceanography, 27(3):413–418, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1982:PDP

[LS82b] John T. Lehman and Craig D. Sandgren. Phosphorus dynamics of the procaryotic nannoplankton in a Michigan lake. Limnology and Oceanography, 27(5):828–838, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Landon:1983:DHC

[LS83] M. Steven Landon and Richard H. Stasiak. Daphnia hemoglobin concentration as a function of depth and oxygen availability in Arco Lake, Minnesota. Limnology and Oceanography, 28(4):731– 737, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1984:MES

[LS84] John T. Lehman and Donald Scavia. Measuring the ecological significance of microscale nutrient patches. Limnology and Oceanography, 29(1):214–216, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Lehman:1985:SSR

[LS85] John T. Lehman and Craig D. Sandgren. Species-specific rates of growth and grazing loss among freshwater algae. Limnology and Oceanography, 30(1):34–46, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1981:SCS

[LSB81] C. J. Lorenzen, F. R. Shuman, and J. T. Bennett. In situ calibration of a sediment trap. Limnology and Oceanography, 26(3): 580–585, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Lindeboom:1985:NBJ

[LSdKvD85] Han J. Lindeboom, Adri J. J. Sandee, and Hetty A. J. de Klerkv.d. Driessche. A new bell jar/microelectrode method to measure changing oxygen fluxes in illuminated sediments with a microalgal cover. Limnology and Oceanography, 30(3):693–698, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Lemmin:1985:IMS

[LSI^{+}85] U. Lemmin, M. Schurter, D. M. Imboden, Th. Joller, and H. Abegglen. An instrument for measuring small bottom currents in lakes. Limnology and Oceanography, 30(5):1116–1122, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Lazar:1983:CSH

[LSK^{+}83] B. Lazar, A. Starinsky, A. Katz, E. Sass, and S. Ben-Yaakov. The carbonate system in hypersaline solutions: Alkalinity and CaCO_{3} solubility of evaporated seawater. Limnology and Oceanography, 28(5):978–986, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1986:OCC

[LSLW86] William M. Lewis, Jr., James F. Saunders III, Suzanne N. Levine, and Franz H. Weibezahn. Organic carbon in the Caura River, Venezuela. Limnology and Oceanography, 31(3):653–656, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Larsen:1981:SIP

[LSM81] David P. Larsen, Donald W. Schults, and Kenneth W. Malueg. Summer internal phosphorus supplies in Shagawa Lake, Minnesota. Limnology and Oceanography, 26(4):740–753, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Legendre:1988:AHB

[LT88] Pierre Legendre and Marc Troussellier. Aquatic heterotrophic bacteria: Modeling in the presence of spatial autocorrelation. Limnology and Oceanography, 33(5):1055–1067, September 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Luckenbach:1986:SSA

[Luc86] Mark W. Luckenbach. Sediment stability around animal tubes: The roles of hydrodynamic processes and biotic activity. Limnology and Oceanography, 31(4):779–787, July 1986. CODEN LI-OCAH. ISSN 0024-3590.

Luecke:1988:PCT

[Lue88] Chris Luecke. Predation by cutthroat trout on Chaoborus flavicans as revealed by a population budget. Limnology and Oceanography, 33(6):1254–1265, November 1988. CODEN LIOCAH. ISSN 0024-

Lund:1984:BRF

[Lun84] J. W. G. Lund. Book review: Fryxell, G. A. [Ed.] 1983. Survival strategies of the algae. Cambridge Univ. Press, New York and London. 144 p. $32.50. Limnology and Oceanography, 29(2): 442, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1988:PAA

[LUP88] Marlon R. Lewis, Osvaldo Ulloa, and Trevor Platt. Photosynthetic action, absorption, and quantum yield spectra for a natural population of Oscillatoria in the North Atlantic. Limnology and Oceanography, 33(1):92–98, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1981:CMS

[LW81] William M. Lewis and Franz H. Weibezahn. Chemistry of a 7.5-m sediment core from Lake Valencia, Venezuela. Limnology and Oceanography, 26(5):907–924, September 1981. CODEN LI-OCAH. ISSN 0024-3590.

LaBaugh:1984:IUH

[LW84] J. W. LaBaugh and T. C. Winter. The impact of uncertainties in hydrologic measurement on phosphorus budgets and empirical models for two Colorado reservoirs. Limnology and Oceanography, 29(2):322–339, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Lorenzen:1983:SRO

[LWCV83] Carl J. Lorenzen, Nicholas A. Welschmeyer, Andrea E. Copping, and Maria Vernet. Sinking rates of organic particles. Limnology and Oceanography, 28(4):766–769, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Lipschultz:1986:NME

[LWF86] Fredric Lipschultz, Steven C. Wofsy, and Lewis E. Fox. Nitrogen metabolism of the eutrophic Delaware River ecosystem. Limnology and Oceanography, 31(4):701–716, July 1986. CODEN LIOCAH. ISSN 0024-3590.

leBWilliams:1982:TMC

[lWJ82] P. J. leB. Williams and N. W. Jenkinson. A transportable microprocessor-controlled precise Winkler titration suitable for field station and shipboard use. Limnology and Oceanography, 27(3):576–584, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Lynch:1986:MCB

[LWL86] Michael Lynch, Lawrence J. Weider, and Winfried Lampert. Measurement of the carbon balance in Daphnia. Limnology and Oceanography, 31(1):17–34, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Lewis:1985:APA

[LWP85] Marlon R. Lewis, Roderick E. Warnock, and Trevor Platt. Absorption and photosynthetic action spectra for natural phytoplankton populations: Implications for production in the open ocean. Limnology and Oceanography, 30(4):794–806, July 1985. CODEN LI-OCAH. ISSN 0024-3590.

leBWilliams:1989:SIP

[lWR89] P. J. leB. Williams and J. I. Robertson. A serious inhibition problem from a Niskin sampler during plankton productivity studies. Limnology and Oceanography, 34(7):1300–1305, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Lyle:1983:BGC

[Lyl83] Mitchell Lyle. The brown-green color transition in marine sediments: a marker of the Fe(III)-Fe(II) redox boundary. Limnology and Oceanography, 28(5):1026–1033, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Lynch:1983:ESS

[Lyn83] Michael Lynch. Estimation of size-specific mortality rates in zooplankton populations by periodic sampling. Limnology and Oceanography, 28(3):533–545, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Mackin:1984:AAM

[MA84] James E. Mackin and Robert C. Aller. Ammonium adsorption in marine sediments. Limnology and Oceanography, 29(2):250–257, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Macan:1980:BRC

[Mac80] T. T. Macan. Book review: Cummins, K. W., and R. W. Morritt [Eds.] 1978. An introduction to the aquatic insects of North America. Kendall/Hunt Publ. Co., Dubuque, Iowa, xiii + 441 p. $21.95. Limnology and Oceanography, 49(6):580–581, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Mackas:1984:SAP

[Mac84] David L. Mackas. Spatial autocorrelation of plankton community composition in a continental shelf ecosystem. Limnology and Oceanography, 29(3):451–471, May 1984. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1985:BRL

[Mac85] Ferren MacIntyre. Book review: Liss, P. S., and W. G. Slinn [Eds.] 1983. Air-sea exchange of gases and particles. NATO ASI Ser. C: Math, and Phys. Sci. 108. D. Reidel Publ. Co., Dordrecht, Boston, and London. 561 p. [Typewriter composition.] $69.50. Limnology and Oceanography, 30(1):236, January 1985. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1986:FMS

[Mac86] Sally MacIntyre. A flow-measuring system for use in small lakes. Limnology and Oceanography, 31(4):900–906, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Malone:1982:PPC

[Mal82] Thomas C. Malone. Phytoplankton photosynthesis and carbonspecific growth: Light-saturated rates in a nutrient-rich environment. Limnology and Oceanography, 27(2):226–235, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Maly:1983:SFO

[Mal83] Edward J. Maly. Some further observations on diaptomid body size and clutch size relationships. Limnology and Oceanography, 28(1):148–152, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Mann:1980:BRC

[Man80] K. H. Mann. Book review: Cole, H. A. [Organizer] 1979. The assessment of sublethal effects of pollutants in the sea. A Royal Society Discussion Meeting. Reprint of Phil. Trans. Roy. Soc. Ser. B. 286(1015), issued in hardcover by The Royal Society, London, viii + 230 p. £18.50, U.K.; £19.10, elsewhere. Limnology and Oceanography, 49(6):577, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Mann:1983:BRH

[Man83] K. H. Mann. Book review: Hobbie, J. E. [Ed.] 1980. Limnology of tundra ponds. US/IBP Synthesis Ser. V. 13. Dowden, Hutchinson and Ross, Inc., Stroudsburg, Pennsylvania. 514 p. $34.00. Limnology and Oceanography, 28(5):1042, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Mann:1987:ICE

[Man87] K. H. Mann. An introduction to coastal ecology. Limnology and Oceanography, 32(4):1015, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Mann:1988:PUD

[Man88] K. H. Mann. Production and use of detritus in various freshwater, estuarine, and coastal marine ecosystems. Limnology and Oceanography, 33(4 (part 2)):910–930, July 1988. CODEN LI-OCAH. ISSN 0024-3590.

Martin:1983:BRE

[Mar83] J. H. Martin. Book review: Eisler, R. 1981. Trace metal concentrations in marine organisms. Pergamon Press, Inc., Elmsford, New York. 687 p. $120.00. (Typewriter composition). Limnology and Oceanography, 28(3):600, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Mayer:1989:EPE

[May89] Lawrence M. Mayer. Extracellular proteolytic enzyme activity in sediments of an intertidal mudflat. Limnology and Oceanography, 34(6):973–981, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1980:EEO

[MB80] Lynda S. Murphy and Rebecca A. Belastock. The effect of environmental origin on the response of marine diatoms to chemical stress. Limnology and Oceanography, 25(1):160–165, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Mackas:1986:PFS

[MB86] David L. Mackas and Karen E. Burns. Poststarvation feeding and swimming activity in Calanus pacificus and Metridia pacifica. Limnology and Oceanography, 31(2):383–392, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Megard:1989:EAS

[MB89a] Robert O. Megard and Tom Berman. Effects of algae on the secchi transparency of the southeastern Mediterranean Sea. Limnology and Oceanography, 34(8):1640–1655, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Morel:1989:SPA

[MB89b] André Morel and Jean-François Berthon. Surface pigments, algal biomass profiles, and potential production of the euphotic layer: Relationships reinvestigated in view of remote-sensing applications. Limnology and Oceanography, 34(8):1545–1562, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Murty:1980:RAT

[MBT80] T. S. Murty, F. G. Barber, and J. D. Taylor. Role of advective terms in tidally generated residual circulation. Limnology and Oceanography, 49(6):529–533, May 1980. CODEN LIOCAH. ISSN 0024-3590.

McDuff:1982:CSG

[MC82] R. E. McDuff and S. W. Chisholm. The calculation of in situ growth rates of phytoplankton populations from fractions of cells undergoing mitosis: a clarification. Limnology and Oceanography, 26(5):783–788, July 1982. CODEN LIOCAH. ISSN 0024-3590.

McAuley:1986:IVU

[McA86] P. J. McAuley. Isolation of viable uncontaminated Chlorella from green hydra. Limnology and Oceanography, 31(1):222–224, January 1986. CODEN LIOCAH. ISSN 0024-3590.

McBride:1988:PRD

[McB88] T. Patrick McBride. Preparing random distributions of diatom valves on microscope slides. Limnology and Oceanography, 33(6 (part 2)):1627–1629, November 1988. CODEN LIOCAH. ISSN 0024-3590.

McClatchie:1986:TSF

[McC86] Sam McClatchie. Time-series feeding rates of the euphausiid Thysanoessa raschii in a temporally patchy food environment. Limnology and Oceanography, 31(3):469–477, May 1986. CODEN LIOCAH. ISSN 0024-3590.

McKnight:1981:CBP

[McK81] Diane McKnight. Chemical and biological processes controlling the response of a freshwater ecosystem to copper stress: a field study of the CuSO_{4} treatment of Mill Pond Reservoir, Burlington, Massachusetts. Limnology and Oceanography, 26(3):518–531, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Morrow:1989:TCD

[MCK89] J. H. Morrow, W. S. Chamberlin, and D. A. Kiefer. A twocomponent description of spectral absorption by marine particles. Limnology and Oceanography, 34(8):1500–1509, December 1989. CODEN LIOCAH. ISSN 0024-3590.

McLaren:1986:SGC

[McL86] Ian A. McLaren. Is “structural” growth of Calanus potentially exponential? Limnology and Oceanography, 31(6):1342–1346, November 1986. CODEN LIOCAH. ISSN 0024-3590.

McLachlan:1989:WFD

[McL89] Anton McLachlan. Water filtration by dissipative beaches. Limnology and Oceanography, 34(4):774–780, June 1989. CODEN LI-OCAH. ISSN 0024-3590.

McManus:1986:BRE

[McM86] Dean A. McManus. Book review: Emery, K. O., and E. Uchupi. 1984. The geology of the Atlantic Ocean. Springer-Verlag, New York. 1,050 p. + 23 separate charts. $98.00. Limnology and Oceanography, 31(3):669–671, May 1986. CODEN LIOCAH. ISSN 0024-3590.

McCauley:1985:PCG

[MD85] Edward McCauley and John A. Downing. The prediction of cladoceran grazing rate spectra. Limnology and Oceanography, 30(1): 202–212, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Manley:1987:MHC

[MD87] Steven L. Manley and Minoo N. Dastoor. Methyl halide (CH_{3}X) production from the giant kelp, Macrocystis, and estimates of global CH_{3}X production by kelp. Limnology and Oceanography, 32(3):709–715, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Marmorino:1980:TFS

[MDM80] George O. Marmorino, Savas C. Danos, and James S. Maki. Temperature fine-structure of Lake Michigan hypolimnion. Limnology and Oceanography, 25(4):680–699, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Madden:1988:FMC

[MDR88] Christopher J. Madden, John W. Day, Jr., and John M. Randall. Freshwater and marine coupling in estuaries of the Mississippi River deltaic plain. Limnology and Oceanography, 33(4 (part 2)):982–1004, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Menzie:1980:NHM

[Men80] Charles A. Menzie. A note on the Hynes method1 of estimating secondary production. Limnology and Oceanography, 25(4):770– 773, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Menzie:1981:PEC

[Men81] Charles A. Menzie. Production ecology of Cricotopus sylvestris (Fabricius) (Diptera: Chironomidae) in a shallow estuarine cove. Limnology and Oceanography, 26(3):467–481, May 1981. CODEN LIOCAH. ISSN 0024-3590.

McManus:1986:BSS

[MF86] George B. McManus and Jed A. Fuhrman. Bacterivory in seawater studied with the use of inert fluorescent particles. Limnology and Oceanography, 31(2):420–426, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Moloney:1989:GAE

[MF89] Coleen L. Moloney and John G. Field. General allometric equations for rates of nutrient uptake, ingestion, and respiration in plankton organisms. Limnology and Oceanography, 34(7):1290– 1299, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Mague:1980:ERC

[MFHM80] T. H. Mague, E. Friberg, D. J. Hughes, and I. Morris. Extracellular release of carbon by marine phytoplankton; a physiological approach. Limnology and Oceanography, 25(2):262–279, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Mayer:1980:BSP

[MG80] Lawrence M. Mayer and Steven P. Gloss. Buffering of silica and phosphate in a turbid river. Limnology and Oceanography, 25(1): 12–22, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Morris:1981:PPM

[MG81] Ian Morris and Hilary Glover. Physiology of photosynthesis by marine coccoid cyanobacteria — some ecological implications. Limnology and Oceanography, 26(5):957–961, September 1981. CO-DEN LIOCAH. ISSN 0024-3590.

Magnien:1983:DCR

[MG83] Robert E. Magnien and John J. Gilbert. Diel cycles of reproduction and vertical migration in the rotifer Keratella crassa and their influence on the estimation of population dynamics. Limnology and Oceanography, 28(5):957–969, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Matter:1980:VDI

[MH80] William J. Matter and A. J. Hopwood. Vertical distribution of invertebrate drift in a large river. Limnology and Oceanography, 25(6):1117–1121, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Marra:1982:PRP

[MH82] John Marra and Kristina Heinemann. Photosynthesis response by phytoplankton to sunlight variability. Limnology and Oceanography, 27(6):1141–1153, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Marra:1984:CBN

[MH84] John Marra and Kristina Heinemann. A comparison between noncontaminating and conventional incubation procedures in primary production measurements. Limnology and Oceanography, 29(2): 389–392, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1985:DAP

[MH85] L. S. Murphy and E. M. Haugen. The distribution and abundance of phototrophic ultraplankton in the North Atlantic. Limnology and Oceanography, 30(1):47–58, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Miller:1986:TRM

[MH86] Kevin M. Miller and Merrill Heit. A time resolution methodology for assessing the quality of lake sediment cores that are dated by 137Cs. Limnology and Oceanography, 31(6):1292–1300, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Maske:1987:QVA

[MH87] Helmut Maske and Hauke Haardt. Quantitative in vivo absorption spectra of phytoplankton: Detrital absorption and comparison with fluorescence excitation spectra. Limnology and Oceanography, 32(3):620–633, May 1987. CODEN LIOCAH. ISSN 0024-

Moran:1989:FBU

[MH89] Mary Ann Moran and Robert E. Hodson. Formation and bacterial utilization of dissolved organic carbon derived from detrital lignocellulose. Limnology and Oceanography, 34(6):1034–1047, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Miller:1984:PCM

[MHB84] Charles B. Miller, Mark E. Huntley, and Elaine R. Brooks. Postcollection molting rates of planktonic, marine copepods: Measurement, applications, problems. Limnology and Oceanography, 29 (6):1274–1289, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1983:CPC

[MHY83] T. P. Murphy, K. J. Hall, and I. Yesaki. Coprecipitation of phosphate with calcite in a naturally eutrophic lake. Limnology and Oceanography, 28(1):58–69, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Miller:1982:BRV

[Mil82] Charles B. Miller. Book review: van der Spoel, S., and A. C. Pierrot-Bults [Eds.] 1979. Zoogeography and diversity of plankton. John Wiley & Sons, Somerset, NJ. 410 p. $69.95. Limnology and Oceanography, 27(1):195–196, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Millero:1984:CDS

[Mil84a] Frank J. Millero. The conductivity-density-salinity-chlorinity relationships for estuarine waters. Limnology and Oceanography, 29 (6):1317–1321, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Mills:1984:BRE

[Mil84b] Eric L. Mills. Book review: Ernst, W. G., and J. G, Morin [Eds.]. 1982. The environment of the deep sea. Rubey Colloq., V.

2. Prentice-Hall, Inc., Englewood Cliffs, N.J. 371 p. $36.95. Limnology and Oceanography, 29(2):449–450, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Millero:1986:PEW

[Mil86] Frank J. Millero. The pH of estuarine waters. Limnology and Oceanography, 31(4):839–847, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Millero:1987:BRP

[Mil87] F. J. Millero. Book review: Practical estuarine chemistry: a handbook (P. C. Head [ed.]). Limnology and Oceanography, 32(5):1179, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Miller:1984:EST

[MJN84] Douglas C. Miller, Peter A. Jumars, and Arthur R. M. Nowell. Effects of sediment transport on deposit feeding: Scaling arguments. Limnology and Oceanography, 29(6):1202–1217, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Michelson:1989:MDA

[MJSM89] A. R. Michelson, M. E. Jacobson, M. I. Scranton, and J. E. Mackin. Modeling the distribution of acetate in anoxic estuarine sediments. Limnology and Oceanography, 34(4):747–757, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Morgan:1984:PUS

[MK84] Mark D. Morgan and Christopher L. Kitting. Productivity and utilization of the seagrass Halodule wrightii and its attached epiphytes. Limnology and Oceanography, 29(5):1066–1076, September 1984. CODEN LIOCAH. ISSN 0024-3590.

McCarthy:1984:CBN

[MKN84] James J. McCarthy, Warren Kaplan, and John L. Nevins. Chesapeake Bay nutrient and plankton dynamics. 2. Sources and sinks of nitrite. Limnology and Oceanography, 29(1):84–98, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Mopper:1982:DDV

[ML82] Kenneth Mopper and Peter Lindroth. Diel and depth variations in dissolved free amino acids and ammonium in the Baltic Sea determined by shipboard HPLC analysis. Limnology and Oceanography, 27(2):336–347, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Mueller:1989:BOP

[ML89] James L. Mueller and R. Edward Lange. Bio-optical provinces of the Northeast Pacific Ocean: a provisional analysis. Limnology and Oceanography, 34(8):1572–1586, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Morel:1987:GRQ

[MLG87] Andre Morel, Luigi Lazzara, and Jacques Gostan. Growth rate and quantum yield time response for a diatom to changing irradiances (energy and color). Limnology and Oceanography, 32(5): 1066–1084, September 1987. CODEN LIOCAH. ISSN 0024-3590.

McKnight:1980:CCS

[MM80] Diane M. McKnight and Francois M. M. Morel. Copper complexation by siderophores from filamentous blue-green algae. Limnology and Oceanography, 25(1):62–71, January 1980. CODEN LIOCAH. ISSN 0024-3590.

MacIntyre:1982:MEA

[MM82a] Sally MacIntyre and John M. Melack. Meromixis in an equatorial African soda lake. Limnology and Oceanography, 27(4):595–609, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Muscatine:1982:DIN

[MM82b] L. Muscatine and R. E. Marian. Dissolved inorganic nitrogen flux in symbiotic and nonsymbiotic medusae. Limnology and Oceanography, 27(5):910–917, September 1982. CODEN LIOCAH. ISSN 0024-3590.

McCaffrey:1980:RBP

[MMD^{+}80] Richard J. McCaffrey, Allen C. Myers, Earl Davey, George Morrison, Michael Bender, Nile Luedtke, Douglas Cullen, Philip Froelich, and Gary Klinkhammer. The relation between pore water chemistry and benthic fluxes of nutrients and manganese in Narragansett Bay, rhode Island. Limnology and Oceanography, 25 (1):31–44, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Madsen:1986:SPM

[MMG86] Eugene L. Madsen, Mark D. Morgan, and Ralph E. Good. Simultaneous photoreduction and microbial oxidation of iron in a stream in the New Jersey Pinelands. Limnology and Oceanography, 31(4): 832–838, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Meise:1985:AFB

[MMH85] Carol J. Meise, Wayne R. Munns, Jr., and Nelson G. Hairston, Jr. An analysis of the feeding behavior of Daphnia pulex. Limnology and Oceanography, 30(4):862–870, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Muscatine:1981:EDC

[MMM81] L. Muscatine, L. R. McCloskey, and R. E. Marian. Estimating the daily contribution of carbon from zooxanthellae to coral animal respiration. Limnology and Oceanography, 26(4):601–611, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Minas:1986:PUA

[MMP86] Hans Joachim Minas, Monique Minas, and Theodore Train Packard. Productivity in upwelling areas deduced from hydrographic and chemical fields. Limnology and Oceanography, 31(6): 1182–1206, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Morin:1987:APS

[MMR87] Antoine Morin, Timothy A. Mousseau, and Derek A. Roff. Accuracy and precision of secondary production estimates. Limnology and Oceanography, 32(6):1342–1352, November 1987. CODEN LI-OCAH. ISSN 0024-3590.

Mazumder:1988:EFP

[MMTL88] A. Mazumder, D. J. McQueen, W. D. Taylor, and D. R. S. Lean. Effects of fertilization and planktivorous fish (yellow perch) predation on size distribution of particulate phosphorus and assimilated phosphate: Large enclosure experiments. Limnology and Oceanography, 33(3):421–430, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Macdonald:1986:SRS

[MMW86] R. W. Macdonald, F. A. McLaughlin, and C. S. Wong. The storage of reactive silicate samples by freezing. Limnology and Oceanography, 31(5):1139–1142, September 1986. CODEN LIOCAH. ISSN 0024-3590.

McCarthy:1986:SNP

[MN86] James J. McCarthy and John L. Nevins. Sources of nitrogen for primary production in warm-core rings 79-e and 81-d. Limnology and Oceanography, 31(4):690–700, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Mobley:1989:NMC

[Mob89] Curtis D. Mobley. A numerical model for the computation of radiance distributions in natural waters with wind-roughened surfaces. Limnology and Oceanography, 34(8):1473–1483, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Molenaar:1989:BRA

[Mol89] Dee Molenaar. Book review: The art of geology (E. M.. Moore and F. M. Wahl [eds.]). Limnology and Oceanography, 34(2):492, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Monismith:1985:WFM

[Mon85] Stephen G. Monismith. Wind-forced motions in stratified lakes and their effect on mixed-layer shear. Limnology and Oceanography, 30(4):771–783, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Moore:1983:OCC

[Moo83] Johnnie N. Moore. The origin of calcium carbonate nodules forming on Flathead Lake delta, northwestern Montana. Limnology and Oceanography, 28(4):646–654, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Moore:1988:DDP

[Moo88] Marianne V. Moore. Density-dependent predation of early instar Chaoborus feeding on multispecies prey assemblages. Limnology and Oceanography, 33(2):256–268, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Morgan:1980:GPG

[Mor80a] Mark D. Morgan. Grazing and predation of the grass shrimp Palaemonetes pugio. Limnology and Oceanography, 25(5):896–902, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Morris:1980:BRH

[Mor80b] Ian Morris. Book review: Hellebust, J. A., and J. S. Craigie [Eds.]. 1978. Handbook of phycological methods. Physiological and biochemical methods. Cambridge Univ. Press, New York and London, xiv + 512 p. $29.95. Limnology and Oceanography, 25(1):197, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Morgan:1984:AHS

[Mor84] Mark D. Morgan. Acidification of headwater streams in the New Jersey Pinelands: a re-evaluation. Limnology and Oceanography, 29(6):1259–1266, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Morse:1988:BRS

[Mor88a] John W. Morse. Book review: Stumm, W. [ed.]. 1987. Aquatic surface chemistry: Chemical processes at the particle–water interface. John Wiley & Sons, Inc., Somerset, New Jersey. 520 p. $69.95. Limnology and Oceanography, 33(2):311, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Mortimer:1988:DTH

[Mor88b] C. H. Mortimer. Discoveries and testable hypotheses arising from Coastal Zone Color Scanner imagery of southern Lake Michigan. Limnology and Oceanography, 33(2):203–226, March 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Morin:1988:EMF

[MP88] Antoine Morin and Robert H. Peters. Effect of microhabitat features, seston quality, and periphyton on abundance of overwintering black fly larvae in southern Québec. Limnology and Oceanography, 33(3):431–446, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Marshall:1989:GPS

[MP89] C. Tara Marshall and Robert H. Peters. General patterns in the seasonal development of chlorophyll-a for temperate lakes. Limnology and Oceanography, 34(5):856–867, July 1989. CODEN LI-OCAH. ISSN 0024-3590.

Millero:1988:DHS

[MPF88] Frank J. Millero, Tinka Plese, and Marino Fernandez. The dissociation of hydrogen sulfide in seawater. Limnology and Oceanography, 33(2):269–274, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Maresca:1980:HFR

[MPsCS80] Joseph W. Maresca, Jr., Robin R. Padden, Ralph Ta shun Cheng, and Erwin Seibel. High frequency radar measurements of tidal currents flowing through San Pablo Strait, San Francisco Bay. Limnology and Oceanography, 25(5):929–935, September 1980. CO-DEN LIOCAH. ISSN 0024-3590.

Mullin:1984:SNI

[MRE84] Michael M. Mullin, Gregory H. Rau, and Richard W. Eppley. Stable nitrogen isotopes in zooplankton: Some geographic and temporal variations in the North Pacific. Limnology and Oceanography, 29(6):1267–1273, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Maestrini:1986:NLB

[MRLD86] Serge Y. Maestrini, Martin Rochet, Louis Legendre, and Serge Demers. Nutrient limitation of the bottom-ice microalgal biomass (southeastern Hudson Bay, Canadian Arctic). Limnology and Oceanography, 31(5):969–982, September 1986. CODEN LI-OCAH. ISSN 0024-3590.

Morel:1984:BRC

[MS84] André Morel and Alain Sournia. Book review: Calkins, J. [Ed.] 1982. The role of solar ultraviolet radiation in marine ecosystems. NATO Conf. Ser. 4: Mar. Sci. 7. Plenum Publ. Corp. New York and London. 724 p. (Typewriter composition.) $79.50. Limnology and Oceanography, 29(4):901–902, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Meyer:1985:MHF

[MS85a] Judy L. Meyer and Eric T. Schultz. Migrating haemulid fishes as a source of nutrients and organic matter on coral reefs. Limnology and Oceanography, 30(1):146–156, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Meyer:1985:TCG

[MS85b] Judy L. Meyer and Eric T. Schultz. Tissue condition and growth rate of corals associated with schooling fish. Limnology and Oceanography, 30(1):157–166, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Morel:1987:BRL

[MS87] André Morel and Alain Sournia. Book review: Light and photosynthesis in aquatic ecosystems (J. T. O. Kirk). Limnology and Oceanography, 32(1):275–276, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Megard:1980:LSD

[MSBC80] R. O. Megard, J. C. Settles, H. A. Boyer, and W. S. Combs, Jr. Light, Secchi disks, and trophic states. Limnology and Oceanography, 25(2):373–377, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Morris:1981:PPP

[MSG81] I. Morris, A. E. Smith, and H. E. Glover. Products of photosynthesis in phytoplankton off the Orinoco River and in the Caribbean Sea. Limnology and Oceanography, 26(6):1034–1044, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Marcus:1986:RIP

[MSG86] Nancy H. Marcus and Jutta Schmidt-Gengenbach. Recruitment of individuals into the plankton: The importance of bioturbation.

Limnology and Oceanography, 31(1):206–210, January 1986. CO-DEN LIOCAH. ISSN 0024-3590.

Mayzaud:1984:SPS

[MTL84] P. Mayzaud, S. Taguchi, and Ph. Laval. Seasonal patterns of seston characteristics in Bedford Basin, Nova Scotia, relative to zooplankton feeding: a multivariate approach. Limnology and Oceanography, 29(4):745–762, July 1984. CODEN LIOCAH. ISSN 0024-3590.

McKnight:1988:RCD

[MTW^{+}88] Diane M. McKnight, Kevin A. Thorn, Robert L. Wershaw, J. Michael Bracewell, and Graeme W. Robertson. Rapid changes in dissolved humic substances in Spirit Lake and South Fork Castle Lake, Washington. Limnology and Oceanography, 33(6 (part 2)): 1527–1541, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Mulholland:1981:FPO

[Mul81] Patrick J. Mulholland. Formation of particulate organic carbon in water from a southeastern swamp-stream. Limnology and Oceanography, 26(4):790–795, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Mullineaux:1989:VDE

[Mul89] Lauren S. Mullineaux. Vertical distributions of the epifauna on manganese nodules: Implications for settlement and feeding. Limnology and Oceanography, 34(7):1247–1262, November 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Murtaugh:1981:IPM

[Mur81a] Paul A. Murtaugh. Inferring properties of mysid predation from injuries to Daphnia. Limnology and Oceanography, 26(5):811–821, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Murtaugh:1981:SPN

[Mur81b] Paul A. Murtaugh. Selective predation by Neomysis mercedis in Lake Washington. Limnology and Oceanography, 26(3):445–453, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Murray:1983:BRD

[Mur83] James W. Murray. Book review: Drever, J. I. 1982. The geochemistry of natural waters. Prentice-Hall, Inc., Englewood Cliffs, N.J. 388 p. $31.95. Limnology and Oceanography, 28(3):600, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Murphy:1984:PPG

[Mur84a] Michael L. Murphy. Primary production and grazing in freshwater and intertidal reaches of a coastal stream, southeast Alaska. Limnology and Oceanography, 29(4):805–815, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Murtaugh:1984:BRM

[Mur84b] Paul A. Murtaugh. Book review: Morgan, M. D. [Ed.] 1982. Ecology of Mysidacea. Developments in Hydrobiology, V. 10. [Reprinted from Hydrobiologia 93(1/2)]. W. Junk, The Hague. 232 p. $55.00. Limnology and Oceanography, 29(2):442–443, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Murchie:1985:DRG

[Mur85] Scott L. Murchie. ^{210}Pb dating and the recent geologic history of Crystal Bay, Lake Minnetonka, Minnesota. Limnology and Oceanography, 30(6):1154–1170, November 1985. CODEN LI-OCAH. ISSN 0024-3590.

Matrai:1988:PTC

[MV88] Patricia A. Matrai and Russell D. Vetter. Particulate thiols in coastal waters: The effect of light and nutrients on their planktonic production. Limnology and Oceanography, 33(4):624–631, July 1988. CODEN LIOCAH. ISSN 0024-3590.

McCarthy:1982:UDN

[MWB82] James J. McCarthy, D. Wynne, and T. Berman. The uptake of dissolved nitrogenous nutrients by Lake Kinneret (Israel) microplankton. Limnology and Oceanography, 27(4):673–680, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Nagata:1988:MPF

[Nag88] Toshi Nagata. The microflagellate-picoplankton food linkage in the water column of Lake Biwa. Limnology and Oceanography, 33 (4):504–517, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Neale:1989:BOI

[NCY89] Patrick J. Neale, John J. Cullen, and Clarice M. Yentsch. Biooptical inferences from chlorophyll-a fluorescence: What kind of fluorescence is measured in flow cytometry? Limnology and Oceanography, 34(8):1739–1748, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Nizan:1986:ATE

[NDS86] Shulamit Nizan, Chanan Dimentman, and Moshe Shilo. Acute toxic effects of the cyanobacterium Microcystis aeruginosa on Daphnia magna. Limnology and Oceanography, 31(3):497–502, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Nehring:1987:TVP

[Neh87] Dietwart Nehring. Temporal variations of phosphate and inorganic nitrogen compounds in central Baltic deep waters. Limnology and Oceanography, 32(2):494–499, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Newbold:1986:BRB

[New86] J. Denis Newbold. Book review: Barnes, J. R., and G. W. Minshall [Eds.]. 1983. Stream ecology: Application and testing of general ecological theory. Plenum Press, New York and London. 399 p. $55.00. Limnology and Oceanography, 31(3):671–672, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Nalewajko:1983:MPS

[NG83] C. Nalewajko and C. Garside. Methodological problems in the simultaneous assessment of photosynthesis and nutrient uptake in phytoplankton as functions of light intensity and cell size. Limnology and Oceanography, 28(3):591–597, May 1983. CODEN LI-OCAH. ISSN 0024-3590.

Newman:1987:MQD

[NG87] Kathleen A. Newman and Philip M. Gschwend. A method for quantitative determination of volatile organic compounds in marine macroalgae. Limnology and Oceanography, 32(3):702–708, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Nicola:1989:MCS

[Nic89] Ricci Nicola. Microhabitats of ciliates: Specific adaptations to different substrates. Limnology and Oceanography, 34(6):1089– 1097, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Nixon:1988:PEI

[Nix88] Scott W. Nixon. Physical energy inputs and the comparative ecology of lake and marine ecosystems. Limnology and Oceanography, 33(4 (part 2)):1005–1025, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Nowell:1984:FSS

[NJF84] Arthur R. M. Nowell, Peter A. Jumars, and Kristian Fauchald. The foraging strategy of a subtidal and deep-sea deposit feeder. Limnology and Oceanography, 29(3):645–649, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Newton:1987:PCS

[NL87] Philip P. Newton and Peter S. Liss. Positively charged suspended particles: Studies in an iron-rich river and its estuary. Limnology and Oceanography, 32(6):1267–1276, November 1987. CODEN LI-OCAH. ISSN 0024-3590.

Napp:1989:NIM

[NL89] Jeffrey M. Napp and Dennis L. Long. A new isotopic method for measurmg diel grazing rates of marine zooplankton in situ. Limnology and Oceanography, 34(3):618–629, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Nijampurkar:1983:STS

[NMM83] V. N. Nijampurkar, J. M. Martin, and M. Maybeck. Silicon-32 as a tool for studying silicon behavior in estuaries. Limnology and Oceanography, 28(6):1237–1242, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Nielsen:1989:DAE

[NO89] Marianne V. Nielsen and Yngvar Olsen. The dependence of the assimilation efficiency in Daphnia magna on the ^{14}C-labeling period of the food alga Scenedesmus acutus. Limnology and Oceanography, 34(7):1311–1315, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Nowell:1984:BRM

[Now84] Arthur Nowell. Book review: McCall, P. L., and M. J. S. Tevesz [Eds.]. 1982. Animal-sediment relations. The biogenic alteration of sediments. Topics in Geobiology, V. 2. Plenum Publ. Corp., New York. 336 p. $42.50. Limnology and Oceanography, 29(2):448–449, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Nyjfeler:1986:RSK

[NSL86] Urs P. Nyjfeler, Peter H. Santschi, and Yuan-Hui Li. The relevance of scavenging kinetics to modeling of sediment-water interactions in natural waters. Limnology and Oceanography, 31(2):277–292, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Novelli:1987:HDM

[NSM87] Paul C. Novelli, Mary I. Scranton, and Robert H. Michener. Hydrogen distributions in marine sediments. Limnology and Oceanography, 32(3):565–576, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Nurnberg:1984:PIP

[Nür84] Gertrud K. Nürnberg. The prediction of internal phosphorus load in lakes with anoxic hypolimnia. Limnology and Oceanography, 29 (1):111–124, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Nurnberg:1987:CIP

[Nür87] Gertrud K. Nürnberg. A comparison of internal phosphorus loads in lakes with anoxic hypolimnia: Laboratory incubation versus in situ hypolimnetic phosphorus accumulation. Limnology and Oceanography, 32(5):1160–1164, September 1987. CODEN LI-OCAH. ISSN 0024-3590.

Nurnberg:1988:SMP

[Nür88] Gertrud K. Nürnberg. A simple model for predicting the date of fall turnover in thermally stratified lakes. Limnology and Oceanography, 33(5):1190–1195, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1985:CTI

[OAI^{+}85] Akira Otsuki, Morihiro Aizaki, Toshio Iwakuma, Noriko Takamura, Takayuki Hanazato, Takayoshi Kawai, and Masayuki Yasuno. Coupled transformation of inorganic stable carbcn-13 and nitrogen-15 isotopes into higher trophic levels in a eutrophic shallow lake. Limnology and Oceanography, 30(4):820–825, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Olson:1987:HAL

[OBP87] Richard Randolph Olson, Isidro Bosch, and John S. Pearse. The hypothesis of Antarctic larval starvation examined for the asteroid Odontaster validus. Limnology and Oceanography, 32(3):686–690, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Olson:1982:IDC

[OC82] Eric J. Olson and Chen-Tung A. Chen. Interference in the determination of calcium in seawater. Limnology and Oceanography, 27 (2):375–380, March 1982. CODEN LIOCAH. ISSN 0024-3590.

ODoherty:1985:SDC

[O’D85] Erin Claire O’Doherty. Stream-dwelling copepods: Their life history and ecological significance. Limnology and Oceanography, 30 (3):554–564, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Oudot:1988:PSD

[OGMG88] Claude Oudot, Robert Gerard, Pascal Morin, and Itaf Gningue. Precise shipboard determination of dissolved oxygen (Winkler procedure) for productivity studies with a commercial system. Limnology and Oceanography, 33(1):146–150, January 1988. CODEN LIOCAH. ISSN 0024-3590. See [Ami88].

Olivieri:1987:VEP

[OH87] Emilia T. Olivieri and Laurence Hutchings. Volumetric estimates of phytoplankton production in newly upwelled waters of the southern Benguela Current. Limnology and Oceanography, 32(5): 1099–1111, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Ohman:1987:ESR

[Ohm87] M. D. Ohman. Energy sources for recruitment of the subantarctic copepod Neocalanus tonsus. Limnology and Oceanography, 32(6): 1317–1330, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Olsen:1986:DRR

[OJR^{+}86] Yngvar Olsen, Arne Jensen, Helge Reinertsen, Knut Yngve Børsheim, Mikal Heldal, and Arnfinn Langeland. Dependence of the rate of release of phosphorus by zooplankton on the p: C ratio in the food supply, as calculated by a recycling model. Limnology and Oceanography, 31(1):34–44, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Orrett:1987:DOP

[OK87] K. Orrett and D. M. Karl. Dissolved organic phosphorus production in surface seawaters. Limnology and Oceanography, 32(2): 383–395, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Oliver:1981:MCD

[OKG81] R. L. Oliver, A. J. Kinnear, and G. G. Ganf. Measurements of cell density of three freshwater phytoplankters by density gradient centrifugation. Limnology and Oceanography, 26(2):285–294, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Olafsson:1980:TSW

[Óla80] Jón Ólafsson. Temperature structure and water chemistry of the caldera Lake Öskjuvatn, Iceland. Limnology and Oceanography, 25 (5):779–788, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Olson:1980:NAU

[Ols80a] Robert J. Olson. Nitrate and ammonium uptake in Antarctic waters. Limnology and Oceanography, 25(6):1064–1074, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Olson:1980:PIC

[Ols80b] Robert J. Olson. Phosphate interference in the cadmium reduction analysis of nitrate. Limnology and Oceanography, 25(4):758–760, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Olson:1987:SCT

[Ols87] Richard Randolph Olson. In situ culturing as a test of the larval starvation hypothesis for the crown-of-thorns starfish, Acanthaster planci. Limnology and Oceanography, 32(4):895–904, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Olsen:1985:ERR

[OØ85] Yngvar Olsen and Kjetill Østgaard. Estimating release rates of phosphorus from zooplankton: Model and experimental verification. Limnology and Oceanography, 30(4):844–852, July 1985. CO-DEN LIOCAH. ISSN 0024-3590.

Orcutt:1983:DVM

[OP83] John D. Orcutt, Jr. and Karen G. Porter. Diel vertical migration by zooplankton: Constant and fluctuating temperature effects on life history parameters of Daphnia. Limnology and Oceanography, 28(4):720–730, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Oren:1983:BMC

[Ore83a] Aharon Oren. Bacteriorhodopsin-mediated CO_{2} photoassimilation in the Dead Sea. Limnology and Oceanography, 28(1):33–41, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Oren:1983:PDH

[Ore83b] Aharon Oren. Population dynamics of halobacteria in the Dead Sea water column. Limnology and Oceanography, 28(6):1094–1103, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Oren:1985:RDB

[Ore85] Aharon Oren. The rise and decline of a bloom of halobacteria in the Dead Sea. Limnology and Oceanography, 30(4):911–915, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Oren:1982:PDD

[OS82] Aharon Oren and Moshe Shilo. Population dynamics of Dunaliella parva in the Dead Sea. Limnology and Oceanography, 27(2):201– 211, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Otsuki:1987:MNG

[OSML87] A. Otsuki, H. Seki, C. D. McAllister, and C. D. Levings. Measurement of net growth rates of herbivorous benthic animals using periphyton labeled simultaneously with ^{13}C and ^{15}N. Limnology and Oceanography, 32(2):499–503, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Oudot:1989:PAD

[Oud89] Claude Oudot. Precision and accuracy of dissolved oxygen measurements. A reply to the comment of Aminot (1988). Limnology and Oceanography, 34(5):965–966, July 1989. CODEN LIOCAH. ISSN 0024-3590. See [Ami88].

Oliver:1984:DER

[OW84] R. L. Oliver and A. E. Walsby. Direct evidence for the role of light-mediated gas vesicle collapse in the buoyancy regulation of Anabaena flos-aquae (cyanobacteria). Limnology and Oceanography, 29(4):879–886, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Prgzelin:1983:PPM

[PA83] Barbara B. Prgzelin and Alice L. Alldredge. Primary production of marine snow during and after an upwelling event. Limnology and Oceanography, 28(6):1156–1167, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Paasche:1980:SCF

[Paa80] E. Paasche. Silicon content of five marine plankton diatom species measured with a rapid filter method. Limnology and Oceanography, 49(6):474–480, May 1980. CODEN LIOCAH. ISSN 0024-

Pace:1986:EAZ

[Pac86] Michael L. Pace. An empirical analysis of zooplankton community size structure across lake trophic gradients. Limnology and Oceanography, 31(1):45–55, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1984:EFF

[Pae84] Hans W. Paerl. An evaluation of freeze fixation as a phytoplankton preservation method for microautoradiography. Limnology and Oceanography, 29(2):417–426, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1985:MFR

[Pae85] Hans W. Paerl. Microzone formation: Its role in the enhancement of aquatic N_{2} fixation. Limnology and Oceanography, 30(6):1246– 1252, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1988:NPB

[Pae88] Hans W. Paerl. Nuisance phytoplankton blooms in coastal, estuarine, and inland waters. Limnology and Oceanography, 33(4 (part 2)):823–843, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1989:BRD

[Pae89] Hans W. Paerl. Book review: Dominance of bloom-forming cyanobacteria (blue-green algae) (W. F. Vincent [ed.]). Limnology and Oceanography, 34(2):490–492, March 1989. CODEN LI-OCAH. ISSN 0024-3590.

Paffenhofer:1984:DPF

[Paf84] Gustav-Adolf Paffenhöfer. Does Paracalanus feed with a leaky sieve? Limnology and Oceanography, 29(1):155–160, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Pedros-Alio:1987:ECP

[PAGG87] Carlos Pedrós-Alió, Josep M. Gasol, and Ricardo Guerrero. On the ecology of a Cryptomonas phaseolus population forming a metalimnetic bloom in Lake Cisó, Spain: Annual distribution and loss factors. Limnology and Oceanography, 32(2):285–298, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Paine:1980:BRN

[Pai80] R. T. Paine. Book review: Newell, R. C. 1979. Biology of intertidal organisms, 3rd ed. Marine Ecological Surveys, Ltd., Kent, ix + 781 p. £22.00. Limnology and Oceanography, 25(2):383–384, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Paine:1984:BRL

[Pai84a] R. T. Paine. Book review: Levinton, J. S. 1982. Marine ecology. Prentice-Hall, Inc., Englewood Cliffs, N.J. 526 p. $35.95. Limnology and Oceanography, 29(2):443–444, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Paine:1984:BRN

[Pai84b] R. T. Paine. Book review: Nybakken, J. W. 1982. Marine biology: an ecological approach. Harper & Row, N.Y. 446 p. $28.50. Limnology and Oceanography, 29(2):443, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Paine:1986:BCW

[Pai86] R. T. Paine. Benthic community–water column coupling during the 1982–1983 El Niño. Are community changes at high latitudes attributable to cause or coincidence? Limnology and Oceanography, 31(2):351–360, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Paine:1989:BRC

[Pai89] R. T. Paine. Book review: Complex interactions in lake communities (S. R. Carpenter [ed.]). Limnology and Oceanography, 34 (2):486–487, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Parsons:1982:NPD

[Par82] T. R. Parsons. The new physical definition of salinity: Biologists beware. Limnology and Oceanography, 27(2):384–385, March 1982. CODEN LIOCAH. ISSN 0024-3590. See reply [Gie82].

Pastorok:1980:EPH

[Pas80] Robert A. Pastorok. The effects of predator hunger and food abundance on prey selection by Chaoborus larvae. Limnology and Oceanography, 25(5):910–921, September 1980. CODEN LI-OCAH. ISSN 0024-3590.

Patalas:1980:CRB

[Pat80] K. Patalas. Comment on “A relation between lake morphometry and primary productivity and its use in interpreting whole-lake eutrophication experiments” (E. J. Fee). Limnology and Oceanography, 25(6):1147–1149, November 1980. CODEN LIOCAH. ISSN 0024-3590. See reply [Fee80].

Patrick:1988:IDF

[Pat88] Ruth Patrick. Importance of diversity in the functioning and structure of riverine communities. Limnology and Oceanography, 33(6): 1304–1307, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Paterson:1989:STC

[Pat89] David M. Paterson. Short-term changes in the erodibility of intertidal cohesive sediments related to the migratory behavior of epipelic diatoms. Limnology and Oceanography, 34(1):223–234, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Parkin:1980:PBP

[PB80a] T. B. Parkin and T. D. Brock. Photosynthetic bacterial production in lakes: The effects of light intensity. Limnology and Oceanography, 25(4):711–718, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Peng:1980:GER

[PB80b] Tsung-Hung Peng and Wallace Broecker. Gas exchange rates for three closed-basin lakes. Limnology and Oceanography, 25(5):789– 796, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Parkin:1981:RPB

[PB81] T. B. Parkin and T. D. Brock. The role of phototrophic bacteria in the sulfur cycle of a meromictic lake. Limnology and Oceanography, 26(5):880–890, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1987:RDA

[PB87] Charles H. Peterson and R. Black. Resource depletion by active suspension feeders on tidal fiats: Influence of local density and tidal elevation. Limnology and Oceanography, 32(1):143–166, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Paul:1984:GMM

[PC84] John H. Paul and David J. Carlson. Genetic material in the marine environment: Implication for bacterial DNA. Limnology and Oceanography, 29(5):1091–1096, September 1984. CODEN LI-OCAH. ISSN 0024-3590.

Paerl:1987:LFC

[PCP87] Hans W. Paerl, Kenneth M. Crocker, and Leslie E. Prufert. Limitation of N_{2} fixation in coastal marine waters: Relative importance of molybdenum, iron, phosphorus, and organic matter availability. Limnology and Oceanography, 32(3):525–536, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1980:AOM

[PD80] Lawrence R. Pomeroy and Don Deibel. Aggregation of organic matter by pelagic tunicates. Limnology and Oceanography, 25(4): 643–652, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Peters:1984:EAZ

[PD84] Robert Henry Peters and John A. Downing. Empirical analysis of zooplankton filtering and feeding rates. Limnology and Oceanography, 29(4):763–784, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Peckarsky:1981:RCS

[Pec81] Barbara L. Peckarsky. Reply to comment by Sell. Limnology and Oceanography, 27(4):982–987, September 1981. CODEN LI-OCAH. ISSN 0024-3590. See [Sel81].

Peters:1981:PAL

[Pet81] Robert Henry Peters. Phosphorus availability in Lake Memphremagog and its tributaries. Limnology and Oceanography, 26(6):1150– 1161, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Peters:1986:RPL

[Pet86a] Robert Henry Peters. The role of prediction in limnology. Limnology and Oceanography, 31(5):1143–1159, September 1986. CO-DEN LIOCAH. ISSN 0024-3590.

Petersen:1986:SPG

[Pet86b] Robert C. Petersen, Jr. In situ particle generation in a southern Swedish stream. Limnology and Oceanography, 31(2):432–437, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1986:EMM

[Pet86c] Charles H. Peterson. Enhancement of Mercenaria mercenaria densities in seagrass beds: Is pattern fixed during settlement season or altered by subsequent differential survival? Limnology and Oceanography, 31(1):200–205, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1988:TFE

[Pet88] Charles H. Peterson. Tidal flat ecology. An experimental approach to species interactions. Limnology and Oceanography, 33(5):1229, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Porter:1980:UDI

[PF80] Karen G. Porter and Yvette S. Feig. The use of DAPI for identifying and counting aquatic microflora. Limnology and Oceanography, 25(5):943–948, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Polishchuk:1981:CTA

[PG81] L. V. Polishchuk and A. M. Ghilarov. Comparison of two approaches used to calculate zooplankton mortality. Limnology and Oceanography, 26(6):1162–1168, November 1981. CODEN LI-OCAH. ISSN 0024-3590.

Priscu:1983:CEA

[PG83] John C. Priscu and Charles R. Goldman. Carboxylating enzyme activity and photosynthetic end products of phytoplankton in the shallow and deep chlorophyll layers of Castle Lake. Limnology and Oceanography, 28(6):1168–1181, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Pan:1988:SVS

[PGB88] D. Pan, J. F. R. Gower, and G. A. Borstad. Seasonal variation of the surface chlorophyll distribution along the British Columbia coast as shown by CZCS satellite imagery. Limnology and Oceanography, 33(2):227–244, March 1988. CODEN LI-OCAH. ISSN 0024-3590.

Porter:1982:EFC

[PGO82] Karen G. Porter, Jeroen Gerritsen, and John D. Orcutt, Jr. The effect of food concentration on swimming patterns, feeding behavior, ingestion, assimilation, and respiration by Daphnia. Limnology and Oceanography, 27(5):935–949, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Peterson:1987:SCN

[PH87] Bruce J. Peterson and Robert W. Howarth. Sulfur, carbon, and nitrogen isotopes used to trace organic matter flow in the salt-marsh estuaries of Sapelo Island, Georgia. Limnology and Oceanography, 32(6):1195–1213, November 1987. CODEN LIOCAH. ISSN 0024-

Patterson:1988:TSL

[PH88a] J. C. Patterson and P. F. Hamblin. Thermal simulation of a lake with winter ice cover. Limnology and Oceanography, 33(3):323– 338, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Price:1988:UUC

[PH88b] Neil M. Price and Paul J. Harrison. Uptake of urea C and urea N by the coastal marine diatom Thalassiosira pseudonana. Limnology and Oceanography, 33(4):528–537, July 1988. CODEN LI-OCAH. ISSN 0024-3590.

Peterson:1983:CFP

[PHCK83] Bruce J. Peterson, John E. Hobbie, Teresa L. Corliss, and Keith Kriet. A continuous-flow periphyton bioassay: Tests of nutrient limitation in a tundra stream. Limnology and Oceanography, 28 (3):583–591, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Poulet:1989:VCM

[PHCT89] S. A. Poulet, A. M. Hapette, R. B. Cole, and J. C. Tabet. Vitamin C in marine copepods. Limnology and Oceanography, 34(7):1325– 1331, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Phillips:1980:CEM

[Phi80] Neal W. Phillips. A comment on energy maximization by deposit feeders. Limnology and Oceanography, 25(6):1143–1145, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Patterson:1984:CDS

[PHI84] J. C. Patterson, P. F. Hamblin, and J. Imberger. Classification and dynamic simulation of the vertical density structure of lakes. Limnology and Oceanography, 29(4):845–861, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Pickard:1982:BRO

[Pic82] G. L. Pickard. Book review: Oceanography: The past. 1980. Proc. 3rd Int. Congr. on the History of Oceanography, Woods Hole, Mass., September 1980. M. Sears and D. Merriman [eds.]. Springer-Verlag, New York, Inc. 812 p. $37.50. Limnology and Oceanography, 27(3):591, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Pickard:1985:BRB

[Pic85] G. L. Pickard. Book review: Bowden, K. F. 1984. Physical oceanography of coastal waters. Ellis Horwood Ser. Mar. Sci. John Wiley & Sons, Inc., Somerset, N.J. 302 p. $68.95. Limnology and Oceanography, 30(2):449–450, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Pimm:1983:BRT

[Pim83] Stuart L. Pimm. Book review: Tilman, D. 1982. Resource competition and community structure. Monogr. Pop. Biol. 17. Princeton University Press, Princeton, N.J. 296 p. $27.50. Limnology and Oceanography, 28(5):1043–1045, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Peter:1989:EDB

[PL89] Hawkins Peter and Winfried Lampert. The effect of Daphnia body size on filtering rate inhibition in the presence of a filamentous cyanobacterium. Limnology and Oceanography, 34(6):1084–1089, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Platt:1986:BRR

[Pla86] Trevor Platt. Book review: Robinson, I. S. 1985. Satellite oceanography: an introduction for oceanographers and remote-sensing scientists. John Wiley and Sons, Inc., Somerset, N.J. 455 p. $59.95. Limnology and Oceanography, 31(5):1167, September 1986. CO-DEN LIOCAH. ISSN 0024-3590.

Pick:1984:NSM

[PLN84] F. R. Pick, D. R. S. Lean, and C. Nalewajko. Nutrient status of metalimnetic phytoplankton peaks. Limnology and Oceanography, 29(5):960–971, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Porter:1984:ECR

[PM84a] Karen Glaus Porter and Robert McDonough. The energetic cost of response to blue-green algal filaments by cladocerans. Limnology and Oceanography, 29(2):365–369, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Preisendorfer:1984:DII

[PM84b] Rudolph W. Preisendorfer and Curtis D. Mobley. Direct and inverse irradiance models in hydrologic optics. Limnology and Oceanography, 29(5):903–929, September 1984. CODEN LI-OCAH. ISSN 0024-3590.

Pelaez:1986:PPP

[PM86] José Peláez and John A. McGowan. Phytoplankton pigment patterns in the California Current as determined by satellite. Limnology and Oceanography, 31(5):927–950, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Palenik:1988:DPH

[PM88] B. Palenik and F. M. M. Morel. Dark production of H_{2}O_{2} in the Sargasso Sea. Limnology and Oceanography, 33(6 (part 2)): 1606–1611, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Palumbo:1987:EDS

[PME87] Anthony V. Palumbo, Patrick J. Mulholland, and J. W. Elwood. Extraction with DMSO to simultaneously measure periphyton photosynthesis, chlorophyll, and ATP. Limnology and Oceanography, 32(2):464–471, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Pick:1984:OMC

[PNL84] F. R. Pick, C. Nalewajko, and D. R. S. Lean. The origin of a metalimnetic chrysophyte peak. Limnology and Oceanography, 29 (1):125–134, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Paasche:1980:ACP

[PO80] E. Paasche and I. Ostergren. The annual cycle of plankton diatom growth and silica production in the inner Oslofjord. Limnology and Oceanography, 49(6):481–494, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Pace:1981:RIP

[PO81] Michael L. Pace and John D. Orcutt, Jr. The relative importance of protozoans, rotifers, and crustaceans in a freshwater zooplankton community. Limnology and Oceanography, 26(5):822– 830, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Page:1984:IAT

[POHMD84] P. Pagé, M. Ouellet, C. Hillaire-Marcel, and M. Dickman. Isotopic analyses (^{18}O, ^{13}C, ^{14}C) of two meromictic lakes in the Canadian Arctic Archipelago. Limnology and Oceanography, 29(3):564–573, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1980:BRK

[Pom80] Lawrence R. Pomeroy. Book review: Kirk, P. W., Jr. [Ed.] 1979. The Great Dismal Swamp. Univ. Press of Virginia, Charlottesville. xii + 427 p. $20.00. Limnology and Oceanography, 25(5):968–969, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Pomeroy:1989:BRP

[Pom89] Lawrence R. Pomeroy. Book review: Postma, H., and J. J. Zulstra [eds.]. 1988. Continental shelves. Ecosystems of the World. V. 27. Elsevier Sci. Publ., Amsterdam. 421 p. $189.50. Limnology and Oceanography, 34(2):489–490, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Pond:1982:BRW

[Pon82] Stephen Pond. Book review: Warren, B. A., and C. Wunsch [Eds.] 1981. Evolution of physical oceanography. MIT Press, Cambridge, Mass. 623 p. $37.50. Limnology and Oceanography, 27(2): 397, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Probyn:1985:NUS

[PP85] T. A. Probyn and S. J. Painting. Nitrogen uptake by sizefractionated phytoplankton populations in Antarctic surface waters. Limnology and Oceanography, 30(6):1327–1332, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Price:1986:CSC

[PP86] Holly J. Price and Gustav-Adolf Paffenhöfer. Capture of small cells by the copepod Eucalanus elongatus. Limnology and Oceanography, 31(1):189–194, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Perry:1989:DCS

[PP89] Mary Jane Perry and Steven M. Porter. Determination of the cross-section absorption coeffcient of individual phytoplankton cells by analytical flow cytometry. Limnology and Oceanography, 34(8):1727–1738, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Price:1983:MCC

[PPS83] Holly J. Price, Gustav-Adolf Paffenhöfer, and J. Rudi Strickler. Modes of cell capture in calanoid copepods. Limnology and Oceanography, 28(1):116–123, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Priscu:1987:PDM

[PPVHW87] John C. Priscu, Linda R. Priscu, Warwick F. Vincent, and Clive Howard-Williams. Photosynthate distribution by microplankton in permanently ice-covered Antarctic desert lakes. Limnology and Oceanography, 32(1):260–270, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Pramer:1980:BRL

[Pra80] David Pramer. Book review: Lynch, J. M., and N. J. Poole [Eds.] 1979. Microbial ecology: a conceptual approach. John Wiley & Sons, New York, xiv + 266 p. $49.50. Limnology and Oceanography, 49(6):580, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Prairie:1990:CNS

[Pra90] Yves T. Prairie. A comment on “Nutrient status and nutrient competition of phytoplankton in a shallow, hypertrophic lake” (Sommer). Limnology and Oceanography, 35(3):778–779, May 1990. CODEN LIOCAH. ISSN 0024-3590. See [Som89b].

Preisendorfer:1986:SDS

[Pre86] Rudolph W. Preisendorfer. Secchi disk science: Visual optics of natural waters. Limnology and Oceanography, 31(5):909–926, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Price:1989:SBK

[Pri89] Holly J. Price. Swimming behavior of krill in response to algal patches: a mesocosm study. Limnology and Oceanography, 34(4): 649–659, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Prieur:1981:OCC

[PS81] Louis Prieur and Shubha Sathyendranath. An optical classification of coastal and oceanic waters based on the specific spectral absorption curves of phytoplankton pigments, dissolved organic matter, and other particulate materials. Limnology and Oceanography, 26 (4):671–689, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Potts:1984:WTI

[PS84] D. C. Potts and P. K. Swart. Water temperature as an indicator of environmental variability on a coral reef. Limnology and Oceanography, 29(3):504–516, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Palmisano:1985:PPC

[PS85] Anna C. Palmisano and Cornelius W. Sullivan. Pathways of photosynthetic carbon assimilation in sea-ice microalgae from McMurdo Sound, Antarctica. Limnology and Oceanography, 30(3):674–678, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Putt:1989:EDC

[PS89] Mary Putt and Diane K. Stoecker. An experimentally determined carbon : volume ratio for marine “oligotrichous” ciliates from estuarine and coastal waters. Limnology and Oceanography, 34(6): 1097–1103, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Psenner:1988:AGS

[Pse88] Roland Psenner. Alkalinity generation in a soft-water lake: Watershed and in-lake processes. Limnology and Oceanography, 33(6 (part 2)):1463–1475, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Priscu:1986:NAH

[PSGD86] John C. Priscu, Robert H. Spigel, Max M. Gibbs, and Malcolm T. Downes. A numerical analysis of hypolimnetic nitrogen and phosphorus transformations in Lake Rotoiti, New Zealand: a geothermally influenced lake. Limnology and Oceanography, 31(4):812– 831, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1983:CER

[PTB83] Hans W. Paerl, Jane Tucker, and Patricia T. Bland. Carotenoid enhancement and its role in maintaining blue-green algal (Microcystis aeruginosa) surface blooms. Limnology and Oceanography, 28(5):847–857, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Paerl:1982:BGA

[PU82] Hans W. Paerl and Joseph F. Ustach. Blue-green algal scums: an explanation for their occurrence during freshwater blooms. Limnology and Oceanography, 27(2):212–217, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Purcell:1981:FER

[Pur81] Jennifer E. Purcell. Feeding ecology of Rhizophysa eysenhardti, a siphonophore predator of fish larvae. Limnology and Oceanography, 26(3):424–432, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Palenik:1987:HPP

[PZM87] Brian Palenik, O. C. Zafiriou, and F. M. M. Morel. Hydrogen peroxide production by a marine phytoplankter. Limnology and Oceanography, 32(6):1365–1369, November 1987. CODEN LI-OCAH. ISSN 0024-3590.

Quay:1980:VDR

[QBHS80] P. D. Quay, W. S. Broecker, R. H. Hesslein, and D. W. Schindler. Vertical diffusion rates determined by tritium tracer experiments in the thermocline and hypolimnion of two lakes. Limnology and Oceanography, 25(2):201–218, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Quay:1986:CCL

[QEQD86] P. D. Quay, S. R. Emerson, B. M. Quay, and A. H. Devol. The carbon cycle for Lake Washington — a stable isotope study. Limnology and Oceanography, 31(3):596–611, May 1986. CODEN LI-OCAH. ISSN 0024-3590.

Quinby-Hunt:1989:PLS

[QHHLS89] M. S. Quinby-Hunt, A. J. Hunt, K. Lofftus, and D. Shapiro. Polarized-light scattering studies of marine chlorella. Limnology and Oceanography, 34(8):1587–1600, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Rahm:1985:TFC

[Rah85] Lars Rahm. The thermally forced circulation in a small, icecovered lake. Limnology and Oceanography, 30(5):1122–1128, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Rahm:1987:OCB

[Rah87] Lars Rahm. Oxygen consumption in the Baltic proper. Limnology and Oceanography, 32(4):973–978, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Richardson:1988:MOP

[RAN88] Laurie L. Richardson, Carmen Aguilar, and Kenneth H. Nealson. Manganese oxidation in pH and O_{2} microenvironments produced by phytoplankton. Limnology and Oceanography, 33(3):352–363, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Riemann:1987:CCP

[RBNF87] Bo Riemann, Peter Koefoed Bjørnsen, Steven Newell, and Robert Fallon. Calculation of cell production of coastal marine bacteria based on measured incorporation of [^{3}H]thymidine. Limnology and Oceanography, 32(2):471–476, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Richman:1983:EFQ

[RD83] Sumner Richman and Stanley I. Dodson. The effect of food quality on feeding and respiration by Daphnia and Diaptomus. Limnology and Oceanography, 28(5):948–956, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Robertson:1989:DAF

[RD89] Alistar I. Robertson and Paul A. Daniel. Decomposition and the annual flux of detritus from fallen timber in tropical mangrove forests. Limnology and Oceanography, 34(3):640–646, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Ross:1982:RCF

[RDE82] Robin M. Ross, Kendra L. Daly, and T. Saunders English. Reproductive cycle and fecundity of Euphausia pacifica in Puget Sound, Washington. Limnology and Oceanography, 27(2):304–314, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Richey:1988:BGO

[RDW^{+}88] Jeffrey E. Richey, Allan H. Devol, Steven C. Wofsy, Reynaldo Victoria, and Maria N. G. Riberio. Biogenic gases and the oxidation and reduction of carbon in Amazon River and floodplain waters. Limnology and Oceanography, 33(4):551–561, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Rassoulzadegan:1981:GRT

[RE81] Fereidoun Rassoulzadegan and Michele Etienne. Grazing rate of the tintinnid Stenosemella ventricosa (Clap. & Lachm.) Jorg. on the spectrum of the naturally occurring particulate matter from a Mediterranean neritic area. Limnology and Oceanography, 26(2): 258–270, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Reeburgh:1982:DSR

[RE82] William S. Reeburgh and Roy E. Erickson. A “dipstick” sampler for rapid, continuous chemical profiles in sediments. Limnology and Oceanography, 27(3):556–559, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Reeve:1981:LCE

[Ree81] M. R. Reeve. Large cod-end reservoirs as an aid to the live collection of delicate zooplankton. Limnology and Oceanography, 26 (3):577–580, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Revsbech:1989:OMG

[Rev89] Niels Peter Revsbech. An oxygen microsensor with a guard cathode. Limnology and Oceanography, 34(2):474–478, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Rutledge:1988:LSC

[RF88] P. A. Rutledge and J. W. Fleeger. Laboratory studies on core sampling with application to subtidal meiobenthos collection. Limnology and Oceanography, 33(2):274–280, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Rutledge:1989:RCB

[RF89] P. A. Rutledge and J. W. Fleeger. Reply to comment by Barnett and Watson. Limnology and Oceanography, 34(6):1143–1144, September 1989. CODEN LIOCAH. ISSN 0024-3590. See [BW89].

Rassoulzadegan:1984:EFS

[RFS84] Fereidoun Rassoulzadegan, Lucienne Fenaux, and Richard R. Strathmann. Effect of flavor and size on selection of food by suspension-feeding plutei. Limnology and Oceanography, 29(2): 357–360, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Rhee:1981:EEFa

[RG81] G-Yull Rhee and Ivan J. Gotham. The effect of environmental factors on phytoplankton growth: Temperature and the interactions of temperature with nutrient limitation. Limnology and Oceanography, 26(4):635–648, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:1983:FST

[RG83] Colin S. Reynolds and Rrian M. Godfrey. Failure of a sedimenttrapping device. Limnology and Oceanography, 28(1):172–176, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Rasmussen:1989:HCL

[RGS89] Joseph B. Rasmussen, Lyse Godbout, and Marc Schallenberg. The humic content of lake water and its relationship to watershed and lake morphometry. Limnology and Oceanography, 34(7): 1336–1343, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Runge:1988:UGP

[RI88] Jeffrey A. Runge and R. Grant Ingram. Underice grazing by planktonic, calanoid copepods in relation to a bloom of ice microalgae in southeastern Hudson Bay. Limnology and Oceanography, 33(2): 280–286, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Riessen:1984:OSC

[Rie84] Howard P. Riessen. The other side of cyclomorphosis: Why Daphnia lose their helmets. Limnology and Oceanography, 29(5):1123– 1127, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Riley:1982:BRS

[Ril82] Gordon A. Riley. Book review: Saltzman, B. [Ed.] 1980. Estuarine physics and chemistry: Studies in Long Island Sound. Advances in Geophysics 22. Academic Press, New York. 424 p. $44.50. Limnology and Oceanography, 27(2):393–394, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Ringelberg:1981:VCC

[Rin81] J. Ringelberg. On the variation in carotenoid content of copepods. Limnology and Oceanography, 27(4):995–997, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Revsbech:1983:PBM

[RJ83] Niels Peter Revsbech and Bo Barker Jorgensen. Photosynthesis of benthic microflora measured with high spatial resolution by the oxygen microprofile method: Capabilities and limitations of the method. Limnology and Oceanography, 28(4):749–756, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Rounick:1984:GCS

[RJ84] J. S. Rounick and M. R. James. Geothermal and cold springs faunas: Inorganic carbon sources affect isotope values. Limnology and Oceanography, 29(2):386–389, March 1984. CODEN LI-OCAH. ISSN 0024-3590.

Revsbech:1981:PPM

[RJB81] Niels Peter Revsbech, Bo Barker Jørgensen, and Ole Brix. Primary production of microalgae in sediments measured by oxygen microprofile, H^{14}CO_{3} — fixation, and oxygen exchange methods. Limnology and Oceanography, 26(4):717–730, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Revsbech:1983:MSP

[RJBC83] Niels Peter Revsbech, Bo Barker Jorgensen, T. Henry Blackburn, and Yehuda Cohen. Microelectrode studies of the photosynthesis and O_{2}, H_{2}S, and pH profiles of a microbial mat. Limnology and Oceanography, 28(6):1062–1074, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Ryan:1987:GMI

[RK87] Douglas F. Ryan and David M. Kahler. Geochemical and mineralogical indications of pH in lakes and soils in central New Hampshire in the early Holocene. Limnology and Oceanography, 32(3): 751–757, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1986:RSR

[RKF86] John W. M. Rudd, C. A. Kelly, and Akira Furutani. The role of sulfate reduction in long term accumulation of organic and inorganic sulfur in lake sediments. Limnology and Oceanography, 31 (6):1281–1291, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1986:MCN

[RKS^{+}86] John W. M. Rudd, C. A. Kelly, Vincent St. Louis, R. H. Hesslein, A. Furutani, and M. H. Holoka. Microbial consumption of nitric and sulfuric acids in acidified north temperate lakes. Limnology and Oceanography, 31(6):1267–1280, November 1986. CODEN LI-OCAH. ISSN 0024-3590.

Reuter:1986:INU

[RLG86] John E. Reuter, Stanford L. Loeb, and Charles R. Goldman. Inorganic nitrogen uptake by epilithic periphyton in a n-deficient lake. Limnology and Oceanography, 31(1):149–160, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Rueter:1981:IBZ

[RM81] John G. Rueter, Jr. and F. M. M. Morel. The interaction between zinc deficiency and copper toxicity as it affects the silicic acid uptake mechanisms in Thalassiosira pseudonana. Limnology and Oceanography, 26(1):67–73, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Risk:1983:PCH

[RM83] Michael J. Risk and Hans R. Muller. Porewater in coral heads: Evidence for nutrient regeneration. Limnology and Oceanography, 28(5):1004–1008, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Riegman:1986:PGP

[RM86a] Roel Riegman and Luuc R. Mur. Phytoplankton growth and phosphate uptake (for P limitation) by natural phytoplankton populations from the Loosdrecht lakes (The Netherlands). Limnology and Oceanography, 31(5):983–988, September 1986. CODEN LI-OCAH. ISSN 0024-3590.

Rodriguez:1986:RBB

[RM86b] Jaime Rodriguez and Michael M. Mullin. Relation between biomass and body weight of plankton in a steady state oceanic ecosystem. Limnology and Oceanography, 31(2):361–370, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Reynolds:1982:SFP

[RMB82] Colin S. Reynolds, Helen R. Morison, and Christine Butterwick. The sedimentary flux of phytoplankton in the south basin of Windermere. Limnology and Oceanography, 27(6):1162–1174, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Revsbech:1986:OPC

[RMJ86] Niels Peter Revsbech, Bent Madsen, and Barker Jørgensen. Oxygen production and consumption in sediments determined at high spatial resolution by computer simulation of oxygen microelectrode data. Limnology and Oceanography, 31(2):293–304, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Robinson:1982:CPF

[RMN82] J. D. Robinson, K. H. Mann, and J. A. Novitsky. Conversion of the particulate fraction of seaweed detritus to bacterial biomass. Limnology and Oceanography, 27(6):1072–1079, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Raspor:1981:VSS

[RNVB81] B. Raspor, H. W. Nürnberg, P. Valenta, and M. Branica. Voltammetric studies on the stability of the Zn(II) chelates with NTA and EDTA and the kinetics of their formation in Lake Ontario water. Limnology and Oceanography, 26(1):54–66, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Runge:1982:SFP

[RO82] Jeffrey A. Runge and Mark D. Ohman. Size fractionation of phytoplankton as an estimate of food available to herbivores. Limnology and Oceanography, 27(3):570–576, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Rodriguez:1987:EPG

[Rod87] Marco A. Rodriguez. Estimating periphyton growth parameters using simple models. Limnology and Oceanography, 32(2):458–464, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Roman:1984:UDC

[Rom84] Michael R. Roman. Utilization of detritus by the copepod, Acartia tonsa. Limnology and Oceanography, 29(5):949–959, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Rosenberg:1980:FOF

[Ros80] Gene G. Rosenberg. Filmed observations of filter feeding in the marine planktonic copepod Acartia clausii. Limnology and Oceanography, 25(4):738–742, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Ross:1981:LCD

[Ros81] Robin M. Ross. Laboratory culture and development of Euphausia pacifica. Limnology and Oceanography, 26(2):235–246, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Roesler:1989:MSP

[RPC89] Collin S. Roesler, Mary Jane Perry, and Kendall L. Carder. Modeling in situ phytoplankton absorption from total absorption spectra in productive inland marine waters. Limnology and Oceanography, 34(8):1510–1523, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Reddy:1989:NDP

[RPL89] K. R. Reddy, W. H. Patrick, Jr., and C. W. Lindau. Nitrificationdenitrification at the plant root–sediment interface in wetlands. Limnology and Oceanography, 34(6):1004–1013, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Roman:1980:CEC

[RR80] Michael R. Roman and Parke A. Rublee. Containment effects in copepod grazing experiments: a plea to end the black box approach. Limnology and Oceanography, 25(6):982–990, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Rivkin:1981:LSC

[RS81] Richard B. Rivkin and Howard H. Seliger. Liquid scintillation counting for ^{14}C uptake of single algal cells isolated from natural samples. Limnology and Oceanography, 26(4):780–785, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Ramcharan:1986:VPM

[RS86a] Charles W. Ramcharan and W. Gary Sprules. Visual predation in Mysis relicta Lovén. Limnology and Oceanography, 31(2):414–420, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Rassoulzadegan:1986:PPI

[RS86b] F. Rassoulzadegan and R. W. Sheldon. Predator-prey interactions of nanozooplankton and bacteria in an oligotrophic marine environment. Limnology and Oceanography, 31(5):1010–1029, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Reimers:1986:RMP

[RS86c] Clare E. Reimers and K. L. Smith, Jr. Reconciling measured and predicted fluxes of oxygen across the deep sea sediment–water interface. Limnology and Oceanography, 31(2):305–318, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Roberts:1988:IGA

[RS88] D. G. Roberts and D. M. Smith. Infrared gas analysis of both gaseous and dissolved CO_{2} in small-volume marine samples. Limnology and Oceanography, 33(1):135–140, January 1988. CODEN LIOCAH. ISSN 0024-3590.

Ramcharan:1989:PRI

[RS89] Charles W. Ramcharan and W. Gary Sprules. Preliminary results from an inexpensive motion analyzer for free-swimming zooplankton. Limnology and Oceanography, 34(2):457–462, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Revsbech:1980:DOM

[RSBL80] Niels Peter Revsbech, Jan Sorensen, Thomas Henry Blackburn, and Jens Peter Lomholt. Distribution of oxygen in marine sediments measured with microelectrodes. Limnology and Oceanography, 25(3):403–411, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Rudd:1980:MOL

[Rud80] John W. M. Rudd. Methane oxidation in Lake Tanganyika (East Africa). Limnology and Oceanography, 25(5):958–963, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Rueter:1983:API

[Rue83] John G. Rueter, Jr. Alkaline phosphatase inhibition by copper: Implications to phosphorus nutrition and use as a biochemical marker of toxicity. Limnology and Oceanography, 28(4):743–748, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Runge:1980:EHS

[Run80] Jeffrey A. Runge. Effects of hunger and season on the feeding behavior of Calanus pacificus. Limnology and Oceanography, 25 (1):134–145, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Runge:1985:REP

[Run85] Jeffrey A. Runge. Relationship of egg production of Calanus pacificus to seasonal changes in phytoplankton availability in Puget Sound, Washington. Limnology and Oceanography, 30(2):382–396, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Rivkin:1987:PPC

[RV87] Richard B. Rivkin and Mary A. Voytek. Photoadaptations of photosynthesis and carbon metabolism by phytoplankton from McMurdo Sound, Antarctica. 1. Species-specific and community responses to reduced irradiances12. Limnology and Oceanography, 32(1):249–259, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Riber:1987:BLI

[RW87] Hans H. Riber and Robert G. Wetzel. Boundary-layer and internal diffusion effects on phosphorus fluxes in lake periphyton. Limnology and Oceanography, 32(6):1181–1194, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Robarts:1989:MTM

[RW89] Richard D. Robarts and Richard J. Wicks. [Methyl-^{3}H]thymidine macromolecular incorporation and lipid labeling: Their significance to DNA labeling during measurements of aquatic bacterial growth rate. Limnology and Oceanography, 34(1):213–222, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Robertson:1987:AHD

[RWB87] K. J. Robertson, P. M. Williams, and J. L. Bada. Acid hydrolysis of dissolved combined amino acids in seawater: a precautionary note. Limnology and Oceanography, 32(4):996–997, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Rzoska:1981:BRB

[Rzó81a] Julian Rzóska. Book review: Baxter, R. M., and P. Glaude. 1980. Environmental effects of dams and impoundments in Canada: Experience and prospects. Can. Bull. Fish. Aquat. Sci. 205. Dept. Fish. Oceans, Ottawa, vi + 34 p. In Canada C$2.50; elsewhere C$3.00. Limnology and Oceanography, 27(4):998–999, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Rzoska:1981:BRW

[Rzó81b] Julian Rzóska. Book review: Ward, J. V., and J. A. Stanford [Eds.] 1979. The ecology of regulated streams. Plenum Press, New York and London, xi + 398 p. (typewriter composition). $35.00. Limnology and Oceanography, 27(4):998, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Schandelmeier:1981:AII

[SA81] L. Schandelmeier and V. Alexander. An analysis of the influence of ice on spring phytoplankton population structure in the southeast Bering Sea. Limnology and Oceanography, 26(5):935–943, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Smith:1983:MBC

[SA83] S. V. Smith and M. J. Atkinson. Mass balance of carbon and phosphorus in Shark Bay, Western Australia. Limnology and Oceanography, 28(4):625–639, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Swanberg:1985:NRT

[SA85] Neil R. Swanberg and O. Roger Anderson. The nutrition of radiolarians: Trophic activity of some solitary Spumellaria. Limnology and Oceanography, 30(3):646–652, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Smith:1986:PTS

[SA86] Craig S. Smith and Michael S. Adams. Phosphorus transfer from sediments by Myriophyllum spicatum. Limnology and Oceanography, 31(6):1312–1321, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Sammarco:1989:HED

[SA89] Paul W. Sammarco and John C. Andrews. The Helix experiment: Differential localized dispersal and recruitment patterns in Great Barrier Reef corals. Limnology and Oceanography, 34(5):896–912, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Sakamoto-Arnold:1986:DHS

[SAJB86] Carole M. Sakamoto-Arnold, Kenneth S. Johnson, and Carl L. Beehler. Determination of hydrogen sulfide in seawater using flow injection analysis and flow analysis. Limnology and Oceanography, 31(4):894–900, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Sakshaug:1989:SSD

[SAK89] Egil Sakshaug, Kjersti Andresen, and Dale A. Kiefer. A steady state description ofgrowth and light absorption in the marine planktonic diatom Skeletonema costatum. Limnology and Oceanography, 34(1):198–205, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Santsch:1984:PFT

[San84] Peter H. Santsch. Particle flux and trace metal residence time in natural waters. Limnology and Oceanography, 29(5):1100–1107, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Santschi:1988:FCB

[San88] P. H. Santschi. Factors controlling the biogeochemical cycles of trace elements in fresh and coastal marine waters as revealed by artificial radioisotopes. Limnology and Oceanography, 33(4 (part 2)):848–866, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Sauer:1980:VLH

[Sau80] Theodor C. Sauer, Jr. Volatile liquid hydrocarbons in waters of the Gulf of Mexico and Caribbean Sea. Limnology and Oceanography, 25(2):338–351, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Smith:1980:BED

[SB80] Raymond C. Smith and Karen S. Baker. Biologically effective dose transmitted by culture bottles in ^{14}C productivity experiments. Limnology and Oceanography, 25(2):364–366, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Somville:1983:MDE

[SB83] Martine Somville and Gilles Billen. A method for determining exoproteolytic activity in natural waters. Limnology and Oceanography, 28(1):190–193, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Smith:1989:EPS

[SB89a] Craig R. Smith and Steven J. Brumsickle. The effects of patch size and substrate isolation on colonization modes and rates in an intertidal sediment. Limnology and Oceanography, 34(7):1263– 1277, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Soli:1989:TDC

[SB89b] Alan L. Soli and Robert H. Byrne. Temperature dependence of Cu(II) carbonate complexation in natural seawater. Limnology and Oceanography, 34(1):239–244, January 1989. CODEN LI-OCAH. ISSN 0024-3590.

Stites:1989:RGR

[SB89c] David L. Stites and Arthur C. Benke. Rapid growth rates of chironomids in three habitats of a subtropical blackwater river and their implications for P : B ratios. Limnology and Oceanography, 34(7):1278–1289, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Santschi:1983:ERC

[SBN^{+}83] Peter H. Santschi, Peter Bower, Urs P. Nyffeler, Antonia Azevedo, and Wallace S. Broecker. Estimates of the resistance to chemical transport posed by the deep-sea boundary layer. Limnology and Oceanography, 28(5):899–912, September 1983. CODEN LI-OCAH. ISSN 0024-3590.

Schroeder:1981:TTA

[SC81a] Lauren Alfred Schroeder and William M. Callaghan. Thermal tolerance and acclimation of two species of Hydra. Limnology and Oceanography, 26(4):690–696, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Sharp:1981:BMC

[SC81b] Jonathan H. Sharp and Thomas M. Church. Biochemical modeling in coastal waters of the Middle Atlantic States. Limnology and Oceanography, 26(5):843–854, September 1981. CODEN LI-OCAH. ISSN 0024-3590.

Sharp:1982:PDS

[SC82] Jonathan H. Sharp and Charles H. Culberson. The physical definition of salinity: a chemical evaluation. Limnology and Oceanography, 27(2):385–387, March 1982. CODEN LIOCAH. ISSN 0024-

Stiller:1984:RDS

[SC84] M. Stiller and Y. C. Chung. Radium in the Dead Sea: a possible tracer for the duration of meromixis. Limnology and Oceanography, 29(3):574–586, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Soluk:1988:VFP

[SC88] Daniel A. Soluk and Douglas A. Craig. Vortex feeding from pits in the sand: a unique method of suspension feeding used by a stream invertebrate. Limnology and Oceanography, 33(4):638–645, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Scavia:1988:RBS

[Sca88] Donald Scavia. On the role of bacteria in secondary production. Limnology and Oceanography, 33(5):1220–1224, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Sharp:1982:CDE

[SCC82] Jonathan H. Sharp, Charles H. Culberson, and Thomas M. Church. The chemistry of the Delaware estuary. General considerations. Limnology and Oceanography, 27(6):1015–1028, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Schindler:1980:EFP

[Sch80] D. W. Schindler. The effect of fertilization with phosphorus and nitrogen versus phosphorus alone on eutrophication of experimental lakes. Limnology and Oceanography, 25(6):1149–1152, November 1980. CODEN LIOCAH. ISSN 0024-3590. See reply [Fee80].

Schanz:1985:VLA

[Sch85] Ferdinand Schanz. Vertical light attenuation and phytoplankton development in Lake Zürich. Limnology and Oceanography, 30(2): 299–310, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Schindlei:1988:COA

[Sch88] D. W. Schindlei. Confusion over the origin of alkalinity in lakes. Limnology and Oceanography, 33(6 (part 2)):1637–1640, November 1988. CODEN LIOCAH. ISSN 0024-3590. See reply [Sto88b].

Smith:1989:BPA

[SCH89] Ralph E. H. Smith, Pierre Clement, and Erica Head. Biosynthesis and photosynthate allocation patterns of Arctic ice algae.

Limnology and Oceanography, 34(3):591–605, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Sosik:1989:CFS

[SCO89] Heidi M. Sosik, Sallie W. Chisholm, and Robert J. Olson. Chlorophyll fluorescence from single cells: Interpretation of flow cytometric signals. Limnology and Oceanography, 34(8):1749–1761, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Sprules:1984:PAB

[SCR84] W. Gary Sprules, John C. H. Carter, and Charles W. Ramcharan. Phenotypic associations in the Bosminidae (Cladocera): Zoogeographic patterns. Limnology and Oceanography, 29(1):161–169, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Smith:1982:PPA

[SE82] Paul E. Smith and Richard W. Eppley. Primary production and the anchovy population in the Southern California Bight: Comparison of time series. Limnology and Oceanography, 27(1):1–17, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Seitzinger:1988:DFC

[Sei88] Sybil P. Seitzinger. Denitrification in freshwater and coastal marine ecosystems: Ecological and geochemical significance. Limnology and Oceanography, 33(4 (part 2)):702–724, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Schelske:1984:MPF

[SEK84] Claire L. Schelske, Brian J. Eadie, and Gerald L. Krausse. Measured and predicted fluxes of biogenic silica in Lake Michigan. Limnology and Oceanography, 29(1):99–110, January 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Sell:1981:CBI

[Sel81] Daniel W. Sell. Comment on “Biological interactions as determinants of distributions of benthic invertebrates within the substrate of stony streams” (Peckarsky). Limnology and Oceanography, 27 (4):981–982, September 1981. CODEN LIOCAH. ISSN 0024-3590. See reply [Pec81].

Serruya:1980:BRL

[Ser80] C. Serruya. Book review: Lerman, A. [Ed.] 1978. Lakes: Chemistry, geology, physics. Springer-Verlag, New York, Heidelberg, Berlin, x + 363 p. $39.80. Limnology and Oceanography, 25(2): 385–386, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Swift:1981:CPC

[SF81] Michael C. Swift and Richard B. Forwarder. Chaoborus prey capture efficiency in the light and dark. Limnology and Oceanography, 26(3):461–466, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Straskrabova:1982:DCN

[SF82] Vera Straskrabova and Josef Fuksa. Diel changes in numbers and activities of bacterioplankton in a reservoir in relation to algal production. Limnology and Oceanography, 27(4):660–672, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Stewart:1984:EIQ

[SF84] Dianne E. Stewart and Franklin H. Farmer. Extraction, identification, and quantitation of phycobiliprotein pigments from phototrophic plankton. Limnology and Oceanography, 29(2):392–397, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Schulze:1989:ECP

[SF89] Peter C. Schulze and Carol L. Folt. Effects of conspecifics and phytoplankton on predation rates of the omnivorous copepods Epischura lacustris and Epischura nordenskioldi. Limnology and Oceanography, 34(2):444–450, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Scavia:1984:SSN

[SFDK84] Donald Scavia, Gary L. Fahnenstiel, Judith A. Davis, and Russell G. Kreis, Jr. Small-scale nutrient patchiness: Some consequences and a new encounter mechanism. Limnology and Oceanography, 29(4):785–793, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Schoener:1981:CBD

[SG81] Amy Schoener and Charles H. Greene. Comparison between destructive and nondestructive sampling of sessile epibenthic organisms. Limnology and Oceanography, 26(4):770–774, July 1981. CO-DEN LIOCAH. ISSN 0024-3590.

Sweet:1984:EFR

[SG84] Stephen T. Sweet and Norman L. Guinasso, Jr. Effects of flow rate on fluorescence in vivo during continuous measurements on Gulf of Mexico surface water. Limnology and Oceanography, 29 (2):397–401, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1986:LRV

[SG86] Ulrich Sommer and Z. Maciej Gliwicz. Long range vertical migration of Volvox in tropical Lake Cahora Bassa (Mozambique). Limnology and Oceanography, 31(3):650–653, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Sharma:1987:SBS

[SGMB87] P. Sharma, L. R. Gardner, W. S. Moore, and M. S. Bollinger. Sedimentation and bioturbation in a salt marsh as revealed by 210Pb, ^{137}Cs, and ^{7}Be studies. Limnology and Oceanography, 32 (2):313–326, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Sicko-Goad:1984:EAI

[SGSS84] Linda M. Sicko-Goad, Claire L. Schelske, and Eugene F. Stoermer. Estim ation of intracellular carbon and silica content of diatoms from natural assemblages using morphometric techniques. Limnology and Oceanography, 29(6):1170–1178, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Stanley:1981:NRN

[SH81] Donald W. Stanley and John E. Hobble. Nitrogen recycling in a North Carolina coastal river. Limnology and Oceanography, 26(1): 30–42, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1983:ECI

[SH83] William G. Sunda and Susan A. Huntsman. Effect of competitive interactions between manganese and copper on cellular manganese and growth in estuarine and oceanic species of the diatom Thalassiosira. Limnology and Oceanography, 28(5):924–934, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Schubauer:1984:ABE

[SH84] J. P. Schubauer and C. S. Hopkinson. Above- and belowground emergent macrophyte production and turnover in a coastal marsh ecosystem, Georgia. Limnology and Oceanography, 29(5):1052– 1065, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1985:RCM

[SH85] William G. Sunda and Susan A. Huntsman. Regulation of cellular manganese and manganese transport rates in the unicellular alga Chlamydomonas. Limnology and Oceanography, 30(1):71–80, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1987:MFC

[SH87a] William G. Sunda and Alfred K. Hanson. Measurement of free cupric ion concentration in seawater by a ligand competition technique involving copper sorption onto C_{18} SEP-PAK cartridges. Limnology and Oceanography, 32(3):537–551, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Sunda:1987:MOM

[SH87b] William G. Sunda and Susan A. Huntsman. Microbial oxidation of manganese in a North Carolina estuary. Limnology and Oceanography, 32(3):552–564, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Suttle:1988:APU

[SH88] C. A. Suttle and P. J. Harrison. Ammonium and phosphate uptake rates, n: P supply ratios, and evidence for N and P limitation in some oligotrophic lakes. Limnology and Oceanography, 33(2):186– 202, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Siver:1989:MSA

[SH89] Peter A. Siver and Josephine S. Hamer. Multivariate statistical analysis of the factors controlling the distribution of scaled chrysophytes. Limnology and Oceanography, 34(2):368–381, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Smith:1989:TBC

[SHDV89] S. V. Smith, J. T. Hollibaugh, S. J. Dollar, and S. Vink. Tomales Bay, California: a case for carbon-controlled nitrogen cycling. Limnology and Oceanography, 34(1):37–52, January 1989. CO-DEN LIOCAH. ISSN 0024-3590.

Sheldon:1984:PGR

[She84] R. W. Sheldon. Phytoplankton growth rates in the tropical ocean. Limnology and Oceanography, 29(6):1342–1346, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Shilo:1985:BRC

[Shi85] Moshe Shilo. Book review: Cohen, Y., R. W. Castenholz, and H. O. Halvorson [Eds.] 1984. Microbial mats: Stromatolites. MBL Lectures in Biology, v. 3. Alan R. Liss, Inc., New York. 508 p.

$88.00. Limnology and Oceanography, 30(4):917–918, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Sorensen:1984:DDS

[SHRW84] Jan Sørensen, David J. Hydes, T. Roger, and S. Wilson. Denitrification in a deep-sea sediment core from the eastern equatorial Atlantic. Limnology and Oceanography, 29(3):653–657, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Schwartz:1985:MSD

[SIH85] Steven S. Schwartz, David J. Innes, and Paul D. N. Hebert. Morphological separation of Daphnia pulex and Daphnia obtusa in North America. Limnology and Oceanography, 30(1):189–197, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Simon:1987:BPS

[Sim87] Meinhard Simon. Biomass and production of small and large freeliving and attached bacteria in Lake Constance. Limnology and Oceanography, 32(3):591–607, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Spigel:1986:MDM

[SIR86] Robert H. Spigel, Jorg Imberger, and Kenneth N. Rayner. Modeling the diurnal mixed layer. Limnology and Oceanography, 31(3): 533–556, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Sirkes:1987:SMI

[Sir87] Ziv Sirkes. Surface manifestations of internal oscillations in a highly saline lake (the Dead Sea). Limnology and Oceanography, 32(1):76–82, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Sorensen:1987:DLD

[SJC^{+}87] J. Sorensen, K. S. Jorgensen, S. Colley, D. J. Hydes, J. Thomson, and T. R. S. Wilson. Depth localization of denitrification in a deep-sea sediment from the Madeira Abyssal Plain. Limnology and Oceanography, 32(3):758–762, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Sand-Jensen:1987:PLA

[SJR87] Kaj Sand-Jensen and Niels Peter Revsbech. Photosynthesis and light adaptation in epiphyte-macrophyte associations measured by oxygen microelectrodes. Limnology and Oceanography, 32(2):452– 457, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Steenbergen:1982:DPM

[SK82] C. L. M. Steenbergen and H. J. Korthals. Distribution of phototrophic microorganisms in the anaerobic and microaerophilic strata of Lake Vechten (The Netherlands). Pigment analysis and role in primary production. Limnology and Oceanography, 27(5): 883–895, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Smith:1983:SPQ

[SK83a] Ralph E. H. Smith and J. Kalff. Sample preparation for quantitative autoradiography of phytoplankton. Limnology and Oceanography, 28(2):383–389, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Smith:1983:CPA

[SK83b] Ralph E. H. Smith and Jaap Kalff. Competition for phosphorus among co-occurring freshwater phytoplankton. Limnology and Oceanography, 28(3):448–464, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Smith:1985:PCA

[SK85a] Ralph E. H. Smith and Jaap Kalff. Phosphorus competition among phytoplankton: a reply. Limnology and Oceanography, 30(2): 440–444, March 1985. CODEN LIOCAH. ISSN 0024-3590. See [SK85b].

Sommer:1985:PNC

[SK85b] Ulrich Sommer and Susan Soltau Kilham. Phytoplankton natural community competition experiments: a reinterpretation. Limnology and Oceanography, 30(2):436–440, March 1985. CODEN LIOCAH. ISSN 0024-3590. See reply [SK85a].

Smith:1986:VFB

[SKW86] S. V. Smith, W. J. Kimmerer, and T. W. Walsh. Vertical flux and biogeochemical turnover regulate nutrient limitation of net organic production in the North Pacific Gyre. Limnology and Oceanography, 31(1):161–166, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Saunders:1987:PUC

[SL87a] James F. Saunders III and William M. Lewis, Jr. A perspective on the use of cohort analysis to obtain demographic data for copepods. Limnology and Oceanography, 32(2):511–513, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Scavia:1987:BLM

[SL87b] Donald Scavia and Gwenyth A. Laird. Bacterioplankton in Lake Michigan: Dynamics, controls, and significance to carbon flux. Limnology and Oceanography, 32(5):1017–1033, September 1987. CODEN LIOCAH. ISSN 0024-3590.

Saunders:1989:ZAL

[SL89] James F. Saunders III and William M. Lewis,, Jr. Zooplankton abundance in the lower Orinoco River, Venezuela. Limnology and Oceanography, 34(2):397–409, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Smith:1983:SCO

[SLB83] K. L. Smith, Jr., M. B. Laver, and N. O. Brown. Sediment community oxygen consumption and nutrient exchange in the central and eastern North Pacific. Limnology and Oceanography, 28(5): 882–898, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Scavia:1986:PPB

[SLF86] Donald Scavia, Gwenyth A. Laird, and Gary L. Fahnenstiel. Production of planktonic bacteria in Lake Michigan. Limnology and Oceanography, 31(3):612–626, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Simard:1985:STG

[SLL85] Yvan Simard, Guy Lacroix, and Louis Legendre. In situ twilight grazing rhythm during diel vertical migrations of a scattering layer of Calanus finmarchicus. Limnology and Oceanography, 30(3):598– 606, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Sathyendranath:1987:VSV

[SLP87] Shubha Sathyendranath, Luigi Lazzara, and Louis Prieur. Variations in the spectral values of specific absorption of phytoplankton. Limnology and Oceanography, 32(2):403–415, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Smith:1980:PCA

[SM80] A. E. Smith and I. Morris. Pathways of carbon assimilation in phytoplankton from the Antarctic Ocean. Limnology and Oceanography, 25(5):865–872, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Scott:1982:MRC

[SM82a] J. M. Scott and J. A. Marlow. A microcalorimeter with a range of 0.1-1.0 calories. Limnology and Oceanography, 27(3):585–590, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Stewart:1982:TCN

[SM82b] Kenton M. Stewart and Patricia J. H. Martin. Turbidity and its causes in a narrow glacial lake with winter ice cover. Limnology and Oceanography, 27(3):510–517, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Scott:1983:SPB

[SM83] Mary Ann Scott and William W. Murdoch. Selective predation by the backswimmer, Notonecta. Limnology and Oceanography, 28(2):352–366, March 1983. CODEN LIOCAH. ISSN 0024-3590.

Setaro:1984:RPE

[SM84] Frank V. Setaro and John M. Melack. Responses of phytoplankton to experimental nutrient enrichment in an Amazon floodplain lake. Limnology and Oceanography, 29(5):972–984, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Schoenberg:1985:RFR

[SM85] Steven A. Schoenberg and Alexander E. Maccubbin. Relative feeding rates on free and particle-bound bacteria by freshwater macrozooplankton. Limnology and Oceanography, 30(5):1084– 1090, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Schoenberg:1984:CDF

[SMH84] Steven A. Schoenberg, Alexander E. Maccubbin, and Robert E. Hodson. Cellulose digestion by freshwater microcrustacea. Limnology and Oceanography, 29(5):1132–1136, September 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Smith:1981:EBS

[Smi81a] Ned P. Smith. Energy balance in a shallow seagrass flat for winter conditions. Limnology and Oceanography, 26(3):482–491, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Smith:1981:HAI

[Smi81b] S. V. Smith. The Houtman Abrolhos Islands: Carbon metabolism of coral reefs at high latitude. Limnology and Oceanography, 26 (4):612–621, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Smith:1982:BRB

[Smi82a] S. V. Smith. Book review: Barnes, R. S. K. 1980. Coastal lagoons. Cambridge Univ. Press, New York. 106 p. $29.50, hardcover; $11.25, softcover. Limnology and Oceanography, 27(2):396– 397, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Smith:1982:BRJ

[Smi82b] S. V. Smith. Book review: James, N. P., and R. N. Ginsburg. 1979. The seaward margin of Belize barrier and atoll reefs. Spec. Publ. 3, Int. Assoc. Sedimentol. Blackwell Sci. Publ., Oxford, London, Edinburgh, Melbourne. $22.95, softcover. Limnology and Oceanography, 27(1):196–197, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Smith:1982:NPD

[Smi82c] Val H. Smith. The nitrogen and phosphorus dependence of algal biomass in lakes: an empirical and theoretical analysis. Limnology and Oceanography, 27(6):1101–1111, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Smith:1982:ZBB

[Smi82d] K. L. Smith, Jr. Zooplankton of a bathyal benthic boundary layer: In situ rates of oxygen consumption and ammonium excretion. Limnology and Oceanography, 27(3):461–471, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Smith:1984:PVN

[Smi84] S. V. Smith. Phosphorus versus nitrogen limitation in the marine environment. Limnology and Oceanography, 29(6):1149–1160, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Smith:1985:MDS

[Smi85] K. L. Smith, Jr. Maerozooplankton of a deep sea hydrothermal vent: In situ rates of oxygen consumption. Limnology and Oceanography, 30(1):102–110, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Smith:1986:SSS

[Smi86a] David F. Smith. Small-scale spatial heterogeneity in dissolved nutrient concentrations. Limnology and Oceanography, 31(1):167– 170, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Smith:1986:BRH

[Smi86b] Stephen V. Smith. Book review: Houghton, J. T. [Ed.]. 1984. The global climate. Cambridge University Press, New York. 233 p. $49.50. Limnology and Oceanography, 31(3):662–667, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Smith:1987:FES

[Smi87] K. L. Smith, Jr. Food energy supply and demand: a discrepancy between particulate organic carbon flux and sediment community oxygen consumption in the deep ocean. Limnology and Oceanography, 32(1):201–220, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Smith:1988:BRB

[Smi88] Val H. Smith. Book review: Burgis, M. J., and P. Morris. 1988. The natural history of lakes. Cambridge Univ. Press, New York. 218 p. $29.95. Limnology and Oceanography, 33(4):647, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Spies:1981:DRE

[SMK81] Robert B. Spies, Kenneth V. Marsh, and James R. Kercher. Dynamics of radionuclide exchange in the calcareous algae Halimeda at Enewetak Atoll. Limnology and Oceanography, 26(1):74–85, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Smith:1989:EPB

[SMP^{+}89] R. C. Smith, J. Marra, M. J. Perry, K. S. Baker, E. Swift, E. Buskey, and D. A. Kiefer. Estimation of a photon budget for the upper ocean in the Sargasso Sea. Limnology and Oceanography, 34(8):1673–1693, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Sherman:1984:TSA

[SMR84] K. M. Sherman, D. A. Meeter, and J. A. Reidenauer. A technique for subsampling an abundant taxon while completely sorting other taxa. Limnology and Oceanography, 29(2):433–439, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Sanchez:1986:FPT

[SMSM86] Arthur L. Sanchez, James W. Murray, William R. Schell, and Laurence G. Miller. Fallout plutonium in two oxic-anoxic environments. Limnology and Oceanography, 31(5):1110–1121, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Seitzinger:1985:ERD

[SN85] Sybil P. Seitzinger and Scott W. Nixon. Eutrophication and the rate of denitrification and N_{2}0 production in coastal marine sediments. Limnology and Oceanography, 30(6):1332–1339, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Scranton:1984:DCH

[SNL84] Mary I. Scranton, Paul C. Novelli, and Peter A. Loud. The distribution and cycling of hydrogen gas in the waters of two anoxic marine environments. Limnology and Oceanography, 29(5):993– 1003, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Scranton:1987:HPN

[SNM^{+}87] Mary I. Scranton, Paul C. Novelli, Anthony Michaels, Sarah G. Horrigan, and Edward J. Carpenter. Hydrogen production and nitrogen fixation by Oscillatoria thiebautii during in situ incubations. Limnology and Oceanography, 32(4):998–1006, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Seitzinger:1984:DNO

[SNP84] Sybil P. Seitzinger, Scott W. Nixon, and Michael E. Q. Pilson. Denitrification and nitrous oxide production in a coastal marine ecosystem. Limnology and Oceanography, 29(1):73–83, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Sheldon:1986:EIF

[SNR86] R. W. Sheldon, P. Nival, and F. Rassoulzadegan. An experimental investigation of a flagellate-ciliate-copepod food chain with some observations relevant to the linear biomass hypothesis. Limnology and Oceanography, 31(1):184–188, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Smith:1987:BSL

[SO87] Richard L. Smith and Ronald S. Oremland. Big Soda Lake (Nevada). 2. Pelagic sulfate reduction. Limnology and Oceanography, 32(4):794–803, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Sodergren:1987:OCS

[Söd87] Anders Södergren. Origin and composition of surface slicks in lakes of differing trophic status. Limnology and Oceanography, 32 (6):1307–1316, November 1987. CODEN LIOCAH. ISSN 0024-

Sommer:1984:PPF

[Som84] Ulrich Sommer. The paradox of the plankton: Fluctuations of phosphorus availability maintain diversity of phytoplankton in flow-through cultures. Limnology and Oceanography, 29(3):633– 636, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1985:CBS

[Som85] Ulrich Sommer. Comparison between steady state and non-steady state competition: Experiments with natural phytoplankton. Limnology and Oceanography, 30(2):335–346, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1988:PSM

[Som88] Ulrich Sommer. Phytoplankton succession in microcosm experiments under simultaneous grazing pressure and resource limitation. Limnology and Oceanography, 33(5):1037–1054, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1989:MGR

[Som89a] Ulrich Sommer. Maximal growth rates of Antarctic phytoplankton: Only weak dependence on cell size. Limnology and Oceanography, 34(6):1109–1112, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Sommer:1989:NSN

[Som89b] Ulrich Sommer. Nutrient status and nutrient competition of phytoplankton in a shallow, hypertrophic lake. Limnology and Oceanography, 34(7):1162–1173, November 1989. CODEN LI-OCAH. ISSN 0024-3590. See comment [Pra90].

Subramanian:1985:MMF

[SP85] A. Subramanian and A. Purushothaman. Mass mortality of fish and invertebrates associated with a bloom of Hemidiscus hardmannianus (Bacillariophyceae) in Parangipettai (southern India). Limnology and Oceanography, 30(4):910–911, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Sathyendranath:1989:CAP

[SP89a] Shubha Sathyendranath and Trevor Platt. Computation of aquatic primary production: Extended formalism to include effect of angular and spectral distribution of light. Limnology and Oceanography, 34(1):188–198, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Shaw:1989:AST

[SP89b] R. D. Shaw and E. E. Prepas. Anomalous, short-term influx of water into seepage meters. Limnology and Oceanography, 34(7): 1343–1351, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Smith:1989:BOM

[SPBB89] R. C. Smith, B. B. Prezelin, R. R. Bidigare, and K. S. Baker. Biooptical modeling of photosynthetic production in coastal waters. Limnology and Oceanography, 34(8):1524–1544, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Sanders:1989:SPB

[SPBD89] Robert W. Sanders, Karen G. Porter, Susan J. Bennett, and Adrienne E. DeBiase. Seasonal patterns of bacterivory by flagellates, ciliates, rotifers, and cladocerans in a freshwater planktonic community. Limnology and Oceanography, 34(4):673–687, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Sweerts:1986:MAF

[SRK86] J. P. Sweerts, John W. M. Rudd, and C. A. Kelly. Metabolic activities in flocculent surface sediments and underlying sandy littoral sediments. Limnology and Oceanography, 31(2):330–338, March 1986. CODEN LIOCAH. ISSN 0024-3590.

Sikes:1980:PCF

[SRW80] C. Steven Sikes, Robert D. Roer, and Karl M. Wilbur. Photosynthesis and coccolith formation: Inorganic carbon sources and net inorganic reaction of deposition. Limnology and Oceanography, 25 (2):248–261, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Stanlaw:1981:GFV

[SRW81] Karen A. Stanlaw, Michael R. Reeve, and Mary Ann Walter. Growth, food, and vulnerability to damage of the ctenophore Mnemiopsis mccradyi in its early life history stages. Limnology and Oceanography, 26(2):224–234, March 1981. CODEN LIOCAH. ISSN 0024-3590.

Sorensen:1984:HTO

[SRW84] Jan Sørensen, T. Roger, and S. Wilson. A headspace technique for oxygen measurements in deep-sea sediment cores. Limnology and Oceanography, 29(3):650–652, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Solorzano:1980:DTDa

[SS80a] Lucia Solórzano and Jonathan H. Sharp. Determination of total dissolved nitrogen in natural waters. Limnology and Oceanography, 25(4):751–754, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Solorzano:1980:DTDb

[SS80b] Lucia Solórzano and Jonathan H. Sharp. Determination of total dissolved phosphorus and particulate phosphorus in natural waters. Limnology and Oceanography, 25(4):754–758, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Sundby:1985:MFB

[SS85] Bjørn Sundby and Norman Silverberg. Manganese fluxes in the benthic boundary layer. Limnology and Oceanography, 30(2):372– 381, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Sherr:1988:RMP

[SS88a] Evelyn Sherr and Barry Sherr. Role of microbes in pelagic food webs: a revised concept. Limnology and Oceanography, 33(5): 1225–1227, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Stockner:1988:RAS

[SS88b] John G. Stockner and Ken S. Shortreed. Response of Anabaena and Synechococcus to manipulation of nitrogen: phosphorus ratios in a lake fertilization experiment. Limnology and Oceanography, 33(6):1348–1361, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Sherr:1982:DOD

[SSB82] Barry F. Sherr, Evelyn B. Sherr, and Thomas Berman. Decomposition of organic detritus: a selective role for microflagellate protozoa. Limnology and Oceanography, 26(5):765–769, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Sherr:1986:SAC

[SSFN86] Evelyn B. Sherr, Barry F. Sherr, Robert D. Fallon, and Steven Y. Newell. Small, aloricate ciliates as a major component of the marine heterotrophic nanoplankton. Limnology and Oceanography, 31 (1):177–183, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Smith:1989:ISS

[SSGB89] S. V. Smith, S. Serruya, Y. Geifman, and T. Berman. Internal sources and sinks of water, P, N, Ca, and Cl in Lake Kinneret, Israel. Limnology and Oceanography, 34(7):1202–1213, November 1989. CODEN LIOCAH. ISSN 0024-3590.

Sigg:1987:VTH

[SSK87] Laura Sigg, Michael Sturm, and David Kistler. Vertical transport of heavy metals by settling particles in Lake Zurich. Limnology and Oceanography, 32(1):112–130, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Schropp:1987:DHF

[SSS87] Steven J. Schropp, Mary I. Scranton, and John R. Schwarz. Dissolved hydrogen, facultatively anaerobic, hydrogen-producing bacteria, and potential hydrogen production rates in the western North Atlantic Ocean and Gulf of Mexico. Limnology and Oceanography, 32(2):396–402, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Stauffer:1980:WTS

[Sta80] Robert E. Stauffer. Windpower time series above a temperate lake. Limnology and Oceanography, 49(6):513–528, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Stauffer:1982:WSE

[Sta82] Robert E. Stauffer. Wind stress effects on chlorophyll distribution in stratified eutrophic lakes. Limnology and Oceanography, 27(1): 66–74, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Stauffer:1985:NIC

[Sta85a] Robert E. Stauffer. Nutrient internal cycling and the trophic regulation of Green Lake, Wisconsin. Limnology and Oceanography, 30(2):347–363, March 1985. CODEN LIOCAH. ISSN 0024-3590.

Stauffer:1985:RBP

[Sta85b] Robert E. Stauffer. Relationships between phosphorus loading and trophic state in calcareous lakes of southeast Wisconsin. Limnology and Oceanography, 30(1):123–145, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Stabel:1986:CPL

[Sta86] H.-H. Stabel. Calcite precipitation in Lake Constance: Chemical equilibrium, sedimentation, and nucleation by algae. Limnology and Oceanography, 31(5):1081–1094, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Schenck:1988:EPI

[STC88] R. C. Schenck, A. Tessier, and P. G. C. Campbell. The effect of pH on iron and manganese uptake by a green alga. Limnology and Oceanography, 33(4):538–550, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Steinhorn:1983:SSP

[Ste83] Ilana Steinhorn. In situ salt precipitation at the Dead Sea. Limnology and Oceanography, 28(3):580–583, May 1983. CODEN LI-OCAH. ISSN 0024-3590.

Steele:1984:BRP

[Ste84] John A. Steele. Book review: Platt, T., K. H. Mann, and R. E. Ulanowicz [Eds.]. 1981. Mathematical models in biological oceanography. Monogr. Oceanogr. Methodol. 7. UNESCO Press, Paris. 156 p. Limnology and Oceanography, 29(1):221–222, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Steinhorn:1985:DLT

[Ste85] Ilana Steinhorn. The disappearance of the long term meromictic stratification of the Dead Sea. Limnology and Oceanography, 30 (3):451–472, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Stemberger:1988:RCH

[Ste88a] Richard S. Stemberger. Reproductive costs and hydrodynamic benefits of chemically induced defenses in Keratella testudo. Limnology and Oceanography, 33(4):593–606, July 1988. CODEN LI-OCAH. ISSN 0024-3590.

Stevenson:1988:CES

[Ste88b] J. Court Stevenson. Comparative ecology of submersed grass beds in freshwater, estuarine, and marine environments. Limnology and Oceanography, 33(4 (part 2)):867–893, July 1988. CODEN LI-OCAH. ISSN 0024-3590.

Stoecker:1984:PPP

[Sto84] Diane K. Stoecker. Particle production by planktonic ciliates. Limnology and Oceanography, 29(5):930–940, September 1984. CO-DEN LIOCAH. ISSN 0024-3590.

Stoddard:1987:ADU

[Sto87] John L. Stoddard. Alkalinity dynamics in an unacidified alpine lake, Sierra Nevada, California. Limnology and Oceanography, 32 (4):825–839, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Stockner:1988:PPO

[Sto88a] John G. Stockner. Phototrophic picoplankton: an overview from marine and freshwater ecosystems. Limnology and Oceanography, 33(4 (part 2)):765–775, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Stoddard:1988:SLD

[Sto88b] John L. Stoddard. Are Sierran lakes different? A reply to the comments of Kelly and Schindler. Limnology and Oceanography, 33 (6 (part 2)):1641–1646, November 1988. CODEN LIOCAH. ISSN 0024-3590. See [Kel88, Sch88].

Stoecker:1989:BRC

[Sto89] Diane K. Stoecker. Book review: Comparative protozoology (O. R. Anderson). Limnology and Oceanography, 34(4):790–791, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Stross:1980:GCN

[Str80] R. G. Stross. Growth cycles and nutrient-limited photosynthesis in phytoplankton. Limnology and Oceanography, 49(6):538–544, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Stromgren:1983:TLG

[Str83] Tor Strömgren. Temperature-length growth strategies in the littoral alga Ascophyllum nodosum (L.). Limnology and Oceanography, 28(3):516–521, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Strayer:1988:LSP

[Str88] David Strayer. On the limits to secondary production. Limnology and Oceanography, 33(5):1217–1220, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Sullivan:1980:SFB

[Sul80] Barbara K. Sullivan. In situ feeding behavior of Sagitta elegans and Eukrohnia hamata (Chaetognatha) in relation to the vertical distribution and abundance of prey at Ocean Station “P”. Limnology and Oceanography, 25(2):317–326, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Sueur:1982:MMC

[SvdBR82] S. Sueur, C. M. G. van den Berg, and J. P. Riley. Measurement of the metal complexing ability of exudates of marine macroalgae.

Limnology and Oceanography, 27(3):536–543, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Schoonmaker:1989:LMO

[SVG89] J. S. Schoonmaker, K. J. Voss, and G. D. Gilbert. Laboratory measurements of optical beams in young sea ice. Limnology and Oceanography, 34(8):1606–1613, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Stewart:1980:FAR

[SW80] Arthur J. Stewart and Robert G. Wetzel. Fluorescence: absorbance ratios — a molecular-weight tracer of dissolved organic matter. Limnology and Oceanography, 49(6):559–564, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Stewart:1981:ARB

[SW81] Arthur J. Stewart and Robert G. Wetzel. Asymmetrical relationships between absorbance, fluorescence, and dissolved organic carbon. Limnology and Oceanography, 26(3):590–597, May 1981. CO-DEN LIOCAH. ISSN 0024-3590.

Sikes:1982:FCF

[SW82] C. Steven Sikes and Karl M. Wilbur. Functions of coccolith formation. Limnology and Oceanography, 27(1):18–26, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Stavn:1989:SFT

[SW89] Robert H. Stavn and Alan D. Weidemann. Shape factors, two-flow models, and the problem of irradiance inversion in estimating optical parameters. Limnology and Oceanography, 34(8):1426–1441, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Swanberg:1983:TRC

[Swa83] Neil R. Swanberg. The trophic role of colonial radiolaria in oligotrophic oceanic environments. Limnology and Oceanography, 28 (4):655–666, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Stoermer:1985:VMI

[SWSC85] E. F. Stoermer, J. A. Wolin, C. L. Schelske, and D. J. Conley. Variations in Melosira islandica valve morphology in Lake Ontario sediments related to eutrophication and silica depletion. Limnology and Oceanography, 30(2):414–418, March 1985. CODEN LI-OCAH. ISSN 0024-3590.

Spinrad:1989:RBA

[SYB^{+}89] R. W. Spinrad, C. M. Yentsch, J. Brown, Q. Dortch, E. Haugen, N. Revelante, and L. Shapiro. The response of beam attenuation to heterotrophic growth in a natural population of plankton. Limnology and Oceanography, 34(8):1601–1605, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Syzdek:1982:CSS

[Syz82] Lawrence D. Syzdek. Concentration of Serratia in the surface microlayer. Limnology and Oceanography, 27(1):172–177, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Taghon:1988:BCD

[Tag88] Gary L. Taghon. The benefits and costs of deposit feeding in the polychaete Abarenicola pacifica. Limnology and Oceanography, 33 (5):1166–1175, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Thode-Andersen:1989:SRF

[TAJ89] Søren Thode-Andersen and Bo Barker Jørgensen. Sulfate reduction and the formation of ^{35}S-labeled FeS, FeS_{2}, and S^{0} in coastal marine sediments. Limnology and Oceanography, 34(5):793–806, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Taylor:1988:APD

[Tay88] Barbara E. Taylor. Analyzing population dynamics of zooplankton. Limnology and Oceanography, 33(6):1266–1273, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Tuschall:1980:CON

[TB80] J. R. Tuschall, Jr. and P. L. Brezonik. Characterization of organic nitrogen in natural waters: Its molecular size, protein content, and interactions with heavy metals. Limnology and Oceanography, 49 (6):495–504, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Thorpe:1983:OSR

[TB83] S. A. Thorpe and J. M. Brubaker. Observations of sound reflection by temperature microstructure. Limnology and Oceanography, 28 (4):601–613, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Tyree:1984:MNP

[TB84] S. Y. Tyree, Jr. and Mary Ann O. Bynum. A method for nitrate and phosphate in saline water. Limnology and Oceanography, 29 (6):1337–1339, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Tramontano:1984:TRE

[TC84] J. M. Tramontano and T. M. Church. A technique for the removal of estuarine seston from Nuclepore filters. Limnology and Oceanography, 29(6):1339–1341, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Tenore:1985:EIR

[TC85] Kenneth R. Tenore and Edward J. Chesney, Jr. The effects of interaction of rate of food supply and population density on the bioenergetics of the opportunistic polychaete, Capitella capitata (type 1). Limnology and Oceanography, 30(6):1188–1195, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Turekian:1982:DSD

[TCN^{+}82] K. K. Turekian, J. K. Cochran, Y. Nozaki, I. Thompson, and D. S. Jones. Determination of shell deposition rates of Arctica islandica from the New York Bight using natural ^{228}Ra and ^{228}Th and bomb-produced ^{14}C. Limnology and Oceanography, 26(5): 737–741, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Tytler:1983:MIC

[TD83] E. M. Tytler and P. Spencer Davies. A method of isolating clean and viable zooxanthellae by density gradient centrifugation. Limnology and Oceanography, 28(6):1266–1268, November 1983. CO-DEN LIOCAH. ISSN 0024-3590.

Threlkeld:1987:EMS

[TD87] Stephen T. Threlkeld and Ray W. Drenner. An experimental mesocosm study of residual and contemporary effects of an omnivorous, filter-feeding, clupeid fish on plankton community structure. Limnology and Oceanography, 32(6):1331–1341, November 1987. CODEN LIOCAH. ISSN 0024-3590.

Tessier:1982:EFL

[TG82] Alan J. Tessier and Clyde E. Goulden. Estimating food limitation in cladoceran populations. Limnology and Oceanography, 26(5): 707–717, July 1982. CODEN LIOCAH. ISSN 0024-3590.

Tessier:1987:CJG

[TG87] Alan J. Tessier and Clyde E. Goulden. Cladoceran juvenile growth. Limnology and Oceanography, 32(3):680–686, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Tenore:1980:ADD

[TH80] Kenneth R. Tenore and Roger B. Hanson. Availability of detritus of different types and ages to a polychaete macroconsumer, Capitella capitata. Limnology and Oceanography, 49(6):553–558, May 1980. CODEN LIOCAH. ISSN 0024-3590.

Torgersen:1981:FGL

[THCP81] T. Torgersen, D. E. Hammond, W. B. Clarke, and T.-H. Peng. Fayetteville, Green Lake, New York: ^{3}H–^{3}He water mass ages and secondary chemical structure. Limnology and Oceanography, 26 (1):110–122, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Tessier:1983:SDE

[THGD83] Alan J. Tessier, Linda L. Henry, Clyde E. Goulden, and Mark W. Durand. Starvation in Daphnia: Energy reserves and reproductive allocation. Limnology and Oceanography, 28(4):667–676, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Thimann:1983:BRP

[Thi83] K. V. Thimann. Book review: Platt, T. [Ed.] 1981. Physiological bases of phytoplankton ecology. (Adv. Study Inst., NATO Sci. Affairs Div.) Can. Bull. Fish. Aquat. Sci. 210. 346 p. In Canada $17.95; elsewhere $21.55. Limnology and Oceanography, 28(1): 198–199, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Threlkeld:1981:RLT

[Thr81] Stephen T. Threlkeld. The recolonization of Lake Tahoe by Bosmina longirostris: Evaluating the importance of reduced Mysis relicta populations. Limnology and Oceanography, 26(3):433–444, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Threlkeld:1988:PPB

[Thr88] Stephen T. Threlkeld. Planktivory and planktivore biomass effects on zooplankton, phytoplankton, and the trophic cascade. Limnology and Oceanography, 33(6):1362–1375, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Tilzer:1983:IFL

[Til83] Max M. Tilzer. The importance of fractional light absorption by photosynthetic pigments for phytoplankton productivity in Lake Constance. Limnology and Oceanography, 28(5):833–846, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Tippett:1987:BRE

[Tip87] Roger Tippett. Book review: The ecology of freshwater phytoplankton (C. S. Reynolds). Limnology and Oceanography, 32(3): 779–780, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Thompson:1989:LLG

[TLH89] Peter A. Thompson, Maurice E. Levasseur, and Paul J. Harrison. Light-limited growth on ammonium vs. nitrate: What is the advantage for marine phytoplankton? Limnology and Oceanography, 34(6):1014–1024, September 1989. CODEN LIOCAH. ISSN 0024-3590.

Timms:1984:PGP

[TM84] R. M. Timms and Brian Moss. Prevention of growth of potentially dense phytoplankton populations by zooplankton grazing, in the presence of zooplanktivorous fish, in a shallow wetland ecosystem. Limnology and Oceanography, 29(3):472–486, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Tilman:1981:CNK

[TML81] David Tilman, Mark Mattson, and Sara Langer. Competition and nutrient kinetics along a temperature gradient: an experimental test of a mechanistic approach to niche theory. Limnology and Oceanography, 26(6):1020–1033, November 1981. CODEN LI-OCAH. ISSN 0024-3590.

Taylor:1983:IMP

[TML83] Craig D. Taylor, John J. Molongoski, and Steven E. Lohrenz. Instrumentation for the measurement of phytoplankton production. Limnology and Oceanography, 28(4):781–787, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Turner:1988:SVD

[TML^{+}88] Suzanne M. Turner, Gillian Malin, Peter S. Liss, Derek S. Harbour, and Patrick M. Holligan. The seasonal variation of dimethyl sulfide and dimethylsulfoniopropionate concentrations in nearshore waters. Limnology and Oceanography, 33(3):364–375, May 1988. CODEN LIOCAH. ISSN 0024-3590.

Tebo:1984:MMM

[TNEJ84] Bradley M. Tebo, Kenneth H. Nealson, Steven Emerson, and Lucinda Jacobs. Microbial mediation of Mn(II) and Co(II) precipitation at the O_{2} /H_{2}S interfaces in two anoxic fjords. Limnology and Oceanography, 29(6):1247–1258, November 1984. CODEN LI-OCAH. ISSN 0024-3590.

Taghon:1984:TBF

[TNJ84] Gary L. Taghon, Arthur R. M. Nowell, and Peter A. Jumars. Transport and breakdown of fecal pellets: Biological and sedimentological consequences. Limnology and Oceanography, 29(1): 64–72, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Top:1987:COB

[Top87] Zafer Top. Comment on “Oxygen budget of a perennially icecovered Antarctic lake”. Limnology and Oceanography, 32(2):520– 521, March 1987. CODEN LIOCAH. ISSN 0024-3590. See reply [WM87].

Thiriot-Quievreux:1986:SMD

[TQS86] C. Thiriot-Quievreux and J. Soyer. A simple method to detect bacterial associations in bivalve gills. Limnology and Oceanography, 31(6):1373–1376, November 1986. CODEN LIOCAH. ISSN 0024-3590.

Taylor:1981:EBD

[TS81a] Barbara E. Taylor and Montgomery Slatkin. Estimating birth and death rates of zooplankton. Limnology and Oceanography, 26(1): 143–158, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Tyler:1981:SRT

[TS81b] Mary Altalo Tyler and H. H. Seliger. Selection for a red tide organism: Physiological responses to the physical environment. Limnology and Oceanography, 26(2):310–324, March 1981. CO-DEN LIOCAH. ISSN 0024-3590.

Titus:1982:PRT

[TS82] John E. Titus and Walter H. Stone. Photosynthetic response of two submersed macrophytes to dissolved inorganic carbon concentration and pH. Limnology and Oceanography, 27(1):151–160, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Tunnicliffe:1983:CMB

[TS83] Verena Tunnicliffe and James P. M. Syvitski. Corals move boulders: an unusual mechanism of sediment transport. Limnology and Oceanography, 28(3):564–568, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Tande:1985:AEH

[TS85] Kurt S. Tande and Dag Slagstad. Assimilation efficiency in herbivorous aquatic organisms — the potential of the ratio method using ^{14}C and biogenic silica as markers. Limnology and Oceanography, 30(5):1093–1099, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Tranvik:1989:EFH

[TS89] Lars J. Tranvik and John McN. Sieburth. Effects of flocculated humic matter on free and attached pelagic microorganisms. Limnology and Oceanography, 34(4):688–699, June 1989. CODEN LI-OCAH. ISSN 0024-3590.

Turner:1988:MCP

[TTF88] Jefferson T. Turner, Patricia A. Tester, and Randolph L. Ferguson. The marine cladoceran Penilia avirostris and the “microbial loop” of pelagic food webs. Limnology and Oceanography, 33(2):245–255, March 1988. CODEN LIOCAH. ISSN 0024-3590.

Tutin:1980:BRM

[Tut80] Winifred Pennington Tutin. Book review: Matter, A. and M. E. Tucker [Eds.] 1978. Modern and ancient lake sediments. Spec. Publ. 2 Int. Assoc. Sediment. Blackwell Sci. Publ., Oxford, London, Edinburgh, Melbourne, vi + 290 p. $32.50. Limnology and Oceanography, 25(2):384–385, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Tipping:1983:SVC

[TW83] E. Tipping and C. Woof. Seasonal variations in the concentrations of humic substances in a soft-water lake. Limnology and Oceanography, 28(1):168–172, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Twombly:1983:SST

[Two83] Saran Twombly. Seasonal and short term fluctuations in zooplankton abundance in tropical Lake Malawi. Limnology and Oceanography, 28(6):1214–1224, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Ullman:1982:DCN

[UA82] William J. Ullman and Robert C. Aller. Diffusion coefficients in nearshore marine sediments. Limnology and Oceanography, 27(3): 552–556, May 1982. CODEN LIOCAH. ISSN 0024-3590.

Urban:1989:AGA

[UB89] Noel R. Urban and Larry A. Baker. Alkalinity generation associated with the formation of ester sulfates. Limnology and Oceanography, 34(6):1144–1146, September 1989. CODEN LIOCAH. ISSN 0024-3590. See reply [KR89].

Urban:1988:RID

[UEB88] N. R. Urban, S. J. Eisenreich, and S. E. Bayley. The relative importance of denitrification and nitrate assimilation in midcontinental bogs. Limnology and Oceanography, 33(6 (part 2)):1611–1617, November 1988. CODEN LIOCAH. ISSN 0024-3590.

Uncles:1981:NSV

[Unc81] R. J. Uncles. A numerical simulation of the vertical and horizontal M_{2} tide in the Bristol Channel and comparisons with observed data. Limnology and Oceanography, 26(3):571–577, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Rhee:1981:EEFb

[uRG81] G ull Rhee and Ivan J. Gotham. The effect of environmental factors on phytoplankton growth: Light and the interactions of light with nitrate limitation. Limnology and Oceanography, 26(4): 649–659, July 1981. CODEN LIOCAH. ISSN 0024-3590.

Vairavamurthy:1985:BDD

[VAI85] A. Vairavamurthy, M. O. Andreae, and R. L. Iverson. Biosynthesis of dimethylsulfide and dimethylpropiothetin by Hymenomonas carterae in relation to sulfur source and salinity variations. Limnology and Oceanography, 30(1):59–70, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Vanni:1986:CZC

[Van86] Michael J. Vanni. Competition in zooplankton communities: Suppression of small species by Daphnia pulex. Limnology and Oceanography, 31(5):1039–1056, September 1986. CODEN LI-OCAH. ISSN 0024-3590.

Vargo:1987:PCP

[VCG^{+}87] G. A. Vargo, K. L. Carder, W. Gregg, E. Shanley, C. Heil, K. A. Steidinger, and K. D. Haddad. The potential contribution of primary production by red tides to the west Florida shelf ecosystem. Limnology and Oceanography, 32(3):762–767, May 1987. CODEN LIOCAH. ISSN 0024-3590.

vandenBerg:1984:OCZ

[vdBD84] Constant M. G. van den Berg and Sirichai Dharmvanij. Organic complexation of zinc in estuarine interstitial and surface water samples. Limnology and Oceanography, 29(5):1025–1036, September 1984. CODEN LIOCAH. ISSN 0024-3590.

vanderLoeff:1984:ATA

[vdLAH^{+}84] Michiel M. Rutgers van der Loeff, Leif G. Anderson, Per O. J. Hall, Åke Iverfeldt, Alf B. Josefson, Bjørn Sundby, and Stig F. G. Westerlund. The asphyxiation technique: an approach to distinguishing between molecular diffusion and biologically mediated transport at the sediment–water interface. Limnology and Oceanography, 29 (4):675–686, July 1984. CODEN LIOCAH. ISSN 0024-3590.

Venrick:1984:WMV

[Ven84] E. L. Venrick. Winter mixing and the vertical stratification of phytoplankton — another look. Limnology and Oceanography, 29 (3):636–640, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Venrick:1987:FDF

[Ven87] Elizabeth L. Venrick. On fluorometric determinations of filterretained pigments. Limnology and Oceanography, 32(2):492–493, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Verity:1985:GRE

[Ver85] Peter G. Verity. Grazing, respiration, excretion, and growth rates of tintinnids. Limnology and Oceanography, 30(6):1268–1282, November 1985. CODEN LIOCAH. ISSN 0024-3590.

VanDover:1987:PHV

[VFB87] Cindy Lee Van Dover, Peter J. S. Franks, and Robert D. Ballard. Prediction of hydrothermal vent locations from distributions of brachyuran crabs. Limnology and Oceanography, 32(4):1006–1010, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Vincent:1980:EAH

[VG80] Warwick F. Vincent and Charles R. Goldman. Evidence for algal heterotrophy in Lake Tahoe, California-nevada. Limnology and Oceanography, 25(1):89–99, January 1980. CODEN LIOCAH. ISSN 0024-3590.

vanGemerden:1985:DCM

[vGMMG85] Hans van Gemerden, Emilio Montesinos, Jordi Mas, and Ricardo Guerrero. Diel cycle of metabolism of phototrophic purple sulfur bacteria in Lake Cisó (Spain). Limnology and Oceanography, 30 (5):932–943, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Vadstein:1989:COC

[VHJ^{+}89] Olav Vadstein, Bjørn Ola Harkjerr, Arne Jensen, Yngvar Olsen, and Helge Reinertsen. Cycling of organic carbon in the photic zone of a eutrophic lake with special reference to the heterotrophic bacteria. Limnology and Oceanography, 34(5):840–855, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Victor:1984:CRF

[Vic84] Benjamin C. Victor. Coral reef fish larvae: Patch size estimation and mixing in the plankton. Limnology and Oceanography, 29(5): 1116–1119, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Vadstein:1988:GPS

[VJOR88] Olav Vadstein, Arne Jensen, Yngvar Olsen, and Helge Reinertsen. Growth and phosphorus status of limnetic phytoplankton and bacteria. Limnology and Oceanography, 33(4):489–503, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Vuorinen:1989:DSF

[VKW89] Ilppo Vuorinen, Matti Ketola, and Mari Walls. Defensive spine formation in Daphnia pulex Leydig and induction by Chaoborus crystallinus De Geer. Limnology and Oceanography, 34(1):245– 248, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Vernet:1987:RAP

[VL87] Maria Vernet and Carl J. Lorenzen. The relative abundance of pheophorbide-a and pheophytin-a in temperate marine waters. Limnology and Oceanography, 32(2):352–358, March 1987. CO-DEN LIOCAH. ISSN 0024-3590.

Vertucci:1989:SRW

[VL89] Frank A. Vertucci and Gene E. Likens. Spectral reflectance and water quality of Adirondack mountain region lakes. Limnology and Oceanography, 34(8):1656–1672, December 1989. CODEN LI-OCAH. ISSN 0024-3590.

Vosjan:1987:MBR

[VN87] J. H. Vosjan and G. Nieuwland. Microbial biomass and respiratory activity in the surface waters of the east Banda Sea and northwest Arafura Sea (Indonesia) at the time of the southeast monsoon. Limnology and Oceanography, 32(3):767–775, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Vadstein:1989:CCP

[VO89] Olav Vadstein and Yngvar Olsen. Chemical composition and phosphate uptake kinetics of limnetic bacterial communities cultured in chemostats under phosphorus limitation. Limnology and Oceanography, 34(5):939–946, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Voss:1989:URD

[Vos89] K. J. Voss. Use of the radiance distribution to measure the optical absorption coefficient in the ocean. Limnology and Oceanography, 34(8):1614–1622, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Vanderploeg:1985:MAC

[VP85] Henry A. Vanderploeg and Gustav-Adolf Paffenhöfer. Modes of algal capture by the freshwater copepod Diaptomus sicilis and their relation to food-size selection. Limnology and Oceanography, 30(4):871–885, July 1985. CODEN LIOCAH. ISSN 0024-3590.

vanRijn:1983:BRN

[vRS83] Jaap van Rijn and Moshe Shilo. Buoyancy regulation in a natural population of oscillatoria spp. in fishponds. Limnology and Oceanography, 28(5):1034–1037, September 1983. CODEN LI-OCAH. ISSN 0024-3590.

vanRijn:1985:CFG

[vRS85] Jaap van Rijn and Moshe Shilo. Carbohydrate fluctuations, gas vacuolation, and vertical migration of scum-forming cyanobacteria in fishponds. Limnology and Oceanography, 30(6):1219–1228, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Valiela:1980:MTE

[VTV^{+}80] Ivan Valiela, J. M. Teal, S. B. Volkmann, C. M. Cogswell, and R. A. Harrington. On the measurement of tidal exchanges and groundwater flow in salt marshes. Limnology and Oceanography, 25(1):187–192, January 1980. CODEN LIOCAH. ISSN 0024-3590.

VanVleet:1980:SSS

[VW80] Edward S. Van Vleet and Peter M. Williams. Sampling sea surface films: a laboratory evaluation of techniques and collecting materials. Limnology and Oceanography, 25(4):764–770, July 1980. CODEN LIOCAH. ISSN 0024-3590.

VanVleet:1983:SPF

[VW83] Edward S. Van Vleet and Peter M. Williams. Surface potential and film pressure measurements in seawater systems. Limnology and Oceanography, 28(3):401–414, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Vincent:1984:SDN

[VWVR84] Warwick F. Vincent, Wayne Wurtsbaugh, Connie L. Vincent, and P. J. Richerson. Seasonal dynamics of nutrient limitation in a tropical high-altitude lake (Lake Titicaca, Peru–Bolivia): Application of physiological bioassays. Limnology and Oceanography, 29(3):540–552, May 1984. CODEN LIOCAH. ISSN 0024-3590.

Wolanski:1989:MAL

[WAI89] Eric Wolanski, Takashi Asaeda, and Jorg Imberger. Mixing across a lutocline. Limnology and Oceanography, 34(5):931–938, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Walker:1980:SMM

[Wal80] Terriss Walker. A simple maximum-minimum salinity sampler. Limnology and Oceanography, 25(6):1122–1123, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Walton:1985:FRR

[Wal85] William E. Walton. Factors regulating the reproductive phenology of Onychodiaptomus birgei (Copepoda: Calanoida). Limnology and Oceanography, 30(1):167–179, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Wangersky:1982:BRD

[Wan82] Peter J. Wangersky. Book review: Duursma, E. K., and R. Dawson [Eds.] 1981. Marine organic chemistry. Elsevier Oceanogr. Ser. No.

31. Elsevier Sci. Publ. Co., Amsterdam, Oxford, and New York. 521 p. $105.00. Limnology and Oceanography, 27(2):398, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Ward:1984:CAI

[War84] B. B. Ward. Combined autoradiography and immunofluorescence for estimation of single cell activity by ammonium-oxidizing bacteria. Limnology and Oceanography, 29(2):402–410, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Westrich:1984:RSO

[WB84] Joseph T. Westrich and Robert A. Berner. The role of sedimentary organic matter in bacterial sulfate reduction: The G model tested. Limnology and Oceanography, 29(2):236–249, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Weidemann:1986:ASC

[WB86a] A. D. Weidemann and T. T. Bannister. Absorption and scattering coefficients in Irondequoit Bay. Limnology and Oceanography, 31 (3):567–583, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1986:PRS

[WB86b] Craig E. Williamson and Nancy M. Butler. Predation on rotifers by the suspension-feeding calanoid copepod Diaptomus pallidus. Limnology and Oceanography, 31(2):393–402, March 1986. CO-DEN LIOCAH. ISSN 0024-3590.

Weidemann:1985:POP

[WBEJ85] Alan D. Weidemann, T. T. Bannister, Steven W. Effler, and David L. Johnson. Particulate and optical properties during CaCO_{3} precipitation in Otisco Lake. Limnology and Oceanography, 30(5):1078–1083, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1985:FLN

[WBF85] Craig E. Williamson, Nancy M. Butler, and Lorraine Forcina. Food limitation in naupliar and adult Diaptomus pallidus. Limnology and Oceanography, 30(6):1283–1290, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Webster:1987:ESP

[WBG^{+}87] J. R. Webster, E. F. Benfield, S. W. Golladay, B. H. Hill, L. E. Hornick, R. F. Kazmierczak, Jr., and W. B. Perry. Experimental studies of physical factors affecting seston transport in streams.

Limnology and Oceanography, 32(4):848–863, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Weber:1983:MSW

[WBS83] Martin E. Weber, Duncan C. Blanchard, and Lawrence D. Syzdek. The mechanism of scavenging of waterborne bacteria by a rising bubble. Limnology and Oceanography, 28(1):101–105, January 1983. CODEN LIOCAH. ISSN 0024-3590.

Worrest:1980:RPP

[WBV80] Robert C. Worrest, Deborah L. Brooker, and Henry Van Dyke. Results of a primary productivity study as affected by the type of glass in the culture bottles. Limnology and Oceanography, 25(2): 360–364, March 1980. CODEN LIOCAH. ISSN 0024-3590.

Wiegand:1981:STT

[WC81] Ronald C. Wiegand and Eddy C. Carmack. Some types of temperature inversion encountered in a freshwater lake with short residence time. Limnology and Oceanography, 26(3):565–571, May 1981. CODEN LIOCAH. ISSN 0024-3590.

Wang:1986:DGP

[WC86] Rong Wang and Robert J. Conover. Dynamics of gut pigment in the copepod Temora longicornis and the determination of in situ grazing rates. Limnology and Oceanography, 31(4):867–877, July 1986. CODEN LIOCAH. ISSN 0024-3590.

Wiegand:1987:IWS

[WC87] Ronald C. Wiegand and Vivian Chamberlain. Internal waves of the second vertical mode in a stratified lake. Limnology and Oceanography, 32(1):29–42, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Wright:1982:FLM

[WCEP82] Richard T. Wright, Richard B. Coffin, Curtis P. Ersing, and Daniel Pearson. Field and laboratory measurements of bivalve filtration of natural marine bacterioplankton. Limnology and Oceanography, 27(1):91–98, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Wilkerson:1987:ULS

[WD87] F. P. Wilkerson and R. C. Dugdale. The use of large shipboard barrels and drifters to study the effects of coastal upwelling on phytoplankton dynamics. Limnology and Oceanography, 32(2): 368–382, March 1987. CODEN LIOCAH. ISSN 0024-3590.

Wodka:1985:PDE

[WED85] Martin C. Wodka, Steven W. Effler, and Charles T. Driscoll. Phosphorus deposition from the epilimnion of Onondaga Lake. Limnology and Oceanography, 30(4):833–843, July 1985. CODEN LI-OCAH. ISSN 0024-3590.

Weiler:1980:PSS

[Wei80] C. S. Weiler. Population structure and in situ division rates of Ceratium in oligotrophic waters of the North Pacific central gyre. Limnology and Oceanography, 25(4):610–619, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Weider:1984:SHD

[Wei84] Lawrence J. Weider. Spatial heterogeneity of Daphnia genotypes: Vertical migration and habitat partitioning. Limnology and Oceanography, 29(2):225–235, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Welander:1988:BRS

[Wel88] Pierre Welander. Book review: Stommel, H. 1987. A view of the sea. Princeton University Press, Princeton, New Jersey. 165 p. $19.95. Limnology and Oceanography, 33(3):488, May 1988. CO-DEN LIOCAH. ISSN 0024-3590.

Wetzel:1982:BRE

[Wet82] Robert G. Wetzel. Book review: Estuarine and wetland processes with emphasis on modeling. 1980. Marine Science, v. 11. P. Hamilton and K. B. Macdonald, Eds. Plenum Publ. Corp., New York. 653 p. (typewriter composition). $69.50. Limnology and Oceanography, 27(2):395–396, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1983:BRB

[Wet83] Robert G. Wetzel. Book review: Barica, J., and L. R. Mur [Eds.]. 1980. Hypertrophic ecosystems. SIL Workshop on Hypertrophic Ecosystems. Developments in Hydrobiology. V. 2. W. Junk Publ., The Hague. 348 p. $87.00. Limnology and Oceanography, 28(5): 1042–1043, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1984:BRF

[Wet84] Robert G. Wetzel. Book review: Furtado, J. I., and S. Mori [Eds.]. 1982. Tasek Bera: The ecology of a freshwater swamp. Monogr. Biol., V. 47. W. Junk, The Hague. 412 p. $79.00. Limnology and Oceanography, 29(2):447–448, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1987:BRMa

[Wet87a] Robert G. Wetzel. Book review: The mangrove ecosystem: Research methods (E. C. Snedaker and J. G. Snedaker [eds.]). Limnology and Oceanography, 32(3):777–778, May 1987. CODEN LI-OCAH. ISSN 0024-3590.

Wetzel:1987:BRMb

[Wet87b] Robert G. Wetzel. Book review: The mangrove ecosystem: Research methods (E. C. Snedaker and J. G. Snedaker [eds.]). Limnology and Oceanography, 32(3):777–778, May 1987. CODEN LI-OCAH. ISSN 0024-3590.

Wheeler:1982:AUI

[WGM82] Patricia A. Wheeler, Patricia M. Glibert, and James J. McCarthy. Ammonium uptake and incorporation by Chesapeake Bay phytoplankton: Short term uptake kinetics. Limnology and Oceanography, 27(6):1113–1119, November 1982. CODEN LIOCAH. ISSN 0024-3590.

Waters:1980:APD

[WH80] Thomas F. Waters and Jay C. Hokenstrom. Annual production and drift of the stream amphipod Gammarus pseudolimnaeus in Valley Creek, Minnesota. Limnology and Oceanography, 25(4): 700–710, July 1980. CODEN LIOCAH. ISSN 0024-3590.

Weger:1989:RLL

[WHFT89] Harold G. Weger, Ronny Herzig, Paul G. Falkowski, and David H. Turpin. Respiratory losses in the light in a marine diatom: Measurements by short-term mass spectrometry. Limnology and Oceanography, 34(7):1153–1161, November 1989. CODEN LI-OCAH. ISSN 0024-3590.

White:1981:MZC

[Whi81] Alan W. White. Marine zooplankton can accumulate and retain dinoflagellate toxins and cause fish kills. Limnology and Oceanography, 26(1):103–109, January 1981. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1986:BRP

[Whi86] M. Whitfield. Book review: Pytkowicz, R. M. 1983. Equilibria, nonequilibria, and natural waters. Wiley-Interscience, New York.

Y. 1: 351 p. $49.95. V. 2: 353 p. $49.95. Limnology and Oceanography, 31(1):235, January 1986. CODEN LIOCAH. ISSN 0024-3590.

Whitmore:1989:FDA

[Whi89] Thomas J. Whitmore. Florida diatom assemblages as indicators of trophic state and pH. Limnology and Oceanography, 34(5):882– 895, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Wood:1985:DBT

[WHM^{+}85] A. Michelle Wood, P. K. Horan, K. Muirhead, D. A. Phinney, C. M. Yentsch, and J. B. Waterbury. Discrimination between types of pigments in marine Synechococcus spp. by scanning spectroscopy, epifluorescence microscopy, and flow cytometry. Limnology and Oceanography, 30(6):1303–1315, November 1985. CODEN LIOCAH. ISSN 0024-3590.

Williams:1980:SRB

[Wil80a] D. Dudley Williams. Some relationships between stream benthos and substrate heterogeneity. Limnology and Oceanography, 25(1): 166–172, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Williams:1980:DDT

[Wil80b] Ernest H. Williams. Disjunct distributions of two aquatic predators. Limnology and Oceanography, 25(6):999–1006, November 1980. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1980:PBM

[Wil80c] Craig E. Williamson. The predatory behavior of Mesocyclops edax: Predator preferences, prey defenses, and starvation-induced changes. Limnology and Oceanography, 25(5):903–909, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Williams:1984:USA

[Wil84] Susan Lynn Williams. Uptake of sediment ammonium and translocation in a marine green macroalga Caulerpa cupressoides. Limnology and Oceanography, 29(2):374–379, March 1984. CODEN LIOCAH. ISSN 0024-3590.

Williamson:1987:PPI

[Wil87] Craig E. Williamson. Predator-prey interactions between omnivorous diaptomid copepods and rotifers: The role of prey morphology and behavior. Limnology and Oceanography, 32(1):167–177, January 1987. CODEN LIOCAH. ISSN 0024-3590.

Winter:1981:EWT

[Win81] Thomas C. Winter. Effects of water-table configuration on seepage through lakebeds. Limnology and Oceanography, 26(5):925–934, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Wyatt:1989:DPL

[WJ89] Philip J. Wyatt and Christian Jackson. Discrimination of phytoplankton via light-scattering properties. Limnology and Oceanography, 34(1):96–112, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Wetzel:1980:BRD

[WK80] Robert G. Wetzel and W. K. Kellogg. Book review: Dykyjova, D., and J. Květ [Eds.]. 1978. Pond littoral ecosystems. Ecological Studies 28. Springer-Verlag, New York, Heidelberg, London. xiv + 464 p. $49.00. Limnology and Oceanography, 25(1):197, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:1986:UIO

[WK86a] Patricia A. Wheeler and David L. Kirchman. Utilization of inorganic and organic nitrogen by bacteria in marine systems. Limnology and Oceanography, 31(5):998–1009, September 1986. CODEN LIOCAH. ISSN 0024-3590.

Winn:1986:DNA

[WK86b] C. D. Winn and D. M. Karl. Diel nucleic acid synthesis and particulate DNA concentrations: Conflicts with division rate estimates by DNA accumulation. Limnology and Oceanography, 31(3):637– 645, May 1986. CODEN LIOCAH. ISSN 0024-3590.

Walls:1989:EPI

[WK89] Mari Walls and Matti Ketola. Effects of predator-induced spines on individual fitness in Daphnia pulex. Limnology and Oceanography, 34(2):390–396, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:1989:DPA

[WKLK89] Patricia A. Wheeler, David L. Kirchman, Michael R. Landry, and Steven A. Kokkinakis. Diel periodicity in ammonium uptake and regeneration in the oceanic subarctic Pacific: Implications for interactions in microbial food webs. Limnology and Oceanography, 34(6):1025–1033, September 1989. CODEN LIOCAH. ISSN 0024-

Ward:1989:BNC

[WKRE89] B. B. Ward, K. A. Kilpatrick, E. H. Renger, and R. W. Eppley. Biological nitrogen cycling in the nitracline. Limnology and Oceanography, 34(3):493–513, May 1989. CODEN LIOCAH. ISSN 0024-3590.

Walsh:1983:PNP

[WL83] P. Walsh and L. Legendre. Photosynthesis of natural phytoplankton under high frequency light fluctuations simulating those induced by sea surface waves. Limnology and Oceanography, 28(4): 688–697, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Welschmeyer:1984:CLP

[WL84] Nicholas A. Welschmeyer and Carl J. Lorenzen. Carbon-14 labeling of phytoplankton carbon and chlorophyll-a carbon: Determination of specific growth rates. Limnology and Oceanography, 29 (1):135–145, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Welschmeyer:1985:CBZ

[WL85a] Nicholas A. Welschmeyer and Carl J. Lorenzen. Chlorophyll budgets: Zooplankton grazing and phytoplankton growth in a temperate fjord and the Central Pacific Gyres. Limnology and Oceanography, 30(1):1–21, January 1985. CODEN LIOCAH. ISSN 0024-3590.

Wurtsbaugh:1985:DMZ

[WL85b] Wayne Wurtsbaugh and Hiram Li. Diel migrations of a zooplanktivorous fish (Menidia beryllina) in relation to the distribution of its prey in a large eutrophic lake. Limnology and Oceanography, 30(3):565–576, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Wafar:1984:SCD

[WLB84] Mohideen Wafar, Pierre Le Corre, and Jean-Louis Birrien. Seasonal changes of dissolved organic matter (C, N, P) in permanently well mixed temperate waters. Limnology and Oceanography, 29(5): 1127–1132, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Wieder:1985:EWC

[WLG85] R. Kelman Wieder, Gerald E. Lang, and Valerie A. Granus. An evaluation of wet chemical methods for quantifying sulfur fractions in freshwater wetland peat. Limnology and Oceanography, 30(5): 1109–1115, September 1985. CODEN LIOCAH. ISSN 0024-3590.

Winter:1988:DUH

[WLR88] T. C. Winter, J. W. LaBaugh, and D. O. Rosenberry. The design and use of a hydraulic potentiomanometer for direct measurement of differences in hydraulic head between groundwater and surface water. Limnology and Oceanography, 33(5):1209–1214, September 1988. CODEN LIOCAH. ISSN 0024-3590.

Wheeler:1982:MUC

[WM82] Patricia A. Wheeler and James J. McCarthy. Methylammonium uptake by Chesapeake Bay phytoplankton: Evaluation of the use of the ammonium analogue for field uptake measurements. Limnology and Oceanography, 27(6):1120–1140, November 1982. CO-DEN LIOCAH. ISSN 0024-3590.

Wharton:1987:RCO

[WM87] Robert A. Wharton, Jr. and Christopher P. McKay. Reply to comment on “Oxygen budget of a perennially ice-covered Antarctic lake”. Limnology and Oceanography, 32(2):521, March 1987. CODEN LIOCAH. ISSN 0024-3590. See [Top87].

Wilkerson:1983:TPC

[WMM83] F. P. Wilkerson, G. Muller, and Parker L. Muscatine. Temporal patterns of cell division in natural populations of endosymbiotic algae. Limnology and Oceanography, 28(5):1009–1014, September 1983. CODEN LIOCAH. ISSN 0024-3590.

Wharton:1986:OBP

[WMSP86] Robert A. Wharton, Jr., Christopher P. McKay, George M. Simmons, Jr., and Bruce C. Parker. Oxygen budget of a perennially ice-covered Antarctic lake. Limnology and Oceanography, 31(2): 437–443, March 1986. CODEN LIOCAH. ISSN 0024-3590. See [Top87].

Whiting:1989:NEB

[WMSW89] Gary J. Whiting, Henry N. McKellar, Jr., John D. Spurrier, and Thomas G. Wolaver. Nitrogen exchange between a portion of vegetated salt marsh and the adjoining creek. Limnology and Oceanography, 34(2):463–473, March 1989. CODEN LIOCAH. ISSN 0024-3590.

Wofsy:1983:SMP

[Wof83] S. C. Wofsy. A simple model to predict extinction coefficients and phytoplankton biomass in eutrophic waters. Limnology and Oceanography, 28(6):1144–1155, November 1983. CODEN LI-OCAH. ISSN 0024-3590.

Wolanski:1987:SEB

[Wol87] Eric Wolanski. Some evidence for boundary mixing near coral reefs. Limnology and Oceanography, 32(3):735–739, May 1987. CODEN LIOCAH. ISSN 0024-3590.

Wood:1980:BRK

[Woo80] R. Brian Wood. Book review: Kalk, M., A. J. McLachlan, and C. Howard-Williams [Eds.] 1979. Lake Chilwa: Studies of change in a tropical ecosystem. Dr. W. Junk Publ., The Hague, xviii + 462 p. $78.95. Limnology and Oceanography, 25(5):964–965, September 1980. CODEN LIOCAH. ISSN 0024-3590.

Woodby:1984:ADM

[Woo84] Douglas A. Woodby. The April distribution of murres and prey patches in the southeastern Bering Sea. Limnology and Oceanography, 29(1):181–188, January 1984. CODEN LIOCAH. ISSN 0024-3590.

Wood:1985:BRS

[Woo85] R. Brian Wood. Book review: Serruya, C., and U. Pollingher. 1983. Lakes of the Warm Belt. Cambridge Univ. Press, New York. 569 p. $89.50. Limnology and Oceanography, 30(4):916, July 1985. CODEN LIOCAH. ISSN 0024-3590.

Weilenmann:1989:PTL

[WOS89] Ulrich Weilenmann, Charles R. O’Melia, and Werner Stumm. Particle transport in lakes: Models and measurements. Limnology and Oceanography, 34(1):1–18, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Wotton:1982:DSF

[Wot82] R. S. Wotton. Does the surface film of lakes provide a source of food for animals living in lake outlets? Limnology and Oceanography, 27(5):959–960, September 1982. CODEN LIOCAH. ISSN 0024-3590.

Wright:1989:CEI

[Wri89] L. D. Wright. Coastal environments: an introduction to the physical, ecological and cultural systems of coastlines. Limnology and Oceanography, 34(4):790, June 1989. CODEN LIOCAH. ISSN 0024-3590.

Wroblewski:1989:MSB

[Wro89] J. S. Wroblewski. A model of the spring bloom in the North Atlantic and its impact on ocean optics. Limnology and Oceanography, 34(8):1563–1571, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Welch:1980:MAA

[WRS80] H. E. Welch, J. W. M. Rudd, and D. W. Schindler. Methane addition to an Arctic lake in winter. Limnology and Oceanography, 25(1):100–113, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Whitfield:1981:HRB

[WS81] Paul H. Whitfield and Hanspeter Schreier. Hysteresis in relationships between discharge and water chemistry in the Fraser River basin, British Columbia. Limnology and Oceanography, 26(6): 1179–1182, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Williams:1980:ASQ

[WST80] J. D. H. Williams, H. Shear, and R. L. Thomas. Availability to Scenedesmus quadricauda of different forms of phosphorus in sedimentary materials from the Great Lakes. Limnology and Oceanography, 25(1):1–11, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Weliky:1983:PAC

[WSU^{+}83] Karen Weliky, Erwin Suess, C. Andre Ungerer, Peter J. Muller, and Kathy Fischer. Problems with accurate carbon measurements in marine sediments and particulate matter in seawater: a new approach. Limnology and Oceanography, 28(6):1252–1259, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Woltemate:1984:CHI

[WWS84] I. Woltemate, M. J. Whiticar, and M. Schoell. Carbon and hydrogen isotopic composition of bacterial methane in a shallow freshwater lake. Limnology and Oceanography, 29(5):985–992, September 1984. CODEN LIOCAH. ISSN 0024-3590.

Yoder:1981:RGS

[YAL^{+}81] James A. Yoder, Larry P. Atkinson, Thomas N. Lee, Hongsuk H. Kim, and Charles R. McClain. Role of Gulf Stream frontal eddies in forming phytoplankton patches on the outer southeastern shelf. Limnology and Oceanography, 26(6):1103–1110, November 1981. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1983:NFC

[Yen83] Charles S. Yentsch. A note on the fluorescence characteristics of particles that pkss through glass-fiber filters. Limnology and Oceanography, 28(3):597–599, May 1983. CODEN LIOCAH. ISSN 0024-3590.

Yen:1985:SPC

[Yen85] J. Yen. Selective predation by the carnivorous marine copepod Euchaeta elongata: Laboratory measurements of predation rates verified by field observations of temporal and spatial feeding patterns. Limnology and Oceanography, 30(3):577–597, May 1985. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1983:FCC

[YHM^{+}83] Clarice M. Yentsch, Paul K. Horan, Katharine Muirhead, Quay Dortch, Elin Haugen, Louis Legendre, Lynda S. Murphy, Mary Jane Perry, David A. Phinney, Shirley A. Pomponi, Richard W. Spinrad, Michelle Wood, Charles S. Yentsch, and Bernard J. Zahuranec. Flow cytometry and cell sorting: a technique for analysis and sorting of aquatic particles. Limnology and Oceanography, 28(6):1275–1280, November 1983. CODEN LIOCAH. ISSN 0024-3590.

Yin:1984:IPA

[YJ84] Chengqing Yin and David L. Johnson. An individual particle analysis and budget study of Onondaga Lake sediments. Limnology and Oceanography, 29(6):1193–1201, November 1984. CODEN LI-OCAH. ISSN 0024-3590.

Yoder:1987:SSC

[YMBO87] James A. Yoder, Charles R. McClain, Jackson O. Blanton, and Lie-Yauw Oeymay. Spatial scales in CZCS-chlorophyll imagery of the southeastern U.S. continental shelf. Limnology and Oceanography, 32(4):929–941, July 1987. CODEN LIOCAH. ISSN 0024-3590.

Yoder:1982:CDP

[YMN82] James A. Yoder, James Martin, and Andrew Nill. Cell division periodicity and the nitrate environment of a marine diatom. Limnology and Oceanography, 27(2):352–357, March 1982. CODEN LIOCAH. ISSN 0024-3590.

Yentsch:1989:BBO

[YP89] Charles S. Yentsch and David A. Phinney. A bridge between ocean optics and microbial ecology. Limnology and Oceanography, 34(8): 1694–1705, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Young:1984:VFD

[YTW84] Stephen Young, Victoria A. Taylor, and Elaine Watts. Visual factors in Daphnia feeding. Limnology and Oceanography, 29(6): 1300–1308, November 1984. CODEN LIOCAH. ISSN 0024-3590.

Zaneveld:1989:ACT

[Zan89] J. Ronald V. Zaneveld. An asymptotic closure theory for irradiance in the sea and its inversion to obtain the inherent optical properties. Limnology and Oceanography, 34(8):1442–1452, December 1989. CODEN LIOCAH. ISSN 0024-3590.

Zaret:1981:LCV

[Zar81] Robert E. Zaret. Lateral compression and visibility in cladocerans. Limnology and Oceanography, 27(4):965–970, September 1981. CODEN LIOCAH. ISSN 0024-3590.

Zevenboom:1982:AFL

[ZBM82] Wanda Zevenboom, Abraham Bij de Vaate, and Luuc R. Mur. Assessment of factors limiting growth rate of Oscillatoria agardhii in hypertrophic Lake Wolderwijd, 1978, by use of physiological indicators. Limnology and Oceanography, 27(1):39–52, January 1982. CODEN LIOCAH. ISSN 0024-3590.

Zlotnik:1989:ELT

[ZD89] I. Zlotnik and Z. Dubinsky. The effect of light and temperature on DOC excretion by phytoplankton. Limnology and Oceanography, 34(5):831–839, July 1989. CODEN LIOCAH. ISSN 0024-3590.

Zuosheng:1989:HDU

[ZEY89] Yang Zuosheng, K. O. Emery, and Xui Yuix. Historical development and use of thousand-year-old tide-prediction tables. Limnology and Oceanography, 34(5):953–957, July 1989. CODEN LI-OCAH. ISSN 0024-3590.

Zimmer-Faust:1989:RBC

[ZF89] Richard K. Zimmer-Faust. The relationship between chemoreception and foraging behavior in crustaceans. Limnology and Oceanography, 34(7):1367–1374, November 1989. CODEN LI-OCAH. ISSN 0024-3590.

Zehr:1988:CBA

[ZFFC88] Jonathan P. Zehr, Paul G. Falkowski, Joanna Fowler, and Douglas G. Capone. Coupling between ammonium uptake and incorporation in a marine diatom: Experiments with the short-lived radioisotope ^{13}N. Limnology and Oceanography, 33(4):518–527, July 1988. CODEN LIOCAH. ISSN 0024-3590.

Zehr:1987:BSL

[ZHO^{+}87] Jon P. Zehr, Ronald W. Harvey, Ronald S. Oremland, James E. Cloern, Leah H. George, and Judith L. Lane. Big Soda Lake (Nevada). 1. Pelagic bacterial heterotrophy and biomass. Limnology and Oceanography, 32(4):781–793, July 1987. CODEN LI-OCAH. ISSN 0024-3590.

Zaret:1980:SST

[ZK80] Robert E. Zaret and W. Charles Kerfoot. The shape and swimming technique of Bosmina longirostris. Limnology and Oceanography, 25(1):126–133, January 1980. CODEN LIOCAH. ISSN 0024-3590.

Zimmerman:1987:ANU

[ZKD87] Richard C. Zimmerman, James N. Kremer, and Richard C. Dugdale. Acceleration of nutrient uptake by phytoplankton in a coastal upwelling ecosystem: a modeling analysis. Limnology and Oceanography, 32(2):359–367, March 1987. CODEN LIOCAH. ISSN 0024-3590.

ZoBell:1980:BRB

[ZoB80] Claude E. ZoBell. Book review: Bitton, G., and K. C. Marshall [Eds.] 1980. Adsorption of microorganisms to surfaces. John Wiley & Sons, Inc., New York, xi + 439 p. $32.50. Limnology and Oceanography, 25(5):969–970, September 1980. CODEN LI-OCAH. ISSN 0024-3590.

Zevenboom:1983:NIM

[ZPVM83] Wanda Zevenboom, Anton F. Post, Udo M. Van Hes, and Luuc R. Mur. A new incubator for measuring photosynthetic activity of phototrophic microorganisms, using the amperometric oxygen method. Limnology and Oceanography, 28(4):787–791, July 1983. CODEN LIOCAH. ISSN 0024-3590.

Zimmerman:1985:ENN

[ZR85] Richard C. Zimmerman and Deborah L. Robertson. Effects of El Niño on local hydrography and growth of the giant kelp, Macrocystis pyrifera, at Santa Catalina Island, California. Limnology and Oceanography, 30(6):1298–1302, November 1985. CODEN LI-OCAH. ISSN 0024-3590.

Zamer:1989:PME

[ZST89] William E. Zamer, J. Malcolm Shick, and David W. Tapley. Protein measurement and energetic considerations: Comparisons of biochemical and stoichiometric methods using bovine serum albumin and protein isolated from sea anemones. Limnology and Oceanography, 34(1):256–263, January 1989. CODEN LIOCAH. ISSN 0024-3590.

Zullig:1981:UCS

[Zül81] Hans Züllig. On the use of carotenoid stratigraphy in lake sediments for detecting past developments of phytoplankton. Limnology and Oceanography, 27(4):970–976, September 1981. CODEN LIOCAH. ISSN 0024-3590.